A Complete Bibliography of Publications in Synthese, 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

07 October 2025 Version 1.18

Title word cross-reference

+ [774]. − [774]. = [774]. ∃ [2577]. ∃□ [2427]. ∀ [2577]. Ω^{−} [2655]. ε [2175].

-bundled [2427]. -operator [2175].

17th [2803]. 19 [2220, 2850]. 1980-2018 [1444].

3D [3050, 3350]. 3D/4D [3050, 3350].

4D [3050, 2313, 3350]. 4E [1596, 3037, 3249].

= [1294].

A-theory [1902]. Abducting [2373]. Abduction [379, 774, 654, 464, 2142, 1985]. Abductive [16, 1849, 129, 2924]. abductivism [2924, 1313]. abilities [702, 811]. ability [148, 3239, 356, 2875]. able [3029]. ABM [1800]. Absence [193, 1340, 2065]. absences [238]. absolute [1429, 368, 703]. absolutely [1961]. absolutism [1343]. Abstract [578, 565, 535, 1018, 2953, 1551, 3034]. abstracta [2273]. Abstraction [2175, 376, 2026, 1863, 559, 560, 558, 2497, 1040]. abstracts [1789]. absurdity [159]. abuse [2334]. academia [1118]. academic [2699]. acceptability [701]. Acceptable [23]. acceptance [2768]. accepting [125]. Access [2310]. Accessibility [351, 983]. accidental [992, 819]. accommodate [3004, 1563]. accommodation [3164, 2711, 2360, 867]. Account [2562, 3197, 1948, 2277, 3140, 2009, 2947, 910, 1566, 86, 1666, 3239, 2524, 174, 677, 1835, 1657, 3042, 3300, 3224, 1223, 2148, 1663, 2214, 1527, 2200, 1341, 2020, 926, 1613, 2008, 134, 2765, 1901, 18, 2835, 2706, 2026, 3274, 3134, 3181, 420, 1982, 1212, 2327, 2602, 1971, 2551, 414, 201, 815, 2371, 454, 2599, 2826, 1004, 1050, 2069, 671, 796, 2561, 818, 1668, 173, 302, 2443, 1638, 2460, 854, 2830, 333, 1066, 2360, 2058]. Accounting [1521, 2489, 3012]. accounts [1371, 3160, 2693, 1759, 956, 3326, 2118, 525, 2283, 760, 743, 2733, 567, 2042]. Accuracy [2316, 765, 2223, 3255, 2804, 1580, 2457, 1942, 3399]. accurate [1953, 1531, 3286]. Achievements [3081, 2365]. Achilles [2829]. Acknowledgments [1983]. Acknowledgments-based [1983]. Acquaintance [1775, 3102, 3052, 1504, 1907, 476]. acquisition [707]. Across [692, 40, 2088, 3025, 3120]. act [2934, 1071, 410]. Acting [2258, 1420, 1220, 2202, 1861]. Action [2077, 276, 1774, 196, 421, 3156, 469, 2099, 117, 3020, 391, 300, 467, 735, 2554, 2991, 1364, 2235, 2489, 357, 1087, 2027, 2705, 3016, 1196, 1788, 2087, 355, 8, 855, 720, 476, 1592, 477, 2202, 478, 1520, 2724, 810, 479, 172, 1218, 473, 2058]. Action-based [276]. Action-Directed [2077]. action-directed-pragmatics [357]. action-guiding [479]. Actions [2957, 470, 67, 925, 2246, 627, 1718]. Active [1719, 1860, 2713, 1255, 225, 3210, 3177, 2211, 1940]. actively [404]. activist [616]. activities [2765, 2044, 2980]. activity [1600]. acts [3076, 3082, 2716, 2738, 2767, 3269, 1315, 1035]. actual [2372, 113]. Actualism [1625]. actually [1939, 738, 2752]. acute [1517]. Ad [2395, 2360, 2470]. Adams [2480]. adapted [3179]. adaptive [1564, 846]. added [1102]. Adding [3375, 540]. addressee [2418]. Addressing [2214, 1792]. adequacy [1015, 652, 2895]. Adjudicating [857]. adjust [1645]. admissibility [2290, 1811]. ado [2869]. adult [3298]. advance [1055]. advanced [232]. advantage [1904, 2184, 1931]. advantages [2607]. adverbial [11, 199, 74]. adverbialism [1854]. adverbialist [619]. adversarial [1038]. advice [1953, 2094, 2715]. Aesthetic [881, 1906, 3276, 2848]. Aesthetics [391]. afar [2377]. Affect [228, 2889, 1087, 104]. Affective [744, 1681, 2681, 1066, 3172, 157, 1532, 897]. affective-evaluative [157]. Affectivism [2053]. afford [280]. affordance [480, 280]. Affordances [1709, 2836, 1517, 472, 1248, 2978, 388, 3320, 1401, 3001]. afoul [3327]. After [937, 3097, 689]. After-effects [937]. again [1911, 1417, 231, 3252, 682].

Against [254, 1825, 989, 1565, 1142, 983, 1144, 289, 3359, 3090, 2247, 1343, 2397, 1951, 3000, 743, 3198, 2447, 422, 2917, 347, 511, 3349, 2969, 215, 970, 780, 753, 2182, 1812, 6, 235, 2487, 1069, 635, 2304, 2198, 2119, 1394, 2332, 1961, 1325, 1011, 1380, 3083, 1853, 3240, 8, 855, 1118, 381, 3395, 686, 2002, 1572, 325, 2483]. age [503, 881]. Agency [2033, 3140, 905, 1566, 256, 2832, 1918, 3289, 2612, 1509, 2489, 2770, 432, 1867, 2231, 1853, 587, 2431, 2530, 1479, 1482, 1481, 2129, 267, 978, 2389, 2481]. agenda [406]. Agent [2399, 3154, 66, 2435, 2531, 2442, 551]. Agent-centered [2399, 3154]. Agential [1748, 2525]. Agentive [3381]. agents [1542, 2270, 2782, 806, 2062, 2033, 591, 2151, 2210]. Aggregating [2151]. aggregation [2741, 271]. aggregative [2350]. AGI [3370]. agnosticism [2277, 3326]. agree [572]. agreement [1913, 3141, 109]. Aharonov [1127]. AI [2023, 2242, 2667, 3172, 2068, 682, 3347, 583, 2513, 3189, 2909, 1482, 1481, 3370]. aims [792, 2420, 3000, 1686]. ain’t [1627]. akratic [1770]. Alan [1, 2361]. Alaskan [737]. aletheic [309]. Alethic [1166, 1839, 1283, 2994, 2659, 2165, 2631]. algebra [2096, 2013, 1813]. algebraic [332]. algebraic-topological [332]. Algorithmic [1030, 2405, 2557, 2510, 3394]. Alia} [2395]. aliefs [967]. aligned [3370]. all-purpose [2978, 475]. allostasis [3001]. allostatic [3190]. allowed [3343]. allowing [3334]. almost [2755]. alone [1420]. along [2872]. Alpha [2250, 1645]. alter [3239]. alternative [1141, 2990, 1540, 152, 1638]. Alternatives [2976, 1184, 2417, 395]. altruism [2868, 1274]. always [1564, 2764]. am [2337, 3034]. Ambient [3001, 3237, 3190]. ambiguities [2555]. Ambiguity [865, 2751, 1106, 1592, 3296, 2133]. ambitious [501]. ameliorative [57, 2113, 1697]. amend [1286]. amid [2059]. among [2915, 1365]. amounts [2495]. ampliative [2289, 322]. analogical [3274]. analogies [155, 1788]. analogue [460]. analogues [2546]. analogy [2358, 2296, 1409, 1053, 1814, 2218, 2048, 2369]. analyses [565, 554, 3339]. Analysis [620, 1089, 1068, 1510, 1452, 1073, 1956, 497, 752, 1236, 50, 1926, 812, 1078, 231, 349, 1669, 803, 1726, 1995, 2803, 3064, 432, 398, 1102, 2490, 1815, 57, 2949, 3145, 1390, 1431, 1950, 2216, 1153, 1530, 1801, 51, 2463, 428, 1039]. analytic [1073, 482, 597, 63, 1403, 1976, 443, 1983]. analyzing [2597, 2516, 3091]. anarchist [608]. Anatomy [2873]. anchoring [3090]. Angst [449, 457]. animal [2834, 2889, 2017, 2874, 2033, 2795, 1780, 2432, 2311]. animalism [2677]. animals [1357, 3383, 1687]. annihilated [904]. anomalous [1615]. anorexia [3116]. Anscombe [2033, 1853]. Anscombean [2231, 2058]. answer [2654]. Answering [2402, 2789]. ant [656]. Ante [2199, 856]. antecedent [1191, 917]. antecedent-contextualism [1191]. anthropology [1015, 1954]. Anti [1712, 131, 1993, 2366, 1062, 1540, 1927, 2337, 147, 1195, 1890, 1520, 2918, 1974, 2699, 1891, 3256, 1988, 2550, 1672, 3288, 2200, 2198, 2119, 1723, 1555, 1414, 3161, 1882, 145, 1443]. Anti-Archimedeanism [1890]. anti-collapse [3256]. Anti-essentialism [1540]. Anti-exceptionalism [1712, 1993, 2366, 1927, 1974, 1988, 3288, 1723, 1555]. anti-expertise [1414]. Anti-Haecceitist [2337]. Anti-individualism [131]. anti-individualist [2200]. Anti-intellectualism [1520, 2198, 2119]. Anti-intellectualist [1062]. Anti-luck [1195]. Anti-luminosity [2918]. anti-physicalism [1891]. anti-psychologism [1882]. anti-realism [2918, 1443]. anti-realists [2550]. anti-reductionist [2200, 3161, 145]. anti-representationalism [1672]. Anti-risk [147]. anti-skepticism [2699]. anticipating [1128]. anticipation [1, 2361, 605]. antirealism [2320]. antisatisfaction [955]. Antoine [1241]. Anxiety [2729, 1751, 1548, 1756, 2103, 897, 2106]. anxious [2003]. any [307]. Anything [1186]. apart [2688]. Aphantasia [2335]. apocalypse [1622, 2407]. apparent [2544]. appearance [2914]. applicability [1618, 3356]. application [3195, 1328, 1310, 860, 2504, 423]. application-introduction [3195]. applications [800, 2424, 3260]. applied [2094, 1352, 2152]. apply [2909]. applying [1402, 3351]. Appraisal [1133, 526]. Appreciating [3299, 1823]. Apprehending [2106]. Approach [1929, 1346, 1463, 1097, 2676, 1158, 2834, 1947, 2415, 315, 2303, 425, 1517, 2686, 2878, 933, 1997, 877, 861, 1812, 2559, 2459, 2800, 3162, 294, 3307, 3402, 2155, 2180, 1377, 2118, 2763, 1887, 3038, 3000, 789, 1299, 1532, 1473, 2626, 856, 2411, 1034, 3268, 541, 1021, 2628, 2573, 546, 3188, 2102, 2577, 1161, 533, 532, 1300, 3249, 178, 1570, 2181, 292, 2907, 561]. approaches [1956, 1696, 2244, 96, 492, 1791, 848, 2408, 1513, 283, 1150, 1239]. Approaching [1666, 1523, 1624, 1892]. appropriate [2952, 1064]. appropriateness [1090]. Approval [587, 1612]. Approval-directed [587]. approximate [2359, 800, 2509, 2804, 3054]. approximation [1353, 1949, 85, 3030]. Approximations [3166]. apriorism [443, 1617]. apt [1406]. Aptness [2352, 2484]. Arational [872, 2413]. Arbitrary [2640]. Arbitrating [2281]. Archaeoastronomy [2218]. Archimedeanism [1890]. architectural [1203]. architecture [945, 3263]. arena [1889]. aren’t [2232, 3350]. Argle [2236]. argue [136]. Arguing [2521, 2395, 983]. Argument [724, 3092, 1616, 1543, 1244, 1290, 215, 970, 2395, 1856, 425, 922, 2832, 780, 175, 82, 2182, 2984, 2385, 2487, 3196, 1656, 2020, 2955, 1824, 635, 2278, 1053, 2247, 1394, 1961, 1049, 3129, 2448, 1235, 3083, 2893, 694, 3072, 8, 855, 1118, 2936, 381, 885, 3292, 2057, 686, 2909, 2644, 675, 1520, 2727, 2352, 2484, 1841, 1207, 3272, 2837]. Argumentation [1913, 2074, 3202, 1551]. Argumentative [1179, 1889]. Arguments [1851, 215, 970, 1082, 3044, 2720, 2684, 693, 592, 3246, 1637, 2472, 3240, 2599, 717, 2457, 431, 3227, 2032, 2576, 858, 325]. Argumentum [2395]. Aristotelian [780, 2771, 598, 1872, 3220, 1490, 2275, 2156, 413, 2807, 3268, 412]. Arithmetic [1571, 2768]. arithmetical [3372, 187, 192]. arm [1773, 1613]. Armchair [61]. Arnauld [1241]. Arne [2450]. arrogance [1477, 2561, 2535]. Art [2404, 3242, 3111, 3165]. artefacts [3268]. artifact [2657, 3175].

Artifacts [472, 2657, 1577, 2221, 1018, 2740]. Artifactual [1905, 525, 2283, 3263, 1762]. Artificial [2562, 2270, 1778, 2389]. Artworks [3268]. ascribing [2893]. ascription [106]. ascriptions [2559, 1549, 2839]. ascriptive [2971]. aspect [2761, 601]. aspects [2393, 33, 2923, 1361, 670]. assembled [2564]. assent [1573]. asserting [746]. Assertion [2043, 1196, 1955, 1809, 18, 1730, 2042, 946]. Assertions [1376, 542, 2656, 3305, 2496, 1107]. assertoric [3082]. Assessing [3044, 2178, 2524, 1885, 789, 1853, 3058]. assessment [2699, 459, 2123, 2848, 2091, 2775]. assessment-sensitive [2848]. assistive [1539]. association [2261]. assumption [2976, 2479]. assumptions [1811, 1037, 502]. assurance [3282, 1451]. asteroids [2690]. astrophysics [3261]. asymmetries [1653, 1208]. Asymmetry [1632, 760, 761, 3304, 764, 678, 2251, 673, 2192, 323, 2397, 2483, 750, 1586]. atemporality [2660]. atmospheric [3278]. Atomic [29, 495, 496]. Atomism [3255]. attempt [2330]. attention [981, 1258, 767, 3233, 795, 2779]. attentional [3104]. attics [66]. Attitude [2839, 2277, 3138, 1109]. Attitudes [1781, 1169, 802, 70, 2450, 2298, 137, 3153, 988, 2901, 744, 247, 791, 2124, 2305, 330]. attractor [3005]. attribution [2795, 953, 1507]. attributions [844, 683, 1782]. attunement [1724]. atypical [1201, 2451]. Audience [577]. audio [684]. audio-visual [684]. auditory [1192]. Augustinian [2699]. austere [2454]. austerity [754, 1763]. Austin [2738]. Austrian [443, 1617, 2312]. Authentication [1828]. Authenticity [2814, 2493, 2615, 2680, 3361]. Authority [3164, 637, 1120, 205, 1216, 2426, 2187, 1749, 2971]. authors [1531]. Autism [1566, 2676, 1366, 3295, 3060, 3037, 2371]. auto [1040]. auto-abstraction [1040]. automorphism [2495]. autonomous [2325]. autonomy [1739, 2437, 150, 3042, 869, 2977, 2211, 987, 600, 3137, 1462]. Autopoietic [300, 282, 2008, 390]. availability [1049]. available [2498, 2621]. averaging [3360]. aversion [219, 865]. Avoiding [1084]. aware [735]. Awareness [988, 70, 468, 1047, 2983, 2489, 674, 1971, 1300, 1607, 1397]. awareness-first [1300]. away [2506, 1067, 3037, 1469]. Axiological [3158]. axiom [40, 68, 2237, 443]. Axiomatic [745]. axiomatization [2243]. Axiomatizing [538]. axioms [2015].

B [2240]. BA [1712]. BA-plan [1712]. babies [2057]. Back [2568, 113, 240, 1192, 2317, 2558, 2338]. Background [135, 3006, 2813]. Bacon [2754]. Bad [3359, 11, 3234, 2799, 2728, 1885, 2941, 74]. badly [421]. balancing [1871]. Bald [2857, 2496]. Bald-faced [2857, 2496]. ballistic [1471]. Bank [2077]. Bank-Type [2077]. bans [609]. Barad [1533]. Barcan [1074]. Bare [2438, 623]. Bargaining [31, 1537]. Barns [260, 2345]. barriers [2665, 3334]. baryon [2655]. base [1284]. Based [1316, 2306, 2175, 425, 1831, 1133, 1480, 3023, 3382, 12, 94, 1306, 2999, 2738, 2372, 276, 1473, 1853, 197, 2871, 1983, 2142, 1985, 2609, 2759, 1761, 2852, 2561, 2585, 2438, 1735, 279, 2577, 2539, 2535, 366, 1606, 1020, 3230, 1014, 3137, 2907]. basic [1993, 1541, 1467, 2161, 3157]. basis [1420, 2040, 2780, 2379]. Battle [1761]. Bayes [803, 3210, 756]. Bayesian [1898, 1013, 2256, 3405, 3287, 1663, 2076, 2990, 591, 1658, 2223, 1761, 2786, 556, 127, 2360]. Bayesianism [1265]. Bayesians [2115, 2316, 3099]. be [1627, 859, 3411, 1616, 2493, 619, 2494, 1564, 3032, 1770, 1531, 2896, 2550, 411, 2894, 1427, 1428, 735, 673, 898, 1421, 24, 2377, 2473, 1737, 685, 2673, 3016, 1846, 1772, 1467, 1717, 329, 1864, 637, 2496, 2422, 3157, 145, 2692, 2880, 1182, 3130, 1881, 3377, 1962, 365, 3211, 2287]. bearers [310, 2428]. bears [737]. Beauty [3276, 690, 2237, 1053]. becauses [2957]. Becher [1649]. become [3145]. bedfellows [450]. been [375, 2174, 3218]. before [913]. begets [928]. begging [63]. beginning [3057]. Begriffsschrift [1101]. behalf [763, 2382]. behaving [421]. behavior [1724]. behavioral [3095, 2978, 2859]. behaviorist [847]. behaviour [2889, 679, 2873, 749, 1402, 1243]. Behavioural [1826, 1402]. behind [975, 2519, 865]. Being [1086, 2374, 3029, 2725, 844, 2547, 2842, 1681, 1756]. Belief [1225, 2860, 252, 3055, 3361, 1753, 2065, 1305, 2291, 2191, 1176, 2579, 668, 2479, 1770, 243, 550, 753, 938, 1778, 2299, 250, 672, 3326, 926, 1614, 453, 3397, 1991, 2034, 3210, 872, 3053, 1423, 2974, 908, 266, 1004, 662, 3148, 1323, 1563, 1526, 1553, 2204, 833, 2254, 2002, 271, 1266, 1883, 1782, 968, 180, 1922, 452]. Beliefs [2993, 1505, 2136, 1392, 2413, 2858, 623, 1519, 2388, 2782, 829, 1509, 2999, 2293, 1090, 695, 135, 1355, 2799, 1085, 1467, 2469, 2417, 3026, 1061, 127, 3089, 271, 2760]. believe [122, 167, 2252, 2545, 2126, 1605, 1392]. believed [1531]. believer [2279, 1877]. believers [1616, 2799]. believing [3139, 1420, 2299, 1912, 2928, 700, 758]. Bell [45]. Bell-type [45]. Belnap [541]. belong [925]. below [913, 159]. Benacerraf [109, 1362]. Benardete [2787]. Bending [3168]. benefit [507]. benefits [2517, 1618, 1863, 730, 2074, 2085]. Benjamin [72]. Bentham [2944]. Bergson [1043]. Bernoulli [1771]. BERTology [2751]. Bertrand [3293]. best [615, 771, 130, 1206, 2190, 1925, 774, 1706, 1638, 2498, 2621]. Beta [1510]. Better [900, 1141, 3002, 1934, 1203, 1300, 395, 1238]. betterness [2648]. Betting [753, 2005]. Between [761, 467, 2885, 1692, 1462, 3154, 1956, 2479, 2358, 2858, 3346, 242, 621, 155, 344, 2944, 673, 616, 3056, 2397, 2483, 1716, 2558, 2928, 1214, 642, 1649, 3079, 882]. Beyond [2988, 3256, 3214, 2638, 1095, 789, 3148, 3227, 1260, 534, 996, 1498, 3064, 2600, 446, 973, 2921, 1690, 2795, 3022, 281, 2461, 3406]. Bias [1319, 2208, 2658, 2016, 2623, 610, 698, 655, 952, 2486, 1030, 1173, 1087, 1842, 2397, 2105, 1292, 1571, 1118, 1767, 2900, 351, 2798, 2507, 2483]. biases [469, 480, 3063, 1030, 1505]. Biasing [479]. bibliography [578]. big [2339, 398, 1990]. bilaterally [531]. bind [594]. Binding [786, 3104, 1780]. bio [1671]. bio-ontologies [1671]. biodiversity [2881]. bioengineering [529]. Biological [3290, 1055, 1719, 2318, 3214, 3289, 3028, 1405, 2706, 954, 2807, 3350, 1262, 1261, 3137]. Biology [1243, 1215, 3266, 2653, 1692, 728, 1599, 1707, 1407, 1060, 445, 738, 3161, 3091, 3146, 2980, 737, 2207, 1233]. biomedical [2315, 3019, 3062]. biorobotics [924]. biosemantics [3363]. biosocial [2318]. biotechnology [504]. bipolar [3147]. birth [2093]. black [2213, 1033]. Blame [1506, 1661]. blaming [1420]. blandsight [3233]. blanket [799, 1462]. Blind [756, 3337, 2005, 2282]. blinding [554]. blindness [831]. Blindsight [3233]. Blindspots [2631]. blingsight [3233]. blue [1282, 1678]. Blumenberg [2885]. Bob [1408]. Bodies [1503, 1558, 2086]. Bodily [688, 468, 2671, 187, 1527, 2914, 674, 750, 2681]. body [688, 3095, 1503, 3085, 5, 1937]. Bohm [1127, 1056]. Bohmian [1791]. Bohr [1533]. Bohume [209]. bold [2115]. Boltzmannian [2292]. Bolzano [2829]. bomb [2877]. bone [1215]. book [1115, 563, 1572]. Books [1084]. Boole [1686]. Boolean [1859, 901, 1442]. Bootstrapping [1424, 3400]. borderline [3245]. boredom [2686]. born [3351]. boson [463]. both [2590, 771]. bots [2608]. bottom [2386]. bottom-up [2386]. bottomless [1852]. boundaries [2790, 799, 3190, 1156]. Boundary [2646]. bounded [2878, 2851, 2122, 2474, 2419]. box [96, 1853]. boxes [1033]. Boydian [1652]. Bracketing [3202]. Bradley [1459, 963]. brain [3017, 933, 66, 1663, 24, 2887, 1397, 1962, 1260]. Brains [3292]. branching [3015]. Brandom [325]. Breaking [3037, 761, 949]. breeds [2311]. Brentanian [2733]. bridge [1522, 2321]. Bridging [3201, 2944]. Bridgman [2877]. brief [2761, 107]. brightspots [2631]. Bringing [240, 1192, 2085]. British [1507]. Brittle [2617]. brittleness [1322]. Broad [957, 2960, 1900, 2462]. Broadening [689]. broadly [3314, 791]. Broken [2760]. brokerage [772]. Brouwer [2300]. brown [2266]. Brownian [1808]. Brutalist [2011]. bucket [3200]. build [3239, 1853]. Building [1602, 574, 3132, 1048, 3356, 1197]. bullet [1931]. Bullshit [3338, 1669]. bundle [346, 1394, 845]. bundled [2427]. Burden [2136, 2072, 1650]. Burgess [3200]. Buridan [2132, 2321, 546]. business [2498, 2621]. bust [2617, 2241]. butterflies [2380].

C [405, 18]. C-word [405]. ca [394]. calculation [2701]. Calculus [1125, 1686]. calibrated [1421]. calibration [2035, 3203]. call [1100, 1001]. called [2374]. Cambridge [849]. Can [1346, 1616, 2494, 3004, 2009, 411, 945, 1421, 2082, 2868, 1737, 39, 3347, 2000, 115, 1373, 1939, 329, 3157, 2880, 268, 1962, 3411, 1100, 1534, 1679, 1183, 655, 1770, 2213, 735, 2920, 81, 272, 1087, 3134, 420, 1599, 1707, 280, 2474, 2419, 104, 1792, 1864, 1563, 380, 2690, 871]. Canberra [1347]. cancelability [1180]. cancelled [2174]. cannot [1632, 460, 3242, 2377]. canonical [537]. can’t [655, 955, 2, 1963, 898, 1121, 1359]. Cantor [3362, 2693]. Cantorian [2175]. capabilities [3072]. capable [1564]. capacities [191, 1776, 431]. capacitist [2473]. capacity [1600]. Capturing [1158]. care [3172, 401, 2951]. Carnap [1744]. Carnapian [1350, 57, 329]. Carnapians [14]. carriers [3166]. Cartesian [3218]. Cartwright [79]. Carving [2964]. case [739, 1201, 3147, 966, 831, 3297, 1848, 3124, 463, 1210, 2240, 679, 2758, 1764, 2292, 2663, 732, 3322, 1886, 427, 397, 2532, 101, 1053, 1595, 3409, 2543, 3122, 2105, 504, 2558, 2289, 446, 1119, 3283, 317, 3031, 1640, 2795, 2091, 1983, 2587, 2682, 1755, 3350, 3045, 3261, 2130, 1008, 2305, 2655, 949, 3236, 2225, 2933, 626, 2895, 2727, 2618, 527, 1014, 990]. case-based [1014]. cases [1729, 2387, 2667, 3163, 2642, 914, 613, 2308, 2820, 2501]. Categorial [2353, 1732]. categorical [911, 789, 2923]. categoricity [1301]. categories [2015, 2270, 1033, 396, 1702, 2745]. categorisations [3062]. categorization [112]. Category [3022, 426]. Catherine [730]. Causal [2208, 2755, 2731, 2256, 1798, 1285, 2531, 3217, 593, 840, 588, 1046, 73, 3365, 2456, 887, 3156, 969, 425, 1345, 585, 3270, 803, 1600, 3398, 3063, 826, 3064, 528, 760, 1874, 1596, 991, 2203, 293, 2273, 2087, 3317, 416, 1796, 671, 2867, 909, 1024, 2585, 838, 103, 1108, 1547, 3394, 1275, 2352, 2484, 400, 399, 3230, 590, 3099, 428, 2172]. causal-historical [3064, 2867]. causal-informational [826]. causal-modeling [1345]. causalist [3042]. Causality [2099, 2231, 2058, 207, 3302, 93, 1945]. Causally [1631]. Causation [1843, 1524, 238, 1349, 1019, 1570, 2456, 1371, 3125, 1641, 2922, 861, 79, 2942, 1309, 1602, 2531, 2442, 1694, 613, 1973, 2827, 2745, 3146, 2269, 319, 805, 2553]. cause [2089, 1351]. causes [972, 3220, 2754, 1235, 2879]. cautionary [1720]. CDT [1115]. censorship [3357]. centered [3154, 2399, 401]. central [992, 819]. centric [2143]. century [3119, 837, 2803, 3187]. cerebral [187]. certain [1133]. certainties [2786]. certainty [3212, 997, 1188, 1098, 123]. certitude [936, 169]. challenge [2402, 1987, 3235, 2947, 1411, 875, 3163, 2043, 1385, 3376, 3203, 2617, 604, 760, 2470, 296, 200, 2603, 1424, 2067, 1123, 2229, 2499, 1362, 1260, 35, 841, 3098, 3034, 3323]. challenges [2214, 1966, 2706, 2144, 1665, 2227, 1844]. Challenging [1393, 62, 1564]. chambers [2191, 3357, 143]. chance [781, 44, 2237, 2287]. chance-credence [2287]. chances [1468, 612, 1122]. change [3050, 2920, 2076, 1821, 1249, 1901, 2738, 1873, 1090, 695, 1895, 553, 304, 2432, 403, 849]. changes [2359, 2018]. changing [1873, 2326]. character [2922, 2905, 2808, 1697]. characterisation [2804]. characterization [3316, 3250, 2207]. characterized [673]. Characterizing [2689, 1255, 3231]. Charting [666]. Charybdis [3275]. cheating [2868]. checkmates [2185]. chemical [3312, 79, 1628]. chemistry [3119, 2240, 444]. chess [2185]. chickens [97]. children [2604]. chimeras [2036]. chimpanzee [1881]. Choice [1459, 2238, 739, 831, 50, 2769, 1076, 2962, 1341, 1395, 2193, 1697]. choice-theoretic [50]. choices [3149]. choose [1683]. Choosing [397]. chronic [1517, 3332]. chronobiology [565]. Chrysippus [3260]. Church [2947, 329]. Cicada [592]. circular [1436, 3049]. circularity [1632, 977, 724, 986, 1326]. circularity/regress [1632]. Circumnavigating [416]. citation [1956, 3145, 2463]. citizen [2085]. claims [2497, 116]. clarifications [199]. clarity [1015, 2652]. Class [1507, 2415]. classical [3303, 3200, 155, 2289, 901, 1442, 1680, 1546, 1735, 2577, 540, 2153, 2164]. Classicality [1442]. classification [505, 584, 3152, 3182, 132]. Classifying [2101, 1255]. Climate [695, 1831, 2883, 1249, 1901, 1102, 1197, 526]. climb [2147]. climbing [1401]. clinical [1615, 394]. clocks [1692]. close [616, 1373, 110]. Closed [1156, 1403]. Closure [453, 451, 3252, 170, 797, 1723, 452, 1703, 2124, 121, 1159, 2444]. clustering [81]. Co [965, 3327, 2803, 1365, 713, 3067, 1224, 2463, 2533]. Co- [965]. co-citation [2463]. co-construction [2533]. co-evolution [3067, 1224]. co-exist [713]. co-hyperintensionality [1365]. Co-intensional [3327]. co-occurrence [2803]. coadaptive [2272]. Coalescent [1382, 1383]. coda [1976]. coded [2109]. codes [2662]. coding [91, 886]. Coercion [2312, 3133, 3012]. Cogito [2542]. Cognition [2060, 284, 1201, 847, 1135, 2720, 2684, 3239, 2232, 3257, 881, 3246, 301, 187, 2874, 75, 2398, 2931, 2855, 2451, 285, 3037, 1407, 1060, 1132, 299, 3374, 149, 184, 2074, 2474, 2419, 1539, 305, 283, 88, 279, 278, 3249, 217, 304, 1570, 164, 190, 162, 163, 303, 291, 882, 2056, 2004]. Cognitive [610, 3156, 969, 2239, 2157, 1272, 1936, 2060, 1981, 143, 871, 768, 3004, 3243, 148, 1564, 191, 1849, 1856, 150, 3020, 2367, 2232, 1131, 290, 1133, 732, 2743, 961, 1568, 3388, 1130, 1385, 945, 285, 165, 189, 1725, 1129, 1132, 46, 276, 1473, 2074, 1203, 2977, 2518, 944, 1136, 2565, 3204, 279, 2646, 2830, 3330, 790, 1845, 282, 2181, 71, 736, 882, 284, 596]. cognitively [479]. cognitivism [294]. cognitivist [276]. Cognizing [3220]. Cohen [3095, 2575]. Coherence [1217, 1490, 1458, 1146, 3284, 2804, 1105, 929, 2352, 2484]. coherent [2372]. coherentism [1013]. coherently [3004]. cohesive [3234]. Coincidence [483, 1478, 992, 819, 1807, 241]. coincidence-free [241]. coincidences [1092, 1091]. cold [2680]. Cole [1668]. collaboration [1038, 2689]. Collaborative [1833, 2392]. collaged [3305]. collapse [2613, 520, 3256, 1235, 197, 874]. Collection [2393, 1892, 2106, 3195, 3082, 2967, 1221, 2981, 2056, 2004, 1228, 1243, 2887, 1893, 1245, 2512]. Collective [2591, 1477, 796, 2561, 2323, 2202, 1320, 2712, 1173, 2551, 978, 1852]. colocation [823]. colony [656]. color [1356, 128, 2549]. colors [1519]. colours [2730, 2039]. combating [3357]. combination [3256]. come [2688, 380]. comes [1843, 393]. coming [501]. comments [963, 88]. Commercial [406]. Commitment [1733, 919, 2620, 2063, 1315, 1833, 322]. commitments [2413, 925, 2963, 3400, 12, 94, 1925, 1965, 2678, 2993, 3158, 2102, 103, 451, 2444]. Common [2722, 2453, 1809, 2904, 1790, 206, 1267, 2789, 2949, 1965]. Common-sense [2722, 2904]. commons [2068]. commonsense [1082]. communicated [1124]. Communicating [1705, 2139]. communication [3155, 2963, 1998, 3209, 2400, 2622, 2778, 2697, 2047, 2401, 3173, 2689]. communicative [2836]. communities [1317, 3234]. Community [3229, 809]. community-oriented [809]. Compact [2690]. compactness [2735]. company [11, 74]. comparability [2101]. Comparative [2509, 2949, 3335, 2139, 3332]. Comparing [1449, 2495, 2597, 2085]. comparison [2075, 2735, 57, 1761, 1649]. comparisons [200]. compatibilism [1244, 161]. Competence [3306, 734, 1662]. competent [81]. Complementarity [26]. complete [2543]. Completeness [2427, 2250]. completion [3031]. Complex [3251, 717, 1782, 1236, 2304, 2497]. complexes [3330]. complexities [1692]. Complexity [928, 2435, 2420, 727, 2221, 2665, 1990, 971]. complicate [396]. complicated [1324]. complicity [3275]. complies [2890]. Components [2768]. composed [1864]. Composing [2708]. composite [1741]. Composition [499, 497, 489, 998, 1525, 1475, 488, 490, 1591, 1093, 495, 496, 153, 493, 491]. Compositional [3348, 214, 133]. Compositionalism [1823]. Compositionality [2963, 1010, 1792]. Compound [2647]. compounds [2647]. comprehensibility [227]. comprehension [877, 3198]. comprehensive [2166]. comprehensiveness [1902]. compression [586]. compulsive [897]. computability [1052]. computation [2314, 1856, 2125, 1822, 1968, 1131, 1721, 1773, 2835, 2149, 2026, 1473, 204]. computation-based [1473]. Computational [2791, 3041, 711, 921, 1893, 732, 1130, 2665, 1299, 1331, 2411, 3211, 1224]. computationally [1198, 2704]. computations [1175]. compute [731]. Computer [2534, 1977, 1673, 2075, 1209, 2455]. computing [1175]. con [2629]. conceivability [1919, 2943, 3085, 5, 1937]. conceive [507, 2669]. conceiving [379, 1339]. Concept [2752, 2393, 3167, 307, 657, 3183, 501, 3014, 2753, 1388, 3133, 1007, 3041, 3382, 1885, 93, 1009, 707, 2490, 2602, 3111, 1267, 3369, 1882, 306]. concept-based [3382]. Concept-formation [3167]. conception [2815, 20, 3195, 2772, 681, 2511, 1247, 1246, 1136, 941]. Conceptions [1273, 883, 913, 2612, 2545, 1984, 1810]. Concepts [112, 1089, 3303, 1855, 1698, 1947, 3215, 3077, 1078, 429, 1253, 2470, 1424, 1508, 3036, 3091, 2973, 309, 2308, 156, 588, 935, 108, 2882, 2843, 228, 3324, 882, 1575]. Conceptual [54, 2555, 1078, 3400, 2121, 3409, 2738, 3320, 3407, 350, 309, 1629, 1844, 1312, 2402, 3303, 160, 1742, 1015, 2762, 3135, 1885, 3410, 2304, 604, 1253, 3152, 3321, 1733, 2746, 157, 1332, 1756, 3072, 2951, 2113, 1939, 3138, 1099, 1890, 1670, 985, 3204, 1857]. conceptualisation [2315]. Conceptualising [2194]. conceptualism [1665, 2702]. conceptualization [2206, 3332]. conceptualize [2389]. Conceptualizing [1600]. concern [2640]. concerning [2210]. conciliation [1095]. Conciliationism [2846, 2884, 1029]. Conciliatory [1952, 1271, 637]. Conclusion [2919]. conclusions [1531, 639, 821, 107]. concrete [954]. concreteness [1394]. condition [1744, 2732, 2864]. Conditional [963, 1275, 3335, 1335, 600, 261]. conditionalization [2480, 765, 676]. Conditionals [1459, 1496, 2356, 1345, 1944, 917, 2806, 3260, 2142, 1985, 2169, 1153]. conditions [2528, 863, 227, 1894, 3043, 561, 759]. conducive [3213, 710]. conduct [2662]. Confabulation [802, 1045]. confabulations [1045]. Confidence [2850, 2559, 1821]. configurations [1019, 3241]. confinement [871]. Confirmation [1478, 1292, 1814, 2301, 463, 464, 1118, 3099]. Confirming [1901]. confirms [3018]. conflict [2969, 2870, 621]. conflicted [2388]. conflicts [1015]. conformant [2435]. conformism [8, 855]. Conformity [912]. confusion [1254]. Conjoined [2501]. Conjoining [1762]. conjunction [1645]. Conjunctive [3002, 1472]. connecting [2015]. connection [1090, 1187, 71]. connectionism [1, 2361]. Connectionist [2953]. connections [1386]. connectives [536]. connectivity [81, 1349]. connexive [3315]. conscious [2998, 628, 2825, 1370]. Consciousness [314, 662, 2925, 2023, 2834, 1264, 1210, 933, 2017, 656, 2984, 352, 2192, 1235, 3240, 197, 416, 1410, 3022, 3383, 2783, 684, 2307, 2907]. consciousness-based [197]. consensual [2334]. consensus [2747, 3205, 568]. consent [2594]. consequence [86, 536, 1160, 2832, 598, 531, 538, 341, 2356]. consequences [1366, 1321, 1741, 2921, 274]. consequentialism [1122]. consequentialist [3213, 1072]. conservation [2881]. conservatism [883, 623, 1441, 2320]. Conservative [545]. conserved [2896]. consideration [396, 1645]. Considerations [3315, 2753, 1674, 1883, 871]. considered [888]. Considering [2169, 3057]. consilience [2360]. consistency [226, 1015, 3276, 2944]. Conspiracy [2849, 3396, 1028, 2022, 3354, 2596, 2110]. conspiracy-theory [3396]. constants [1442]. constituents [56]. constitution [523, 3182]. Constitutive [1556, 227, 446, 3289, 991, 671, 1557]. constrained [221, 1954]. Constraint [2966, 2959, 2314, 1529]. constraints [2127, 3028, 2131, 2682, 1349, 2585, 2002]. construct [1587, 1620, 3027]. ’constructed [1366, 1007]. constructing [529, 965]. construction [2771, 3094, 386, 1702, 2271, 71, 871, 2533]. constructionism [707]. constructions [2563]. constructive [2670]. constructivism [1668]. construing [443]. Consumer [2822, 225, 3363]. Consumer-side [2822]. Contact [1918]. containing [542]. contamination [3232]. contemporary [3004, 1869, 430, 126, 432, 1299, 2021, 2810, 2216, 1014]. Content [1099, 862, 1312, 40, 3097, 3160, 1978, 967, 63, 1403, 2983, 2903, 293, 157, 694, 280, 767, 2047, 1643, 2564, 909, 116, 1435, 3204, 2925, 127, 2671, 937, 190, 1294]. Contents [1236, 69, 1903, 740]. Contested [2619, 1947]. Context [1493, 3309, 2613, 1709, 459, 83, 820, 2104, 1549, 22, 3141, 2476, 1932, 1805, 1196, 774, 1348, 1106, 3036, 3264, 2541, 848, 962]. Context-dependence [3309]. Context-sensitivity [1493]. contexts [1173, 1383, 1382]. Contextual [1691, 2524, 1705, 2181]. contextualising [1238]. contextualism [2384, 1302, 698, 2317, 1191, 244, 2100, 712, 1423]. contingency [1981, 822, 110]. Contingent [1369, 490, 547, 2938, 1613, 1817]. contingentism [2188]. contingents [2009, 251]. Continuants [171]. continue [1738]. continued [2965]. continuity [2995, 390, 2923]. continuum [39]. Contra [1172, 1962]. contract [2776]. contraction [709]. contradict [1860]. Contradiction [1478]. contradictions [3411, 1835, 755, 1311]. Contrarianism [3354]. Contraries [1835]. contrariety [1835]. Contrasting [848]. Contrastive [319, 67, 1437, 474]. contrastivism [2317]. contrastivity [977]. contribution [3398]. contributions [301, 914]. contributory [2613, 1446]. control [1263, 2969, 1445, 2239, 1432, 3381, 2325, 945, 2492, 1034, 1774, 3392, 3190]. controlled [2642]. controversies [1603]. controversy [2541, 110]. convention [3095]. Conventionalism [1802, 2954, 1537]. conventionalist [2773]. Conventions [1837, 2403, 1360, 3269, 1820, 2981, 2538, 1903]. converge [2700]. convergence [2824, 872, 1649]. conversation [3180, 1755]. Conversational [3209, 3359, 3228]. conversations [396]. conversion [3148]. convert [614]. Convincing [3399]. convolutional [2766]. Cook [2055]. Cookie [1381]. cookies [1381]. cooks [2090]. cooled [2412]. cooped [97]. cooperation [2025]. Coordinated [2806]. Coordination [1599, 2382, 1869, 2047, 1461, 1460]. coping [2005]. Cops [3333]. coral [1399]. core [311, 2889, 1331, 2498, 2621, 1344]. corpora [2021]. corporate [806, 2535]. corporations [2452]. Corporeal [1093]. corporeity [2914]. corpus [2514]. correct [562]. Correcting [612]. Correction [229, 3108, 2739, 970, 1679, 419, 1297, 2462, 2522, 1330, 2257, 2720, 2096, 2361, 43, 1557, 2125, 1653, 2455, 3085, 1428, 277, 2507, 893, 992, 1963, 235, 1079, 2931, 2854, 1678, 258, 2828, 3329, 942, 94, 2842, 186, 2275, 2198, 1937, 74, 3100, 2283, 1515, 1932, 2483, 2279, 496, 1407, 183, 1247, 2718, 2056, 2007, 2322, 825, 2481, 855, 2474, 2142, 3170, 2164, 1383, 2628, 1461, 1482, 1984, 533, 1262, 2621, 3006, 1092, 2882, 2440, 2484, 400, 1635, 2831, 1251, 2444, 556]. correctly [280]. Correctness [2528, 1290]. correlate [2566]. correlates [101]. correlational [3022]. correlationism [2947]. Correspondence [2698, 419, 418, 529, 1037, 3109]. corresponding [1585]. Corroboration [3193]. cosmic [992, 819]. cosmological [3375, 3248]. cosmology [415]. cost [3014, 1745, 1723]. cost-free [1745]. costly [2028]. costs [2493, 1756, 161, 1489]. could [2894, 1525, 1456]. count [1262, 1261]. counter [2674]. counterevidence [3281]. counterexamples [79, 259]. counterfactism [1172]. Counterfactual [466, 1191, 640, 1345, 1230, 2465, 678, 760, 1054, 613, 2131, 1205]. Counterfactuals [1058, 1830, 1688, 207, 1067, 2252, 1540, 2476, 1796, 1633, 110, 1924, 99]. Counterpart [2162, 233]. Counterpossibles [647, 2331, 1492, 1125, 2759]. counting [675]. coupling [1685]. covertly [2109]. COVID [2220, 2850]. COVID-19 [2220, 2850]. Crabbé [1155]. Crabbé-Tennant [1155]. crazy [607]. create [2048]. creation [1018, 3128, 849]. Creativity [1685, 2235, 3137]. Credal [1942, 1580]. Credence [223, 107, 3405, 1225, 252, 453, 1493, 2287, 2907]. credence-based [2907]. credences [1953, 612, 1084, 590]. credentiam [3026]. credibility [2028]. Credible [1644]. Credo [3026]. creedal [2293]. crimes [2217]. criminal [2251, 264]. crisis [448, 2583, 1967]. Criteria [1857, 2495, 160, 4, 182, 2707, 1083, 2255, 399, 400]. Criterion [1750, 176, 3287]. Critical [2406, 2508, 395, 1452, 920, 2248, 1142, 126, 101, 3272, 1617, 2091, 2166, 502, 645]. critics [3244, 1999, 288]. critique [2387, 2720, 2684, 2625, 2086, 3074, 55, 1890, 567, 389]. Cronus [3260]. cross [636, 2383, 3090, 2497, 2463, 3241]. cross-disciplinarity [2463]. cross-disciplinary [2383]. cross-fertilization [3241]. cross-linguistic [636]. cross-sortal [2497]. cross-world [3090]. crosscutting [928]. crossroads [2699]. crowds [1648]. Cryptomnesia [2327]. Cultural [714, 2056, 2004, 1564, 1425, 187, 2398, 2170, 839, 1242, 194, 971, 1167, 1224]. Culturally [1176]. Culture [2244, 783, 873, 276, 1278]. cultures [575]. cumulative [1278, 2169]. curative [1840]. Current [2667, 2196, 2237]. curry [2356, 1037]. curve [2362]. curved [3058]. customizing [2383]. Cut [1877, 2633, 1154, 1158, 2279]. Cut-free [1158]. Cut-off [1877, 2279]. cycle [3231].

D. [230]. damages [2139]. damaging [1999]. dance [1330, 1329, 3245]. dance-movement [3245]. dancers [1503]. dancing [3180]. dark [1562, 3035, 3087]. Darwinian [3171]. Data [1104, 728, 2271, 562, 1236, 2248, 2339, 1908, 586, 1454, 3152, 398, 789, 2949, 1866, 2895, 1168]. data-driven [1908, 789]. Davidson [1749]. days [2021]. de-idealisation [2933]. dead [1989, 599]. deal [1212]. Death [1267, 1759, 2607]. debate [65, 2847, 3050, 712, 1053, 1716, 1851, 2446, 3205, 2280, 3146, 3385, 3309, 876]. debates [65, 1252, 84, 1238]. debiasing [655]. debunkers [1499]. Debunking [2329, 1028, 2293, 1082, 635, 1695, 2472, 2599, 2032, 3067, 858]. decades [1299]. deceit [606]. deceitful [2496]. deceive [1909]. deceived [2496]. deceiver [2606]. Decentring [2242]. deception [2591, 2082, 2857, 2872]. decidability [2735]. decidedly [2555]. Decision [1316, 3183, 2712, 584, 1480, 1083, 2828, 2797, 397, 1455, 1875, 2557, 963, 896, 914, 1815, 1194, 587, 840, 3370, 890, 2391, 3236, 2352, 2484, 589, 583]. Decision-Based [1316]. decision-making [2712, 397, 2557, 914, 2391, 3236]. decisions [801, 1781, 1664, 2438, 1397, 2798]. decoherence [2040]. decomposition [620]. Deducing [246]. Deduction [10, 1039, 2015, 2356, 451, 2444, 277]. deductive [2385, 3272, 2602, 1313]. deed [297]. deemed [811]. Deep [2752, 1418, 2766, 3025, 2010, 248, 3404, 2415, 2138, 665, 3212, 1975, 1033, 3180, 3121, 3148, 2049, 670, 2830, 2355, 3389, 3167, 2545]. Deepfakes [2297, 2967, 2407, 2661, 2345, 2401, 2569, 2555, 1747, 2098, 2347, 2334]. default [1651, 636, 669]. defeasibility [363]. Defeasible [76, 636, 1952, 1467]. defeat [205, 174, 2588, 2786, 273]. defeated [3117]. Defeaters [1766]. defeating [2329, 2805]. Defective [1698, 3141]. defence [1838, 2454, 561, 821, 2069, 2461, 1779, 2343, 2498, 2621, 2576]. Defending [2905, 149, 2867, 1573, 135]. defense [1769, 2713, 2790, 3287, 893, 892, 1872, 2979, 3252, 2167, 2759, 818, 1044, 204, 2810, 3393, 2172]. defensible [2815]. defensive [801]. deference [1120, 2480, 1323]. Deferred [3012]. deficit [1926]. definability [10, 277]. Defining [448, 2906]. Definite [2762, 2627]. definites [1383, 1382]. definition [1957, 349, 376, 2332, 3071, 886]. definitionism [152]. Definitions [1055, 2552, 1151, 643, 789, 2443, 1270]. Deflating [1670]. Deflationary [15, 3150, 933, 1886, 2800, 524]. Deflationism [371, 729, 1363, 308, 1647, 2659, 1984, 1810]. deflationists [2402]. Deformation [1064]. degradation [1399]. degree [927, 2719]. Degrees [1675, 1041, 1677, 2224, 1843, 3398, 1057, 1563, 322, 1526, 271, 540]. Deliberation [1821, 1521, 792, 1781, 216]. deliberative [1534]. delusion [1247, 1246]. Delusional [1447]. Delusions [2002, 1615, 2993, 3089]. Delving [3077]. demand [1747, 1827, 1098, 2568]. demands [2770, 1276, 488]. Demarcating [285, 2652]. demarcation [1336]. dementia [3172]. democracy [1534, 1217, 2603, 3070]. democratic [2664, 934]. Demographic [801]. Demography [971]. demon [1022]. Demonstrations [2246, 2754]. demonstrative [863, 381]. demonstratives [717, 1782]. demystified [2335]. Demystifying [1854]. deniability [2892]. denial [695]. denialism [569]. denials [542]. denied [2310]. Dennett [3383]. deontic [978]. Deontological [896]. deontology [2351]. depend [52]. dependence [1052, 148, 140, 150, 1675, 607, 1569, 1577, 1016, 143, 3189, 149, 147, 336, 17, 2380, 3309]. dependency [3328]. Dependent [1824]. depends [132]. depiction [2401]. depression [3032, 3186]. deprivation [1756]. deprivationist [1759]. depth [3077, 974]. derealization [3110]. deregulation [2883]. derivation [1151, 1674]. derived [3179, 190]. deRosset [875]. Descartes [2606, 2652]. descendants [2761, 1190]. description [1411]. descriptions [2627, 2839]. Descriptive [1908, 2589]. descriptivism [3064]. desert [3067]. design [2534, 1778, 841]. Designed [2334]. designs [394]. Desirability [589, 963]. desires [3368]. despite [1084]. detach [2, 1963]. detectability [703]. Detecting [2608]. detection [3035]. deter [2868]. determinable [2997]. Determinables [773]. determinacy [2125, 1822, 35]. determinants [404]. determinate [2762]. determination [2099, 607, 119, 2231, 2425, 3261, 2058]. determine [799]. deterministic [1808, 1666, 781, 1523, 538]. develop [3239]. Developing [2952, 3237, 579]. development [3092, 577, 1855, 696, 155, 783, 2604, 276, 3306, 2879, 1628, 2225]. developmental [470]. developments [2300, 462]. deviance [899]. deviant [1424, 669]. deviant-interpretation [1424]. deviation [3214]. deviations [3343]. devoid [1279]. devotion [120]. Dewey [301]. Diachronic [1789, 991, 913, 984, 222, 997, 3350, 3021]. diagnoses [1366, 2707]. diagnosis [985]. diagnostic [2707]. diagnostics [552]. diagonal [1979]. diagonalization [958]. diagram [585]. diagrammatic [2250]. diagrams [1334, 1865, 598, 708, 2079]. dialectical [2699, 2459, 137, 741]. dialectics [3210, 2457]. Dialetheism [2645, 47]. dialogical [519, 2602]. Dialogue [2602]. dialogues [1658]. Diaphonic [2177]. Dictum [2487]. did [1434, 762]. diegetic [2948]. dietary [3140]. difference [1941, 344, 1655, 927, 642]. differences [596, 1260, 3241]. differend} [2138, 3121]. different [2707, 1321, 112, 2151, 366]. differential [1328, 2819, 641, 1924, 270]. differentiation [786, 2864]. difficult [1497]. difficulties [3180]. difficulty [2567, 2913]. diffusion [2334]. digging [1780]. Digital [2107, 2597, 3352, 1917, 2783]. Digitally [3096]. dilemma [3378, 1837, 2743, 3308, 3329, 2316, 260, 1395, 1916, 424, 1489, 1414, 3370, 361]. Dilemmas [1777, 1948]. diligence [3119]. dimension [2206, 2128]. dimensional [958, 2006, 1004]. dimensionalism [16, 1759]. dimensions [187, 2777]. Diodorus [3260]. Direct [690, 962, 1943, 1244, 848, 1749].

Directed [2077, 1542, 1068, 357, 3016, 587]. directedness [2482]. direction [1802, 2168, 775]. directionality [2891]. directions [384, 367]. directives [2656]. disablement [1566]. disadvantage [839]. disagree [572]. Disagreement [574, 566, 1799, 2183, 569, 3279, 144, 1986, 2349, 1250, 1841, 3378, 2138, 2030, 974, 1899, 957, 3212, 134, 1824, 32, 3025, 1102, 730, 3180, 570, 3121, 2177, 857, 222, 1933, 8, 855, 2812, 1827, 3148, 360, 248, 572, 567, 568, 638, 103, 3389, 3039, 1635, 1634, 2668, 1251]. disagreements [3404, 806, 1952, 2049, 1890, 571, 3167]. disambiguating [1755]. Disambiguation [1755]. disappearance [1144]. Disbelief [1109]. discernibility [966]. disciplinarity [2463]. disciplinary [2383]. discipline [574, 1594]. discipline-building [574]. disciplines [772]. discontents [1111]. discourse [1938, 2253, 2502, 3351]. discoverer [2242]. discoveries [529]. discovering [2567]. discovery [3298, 2242, 4, 182, 1039, 774, 2814, 1841]. discrepancy [242]. discrimination [2658, 1776, 1059, 2269]. discriminatory [2613]. discussion [2583]. discussions [2443]. disease [2663, 1351, 3332]. Diseases [2792]. disinformation [854]. disintegration [1826]. Disjunction [909, 1645]. Disjunctive [3250, 917]. disjunctivism [2984, 902, 3412, 292]. disjunctivist [450]. dismiss [2596]. Disorder [3174, 3147, 3267, 2517, 3110, 3185, 1007, 3041, 24, 3245, 2371, 3036, 2898, 897, 1610, 2760]. disorders [3140, 295]. disparities [404]. dispensability [1741]. displacing [2503]. displays [2028]. displeasure [3131]. disposition [1089]. Dispositional [2001, 2050, 1704, 1527, 988, 1563, 1684, 2440, 1935]. Dispositionalism [1288, 2163, 947, 1257]. dispositionalist [3160]. Dispositionality [1301]. Dispositions [1752, 207, 803, 829, 239, 3247, 3213, 198, 2089, 742, 1169, 2475, 428]. dispute [1986, 1416]. disputes [806]. disruption [3212]. dissent [2663, 2105]. dissertation [2803]. dissimilar [1190]. dissimilarity [834]. dissolves [214]. distality [909]. distance [1017, 2372]. distances [922]. distinct [1543, 1109, 2647]. distinction [2915, 1587, 2026, 559, 1784, 62, 2330, 21, 2095]. distinctions [1924]. Distinctive [2505]. Distinctively [2838, 1972]. distinctly [2812]. distinctness [2652]. Distinguishing [1399, 3060, 1024, 642]. distress [3174]. distributed [2645, 1833]. distributes [608]. Distributing [2139]. Distribution [706, 768, 2310]. distributions [1789]. distributive [2876, 3313, 857]. disturbances [2578]. disvalue [825, 824]. Ditching [607]. divergent [1383, 1382]. Diverse [2136]. Diversifying [2085]. Diversity [50, 2136, 1534, 3168, 1736, 2244, 2507, 2486, 2262, 1981, 2074, 1118, 2977, 2565, 2100, 649, 2341]. divide [692]. divine [3403]. division [144, 2518]. divisional [31]. divorce [3409]. Do [785, 3396, 2550, 1480, 2545, 1289, 2135, 650, 1314, 2057, 1469, 1941, 2751, 3368, 1411, 1886, 1747, 2043, 3153, 3295, 3213, 2750, 112, 2490, 2347, 1456, 1738, 1860, 1945, 3393]. doctor [3173]. Does [2301, 1281, 1988, 731, 2017, 764, 2700, 902, 2909, 516, 1098, 1879, 123, 1386, 2581, 1941, 802, 2772, 807, 2040, 3252, 24, 393, 1087, 1842, 1146, 3016, 396, 738, 3343, 1806, 1945, 526]. doesn’t [3238, 615, 833]. Dog [2109]. dogmas [1140]. dogmatic [2844]. dogmatism [3279]. dogmatist [3108, 3066]. dogmatists [364]. Dogramaci [1363]. dogwhistling [1982]. Doing [3334, 302, 172]. Domain [3274, 1739, 1330, 1329, 2513, 1010, 1870]. domains [2176, 1384, 2223]. domestication [1896, 921]. dominance [2488, 2889, 2237, 2352, 2484]. domination [641]. done [2377]. Don’t [2943, 1846, 2496, 828, 2784, 746, 2904, 2688, 2703, 2831]. doombehaviour [2921]. doomchecking [2921]. doomed [2976]. dooms [928]. doomscrolling [2921]. doomsurfing [2921]. door [1373]. Dormant [1940]. double [1147, 2582]. double-preventers [2582]. doubly [2760]. Doubt [456, 1265, 1206, 3093, 897, 2003]. Doubting [815]. down [2386, 390, 716, 1351, 1878, 2412, 2545]. downscaling [1249]. Doxastic [2089, 3230, 792, 1483, 256, 956, 1997, 223, 2489, 3038, 1004, 247, 161, 1109, 2210]. doxasticism [2084]. drain [1469]. drawing [821]. dream [1281, 2916, 366]. dream-based [366]. Dreaming [477]. Dreams [267, 1456, 3098]. Dretskean [637]. Dreyfus [2614, 1896]. Driftability [3094]. driven [2845, 1908, 789, 2216, 2432]. drives [275]. drug [2986]. Dthats [721]. dual [2761, 772, 1894, 3134, 651]. dual-aspect [2761]. dual-individuation [3134]. Dualism [3159, 1878, 1667, 1254]. dualists [2780]. dualities [3387, 25]. duality [2096, 2013, 722, 817, 3263]. Duhem [28]. Dummett [880]. Dunn [541]. duration [1005]. during [1645]. Dutch [3119, 2862, 1115, 1084, 1572]. duty [3016]. Dynamic [1459, 704, 1551, 1412, 2171, 3307, 2516, 2682, 2014]. dynamic-mechanistic [2682]. dynamical [1259, 861, 3289, 1272, 1972, 1901, 154, 776, 241]. dynamicism [1140]. dynamics [1521, 1213, 2525, 31, 463, 1622, 2539, 304, 1209]. dynamism [2578]. dysmenorrhea [2477].

each [3411, 218]. early [1956, 1720, 2803, 2604, 2021, 1368, 881]. Earning [2775]. earth [2955, 862, 2632, 2679, 371]. earthborn [1555]. earthquake [2139]. Eastern [117]. easy [1326, 2576, 870]. eating [3140, 3381]. Ebola [394]. echo [2191, 3357, 143]. Ecological [3364, 2371, 2439, 679, 139, 3219, 1083, 745, 2598, 294, 658, 1609, 292, 2533]. Ecological-enactive [2371, 2533]. ecologies [3172]. ecology [2386, 1003, 2432]. Econometric [1964, 2007]. economic [2429, 3300, 2701]. economics [2653, 442, 441, 443, 1617, 1811, 2439, 2312]. economy [2453, 1137, 2779, 736]. ecosystem [2398, 402]. ecosystems [1399]. Ecumenical [705, 539]. Edelberg [2143]. Editor [313]. Editorial [1208, 583, 2107, 306, 3187, 3223, 1653]. Educating [940]. education [1321, 1539, 1507]. effect [391, 1127, 1932, 1805, 274, 2161, 2174, 2210, 132, 2748, 2210]. Effective [3178, 3132, 1123, 3387, 2770, 886]. effectively [3357]. effectiveness [2095]. effects [1651, 616, 528, 2706, 937, 649, 2210]. efficacy [350, 1547]. efficiency [2516]. Effingham [1565]. effort [782, 3313]. efforts [2913]. egalitarian [49, 2603, 1118]. egalitarianism [1593]. ego [2623]. eigenstate [1654]. eigenstate-eigenvalue [1654]. eigenvalue [1654]. Eight [1299]. eighteenth [3119]. Einstein [1992, 1434]. Ekman [1155]. elaboration [649]. electrode [2225]. electron [1116]. electronic [1019]. elegant [237]. Elements [444, 3255, 446, 1435, 1628, 1241]. Elenchus [3226]. eleven [2224]. Elgin [807, 809, 730, 1221]. eliminability [1154]. eliminate [3005]. eliminated [3078]. Eliminating [1732]. elimination [2583]. eliminative [1676]. eliminativism [989, 2104, 2130, 2172]. else [1121]. Elsewhere [2748]. embedded [1176, 2286, 245, 766, 525, 2283, 1153, 288]. Embedding [2383, 1071, 1587, 2289]. Embodied [1364, 847, 290, 1813, 1407, 1060, 1132, 7, 183, 1359, 1532, 2474, 2419, 1529, 217, 284]. embodiment [879, 384, 1407, 1060, 1339, 1134]. embodiments [2710, 2840]. embracing [2356, 3241]. Emergence [979, 1518, 3199, 1938, 3201, 3200, 913, 984, 1291, 1145, 1033, 339, 839, 564, 976, 1513, 990, 1708]. Emergent [2582, 2442, 489, 1692]. emergentism [1264]. Emotion [2969, 1487, 2092, 2299, 7, 183, 1681, 3045, 2526, 2681]. emotion-regulation [2092]. emotional [2471, 391, 844, 1940]. Emotions [1256, 1689, 108, 2952, 1934, 1757, 2469, 1582]. empathetic [1590]. Empathy [1590, 905, 2680, 1216]. Empirical [1147, 363, 3308, 3097, 615, 2393, 459, 425, 2450, 2858, 3077, 2123, 1015, 2243, 3135, 2509, 1252, 1454, 1741, 1621, 1146, 652, 1603, 2144, 1273, 1889, 1731, 626, 3329]. empirically [2463]. empirically-informed [2463]. empiricism [2044, 1649, 841]. empiricist [2524, 962]. empricist [1560]. emption [80, 613]. empty [1651, 272, 1284, 611]. empty-base [1284]. enable [2109]. Enabling [575]. enact [2910]. Enacted [2888]. enaction [1813]. Enactive [1401, 1517, 3155, 3162, 2371, 3369, 305, 302, 2533]. Enactivism [286, 295, 847, 480, 1911, 300, 294, 390, 711, 296, 1407, 1060, 293, 297, 2702, 279, 3227, 282, 289]. enactivist [301, 283, 3098]. Enactment [71]. encoding [856]. encroachment [2579, 1207, 1225, 1574, 3279, 1195, 136, 2917]. enculturated [187, 732]. Enculturating [1211]. Enculturation [1575, 184, 192]. Encultured [2170, 177]. End [2562, 1622, 2610, 2352, 2484, 599]. end-Permian [1622]. ended [911]. endogenization [1599, 445]. Endogenous [573]. endorse [1029]. endorsement [807, 3326]. ends [3028]. endurantists [823]. Enduring [2070, 240, 2654]. energy [1719, 1065, 447, 4, 182, 3405, 1142, 1165, 711, 2890, 2587, 1262, 1261, 1516, 2430, 3087]. engage [404]. engaged [791]. engagement [504, 2702]. Engel [3378]. engineered [142]. Engineering [1253, 54, 1947, 1015, 1885, 604, 3152, 1733, 3320, 350, 2113, 1939, 3138, 1099, 1670, 309, 1629]. engineers [2951]. engram [1588]. Engrams [3086]. Enhanced [724, 1198]. enhancement [150]. enhancing [2028]. enough [415, 1699, 1221, 1815, 710, 1499]. enquiry [3264]. enrichment [2738]. ensemble [1831]. entail [1988]. Entailed [3228]. entailment [534, 539, 546, 533, 532, 530]. entails [340]. entangled [7, 183, 1581]. Entanglement [1017, 1111, 207, 212, 2504, 414, 886]. entanglements [616]. Entering [2018]. entities [1647, 1915, 921]. entitlement [245, 1763]. entity [1112]. Entropic [2748]. Entropy [2398, 2425]. enumerative [1571]. environment [398]. Environmental [2817, 3283]. environmentalism [3310]. environments [3237, 3239, 142, 388, 201, 3001]. environs [1635, 1634]. epidemiologic [396]. epidemiological [401, 395, 3390]. Epidemiology [402, 2261, 405, 397, 393, 403, 404]. Epigenetic [2597, 3137]. epigenetics [1407, 1060]. epilepsy [2682]. epiphobia [1293]. Episodic [1011, 1888, 3255, 2615, 3142, 2155, 2868, 1784]. Epistemic [2351, 48, 3147, 148, 3183, 1000, 449, 205, 623, 2517, 869, 3133, 2486, 1470, 1980, 44, 1374, 432, 2027, 797, 1430, 1497, 2899, 2103, 888, 1763, 457, 2530, 195, 452, 1866, 1303, 1823, 2305, 3286, 710, 2187, 359, 1585, 1055, 2635, 3378, 2116, 2297, 3125, 2930, 3401, 863, 2310, 3119, 1534, 3168, 1838, 1120, 668, 910, 140, 97, 1002, 505, 1751, 2160, 160, 150, 1566, 1926, 1675, 2591, 2870, 2769, 174, 660, 832, 28, 1997, 1076, 3346, 2270, 3354, 1388, 2244, 1366, 2915, 2399, 344, 1237, 1574, 3400, 957, 2407, 144, 2295, 843, 616, 245]. epistemic [1317, 1464, 2665, 328, 1509, 134, 2179, 2622, 2348, 1216, 406, 2701, 1544, 1173, 1836, 2476, 1343, 3192, 181, 1725, 3037, 1618, 1863, 3074, 420, 730, 3000, 3284, 1402, 1051, 1355, 1276, 3278, 878, 265, 808, 2452, 3283, 2551, 3341, 2603, 750, 2849, 2789, 2381, 2921, 1174, 143, 2098, 1933, 3189, 149, 2871, 274, 2648, 759, 2481, 2431, 1693, 2074, 1489, 1706, 1373, 1662, 2409, 147, 835, 1551, 2808, 2038, 2775, 434, 2628, 2573, 1529, 2926, 642, 2422, 705, 2820, 828, 2368, 742, 3027, 649, 2216, 268, 2980, 1181, 2210, 2499, 1570, 2189, 1362, 351, 2432, 1883]. epistemic [372, 261, 394, 897, 1324, 3010, 3144, 573, 132, 2139, 2507]. Epistemically [142, 1085, 1661, 3040]. Epistemicism [1569]. Epistemological [2451, 1422, 812, 1591, 2248, 320, 1247, 1246, 1204, 2227, 370, 1841]. epistemologically [218]. epistemologies [3154, 1765]. Epistemologists [2721, 1711, 2739]. Epistemology [1316, 1232, 2817, 1179, 3154, 379, 3103, 1679, 1183, 2862, 2998, 2916, 450, 2769, 701, 1240, 2851, 1786, 2962, 1184, 430, 218, 455, 957, 1317, 749, 1655, 3339, 119, 1970, 3310, 453, 263, 253, 1789, 1695, 681, 59, 3213, 367, 1119, 1788, 1327, 1287, 222, 3189, 149, 2196, 1373, 2409, 147, 1195, 360, 248, 321, 431, 3233, 452, 1665, 438, 651, 268, 1187, 372, 2512, 836, 3332]. EPR [1750]. equal [82]. equality [3380, 1377, 1589]. Equalized [2405]. equals [1365]. equation [2601, 1707]. equations [2819, 1924]. equilibration [3245]. equilibrium [1346, 939, 2824, 2815, 54, 1947, 3195, 2858, 2781, 2700, 2076, 2876, 1216, 2747, 2794, 2051, 2906, 838, 3039, 1727, 2788]. Equivalence [3379, 2019, 3050, 817, 2423, 1578, 664, 1851, 1723, 2694, 2145, 3350, 176]. era [2896]. Erich [1649]. Ernest [253]. Eros [2729]. erosion [2674]. Erratum [182]. Error [698, 933, 618, 3023, 1011, 46, 848, 2064, 1771]. errors [3095, 3253]. ersatzing [2830]. Ershov [2359]. Essence [930, 631, 376, 374, 3353, 1600, 611, 335, 2730, 337, 347]. essences [2915, 427, 2288, 2001, 3247]. Essential [346, 1307, 661]. essentialism [1540, 2996, 171, 1935]. Essentialist [377, 3391]. established [1360]. Establishing [399, 400]. Estimating [3398]. Eternal [465]. eternalism [1429, 1759, 2912, 151]. ether [1434]. Ethical [2108, 1476, 1664, 3068, 433, 1900, 2462, 569, 557]. ethico [218]. ethico-political [218]. ethics [3151, 1310, 430, 2346, 3409, 3321, 3361]. ethnic [880]. Ethno [3062]. Ethno-racial [3062]. ethnography [1015]. Etiological [2732, 743]. etiology [680]. etiquette [3279]. Euclidean [1334, 2245]. evaluability [712, 2798]. Evaluating [660, 1266, 596, 400, 399]. evaluation [2008, 2190, 1901, 1374, 2642, 1811, 3343, 968, 359]. evaluations [194]. Evaluative [1422, 3131, 157, 228]. Evans [1015]. Evans-Pritchard [1015]. evasiveness [2409]. even [2232, 1529]. Event [3007, 3031]. events [2710, 1280, 1870, 2960]. ever [1340, 2478]. Everett [3315]. Everettian [3304, 874]. Everybody [1121]. Everyday [2003]. everyone [632, 233]. Everything [1379]. Evidence [2348, 2218, 2648, 1097, 2291, 1561, 1340, 1330, 1329, 2256, 1699, 1256, 2108, 2296, 1745, 1447, 2559, 2251, 2956, 3196, 2346, 2557, 252, 2424, 3202, 1057, 3053, 1981, 993, 1603, 813, 2401, 541, 2438, 2269, 2254, 2785, 1771, 549, 626, 1608]. evidence-responsiveness [1447]. evidence-seeking [2559]. Evidential [1561, 2100, 2986, 951, 1736, 2237, 1950]. Evidentialism [908, 833, 135]. evidentialist [454]. Evidentially [245]. Evil [642, 1759, 896, 1022]. Evolution [2056, 2004, 1050, 752, 1564, 1425, 2914, 3337, 1242, 2025, 2307, 3067, 1224]. evolutionarily [1278]. Evolutionary [1082, 1778, 2472, 2482, 1106, 858, 714, 1695, 1599, 1707, 201, 445, 2032, 2146, 1499]. evolvable [553]. evolving [3041]. exact [848, 327]. examination [1729, 3272]. examined [300]. examining [2799, 356, 219]. example [1217, 1869, 2536]. examples [565]. exception [833]. exceptionalism [1712, 1988, 1993, 2366, 3288, 1927, 1723, 1555, 1974]. exceptionalist [1970]. exceptions [2169]. exchange [2312]. exclude [123]. excluded [2520]. exclusion [653, 885, 1108, 1046]. exclusions [2930]. excusability [812]. Exercising [702]. exile [2044]. exist [2904, 713]. Existence [806, 27, 1074, 1387, 3294, 1852]. existences [1543]. existential [3361, 2375, 1510, 2250]. Exit [2031]. exoplanetary [3278]. Exorcising [2084]. expansion [1925, 3379]. expectation [1826, 1581]. expectations [471, 713, 3286]. Expected [1194, 2540, 1774]. Expecting [2685]. Experience [62, 981, 3154, 2651, 2594, 1128, 2707, 1568, 628, 2825, 1356, 1601, 276, 157, 1772, 2912, 1615, 1784, 2564, 944, 818, 1149, 2780, 2477, 2527, 1061, 897, 1066, 795, 2514]. experience-first [3154]. experiences [1422, 1509, 1466, 2607, 791, 2574, 629, 1026]. experiencing [603]. experiential [2404, 20, 2728]. Experiment [2845, 2338, 2876, 3356, 2785]. experiment-applicability [3356]. Experiment-driven [2845]. Experimental [385, 693, 636, 1587, 528, 1371, 1001, 554, 1038, 1812, 2406, 2854, 2816, 2722, 2331, 2811, 1496, 3176, 3079, 3392]. experimentalist [3163]. experiments [655, 460, 1440, 2213, 2677, 2521, 1970, 2000, 669, 2181, 751]. Expert [2480, 1323, 1223, 2339, 2094, 181, 1971, 2187]. Expertise [2883, 2368, 1921, 60, 1812, 298, 2235, 1322, 759, 1414, 568, 2882, 2843]. experts [1793, 1996, 1576, 2707, 3056, 3341, 143]. experts-by-experience [2707]. explain [1346, 3368, 2168, 1886, 945, 2082, 1842, 1314, 2006, 268, 2475]. Explainable [2562]. explained [1200]. Explaining [565, 2028, 679, 1328, 1622, 826, 1021, 850, 2133, 1274, 3221, 933, 596, 2854, 2816]. Explanation [1284, 1295, 1593, 132, 1100, 777, 1495, 2306, 615, 1636, 974, 592, 1622, 875, 3322, 2357, 2554, 2272, 1405, 461, 581, 1438, 1206, 2190, 323, 2809, 760, 1925, 158, 728, 2289, 3290, 774, 1615, 2607, 274, 720, 1706, 1220, 1767, 593, 1024, 976, 127, 41, 1606, 337, 999]. explanationist [135, 2599]. Explanations [1129, 2434, 2731, 969, 67, 1135, 2873, 1436, 1718, 3002, 2766, 1936, 2121, 1972, 1800, 32, 2543, 2167, 650, 2012, 2682, 838, 2838]. Explanatory [2516, 3184, 1257, 600, 3068, 2384, 3266, 480, 1911, 226, 3042, 1137, 2641, 643, 1621, 2458, 1243, 1674, 476, 2211, 1779, 2585, 488, 790, 2095, 1443, 858]. explanatory/pragmatic [2095]. explanatory/pragmatic-trial [2095]. Explicating [1388]. Explication [58, 54, 56, 57, 55, 329]. explicit [1767]. exploitation [2244]. exploration [741]. Exploring [2105, 1061, 2830, 2980, 390, 1880, 504]. expounding [2830]. Expressing [42, 1899, 2073, 43]. expression [2569, 1827]. Expressions [2526]. expressive [3155, 1863, 1710]. Expressive-relativism [1710]. expressives [1710, 1001]. Expressivism [3373, 2183, 1910, 3280, 2165, 2760]. expressivist [1389, 1966]. expressivists [169]. Expressivity [978]. Extended [2874, 2060, 111, 1539, 190, 2971, 3380, 2720, 2684, 3246, 2286, 3080, 3078, 3096, 525, 2283, 149, 3351, 3227, 2114, 2724, 3190]. extenders [3388, 3172]. Extending [3369, 2375, 1954, 203, 132]. extension [768, 3020, 996, 1675, 1599, 1864, 541]. extensional [2923, 1383, 1382]. extensionalism [2410]. Extensionality [331, 487]. extensionless [1864]. extensions [2750]. extensive [3227]. extensiveness [299]. external [242, 1587, 190]. Externalism [1979, 645, 694, 281, 2067, 3292, 261]. externalist [1554, 3064, 604]. externalists [2127]. extinction [1622, 3262]. extra [187, 1936, 2756]. extra-cerebral [187]. extra-mathematical [1936]. extra-scientific [2756]. extrapolation [719]. extreme [443, 1617]. extrinsic [2932]. Eyewitness [2190].

fabric [1216]. face [2994, 963, 8, 855, 64, 3039]. faced [2857, 2496]. faces [1077]. facie} [1244]. facilitate [1855]. facing [2295, 1260]. fact [1524]. factionalization [2782, 573]. Factive [1606, 2214, 853, 1266]. factives [3377]. Factivity [2902, 807]. factor [32, 2327, 2599]. factors [2517, 956, 1174, 2038, 2181]. factory [79]. facts [1100, 2858, 1387, 1338, 950, 778, 2838]. factualist [315]. faculty [432]. fail [3252, 2262, 465, 3377]. failed [2057]. fails [1235, 3350, 709, 1909, 841]. failure [2169]. failures [2434]. fair [554, 3062]. Fairness [3236, 2405, 2510, 3394]. Faith [3126, 3284, 1423]. Fake [2635, 2345]. fallacy [719, 3301, 219]. fallibilism [763, 358, 677, 441]. fallibilist [683]. fallibility [946]. Fallible [763]. False [1991, 1176, 2904, 672, 1085, 246]. False-belief [1991]. falsehoods [895, 830]. falsifiability [1677]. Falsity [2321, 2581].

Fame [933]. Familiar [2574, 531, 1419]. familiarity [1383, 1382]. Fanaticism [3277, 2494, 2941]. far [1534, 110]. fast [1564, 624]. faultless [134]. favoured [2115]. fear [2035, 2530]. feasibility [2607]. features [565, 468, 2006, 3157, 1026]. feeding [291]. feel [1281]. Feeling [2299, 1934, 1888, 2375, 897]. feelings [1470]. feels [1223]. felicitous [895, 830]. feminism [3411]. Feminist [3314, 3411, 3243, 1829]. Fernández [1011]. Fernparallelismus} [1992]. ferromagnetism [949]. fertilization [3241]. few [3268]. Feyerabend [2220, 1316, 1562, 568, 608]. Feyerabend-inspired [2220]. Feynman [2079]. fiat [3349]. Fichte [2537]. fiction [2215, 1754, 3074, 751]. Fictional [842, 935, 2253, 525, 2283, 1311]. fictionalism [2695, 3060]. fictions [954, 1314, 1762]. Fidelity [1425]. field [3033, 471, 1517, 2858, 155, 1992, 393, 388, 45, 1056, 404, 3169, 109, 1362]. fields [2324, 1956, 1813, 616, 1983, 1116, 2159]. Fighting [1514, 1515]. Files [2229, 3086]. filing [2965]. financial [579]. find [1301]. Finding [3304, 1348, 210, 224, 3079]. Fine [2020, 565, 2956, 3153, 3248]. fine-grained [565, 3153]. Fine-tuning [2020, 2956, 3248]. Fineness [2527]. Finetti [1105]. finetuning [461]. finite [2790, 3135]. finkishness [663]. First [2676, 447, 441, 1365, 401, 2682, 1481, 3092, 3154, 2755, 2478, 3382, 444, 440, 2041, 2155, 2179, 2198, 2119, 3187, 2332, 373, 1054, 2154, 944, 2577, 1300, 2426, 757, 2427, 3391, 758, 1749, 1482]. first-order [2577, 2427]. First-person [2676, 2682, 1481, 944, 2426, 1749, 1482]. fish [737]. Fisherian [848]. Fission [1128]. Fitch [2947]. fitness [2746]. Fitting [1548, 2362, 634]. fittingness [1757, 2733]. Five [2386, 3018]. Fixed [644, 1353, 1803, 673, 19, 2683]. Fixed-point [644, 1803]. fixing [79, 2867]. fixity [2757]. flawless [1085]. Fleck [2344]. Fleckian [3332]. flexibility [2221, 423]. Flexible [554]. flip [1769]. flip-flopping [1769]. flocks [1692]. flooding [2635]. flopping [1769]. fly [2282]. fMRI [943]. focal [90]. focus [1932, 1805]. Fodor [1404]. fold [2389]. folding [580]. Folk [275, 953, 657, 1855, 844, 2388, 829, 2612, 1009, 1211, 396, 3111, 1390, 1860, 1239]. folk-psychological [844]. following [926]. food [1698]. force [2253]. Forces [639, 714, 139, 3247]. forensic [1967]. Forever [751]. forgeries [3354]. forget [828]. forgetting [828]. forgotten [1966]. fork [1586]. form [415, 969, 1289, 407, 1795]. Formal [2249, 155, 555, 3070, 2093, 1956, 752, 2180, 1789, 2024, 2537, 999]. formalism [2018, 36]. formalization [3195, 2823]. Formalizing [3260]. formation [2393, 2171, 2299, 3167]. formats [478]. forming [67, 2704]. Forms [2802, 942, 437, 2807, 2055, 3389]. formula [2351, 1111]. formulas [1074]. formulation [3084, 676]. formulations [28]. forward [499, 1680, 1774]. fossil [562]. foster [1055]. found [1175]. foundation [2746, 2078]. foundational [2157]. foundationalism [362]. Foundations [334, 1110, 1352, 2163, 2693, 2248, 2894, 3219, 2944, 461, 3321, 2694, 412, 589]. foundherentism [2390]. foundherentist [1663]. Four [1759, 1339, 721, 1777, 736, 3220, 545, 1716, 1303, 2389]. Four-dimensionalism [1759]. four-fold [2389]. four-valued [545]. fragment [2427]. Fragmented [2388, 3367]. fragments [2725]. frame [2637, 12, 94, 2738, 1139, 1434]. frame-based [12, 94, 2738]. Framed [2364]. Framework [464, 1330, 1329, 129, 933, 2978, 1184, 2179, 1316, 2516, 704, 3091, 1479, 999]. frameworks [40, 1350]. Framing [301, 1651, 2364]. fraternity [1589]. Fred [2080]. Free [677, 627, 3280, 1034, 2662, 1158, 1739, 2881, 3044, 97, 2099, 447, 2776, 1745, 868, 1165, 711, 3347, 1783, 297, 1853, 1683, 2890, 3081, 2587, 2977, 3229, 1010, 2193, 3406, 1262, 1261, 1516, 1604, 241, 1279, 2058, 1623, 2430]. free-energy [447, 2890, 2430]. Freedom [2569, 48, 677, 2265, 220, 1041, 1841]. Frege [2966, 727, 1457, 837, 1101, 2541, 1714, 2497, 3128, 1686, 762]. Fregean [178]. Fregeans [1096]. French [1105]. frequency [1943, 1514, 1515]. frequentism [1400]. Frequentist [1868, 3122, 1761]. fresh [732, 3268]. friction [372]. Friedman [249]. Friend [3047, 3170]. friendship [1276]. Friston [1294]. front [2558]. frontiers [335]. frontline [2196]. frontloading [383]. fulfilling [1355]. fulfilment [2947]. full [1651, 2747, 1335]. function [1179, 2416, 1949, 3410, 2706, 1003, 435, 743, 2893, 2732, 210, 2549, 1260]. function-structure [1260]. Functional [2752, 581, 2802, 266, 1475, 1142, 1773, 442, 602, 2795, 1670, 3091]. Functionalism [3266, 2827, 2783, 1700, 1141, 1147, 2285, 1138, 1142, 1143, 1145, 1146, 111, 1139, 2245, 361]. functionalist [3197, 1264]. functioning [867]. functions [1305, 1920, 680, 826, 2657, 1885, 3290, 3091, 3179, 2254]. Functionspaces [2362]. Fundamental [1296, 726, 866, 2456, 3003, 2314, 93, 443, 1297]. fundamentalism [2011]. Fundamentality [1433, 1536, 338, 1798, 318, 2159, 1020]. funds [608]. fungal [3374]. fungi} [3374]. further [80]. fusions [487]. Future [1087, 1627, 2623, 122, 2009, 497, 1376, 2667, 673, 1842, 2397, 583, 1342, 1497, 102, 748, 251, 113, 2483]. future-bias [2623, 1842, 2397, 2483]. Futures [1644, 2691]. fuzzy [2359, 2224].

gadget [783, 1164]. gadgets [1564]. Galilean [1528]. Game [2296, 1837, 1926, 2185, 3134, 3333, 2981, 2098, 2690, 999]. game-theoretic [1926]. games [1938, 2467, 3348, 3208, 90, 3134, 2516, 2602, 1106, 982, 1461, 1460, 2880, 1701]. Gao [646]. gap [596, 630, 2944, 1069, 435, 1146, 1889, 1278, 476, 1522]. gapless [297]. Gappy [1659]. Gaps [2145, 23, 2680, 2325]. garbage [146]. Gatekeeping [2896]. gauge [462]. Geach [727]. Geachianism [1486]. Gender [2426, 596, 1889, 3106, 1238]. Gendering [1357]. genealogical [59]. genealogies [104]. genealogy [1358]. general [1138, 1764, 1148, 3058, 1386]. general-relativistic [1764]. generalisation [1618, 1863, 2530]. generalism [511]. generality [175, 2026, 1766, 1792]. generalizability [2095]. generalization [2638]. generalizations [125]. generalized [1538, 538, 3198]. Generalizing [652, 3241]. generating [3335]. generation [3006, 2813]. generative [1333, 864, 1008]. generativity [2736]. Generic [3142, 880, 127, 2838]. Generics [1713, 2354, 3345, 2532, 2970, 3290, 2949]. generosity [1178]. genes [3375]. genesis [1978]. genetic [2968]. Genetics [3345, 2289, 1407, 1060]. genidentity [2079]. genuine [1616, 232]. genus [1835]. geohistory [1622]. geometric [1142]. geometrical [2754]. géométrie [1241]. Geometrization [1992]. geometry [2096, 2013, 1457, 837, 1813, 1375, 2652]. Gerhard [88]. German [2803]. Gestalt [383, 392, 386]. get [955, 961, 1423, 2831, 2784]. Gettier [1729, 669, 2820]. Getting [877, 1796, 948, 669, 1703, 3107]. ghost [847]. Gibson [386]. Given [1500, 2084, 1690, 1626, 1607, 1642, 1762, 1960, 1450]. glappy [1659]. glasses [1651]. Global [2165, 254, 215, 970, 3299, 3000, 3036]. glut [1069]. gluts [2145]. glutty [1659]. go [996, 2224, 2506, 1473, 871]. goal [2349, 2482]. Goals [470, 2024, 3140]. God [2606, 2020]. Gödel [93]. goes [2599]. going [3022]. golden [1771]. Golemon [3272]. gone [2703]. Good [3152, 2544, 3237, 2357, 1885, 914, 1973, 1876]. Goodness [1049]. got [2636, 2703, 3393]. governing [2020]. Graber [3272]. gradability [2875]. gradable [2224]. Graded [3225, 1688, 1209]. grades [3328, 736]. Gradual [2933]. grain [1425, 2527]. grained [565, 3153]. grammar [1419]. grandfather [1951]. granularity [1524]. Graphs [2250, 1510]. Grasping [3344]. grasshoppers [3313]. graves [2297]. gravitation [3200]. gravitational [1142]. gravity [1639, 448, 1138, 1144, 1148, 1009, 1145, 1146, 3058, 102]. great [692, 88, 1507]. Greco [2384, 2453, 2052]. Greta [1837]. Gricean [1479, 2759]. Gricean-style [2759]. grief [2703, 3045]. grips [501]. Grothendieck [2096, 2013, 2885]. Ground [664, 2332, 1809, 961, 2915, 206, 2445, 2809, 339, 2789, 335, 1522, 412, 341, 337]. Ground-theoretic [664]. grounded [1409, 121]. Grounding [1398, 340, 5, 915, 336, 17, 1665, 1081, 2726, 1040, 3197, 1369, 160, 2972, 1433, 1536, 2891, 860, 2932, 2102, 105, 1108, 436, 1020, 2553, 3085, 1937]. grounding-based [1020]. grounding-reductive [2102]. grounds [615, 2680, 2655]. Group [894, 1848, 263, 1662, 1922, 2579, 869, 3346, 550, 155, 806, 1989, 2062, 2435, 2551, 662, 2417, 2535, 2210, 2210]. group-based [2535]. Groups [1042, 2282, 1521, 1616, 907, 81, 2336, 2062, 591, 2949, 788]. Growth [772, 988, 3006, 2813]. guaranteed [1045]. guessing [255]. guidance [2272, 478]. guide [1404, 781, 2809, 755, 1064]. guided [469]. guiding [479]. guilty [1450]. guise [2519]. gullibility [2818]. gunk [1418, 491].

habit [1080, 3369]. habits [3237, 2122, 1609]. Haecceitist [2337]. Haeccities [3222]. haka [1359]. Hale [1408, 675]. half [1651]. hallucination [2083, 2902, 3129]. hand [2596]. handle [79]. Hans [2752, 1595, 1715, 2044, 1649, 2079]. Hanson [818]. Happily [7, 183]. happiness [1421]. happy [1846]. Haptic [2592]. Hard [964, 1145, 909]. hardier [544]. Hardy [446]. Hardy-Weinberg [446]. harm [1747, 3208, 359]. Harman [364]. harmful [2219]. Harmony [2276, 3232, 333]. harms [2622]. Haslanger [2113]. Having [3014, 1827, 756, 1128]. Hayekian [836]. head [2780]. Health [2307, 3147, 3162, 398, 2519, 1254, 2936, 404, 2898]. healthcare [3209]. heaps [1558]. Hearing [2036, 820]. heavy [2834, 2907]. heavyweight [1891]. Hedging [1426, 1955]. Hegel [3210]. Heidegger [1896, 3361]. Heisenberg [2633]. Heisenberg-like [2633]. help [2700, 1146, 1599, 1175]. helped [2048]. helps [2242]. Hempel [3308, 3329]. here [256]. here-and-now [256]. hereditary [2146]. heredity [2146]. Heritability [887, 692]. hermeneutic [696, 502]. hermeneutical [2303, 3174]. Heterogeneous [1337]. heuristic [1341, 2186, 2458, 1674, 2425]. heuristic-instrumentalist [1341]. heuristics [1787, 2524, 2759]. hidden [717]. hierarchies [1679, 1183]. hierarchy [1131, 928, 2147, 2149, 2026, 2212, 2359]. Higgs [463, 1685]. High [1487, 3366, 4, 182, 1602]. high-impact [3366]. High-level [1487]. Higher [3059, 3326, 273, 2834, 2791, 2946, 1952, 3202, 1605, 2827, 1098, 268, 1547, 821]. higher-level [1547, 821]. Higher-order [3059, 3326, 273, 2834, 2791, 2946, 1952, 3202, 1605, 2827, 1098]. highlight [1386]. Hilbert [2541]. Hinge [2413, 455, 1965, 2678, 450, 2030, 3400, 453, 815, 1373, 2993, 3158, 248, 452, 451, 2444]. Hinges [2453, 454]. Hintikka [2511]. historians [1977, 2455]. historical [427, 3064, 2867, 1575]. historicity [299]. Historiographic [626]. historiography [2550]. History [749, 689, 2529, 110]. HL [1685]. HL-LHC [1685]. hoc [2470]. hocness [2360]. Hoffman [2329]. hole [2213, 2236, 2837]. hole-linings [2236]. holes [2236]. holism [206, 208, 1829, 876]. Holmes [1039]. Holmesian [1039]. Home [3172]. homoeopathy [2173]. homotopy [43, 42]. honesty [2025]. Hoops [260]. hope [358, 2884, 1203]. horizonality [1068]. horizontal [653, 1101, 2026]. Horwich [105]. Hot [2680]. HOTT [2834, 41]. Howard [1037]. HPC [1112]. hue [2284]. Human [2221, 2056, 2004, 3156, 2099, 120, 2832, 2994, 2398, 2612, 2272, 945, 963, 2914, 3260, 2033, 1203, 1274, 2907]. Humanistic [1229]. humanities [2597]. humanity [2351, 3262, 2914]. humans [3239]. Hume [1200, 754, 1190, 597, 2487, 1485, 1188]. Humean [236, 977, 1835, 1404, 1240, 1436, 992, 819, 2979, 2020, 3412, 3273, 2560, 2869, 1744, 1220, 232, 1188]. Humeanism [1639, 868, 986, 200, 1199, 1287]. Humeanisms [1204]. Humeans [852]. humility [3119, 1477, 360, 1415]. hungry [291]. hunt [692]. Husserl [2083, 2947, 3408, 3353, 2679, 3113, 1188, 1113, 1450]. Hybrid [1320, 1135, 1660, 2627, 2765, 69]. hygiene [2639]. hyloenergeism [410]. hylomorphic [411, 414]. Hylomorphism [408, 415, 320, 407, 416, 2527]. HYPE [1958]. hyper [2692]. hyper-inferentialist [2692]. hypergraphs [2350]. Hyperintensional [632, 178, 2628, 2573]. Hyperintensionality [2853, 1365, 3165]. hypodoxes [2634]. hypothesis [2540, 1189, 474, 39, 839, 662, 3018, 2950]. hypothetical [2064, 2227].

ICM [1203]. idea [2437, 1564, 1161, 2810, 3330, 2084]. ideal [2662, 2881, 3044, 1127, 2776, 2664, 2626, 2977, 3229, 2948, 3406]. idealisation [2933, 1233]. idealisations [561, 1175]. idealised [2835]. idealism [202, 2461]. Idealization [559, 560, 778, 558]. idealizations [3112, 1640, 1123]. idealized [154]. ideals [2903]. ideas [3312, 246, 2996, 395, 1628]. identification [16, 1601, 2551, 2098]. identify [498]. Identifying [993]. identities [396, 317, 3369]. Identity [334, 160, 481, 2696, 561, 499, 497, 489, 2050, 494, 490, 425, 2015, 3222, 1752, 1494, 2255, 2143, 2983, 1104, 491, 492, 1393, 495, 998, 1630, 2000, 111, 3083, 2849, 2098, 2047, 1525, 2461, 176, 2497, 2271, 661, 171, 493, 496]. Ideological [2195]. Ideology [633, 918]. idling [2387]. If [2904, 1235, 2232, 2806]. Ignorance [920, 878, 836, 768, 3103, 831, 746, 812, 1705, 891, 1189, 2590, 853, 931, 975, 3283, 1396, 759, 940, 1032, 706, 796, 638, 865, 934, 736, 1426]. ignorance-involvement [736]. ignorant [921]. ignore [557]. II [1863, 652, 1442]. ill [1989, 2760]. illness [3296]. Illocution [2536, 867]. Illocutionary [2593, 1465, 3164, 3269]. illusion [2943, 628, 46, 1370]. illusionism [2017]. Illusory [2185]. image [3283]. imagery [735, 1370, 2681, 3194]. images [1368, 2334]. imaginary [1762]. imaginary-object [1762]. Imagination [1898, 525, 3074, 1581, 2309, 2404, 1590, 378, 1488, 1728, 1367, 1737, 1880, 3060, 1601, 434, 1529, 1858, 1318, 651, 2230, 2283]. Imaginative [1483, 1312]. imagine [2943, 1683]. imagined [954, 2681]. imaging [552]. Imagining [1498, 2991, 1501, 1772, 1311, 2670]. immanence [2771]. immanent [200]. immersion [1483]. immodest [2237]. Immodesty [2045]. immune [1011, 1182]. Immunity [1269, 618, 3023, 1036]. immunology [2758]. impact [3366, 873, 1384, 2775]. impacts [504]. impenetrable [2665]. impermissibility [125, 1574]. implementation [3409, 2149, 2951]. implication [915]. implicational [3173]. Implications [2302, 1281, 2667, 3110, 945, 3263, 2898]. implicature [3359, 3173]. implicatures [1651, 3228]. Implicit [2658, 3073, 70, 469, 2467, 655, 2828, 2797, 1030, 2063, 1767, 389, 1516, 2305, 351]. imply [2017, 516]. import [2634]. importance [1918, 2707, 2339, 280, 828]. imposition [2194]. impossibilities [2434, 2744]. impossibility [2728, 2701, 3335]. impossible [2938]. impotence [2029]. impoverishment [2856]. imprecise [2254, 1182, 1270]. Imprecision [1998, 1335]. improve [1045, 3343, 3392]. improvement [2505, 1241]. Improving [2059, 1014]. Improvisation [1918]. impugned [2141]. Impure [156, 2445, 1374]. Impurism [1207]. inaction [2202]. inadequacy [97]. Inappropriate [1038, 1934]. incarceration [1837]. inclusion [3106]. incoherence [1147]. incommensurable [282]. incommensurate [3215]. Incompatibilism [2033]. Incompatibility [1530, 694]. incompleteness [211]. inconsistency [2183, 975, 540, 1530]. inconsistent [2197, 938, 1084]. Incorporating [2587]. increase [2262]. increasing [2603]. Incubating [102]. Indecision [2132, 2441, 1417]. indeed [2202]. independence [2877, 1057, 1458, 2974, 1923]. independent [2740]. Indeterminacy [2029, 1807, 2009, 2274, 773, 1417, 665, 1654, 1743, 1721, 1753, 1283, 1830, 108, 211]. indeterminate [1342, 37, 1689, 590].

Indeterminism [1746, 3325, 1687, 1853]. indeterminism-based [1853]. Indeterministic [2972, 1580]. indexical [1307]. Indexicals [2836, 2604]. Indicating [3323]. indication [2976]. Indicative [1345, 2169]. indicatives [1787]. indicator [927]. Indigenous [396]. indiscernibility [1968, 980]. indiscernibles [3222]. Indispensability [724, 592, 1915, 3179, 779]. Indistinguishability [2959]. Individual [2990, 2194, 2877, 1645, 935, 2535, 1260]. individualism [131, 876]. individualism-holism [876]. individualist [2200, 1800]. individuals [3171, 425, 2511, 2315, 738]. individuate [2604]. Individuating [982, 1354]. individuation [3115, 693, 1773, 2736, 3134]. Induction [643, 2249, 3276, 3308, 3329, 1571, 1485, 1638, 2655]. Inductive [2488, 2485, 2990, 1523, 1435, 1649]. indulgent [2081]. industrial [2507, 2486]. inelasticity [2122]. inequality [2208, 1238]. inequivalence [949]. inertial [1141, 1139]. inevitability [110]. inevitability/contingency [110]. Inevitable [812]. inevitably [1800]. Inexact [1412]. inexistence [1502]. infallibilism [1023]. infelicity [3224]. infer [655]. Inference [323, 1925, 1943, 916, 668, 615, 1849, 129, 631, 690, 1206, 2190, 3210, 2111, 3274, 774, 1706, 1313, 3071, 1860, 3394, 1868]. inferences [1337, 2987, 3256, 1831, 636, 1038, 1765, 2595]. Inferential [3232, 3216, 1804, 324, 880, 534]. Inferentialism [3288, 3305, 322, 321, 325, 3258, 3188, 1606]. Inferentialist [2137, 3402, 3181, 2644, 2692, 3309]. inferred [1117]. Inferring [353]. Infinite [1943, 1612, 1305, 10, 277, 808, 2223, 1123, 2787, 3057]. Infinite-population [1612]. infinitely [914, 540]. infinites [2693]. infinitesimal [785, 625]. Infinitism [2789]. infinity [2693, 2316, 2610, 3293]. inflation [465]. inflationary [3150, 1886]. Influence [2515, 394, 887, 2696, 585, 579, 2642, 1384, 2756]. inform [3320]. informal [2584]. Information [2015, 3287, 649, 1279, 1777, 3178, 314, 1264, 2122, 2065, 2602, 416, 796, 2368, 854, 2968, 26, 945]. Information-devoid [1279]. informational [1346, 2930, 969, 826, 1152, 533, 532, 2259]. Informative [139]. informativity [1031, 1079]. informed [2594, 2463]. infra [578]. infra-structures [578]. ingredients [1970]. inherently [1276]. inheritance [3235, 3388, 2056, 2004]. initial [1586]. injustice [2613, 2930, 1566, 2517, 844, 2171, 1544, 195, 3173, 1866]. injustices [3174]. innateness [1136]. inner [1643, 1607]. innocence [985, 2195]. innocuous [3040]. innovation [507, 501, 2637, 504, 192, 502, 508, 503]. input [1664]. input-selection [1664]. inquire [1237]. Inquiry [2911, 354, 3011, 247, 3371, 2833, 2676, 1419, 66, 2364, 3395, 2189, 1426]. inquisitive [196]. insecure [2750]. insertion [3038]. Insight [1823, 905]. insights [2456, 2021]. insignificance [2264]. insignificant [1885]. insignificant-assessing [1885]. inspire [1114]. inspired [2220]. instabilities [154]. instantial [2199]. instantiation [2736]. institutional [1387, 2262, 1668, 849, 836]. institutionalization [772]. institutions [680, 2857, 2859, 2888, 1701]. Instructions [2563, 1713]. instructive [1520]. instrumental [2272, 650, 836]. instrumentalism [1002, 832, 705]. instrumentalist [1341, 137, 1380, 2430]. instrumentalists [2422]. Instruments [2270]. insults [996]. integer [1424]. integrate [2793]. Integrated [1264, 976, 2605, 416, 2015]. Integrated-structure [976]. Integrating [1131, 3342]. Integration [2605, 3178, 314, 505, 281, 1262, 1261, 790, 841]. integrative [2676, 3178, 679, 1119, 1066, 841]. integrity [2662, 2783, 3311]. Intellectual [150, 1178, 360, 835, 1415, 2535, 1848, 1751, 770, 1477, 851, 681, 1845, 1598, 273]. intellectualism [3297, 2198, 2119, 2447, 1520, 356]. intellectualist [1062]. Intellectualizing [667]. Intelligence [1921, 2562, 1911, 2270, 3080, 945, 586]. Intelligent [478]. Intensional [480, 332, 2468, 3327, 333]. Intensionality [326, 334, 1911, 1372]. intensity [2490, 1879]. intensive [886]. intent [1909]. intention [1044, 1397]. Intentional [1511, 3020, 2033, 1218, 421, 2143, 1982, 1430, 1541, 1502, 1479, 1520]. intentionalism [782, 3269, 1888]. Intentionality [1294, 619, 1068, 1320, 2083, 290, 1894, 2272, 3024, 287, 1532, 1231, 283, 1607, 3194]. Intentions [3155, 2239, 784, 788, 2538]. Inter [3322]. Inter-level [3322]. interaction [3375]. Interactionist [2268]. interactions [3166, 3155, 3056]. Interactive [2525, 924, 298, 421, 2215]. Interdisciplinarity [2803, 2432]. Interdisciplinary [2369, 3184, 2518, 2887]. interest [2870, 737]. interest-relative [737]. Interests [1820, 2008, 406, 743, 2893]. interface [2329]. interkinaesthetic [1918]. Internal [2050, 2127, 3017, 1587, 2989]. internal-external [1587]. internalism [1740, 2905, 2448, 694, 161, 1099, 281]. Internalist [2414, 269]. internally [850]. Interpersonal [2974]. interpretability [724]. interpretable [999, 1844]. Interpretation [2423, 666, 2471, 242, 1455, 2040, 658, 1152, 2746, 1229, 1424, 2814, 2948, 567]. interpretational [2611]. Interpretationism [1016, 1616, 2609]. interpretations [2301, 2228, 3154, 669]. Interpreting [2228, 2513, 3220, 1362]. interpretive [961]. interrelations [2858]. interrogatives [793]. Intersubjective [2263]. intersubjectivity [2301, 392]. intertheoretic [800]. Interval [2539]. Interval-based [2539]. intervention [3214, 2657, 2498, 2621]. interventionism [2942, 2827]. Interventionist [1633, 1371]. interventions [3396, 1840]. interview [2552]. intimate [3180, 2334]. intragroup [1521]. intransparency [1673]. intrapersonal [2955]. Intrinsic [922, 923, 3067]. intrinsicality [930, 1914]. intrinsically [2728]. intrinsics [240]. introduced [748]. Introducing [2274, 2389]. Introduction [1239, 34, 1193, 2420, 348, 2393, 3082, 117, 449, 75, 430, 2975, 440, 382, 2424, 393, 1880, 95, 1221, 230, 407, 2056, 2004, 2887, 2512, 1611, 2163, 548, 140, 3195, 509, 2967, 313, 566, 53, 467, 1976, 2451, 253, 3187, 2511, 575, 367, 2981, 1228, 1243, 1892, 335, 3223, 184, 2107, 558, 2038, 2565, 13, 321, 1513, 1893, 2497, 2100, 278, 2106, 1245, 530, 306, 1150, 417]. introspection [802, 1412, 1022]. Introspective [1907]. Intuition [586, 1457, 1667, 1223]. intuitionism [2300]. Intuitionistic [3401, 484, 1746].

Intuitions [2346, 1912, 1697, 951, 3203, 2053, 1390, 687, 3067, 2412]. intuitive [2210]. Invariance [1578, 1914, 1768, 2333]. Invariances [1291]. invariant [398, 1546, 2164]. Invasive [1003, 569]. investigate [2612]. Investigating [3261, 2623]. investigation [1001, 3077, 2144]. investigative [1826]. invitation [2954, 800]. Involuntarism [2141, 161]. involve [807]. involvement [736]. involves [1911]. involving [2707, 1276, 127]. Inwagen [621]. iota [2376]. ir [891]. irony [877]. irrational [3032, 219]. Irrationality [2441, 127]. irrealist [2380]. irreducible [2922]. Irreducibly [1852]. irrelevance [600]. irreplaceability [634]. Irreplaceable [2821]. is-ought [630]. Isaac [3030]. Isaacson [3372, 2790]. islands [143]. isn’t [1395, 2953, 1676]. isolation [2031]. isomorphism [2837]. isomorphisms [561]. Issue [393, 230, 2191, 348, 140, 480, 1911, 117, 449, 566, 653, 53, 75, 467, 367, 407, 417, 184, 13, 321, 1513, 88, 278, 530, 1150]. issues [3017, 301, 2451, 2424, 1252, 2196, 2280]. Istanbul [2044]. iterative [2772]. itself [3004].

J [3315]. J. [2450]. Jan [1978]. Jean [546]. Jeffrey [926, 1194, 107]. Jeffreys [4, 182]. Jeremy [1710]. jerk [1001]. Jerzy [3122]. job [1411, 1423]. John [2052]. Johnston [1946]. Joint [981, 2365, 925, 767, 810, 298, 3016, 271, 805]. joints [2964, 1271]. Joongol [74, 11]. Journalism [3009]. journals [1299, 2717]. Joyce [216]. JuDAS [250]. judge [1309]. judgement [1016, 1173, 267]. judgement-dependence [1016]. Judgementalism [890]. judgements [165, 194]. Judging [670]. judgment [255, 435, 249, 1867, 1878]. judgmental [266]. judgments [2663, 2872, 1485, 275, 2210, 2898]. judicial [2438]. Judy [72]. Julian [1668]. junk [2943, 1418]. jurisprudence [1890]. jurisprudential [3406]. jury [3340]. Just [3036, 2582, 844, 1223, 2759, 2425, 1876, 3010]. Justice [3313, 1904, 2876, 3037, 3283, 1537, 2981, 1939, 2184, 1931]. Justification [1504, 2431, 20, 348, 1993, 2870, 363, 983, 3214, 1663, 843, 440, 1219, 2348, 2770, 2394, 2902, 2901, 694, 3367, 3071, 2089, 1735, 1593, 3027, 351, 946, 180, 804, 2481]. justifications [3070]. Justified [2291, 2579, 1531, 2204]. justifiers [1689]. justify [2772]. Justifying [951, 1341]. juxtaposition [2987].

Kahneman [2122]. Kant [2351, 433, 3312, 430, 3135, 942, 437, 3247, 432, 439, 1485, 2461, 434, 1113, 2996, 1823, 164, 162, 436, 3391, 163]. Kantian [3052, 227, 3220, 435, 265, 431, 687]. Katz [2636]. Keep [97, 1730]. ken [1971]. key [1592]. Kim [11]. kind [3298, 989, 1536, 627, 426, 3339, 3064, 3175, 2112, 423, 927, 234, 2288, 424, 2867, 422, 737, 1304, 366]. kindhood [1828]. Kinds [421, 2500, 1897, 419, 418, 425, 2653, 1696, 2219, 928, 3060, 2394, 420, 2062, 100, 2117, 417, 3358, 2087, 424, 1112, 1923, 108, 790, 2311, 1610, 428]. kinematics [1213]. King [839]. kink [297]. kink-free [297]. kinship [1954]. KK [170]. knew [2872]. Knobe [2210]. Knot [1928, 544, 1488]. knots [2624]. Know [2506, 2119, 355, 746, 2753, 667, 938, 1047, 807, 2043, 2934, 1991, 3100, 3069, 2703, 3206, 380, 356, 2198, 2753]. Know-how [2506, 355, 807, 1991]. Know-how-first [2119, 2198]. knowability [3307, 640, 438]. knowers [1585]. Knowing [3029, 3384, 2938, 1816, 375, 3114, 1193, 761, 827, 353, 2112, 1550, 2170, 2544, 648, 1660]. Knowledge [3092, 358, 956, 2753, 179, 1403, 3382, 672, 1875, 1103, 2588, 663, 766, 165, 3213, 1227, 181, 1180, 1402, 2345, 274, 1619, 549, 757, 180, 758, 166, 1507, 1055, 3095, 1775, 2404, 2291, 2310, 3073, 348, 2488, 2614, 117, 255, 3020, 1412, 2453, 2505, 1920, 772, 3281, 869, 349, 938, 175, 1480, 1062, 2559, 2984, 1047, 1907, 2214, 146, 455, 2200, 2339, 2418, 1790, 683, 2155, 12, 94, 3277, 262, 2179, 1549, 3096, 354, 374, 1595, 432, 1023, 2974, 1730, 2435, 264, 2055, 3102, 1276, 439, 2170, 1359, 1406, 2912, 33, 2466, 1833, 824, 2154, 1965, 1592, 637]. knowledge [788, 1818, 1553, 2323, 2616, 1300, 1942, 3371, 364, 889, 123, 1817, 1169, 2392, 1218, 121, 269, 163, 2258, 1749, 2875, 2263, 325, 622, 1107, 825, 2052]. knowledge-attributions [683]. knowledge-based [1480]. Knowledge-conducive [3213]. Knowledge-first [3382, 758, 2155, 2179, 1300]. Knowledge-how [672, 2588, 663, 2258]. knowledge-involving [1276]. knowledge-that [2614]. known [2179]. knows [2350]. Knox [1141]. Koffka [388]. Korta [2983]. KR [725]. Kripke [1850, 2071, 3365]. Kripkenstein [2067]. Kuhn [2550, 567, 30]. Kuhnians [2550]. Kuleshov [391]. Kürbis [2577].

L. [2450]. lab [1022]. labor [144, 2518]. laden [2756, 2760]. ladenness [943, 1731, 78]. lag [1656]. Lagrangian [2861]. Lakatos [1133]. lambda [2376]. land [201]. landscape [666, 388, 1512]. landscapes [1179]. Language [1938, 317, 2387, 2751, 15, 996, 2450, 2134, 3246, 2894, 636, 1038, 2406, 821, 3024, 126, 3118, 965, 2641, 276, 417, 2890, 563, 2529, 194, 1214, 2222, 2133, 2863, 1903, 1224]. languages [2093, 966, 2088, 89]. Large [3024, 2751, 2467, 2894, 3118, 1950, 219]. large-N [1950]. Laser [1491]. last [229, 168]. later [755]. Law [2585, 3235, 3103, 2296, 2520, 1385, 1613, 264, 1674, 2396, 489]. Law-based [2585]. Lawful [1436]. Laws [1819, 2600, 2238, 3391, 731, 977, 992, 819, 2891, 1579, 1058, 2979, 2020, 1613, 1438, 986, 1783, 1212, 1624, 2097, 99]. lay [3056, 2898]. lead [3238]. leads [529, 3276, 3159, 3079]. learn [460, 2213, 1711, 2945]. learnable [1379]. learned [1588, 3157]. Learning [3215, 327, 2945, 1567, 2731, 3366, 2239, 1076, 585, 3132, 2638, 1778, 2339, 1249, 90, 3053, 504, 1033, 2170, 2010, 2411, 1229, 1032, 1774, 2193, 670, 3027, 1604, 2830, 2355, 1167, 1275, 999, 1844, 2935]. least [2554]. Legacy [230, 1538, 30]. Legal [2843, 2256, 2800, 2908, 2069, 3263, 1890, 2882, 2882]. legitimacy [2664, 2438]. legitimate [2595, 895]. Leibniz [489]. lens [1048]. less [418, 81, 1209, 419]. lesser [896]. lesser-evil [896]. Lessons [2046, 213, 3178, 3050, 689, 2386, 1786, 2241, 1784, 571]. Let [572]. Let’s [1853]. letters [1957]. level [2158, 3322, 1602, 2830, 1547, 502, 1487, 821].

Levelling [1205]. levels [534, 3107, 876]. Lewis [235, 231, 242, 6, 230, 234, 1572]. Lewisian [3270, 244]. Lewontin [1050]. Lexical [3153, 1552, 3216]. lexicalized [2137]. LHC [1679, 1183, 1685]. liar [2961, 2634, 546]. liberalised [841]. Liberalism [2741]. liberation [1595]. liberatory [2494]. Lie [1909]. lies [159, 2496, 1209, 1734]. Life [152, 2923, 1055, 3017, 3140, 1724, 447, 3084, 3231, 285, 2807, 789, 3081, 244, 3161, 2443, 3389, 3137]. Life-mind [2923]. lifeness [789]. light [1639, 300, 1033]. Lightweight [1891]. like [1519, 1928, 1831, 1366, 2633, 408, 587, 2436, 116, 1881, 3130]. likelihood [2217]. likely [1467]. Lima [1470]. Limitation [1053, 1867]. limitations [1139]. limitative [339]. limited [2081]. Limiting [515]. Limits [1950, 876, 3033, 1590, 2614, 2572, 1584, 2588, 1880, 2105, 1601, 1763, 2098, 1318, 2308, 330, 3390]. Lindley [4, 182]. line [2610, 1864]. lineage [2272]. linear [1095, 3191]. Linguistic [2464, 2940, 2206, 951, 1998, 636, 1812, 3024, 1094, 287, 1051, 2205, 987]. linguistically [1356]. Linguistics [1137, 2450, 1333, 1008]. linings [2236]. link [1654, 1214]. links [1496, 1573]. literacy [395]. literature [2328]. little [1703]. Live [1252, 3149]. Lives [1544]. living [2416, 1503]. loaded [804]. Local [1641, 922, 2350, 3000, 241]. localist [2987]. locality [887]. localized [2411]. locate [2780]. located [151]. locating [280, 2887]. location [773, 2709]. locations [2566]. Locke [1926]. Logic [1073, 1760, 1453, 1594, 543, 3411, 3401, 1567, 1988, 509, 3124, 2061, 1825, 338, 2350, 1682, 1427, 1428, 1189, 3315, 2366, 2962, 958, 3288, 1039, 1966, 2627, 2946, 2999, 1952, 1637, 545, 853, 797, 1550, 3260, 2924, 3314, 1152, 1927, 2794, 2381, 2162, 2871, 2608, 1555, 1010, 541, 1882, 2468, 929, 2537, 1735, 2204, 1113, 1984, 1810, 2820, 2368, 1974, 2014, 540, 3144, 2427, 3230, 648, 196, 2833, 821]. Logical [513, 2924, 1032, 1313, 2044, 1795, 3378, 542, 2947, 1151, 518, 1993, 904, 28, 515, 1899, 519, 1419, 3143, 1994, 1966, 510, 993, 1723, 1560, 517, 88, 1649, 516, 1984, 1810, 798, 1686, 1530, 1608, 511, 1697, 75]. Logicality [2863, 331]. logician [1039]. logicism [1286]. logicization [1978]. Logico [1530, 2490]. logico-mathematical [2490]. Logics [2735, 520, 2224, 3257, 3355, 1743, 2445, 545, 901, 1442, 1551, 2054, 1735, 632, 978]. Long [1773, 1815]. Long-arm [1773]. long-term [1815]. longitudinal [2642]. longtermism [2499]. look [732, 756, 128, 2748]. looks [1678, 1282]. loophole [2829]. loops [2612]. loose [788, 2539]. Lord [2217]. Lorentz [1528]. lose [2862, 938, 775]. losing [2862, 3324]. loss [3381, 1182]. Lost [1756, 2493, 3107, 1251, 1250]. Lotka [2653]. Lotteries [938]. loudness [1879]. love [1997, 634, 1767]. Loving [1997]. low [3405, 1602, 2711]. low-quality [2711]. Luck [2400, 657, 910, 1785, 3102, 355, 1195, 123, 51]. Ludwik [2344]. Łukasiewicz [1978, 1882, 2054]. Luminosity [352, 2918]. luminous [1646]. lunch [1739, 1604]. Lying [1955, 2805, 2816, 827, 1315, 1209, 624, 2857, 3181, 1903, 2581, 2854]. Lyme [2663]. Lynch [3197]. Lyotard [2138, 3121].

Machine [2355, 2731, 3366, 1076, 2638, 2339, 1249, 1664, 2021, 2411, 1229, 2193, 3027, 999, 1844]. machines [2325, 670]. MacIntyre [1310]. macro [2292, 210, 3107]. macro-object [210]. made [436]. madmen [2002]. magic [485, 6, 235, 234]. mainstream [2673]. maintenance [2409]. majorities [1824]. Make [1135, 1941, 2656, 2790, 2920, 81, 2252, 1783, 2066, 2322, 833, 2166, 1392]. make-beliefs [1392]. make-believe [2252]. maker [2147]. makers [843]. makes [3175]. Making [3342, 3239, 960, 1998, 1655, 130, 682, 3038, 1529, 522, 1953, 1855, 2712, 397, 2041, 2557, 1017, 390, 141, 914, 299, 2872, 1702, 521, 3370, 2391, 1392, 3236, 2888, 589]. malfunction [3242]. mammalian [3298]. managing [2663, 811]. Mandevillian [2982]. manifest [1752, 3048]. manifestation [3063]. manifestations [734]. manifold [3132]. manipulability [1557, 1556, 3063]. manipulation [2857, 1658]. Manipulationist [3065]. manners [3359]. manufactured [759, 568]. Many [1859, 1308, 1637, 3159, 500, 536, 3304, 2714, 2104, 2040, 2090, 3101, 492, 538, 914, 1298, 1876, 89, 540, 1413]. many-one [492]. Many-Valued [1859, 1637, 536, 538]. many-worlds [2040]. Manyism [3331]. Maori [1359]. Mapping [1975, 2641, 3204, 1547, 1983, 1260]. maps [1337, 3204]. marginalization [1934, 1725]. Mark [2060]. markers [1214]. Markov [79, 799, 1462]. Markovian [1452]. marks [3113, 282]. marriage [1045]. Martin [1732]. masculinity [2630]. masking [2622, 663, 1439]. masks [2622, 1684, 2440]. mass [1065, 1622]. masses [703]. massive [945]. master [3129, 118]. masters [2650]. matching [2127]. Material [3219, 188, 3008, 580, 1540, 2595, 864]. material-origin [1540]. materialism [1676]. Materializing [2958]. Materials [2429, 1255]. Mates [2212, 3402]. Mathematica [1286]. Mathematical [2639, 462, 2273, 184, 1758, 1181, 1443, 1611, 739, 23, 2567, 577, 3115, 916, 1832, 1495, 2495, 615, 2656, 2303, 3084, 1865, 639, 732, 1405, 1936, 2186, 1972, 188, 579, 2825, 189, 2111, 158, 2490, 575, 1807, 2694, 650, 1175, 246, 517, 2585, 3128, 2785, 190, 2575, 35, 1449, 2078, 2785]. mathematician [2018, 1925]. Mathematicians [576, 109]. Mathematics [1623, 1310, 1813, 563, 1202, 578, 592, 2790, 2637, 2885, 1746, 581, 1321, 1352, 1241, 643, 1621, 2558, 326, 2458, 1226, 2152, 1178, 195, 1415, 1035, 1668, 2575]. Mathematizing [1321]. Matraversian [3075]. matrices [2456]. Matter [1544, 3166, 2275, 2156, 1087, 1562, 2211, 110, 3035]. matters [2101, 3008, 1317, 1322, 3152, 1635, 1634]. Matthen [2564]. maxim [1162, 1559]. Maximal [2156, 2275]. maximalism [87]. Maximality [40, 1230]. Maximin [49]. maximum [2425]. may [2665, 613, 3306]. McDowell [3097, 2984, 1023]. McKinsey [1817]. me [2490, 791, 1909]. mean [3046, 3016, 72]. Meaning [2920, 202, 3118, 1873, 2067, 2206, 2471, 2947, 1803, 1750, 3012, 3141, 2795, 3081, 2205, 2644, 41, 2267]. Meaningful [1248, 682]. Meaningfulness [2703]. meanings [2209, 820, 2137, 3216, 2257]. Means [2562, 2, 1963, 1050, 2024, 2352, 2484]. means-end [2352, 2484]. meant [2262]. Measure [696, 785, 1020]. Measurement [3264, 3342, 1217, 9, 2633, 646, 1685, 1116, 2382]. measures [2208, 1421, 2372, 401]. Measuring [2265, 1005, 541, 3176, 1281]. mechanical [1692]. mechanics [666, 2301, 1213, 3154, 3304, 2292, 1533, 155, 2040, 1791, 2178, 874]. mechanism [3239, 447, 2459, 1596, 1642, 1708]. mechanism-the [1642]. Mechanisms [221, 864, 1561, 1719, 565, 3256, 689, 2157, 2791, 2985, 1975, 2131, 2587]. Mechanist [1233]. Mechanistic [1439, 2873, 1131, 2766, 1130, 2835, 2543, 2026, 728, 1603, 2607, 2682, 1950, 976, 2181]. mechanistic-computational [1130]. mechanists [389]. media [2613, 2010, 2949, 2526]. mediation [1483]. Medical [1900, 2462, 3116, 1840, 2519, 2196]. Medicine [1336, 2594, 719, 2196, 2498, 2621, 1254]. medium [962]. meet [1411, 398, 2439]. meets [2569, 1956, 2318, 2851, 2927, 2376, 3210, 279, 2512, 306]. Megarian [1625]. Meinongians [1834]. melodies [1819]. Memberless [2336]. members [3346]. memorial [983]. memoriam [1408]. memories [3142]. Memory [2777, 3023, 293, 2852, 1586, 1063, 2822, 2862, 3255, 1329, 2615, 618, 66, 2155, 2868, 1011, 1501, 3031, 2912, 1784, 1888, 69, 3075, 2681, 1330]. Memory-based [3023, 2852]. Mental [1213, 2942, 2965, 2489, 1370, 805, 3086, 2897, 2582, 3267, 2080, 2517, 2922, 782, 3209, 3185, 1007, 3162, 3041, 2991, 3078, 3113, 3101, 295, 115, 767, 3036, 909, 116, 3204, 319, 2859, 2229, 2898, 740, 1610, 3194, 2760]. mentalistic [646]. mentality [3003, 1297, 1296, 2314, 2971]. mentalizing [3382, 3306, 720]. mere [3136, 849]. Merely [349, 3141, 1817, 65, 3382]. Mereological [236, 1565, 1957, 1394]. mereologicism [3331]. Mereology [918, 481, 485, 2943, 484, 412, 629, 1727]. merge [1833]. merging [2065]. merits [2203]. Merleau [1813, 387, 389]. Merleau-Ponty [1813, 389]. merry [1199]. MES [592]. mesoscale [1215]. message [2564]. Meta [534, 2695, 497, 2123, 3308, 3329, 2278, 2417, 1435, 1648, 2717]. meta-empirical [2123]. Meta-fictionalism [2695]. meta-induction [3308, 3329]. meta-inductive [1435]. meta-metaphysics [2417]. meta-ontological [497]. meta-ranking [2717]. meta-wisdom [1648]. metacognition [1470, 1888, 882]. metacognitive [882]. metaconceptualism [2363]. metaethical [2487, 114]. metaethics [32, 1890, 3301, 2918]. metalanguage [512]. Metalinguistic [510, 2088, 1416, 2354, 2619, 1984, 1810]. metametaphysical [1986]. metametaphysics [1597]. metaontology [621, 1271, 13]. metaphilosophy [1316]. Metaphor [891, 2386, 3060, 189]. metaphorical [2693]. metaphors [2885, 1581]. Metaphysical [368, 2390, 2891, 369, 65, 3050, 2367, 1636, 639, 2915, 344, 96, 1671, 2333, 2072, 343, 3240, 1254, 2006, 2280, 3350, 1204, 380, 211, 2078]. metaphysically [158]. metaphysician [1605]. metaphysicians [2503]. Metaphysics [1669, 3318, 2116, 3238, 2016, 2636, 2358, 2998, 97, 315, 3408, 15, 3050, 1486, 3061, 764, 633, 6, 875, 430, 96, 212, 3319, 2503, 2985, 2150, 95, 704, 373, 2531, 2866, 317, 417, 102, 2417, 2675, 1649, 2859, 2207, 172, 2153, 213, 436, 543, 3349, 2425, 235]. Metarepresenting [2697]. metasemantic [35]. Metasemantics [1652]. Metastable [1724]. metatheoretical [2446]. Method [3226, 1511, 2387, 254, 3297, 2450, 6, 235, 430, 3163, 1341, 59, 432, 1331, 55, 1950, 2906, 2308, 1686]. Methodological [433, 2358, 1366, 1130, 2876, 686, 1594, 1974, 2305, 876, 985]. methodologies [1133, 1761]. methodology [1133, 1341, 3122]. methods [1956, 53, 155, 2597, 2007, 1964]. Metrics [2881]. micro [2225, 3107]. micro-electrode [2225]. micro-macro [3107]. microaggressions [1574]. microbiome [679, 2216]. microscope [60]. middle [1215, 2520, 3274]. middle-out [1215]. middle-range [3274]. might [1427, 1428, 375, 2759]. migraine [2682]. Mill [3208]. Miller [2737]. Millianism [2072]. Millikan [3363]. Millikanian [1929]. Mills [462]. Milne [2577]. Mind [794, 2286, 3374, 1278, 1214, 2302, 1511, 1929, 2359, 1135, 1404, 3085, 2234, 3080, 430, 2104, 203, 965, 1577, 5, 1937, 799, 3172, 141, 2740, 7, 183, 111, 1242, 2056, 2004, 387, 1126, 2923, 662, 3351, 3227, 2114, 810, 1923, 1066, 71, 2463, 3034, 3078]. mind-body [3085, 5]. mind-changes [2359]. mind-dependence [1577]. mind-independence [1923]. mind-independent [2740]. mind-modelling [1242]. mind-wandering [1511]. mind-world [71]. minded [1609]. mindful [1609]. mindless [1609]. Mindreading [2092, 961, 1164, 873, 953, 1224]. Minds [3159, 3358, 3380, 3135, 1558, 806, 295, 1792, 1262, 1261, 1780, 177, 1117, 3190, 2935]. mindshaping [953]. mine [1325]. Mini [162]. Mini-symposium [162]. Minimal [1929, 3061, 1903, 3318, 972, 1894, 1162, 1272, 2275, 2156, 2211, 1136]. Minimalism [19, 712]. minimalist [1591, 98]. minimality [3256]. minimising [2756]. minimization [933]. minimize [2102]. mining [2785]. Minkish [239]. Minnesota [1405]. minority [839]. minute [3018]. Miracle [3272]. miracles [215, 970, 1075, 2385, 2909]. mirror [1707]. Mirrors [699, 1788]. mis [1339]. mis- [1339]. misalignment [2667, 2216]. misappropriation [2613]. misconceptions [920]. misfunctioning [3089]. misguided [3370]. misidentification [618, 3023, 1269, 1011]. misinformation [3357, 3059, 3407, 854, 3390]. misleading [1720, 3253]. mismatch [707, 201]. misrepresentations [2059]. misrepresenters [2059]. Missing [1375, 716, 692, 1214, 527]. mistake [3396]. Mistakes [734]. mitigating [1173]. mixed [2987, 3256, 922, 478]. Modal [3285, 1765, 381, 232, 485, 3401, 1068, 2378, 379, 1290, 910, 2353, 2995, 2467, 2350, 1240, 2943, 377, 1464, 1540, 374, 545, 435, 647, 373, 1059, 2162, 1174, 2069, 1027, 2054, 2460, 3006, 2813, 2379, 380, 372, 370, 2427, 841, 2512]. modalities [2464, 2500, 1026]. Modality [2540, 3136, 2721, 2739, 930, 3353, 231, 346, 1295, 1754, 345, 367, 343, 335, 372, 347, 99]. modalizing [232]. modally [748, 1362]. Modals [2465]. Model [1831, 1306, 2504, 1811, 2046, 2324, 1120, 463, 2653, 2591, 3405, 2465, 837, 535, 2272, 2076, 397, 2983, 2876, 2278, 1901, 3053, 2519, 197, 2006, 1048, 3356, 2775, 2369, 2989, 1606, 1197, 1224]. Model-based [1831, 1306, 1606]. Model-theoretic [1306, 2278]. Modeling [1552, 2313, 814, 3317, 2211, 1479, 1215, 1679, 1183, 3178, 2429, 1345, 1831, 2465, 1908, 2383, 2500, 2767, 2433, 1299, 2021, 560, 1893, 2227, 2396, 2512]. Modelling [665, 1516, 2456, 2969, 3151, 1618, 1863, 3074, 1242, 564, 1685, 708, 2379]. Models [1621, 524, 1463, 3112, 1179, 562, 3178, 2751, 2249, 529, 2256, 639, 2465, 139, 2894, 3270, 1184, 3024, 1130, 3118, 2605, 1272, 2650, 525, 2283, 2543, 85, 1102, 2203, 1851, 3278, 3217, 555, 2850, 1833, 1175, 2587, 2430, 2682, 558, 2211, 1774, 3070, 3075, 1860, 2222, 3259, 217, 2830, 864, 1275, 2475, 3365, 3390, 2935, 196]. Moderate [1555, 2011]. modern [578, 2954, 2803, 2903]. modernity [1826]. modes [2852]. modesty [747]. modular [2492]. modularity [945]. modulation [3104]. molecular [1719, 552, 2433, 728, 3091]. Monadic [2774]. MOND [2123]. Money [2572, 740, 2862]. monism [2294, 2761, 3199, 1452, 236, 1825, 3294, 2100]. Monist [3129]. monitoring [3376]. Monk [753]. Monothematic [3089, 1615]. monotonicity [1285]. Monster [1381]. Monstrous [2919]. Montessori [770]. mood [1681]. moods [1090, 906]. Moore [2542, 2687]. Moorean [1389, 1734]. Mooreanism [1476]. Moral [1987, 3258, 1574, 3212, 1241, 1912, 2064, 638, 3389, 3404, 1590, 777, 1422, 2233, 2160, 2720, 2684, 2858, 2555, 3275, 2883, 3061, 1647, 1225, 3373, 3143, 2612, 3279, 783, 3025, 2472, 1072, 3180, 3337, 3065, 2513, 603, 2793, 889, 1652]. Morality [3226, 783]. morally [2194]. morals [1837, 1027]. morass [1254]. morning [689]. morning-after [689]. most [2304]. motion [1808, 1719, 731, 2275, 2156]. Motivated [934, 3368, 2931, 2855, 2618]. Motivating [25]. Motivation [2131, 2869, 787]. Motivational [2855, 228, 2931]. motives [3399]. Motor [1861, 1062, 1813, 2875]. motors [1719]. movement [1330, 1329, 3245]. movements [2171, 1976, 616]. Moving [2623, 2859, 3037, 2558, 2673]. much [1928, 2115, 2490, 914, 2869]. Multi [1056, 66, 392, 2435, 564, 1126, 1950, 2054]. multi-agent [66, 2435]. Multi-field [1056]. multi-method [1950]. multi-person [392]. multi-scale [564]. multi-storey [1126]. multi-valued [2054]. multilayer [2543]. Multilevel [2601]. multileveled [2315]. multilingual [2021]. Multiple [2285, 3053, 133, 1148, 995, 1645, 549, 1960]. multiplexes [2313]. multiplicative [3250]. Multiplicity [980]. Multiscale [281, 1908, 2211, 1262, 1261]. multisensory [468, 786]. multiset [1160]. Multivariate [1726]. multiverse [2956, 2020, 1512, 3248, 642]. Musical [2740, 298]. Must [3032, 2493, 365]. mutilation [1162]. Mutual [2019, 1557, 1556, 1904, 2184, 1931]. My [1238, 2953, 1456]. Myth [1500, 2084, 1854, 719, 1690, 1626, 1450].

N [1950]. Naïve [3129, 1163, 3271, 2284, 1158, 776]. Naivety [3282]. names [272, 3012, 3251, 2370, 935]. narrative [761, 2720, 2684]. narratives [1321, 626]. narrow [957]. narrowmindedness [289]. nasty [544]. Native [2250]. Natural [2471, 2134, 2890, 1923, 2396, 1610, 428, 3375, 419, 418, 2393, 989, 1696, 627, 2520, 1778, 821, 928, 3064, 2754, 1003, 1319, 417, 822, 1050, 2867, 2425, 3351, 422, 108, 1304, 858]. natural-born [3351]. naturalised [316]. Naturalism [846, 2517, 2487, 390, 1595, 2807, 602, 2749, 659, 1652]. naturalist [2763]. Naturalistic [3238, 777, 2408]. naturalists [1358]. naturalize [2474, 2419]. Naturalized [2503, 3288, 61]. Naturalizing [3185, 1446, 3289, 2446]. Naturalness [2926, 461, 420, 1857]. Nature [2238, 1312, 2416, 1926, 1367, 3135, 2449, 1613, 2031, 1595, 1836, 986, 1023, 2913, 1271, 563, 2585, 533, 532, 2968, 1386, 1635, 1634, 2760, 783]. Natures [2996]. Navigating [2663]. Near [2607, 2623, 2397, 2483]. near- [2397, 2483]. near-bias [2623]. Near-death [2607]. nearness [1633]. necessary [1100, 863, 1074, 16, 2938, 1613, 371, 158, 3394, 1386]. necessitarianism [1285]. necessitation [2099, 1600]. Necessity [3135, 2556, 1919, 344, 369, 1945]. need [2301, 1371, 1736, 925, 1531, 1480, 2633, 1289, 1374, 1358]. needed [3281]. needs [415, 2492, 2161]. negation [911, 540]. Negative [759, 18, 147, 2831, 2784]. negativity [2949]. neglected [2493]. Negotiating [2088]. negotiation [510, 2354, 2619]. negotiations [2088]. Neighborhood [1550, 2871]. neighbors [2187]. Neither [1609, 1117]. Nelson [3315]. Neo [413, 3412, 244, 1479, 412, 1096, 14]. Neo-Aristotelian [413, 412]. neo-Carnapians [14]. neo-Fregeans [1096]. neo-Gricean [1479]. neo-Humean [3412]. neo-Lewisian [244]. Neopragmatism [2632]. neopragmatist [3185]. nervosa [3116]. nested [1401]. net [803]. network [2415, 2591, 1847, 2171, 2076, 3339, 2543, 3360]. Networks [2752, 1542, 1679, 1183, 2256, 679, 2766, 81, 1975, 1983, 1378, 2940, 912, 590]. Neural [2752, 910, 2743, 101, 1948, 2415, 1488, 1131, 2157, 1726, 2766, 186, 185, 1975, 3123, 1378, 1397]. neurocognitive [1213, 3114]. neurodivergence [2451, 1725]. neurodiversity [1725]. neurodynamics [1279]. ‘NeuroGate’ [3080]. neurogenesis [3298]. Neuroidealism [3052]. neuroimaging [1431]. Neuromedia [1845]. neuromorphic [3080]. neuronormativity [3037]. neurophilosophy [1289]. neuropsychoanalysis [3322]. neurorights [3080]. Neuroscience [1277, 3243, 1236, 1849, 1198, 1720, 2592, 2743, 1259, 1908, 2641, 2682, 1349, 279, 2830, 2225, 2181, 795, 3211, 1962]. neuroscientific [3017, 2978]. neurotic [3108, 3066]. neutral [512]. neutrality [2761, 2367, 2386, 2598, 2485]. never [2346, 3218]. Newcomb [584, 1395, 587]. Newcomb-like [587]. Newness [1807]. news [2635]. Newton [3030]. Newtonian [464, 703]. Neyman [3122, 848]. niche [3094, 71, 871]. Niches [2598, 768]. Nicholls [2217]. Nicole [1241]. Nieuwentijt [246]. nihilism [904, 1994, 214, 1852]. nineteenth [837]. No [1739, 2884, 919, 1338, 2149, 491, 420, 2193, 1522, 742, 2174, 1883, 1371, 215, 970, 955, 1075, 3281, 2385, 1989, 808, 1971, 3189, 1489, 2909, 1604, 110, 874, 2189, 1491, 1878, 3034, 2314, 3272, 452]. No-Belief [452]. no-collapse [874]. no-free-lunch [1604]. No-Miracle [3272]. no-miracles [1075]. no-platforming [1489]. Nobody [126]. noetic [1657]. noise [884, 3079]. nomic [3318, 3135, 85, 1523, 2372, 3391]. Nominalism [1565, 822, 2182, 659, 2865]. Nominalist [1935]. nominalists [2528, 2041]. Non [777, 2578, 3082, 3008, 901, 1442, 8, 1546, 1513, 1429, 2416, 65, 668, 1848, 459, 3124, 691, 2695, 3042, 956, 1230, 1328, 1387, 1076, 1127, 992, 819, 2487, 2828, 2797, 2979, 2020, 1317, 453, 2304, 1289, 760, 2021, 538, 157, 2626, 3271, 1313, 1383, 1382, 2038, 726, 3204, 2202, 1984, 1810, 156, 866, 638, 2708, 2166, 394, 2334, 190, 2907, 288, 855, 2164]. non-accidental [992, 819]. non-action [2202]. Non-assertoric [3082]. non-belief [453]. Non-Boolean [901, 1442]. non-causal [760]. non-causalist [3042]. Non-classical [1546, 2164]. non-conceptual [2304, 157, 3204]. non-conformism [8, 855]. non-consensual [2334]. non-deductive [1313]. non-derived [190]. non-deterministic [538]. Non-dynamism [2578]. non-empirical [459]. non-epistemic [668, 1076, 1317, 2038, 394]. non-eternalism [1429]. non-existence [1387]. non-extensional [1383, 1382]. non-fundamental [726]. non-governing [2020]. non-human [2907]. non-ideal [1127, 2626]. non-living [2416]. non-maximality-solution [1230]. non-metalinguistic [1984, 1810]. non-moral [638]. non-naturalism [2487]. Non-naturalistic [777]. non-paradoxical [2166]. non-parallel [2021]. non-philosophers [2828, 2797]. non-present [2695]. non-qualitative [691, 156, 866]. Non-rational [8, 855]. non-reductive [1289]. non-representational [288]. Non-spatial [3008]. Non-standard [1513]. non-substantiality [3124]. non-substructural [3271]. non-summativism [1848]. non-symmetric [1328]. non-traditional [956]. non-universal [2979]. non-verbal [65]. non-words [2708]. noncausal [1034]. nonclassical [629]. noncommutative [1375]. nonconceptuality [2885]. nondescriptivist [3267]. none [2714]. nonexistent [713]. nonhuman [3383]. nonlinguistic [3072]. nonmonotonic [1157]. nonphilosophers [1941]. nonsense [2454]. nonseparability [206]. nonsingular [272]. nontransitive [1157]. nor [1609, 1117]. norm [668, 3214, 175, 2502, 1464, 2034, 1730, 1815, 1406, 2409, 1592, 3395, 324, 1266, 1426, 1107]. norm-maintenance [2409]. Normal [2844]. normality [1103]. Normalizability [1154]. normalization [1680, 2577]. Normative [2927, 2234, 169, 2733, 889, 103, 657, 1739, 936, 2393, 2720, 2684, 1427, 1428, 2272, 2955, 2075, 2877, 2397, 2483, 3352, 1532, 1327, 1220, 2443, 76, 890, 644, 3286, 2267]. normativism [3267]. normativists [647]. Normativity [2118, 2644, 513, 256, 2625, 3214, 3185, 1905, 3162, 1165, 1971, 299, 813, 2131, 3259, 2833]. Norms [2476, 1955, 31, 2281, 1997, 1317, 2179, 3192, 3184, 2910, 3352, 2799, 2951, 888, 2561, 452, 987, 2189, 946, 710, 2888, 2287, 836]. norms-based [2561]. Norwood [818]. notations [1052]. note [1142, 3191, 517, 105, 2820]. notebook [794]. Notes [1324]. nothing [229, 168, 1494, 1678, 1282, 2869]. notion [804, 2837, 539, 1714, 105]. notions [531, 343, 208, 1516]. Noumenal [3247]. nourish [736]. Nouveaux [1241]. Novel [2711, 468, 1848, 2270, 1377, 1836, 2021, 3189, 1026, 3034]. novelties [969]. novelty [3366]. Nudging [2326, 3286]. number [2865, 1575]. Numbers [2643, 2520, 3, 11, 199, 791, 327, 2497, 219, 74]. Numerals [185, 191, 328, 186]. numerical [1052, 2461, 3030, 303]. nurture [783]. nutrition [406].

obedience [987]. obesity [2519]. Object [2461, 1068, 786, 2766, 2143, 3328, 733, 210, 1762, 1960, 3034]. object-centric [2143]. object-directed [1068]. object-seeing [733]. objecthood [1457]. Objectification [1368]. Objection [1904, 1812, 1285, 11, 74, 1853, 2460, 370, 1931]. objections [1266, 2184]. Objective [1122, 1168, 1007, 44, 1086, 197, 2724, 3221, 2581, 886]. Objectivity [1173, 2796, 1185, 1270, 311, 2080, 1217, 1587, 1252, 1691, 1348, 1238]. Objects [733, 3318, 3115, 3008, 1192, 1142, 2604, 1394, 1741, 2087, 2953, 2646, 661, 1245, 1449]. objectual [2901]. obligation [1900, 2462, 557]. Obligations [3355, 1933]. obliviousness [2572]. obscuring [2517]. observables [3336]. observation [1731]. Observer [618]. Observing [1484]. obsessive [897]. obsessive-compulsive [897]. obsolescence [2382]. obstacle [265]. obstruction [2310]. Obtaining [540]. occlusion [2864]. occupies [2938]. occurrence [2803]. Ockham [3327]. octopus [1135]. odds [2405]. off [1538, 961, 1877, 2279]. office [2938]. offices [2938]. offloading [3156, 1845]. offs [2351]. ok [1001, 2606]. old [2555, 84, 2839, 2810]. oldest [986]. olfaction [3021, 1354]. olfactory [1691, 2117, 629]. omission [1511, 2202]. omissions [1541]. Omnipresence [3403]. omniscient [2801]. on-shell [3048]. Once [2948, 365, 2386, 2764, 1379, 231]. One [1876, 2158, 1576, 3002, 492, 141, 1126, 1772, 594, 1960, 500, 2612, 2461]. one-level [2158]. ones [954]. online [1793]. only [152, 563]. ontic [1398, 2985, 3294, 2024, 1064]. Ontological [12, 1290, 497, 2121, 320, 1925, 3263, 336, 2102, 94]. ontologically [2673]. ontologies [2318, 1671]. ontologist [685]. Ontology [1378, 1777, 1463, 40, 2324, 1065, 3051, 411, 3265, 98, 2234, 621, 2157, 2800, 715, 2722, 2449, 2180, 2763, 413, 2589, 1596, 998, 1271, 2087, 2006, 2811, 29, 1732, 1326, 2508, 3054, 2408, 1112, 2568, 73, 71, 990, 870]. Opacity [1673]. opaque [3189]. Open [3120, 2120, 911, 2896, 673, 683, 1342, 3091]. open-ended [911]. openness [1900, 2462]. operationalism [1629]. operative [3346]. operator [2175, 2727]. operators [3401]. opinions [2151]. opportunity [2208]. oppositions [1835]. oppression [2092, 3133]. optimal [201]. optimal-environments [201]. optimality [82, 3208]. optimality-argument [82]. optimistic [3308, 3329]. optimization [1435]. options [896, 2032]. oracles [3189]. orbit [3261]. Order [3084, 2834, 2791, 3059, 3326, 2946, 1952, 3202, 1605, 1114, 2827, 2577, 1098, 2427, 273]. ordered [506, 608]. ordering [588]. Ordinary [2649, 2387, 15, 2450, 2832, 3077, 636, 1038, 2406, 126, 2673, 2513, 244, 2529, 2514, 2718]. organic [3006, 2813, 1351]. organicism [447]. organism [2873, 3220]. organisms [1462]. organization [3084]. Organizational [2625, 1990]. Organized [2704]. orgasm [1270]. oriented [809]. origin [1540, 767]. original [617]. Origins [3028, 497, 837, 961, 1575]. orthodox [2301, 2183]. orthonasal [1354]. Ostrich [1797]. Oswald [1932, 1805]. Other [1117, 307, 3411, 470, 218, 96, 295, 1267, 3268]. Ought [1302, 1372, 784, 77, 630, 2, 2179, 1963, 2129]. Ought-contextualism [1302]. ought-statements [784]. oughts [1585]. our [689, 2650, 776, 1368, 1516, 164, 3324]. ourselves [1516, 3399]. outcomes [3366, 504]. outer [2572]. outline [3220]. Outlines [3412]. outright [1526]. Outward [2295]. Outward-facing [2295]. Overcoming [1823]. overconfidence [983]. overdiagnosis [3174]. overfitting [2853]. Overlap [2928, 1957]. Overlapping [2747]. overlooked [2984, 2446, 944]. oversight [3173]. overt [1982]. overview [2366]. owe [218]. own [393, 2135, 141]. ownership [2899].

P [405]. P-word [405]. PA [2629, 2629]. package [1212]. Pain [1517, 1340, 1527, 92, 2685, 3332]. pairing [416]. Paleoclimate [2546]. paleodiversity [562]. paleoecological [2895]. Pandemics [2220]. Pandispositionalism [2150]. panicky [2531]. Panpsychism [2023, 2774, 2091]. panpsychist [2276]. paper [580, 928]. papers [88]. paraconsistency [1097, 2228, 1361, 1472]. paraconsistent [1097, 540]. paradigm [534, 2988, 2306, 2037, 502, 1413]. paradigm-based [2306]. paradigm-level [502]. paradigms [276]. Paradox [1493, 2947, 1151, 87, 1862, 785, 4, 182, 1819, 2737, 1528, 2556, 1951, 1027, 546, 1722, 1161, 438, 1154, 1817, 365, 1878, 1150, 1859, 2687, 3271, 3293]. Paradoxes [330, 3316, 2634, 1389, 52, 2929, 1194, 2787]. paradoxical [2437, 2166]. Paradoxicality [1850]. parallel [2021]. parallelism [2899]. Parameterization [2433]. parameters [398]. paraphrases [1383, 1382]. Paraphrasing [1067]. parentheticals [1755]. Pareto [3208]. parity [3238]. parsimony [1650]. Part [412, 1510, 486]. parthood [494, 483]. Partial [830, 2570, 86, 349, 753, 90, 652]. participants [3062]. participatory [2888]. particle [3112, 1065, 155, 3048]. Particles [2324, 1116, 1808, 966, 3166]. particular [3312]. particularism [2226]. particularity [2757]. particulars [1690, 2229]. Partisan [2664, 2618]. partly [2050]. parts [785, 2718, 2649, 498]. passage [1902, 1842]. Past [748, 1063, 2757, 673, 1087, 1501, 2787, 3057, 2683, 2950]. past-perception [1063]. pastness [1888]. patchwork [2995, 2787]. paternalism [3275, 3116]. path [160, 2616]. Pathological [1351, 3116]. paths [1687, 1149]. Pathways [1384]. Patience [3119]. patient [401, 3173]. patients [505]. pattern [1726]. Patterns [2046, 2211, 779]. Paul [2575]. Paying [2779]. pays [2025]. PDA [1212]. Peano [2093]. Pearson [848]. pebbles [1462]. peculiarities [2911]. peer [222, 2668, 2846]. Peer-undermining [2846]. peerhood [730]. Peirce [620, 2250, 300, 1419, 2054]. Peircean [1729]. pejorative [996]. pejoratives [880]. penetrable [1356, 479]. people [1509, 1061]. perceive [826, 115]. perceived [1582]. Perceiving [899, 2928]. Perception [316, 475, 2492, 1487, 1063, 193, 2998, 3408, 480, 1488, 1075, 391, 1192, 1258, 362, 1047, 902, 786, 3328, 1691, 750, 280, 2808, 1906, 1021, 3233, 1187, 1780, 3393, 473, 962]. Perceptual [1500, 1663, 1776, 2770, 2901, 1396, 2864, 862, 981, 3052, 286, 1768, 902, 2912, 2564, 906, 2526, 3157, 937, 2810, 1245, 1026]. Perceptualism [1327]. perceptually [1960]. perdurantism [1072]. perfection [120]. Performance [256, 1503, 1918, 1367, 81, 1364, 298, 1359, 1971, 1532, 3306, 1456, 1609, 1506, 268]. Performative [2586, 2767]. performatives [547]. perils [3054, 2216, 1270]. period [101]. periodic [593, 1019, 1628]. Peripersonal [473]. periplanone [2240]. Permian [1622]. permissible [2194]. Permission [2126]. permissions [3355]. permissive [1871, 1271]. permissivism [2081, 2045]. perpetration [2330]. Perry [2983]. perseverance [1845]. persist [907]. Persistence [3043, 345, 3050, 2657, 275, 1044]. Person [2136, 2676, 392, 401, 2682, 944, 1482, 1481, 2426, 1749]. personal [242, 1454, 2000, 111, 2267]. personality [1941, 596, 3245, 1169, 953]. persons [2719]. Perspectival [3122, 2925, 3364, 439]. Perspectivalism [2571, 3285]. perspective [348, 2220, 1110, 585, 770, 2255, 2556, 3220, 2985, 3360, 1143, 1132, 222, 1796, 369, 2887, 2368, 3309, 1608, 882, 3137]. Perspectives [1036, 2676, 1590, 509, 117, 75, 1886, 430, 1338, 276, 2231, 13, 3128, 88]. perspectivism [1104, 3364, 570, 2396]. perspicuity [1914]. persuasion [1827]. perturbations [878]. Perturbative [3302]. pessimism [2931, 2855]. pessimistic [1086]. pharmaceutical [3168]. phase [910]. phenomena [565, 2831, 2784]. Phenomenal [1527, 1646, 2808, 198, 2114, 1607, 619, 1068, 905, 2127, 3196, 2905, 3194, 1878]. phenomenality [289]. phenomenological [20, 46]. Phenomenology [390, 1410, 385, 1422, 782, 383, 735, 3381, 1842, 46, 3093]. phenomenon [357]. Phi [2470]. Philosopher [1812, 1715]. philosophers [1941, 1711, 2828, 2797, 3295, 112, 791]. Philosophical [53, 60, 3110, 1917, 2228, 2300, 40, 2387, 2693, 3297, 693, 2030, 2264, 596, 1669, 430, 1660, 3203, 1824, 3202, 2673, 2642, 1331, 1858, 670, 2880, 2514, 2898]. Philosophicus [1530]. Philosophy [596, 1568, 1941, 334, 193, 3338, 2387, 1073, 1956, 1978, 2138, 2450, 2414, 1794, 2954, 1847, 1799, 1977, 2455, 3061, 636, 1038, 1812, 2406, 1976, 2915, 430, 126, 1966, 1006, 2424, 393, 2754, 2349, 1444, 58, 93, 3122, 704, 61, 1299, 2021, 1226, 1331, 3352, 1829, 814, 3374, 1983, 3148, 55, 1384, 2529, 3145, 738, 3343, 321, 3406, 1893, 88, 659, 404, 2575, 2512, 2078, 1014, 2463, 2717, 1889]. phylogenetic [1640]. Physical [3030, 217, 99, 1297, 1296, 3231, 639, 1968, 242, 466, 2149, 2026, 1851, 2913, 2948, 2780, 642, 2369, 949, 3221, 2838]. Physicalism [2294, 3003, 1650, 1891, 2314, 3308, 3329, 381, 3323]. physicalist [1108]. physicalists [2158]. physicality [2080]. physicist [2877]. Physics [2802, 2456, 1346, 3112, 1215, 4, 182, 1653, 1208, 227, 458, 155, 1746, 2554, 461, 3048, 2275, 2156, 446, 1707, 1199, 1380, 1222, 850, 2079, 2153, 213, 2420]. Piagetian [1608]. pictorial [1258]. picture [3136, 2863]. pictures [2047]. Pierre [1241]. pill [689]. PKF [2343]. place [1317, 659, 164]. placed [395]. placeholder [1037]. Plain [2994]. plains [1405]. Plan [1347, 1712]. Planetary [3248]. planning [2435, 1401]. plant [2232, 749, 2705]. plants [2232]. platforming [1489]. Platonic [681, 1260]. Platonism [615, 2273, 3128]. plausibility [135]. plausible [1278, 2892]. play [3134, 2302]. playing [2858]. plea [754, 3151, 725, 2589, 3352, 1924]. pleasure [3131]. Plenitude [3273, 3007, 342]. Pleonastic [64]. Plural [821, 998]. Pluralism [2595, 2986, 520, 1561, 2713, 513, 2220, 518, 2994, 679, 1166, 514, 515, 519, 1088, 2176, 2104, 1839, 2659, 510, 1465, 2531, 692, 1129, 2177, 1243, 1717, 3036, 2165, 517, 1950, 2100, 516, 2698, 2428, 1974, 737, 1324, 2631, 836]. Pluralist [771, 893, 892, 2747, 2024]. Pluralistic [1239, 509, 2878, 829, 3078, 749, 985, 404]. pluralists [1717]. pluralities [1042, 1525, 487]. plurality [1122]. Plurivaluationism [89]. pluriverse [1067]. plus [874]. poietic [1430]. point [1803, 1375, 2610, 644, 727]. pointillism [1630]. Pointless [3192]. Points [439, 1649, 1353, 2234, 90, 2279, 1877, 19, 1864]. Poiseuille [1674]. polarisation [1614]. Polarization [591, 3040, 2191, 2741, 908, 3234]. Poldrack [1948]. poles [1332]. policing [2510]. policy [2220, 397, 1402, 1815, 568, 103]. policymaking [1402]. politeness [2940]. Political [2590, 1318, 3061, 218, 2502, 1118, 2933]. politics [934]. pollution [3232]. Polycratic [1183, 1679]. polyhedra [739]. polysemous [2134]. polysemy [2753, 486]. Ponty [1813, 387, 389]. pooled [788]. pooling [90]. Popper [2737, 3343]. popular [2568]. population [1612, 3062]. portrait [791]. poses [3189]. Position [261, 1806]. positions [1710]. positive [2179, 3126, 2628, 2573, 2943]. Positively [3253]. Possessing [1300]. Possibilities [244, 2861, 3015]. Possibility [1818, 1429, 698, 50, 368, 2505, 3008, 634, 2628, 2573, 642, 1362]. Possible [2736, 3097, 2188, 2764]. post [227, 294, 293, 2872, 3351]. post- [3351]. post-cognitivism [294]. post-Kantian [227]. post-self-deception [2872]. posteriori [1988, 371, 62, 3323]. postgenomic [1407, 1060]. posthuman [3351]. posthumanist [2286]. postulate [874]. postulates [2015, 85]. potency [410]. Potential [2712, 301, 2149, 2705, 3260, 886]. potentials [2705, 3216]. poverty [1841]. power [2186, 1863, 2458, 1325, 2951, 2475, 177]. Powered [2995]. powerful [960, 1631, 2647]. powers [1884, 1288, 1600, 2150, 1596, 2231, 2647, 1081, 342, 73]. Practical [1871, 926, 1268, 1508, 2154, 1526, 2810, 1237, 1464, 973, 1406, 2899, 2669, 1883, 1397, 3167, 2139]. practical/epistemic [2899]. practicalism [3114]. practice [739, 1371, 1856, 1190, 2552, 2876, 137, 188, 2018, 579, 1321, 504, 1851, 1999, 1273, 1758, 194, 2785, 1181, 2575, 2533]. Practices [1329, 1611, 2613, 2028, 961, 1587, 2500, 1990, 1640, 192, 2109, 1330]. practicing [1999]. Pragmatic [2579, 1214, 2699, 1463, 1207, 2847, 2522, 2491, 2257, 2209, 129, 3224, 2633, 3306, 1706, 1195, 136, 2443, 1358, 990]. pragmatic-trial [2095]. Pragmatics [2077, 2753, 357, 21]. Pragmatism [294]. Pragmatist [2788, 2962, 212, 2364, 1896]. pragmatists [1839, 2027]. prairies [1405]. Praise [2724]. praxeology [443]. Pre [613, 80, 1397]. pre-awareness [1397]. Pre-emption [613, 80]. precautionary [1552]. Precedent [2824]. precipice [159]. precise [1998]. precise-Imprecision [1998]. predetermination [840]. predicate [2556]. Predicates [1454, 2267, 2224, 1704, 3153, 27, 310, 2848]. Predication [429]. predicative [1710, 1995]. predictability [1819]. predicted [3149]. Prediction [1198, 2094, 2988, 933, 2398, 216, 728, 7, 183, 2711, 1351]. predictions [2131]. Predictive [1672, 362, 2510, 1792, 91, 3052, 2686, 933, 877, 1294, 3319, 2082, 2241, 3104, 2131, 2587, 2669, 1021, 787, 1066, 2060]. predictivism [3105]. preemption [205]. Preface [1493, 365]. preference [2191, 1466]. preferences [753, 2946, 2381]. preferred [2040]. prefix [1492]. Pregnancy [2758]. preliminary [1715]. Preparedness [1167]. Preregistration [3105, 3343]. Prescriptive [3296]. presence [1068, 1281, 286]. Present [2950, 1281, 2695, 1005, 1086, 1342, 748, 2690]. presentation [1500, 386, 2852]. presentations [1960]. Presentism [2340, 151, 964, 2904, 224, 1344]. preserving [523, 521]. pressures [3209]. prestige [952]. presumption [1451]. presumptions [985, 867]. Presumptuous [985]. presupposition [717, 1714]. presuppositions [1402]. pretend [2302]. Preteriception [1063]. preventers [2582]. Preventive [1840]. Price [2601, 1707]. prima [1244]. primacy [1950]. primer [2761]. primitive [2197, 3335]. primitivism [128, 1149, 370]. Prince [238]. Principia [3030, 1286]. Principle [2132, 1552, 754, 1075, 447, 2995, 2594, 2554, 3208, 442, 1165, 3220, 3012, 711, 3313, 2203, 446, 2007, 1964, 2890, 2587, 546, 176, 2425, 1262, 1261, 1516, 2989, 597, 2430]. principles [1943, 2755, 1993, 447, 448, 3231, 227, 444, 440, 441, 1578, 915, 2787, 1923, 3010, 1207]. principles-a [3231]. Prior [2468]. priori [16, 958, 442, 2394, 523, 547, 62, 1665, 2373, 1817, 1485, 2309]. priori/a [62]. priorities [2414]. prioritised [553]. prioritization [394]. Priority [3199, 875, 2769, 3000]. priors [2959]. Pritchard [1015, 1675, 450, 815]. privacy [2271]. private [3017]. privilege [637, 1806]. privileged [2673, 1735, 2690]. Pro [2136]. Pro-Diversity [2136]. Probabilifying [2781]. probabilism [765, 2223, 2457, 1559]. Probabilistic [1353, 2871, 1435, 1567, 2471, 1666, 701, 2922, 1579, 2942, 1523, 1624, 1892, 909, 541, 889, 2079]. probabilistically [2782]. probabilities [465, 1546, 625, 2164, 3335, 2254, 1771, 3221, 1335]. Probability [1305, 1070, 579, 1435, 874, 551]. probably [2, 1963, 1842, 1876]. Probing [1440]. problem [2301, 1913, 3125, 2987, 3256, 910, 286, 1768, 2415, 1632, 2365, 1740, 1425, 2239, 631, 1075, 1591, 1803, 28, 653, 690, 3085, 2183, 3116, 732, 9, 2741, 1470, 2633, 2040, 2094, 724, 272, 945, 2932, 1018, 453, 1441, 5, 1937, 2856, 986, 3101, 2149, 1145, 1395, 1146, 707, 1766, 1664, 1502, 200, 808, 210, 646, 416, 2024, 1010, 1512, 2839, 1758, 909, 3261, 1638, 3334, 840, 1108, 89, 2711, 1092, 1091, 2124, 251, 644, 372, 2683, 1960, 1439, 3106, 1413, 2546, 3087, 3107, 2752, 2846, 261]. problem-solving [1758]. Problems [3160, 3316, 1731, 2101, 584, 290, 585, 1669, 2157, 2632, 2226, 1308, 80, 663, 2026, 326, 2000, 143, 587, 3189, 1680, 2518, 2792, 2865, 508]. procedural [2453, 2122]. procedure [2577]. Process [2046, 2456, 529, 1094, 1685, 651, 549]. Processes [2710, 2378, 2840, 80, 574, 1631, 1430, 3317, 1368, 601]. Processing [2060, 3052, 2686, 877, 1672, 362, 386, 3319, 2082, 2241, 2587, 1021, 2704, 787]. processors [2669]. processual [2315]. production [2425]. Productive [943, 826, 878, 1508]. professional [3209]. profile [2378, 3006, 2813]. profundity [3077]. progeneratively [1954]. program [1978, 1307]. programmes [1133]. progress [1657, 1786, 2509, 2349, 1273, 2443, 2933, 436]. progression [164]. progressive [601]. prohibitive [1871]. project [3140, 2406, 126, 2113, 3022]. projective [1457]. projectivism [1200]. Prolegomena [1226]. promiscuous [2822]. promise [3300, 2742, 1668]. promises [3017, 1190]. Promising [3087]. promote [2868, 1939, 2341]. promoting [2778]. Proof [514, 2073, 333, 577, 3316, 617, 1170, 2332, 264, 1268, 2908, 2120, 2523, 3362]. Proof-theoretic [514, 2073, 333, 3316, 617, 2332, 2523]. proofs [23, 2567, 2250, 2342, 2563]. propensities [2879]. propensity [2746]. Proper [3321, 3012, 2732, 676, 3179, 361]. properly [775]. Properties [903, 485, 489, 2582, 2995, 1752, 735, 2997, 1646, 1067, 2180, 3048, 2304, 611, 1605, 2516, 1393, 1212, 234, 979, 603, 1722, 309, 156, 866, 3327, 2647, 2643]. property [2528, 1632, 312, 2041, 1152, 306]. prophecies [2805]. Proportional [1704]. proportionality [1870, 600]. proposal [1898, 747, 1612, 3257, 98, 1894, 79, 2364, 2876, 799, 1563, 2153]. proposals [3155]. proposed [2583]. proposition [3344]. Propositional [727, 2188, 1580, 802, 2014, 330]. Propositions [1391, 1979, 1886, 272, 1365, 2445, 815, 246, 2151, 153, 1071, 64]. Propping [1874]. Proprietary [2205]. Propter [2754]. Prospecting [1367]. Prospects [283, 507, 956, 923, 455, 326]. Prosthetic [879]. protection [3263]. Prototypes [1332]. provability [2548, 2584]. proved [762]. provide [1381]. Proving [974]. provisional [1431]. Proxies [2752]. proximate [3146]. proximity [1070]. proxy [860, 1959, 1431]. prudent [1548]. prudential [2921]. pseudo [3395]. pseudo-questions [3395]. pseudoscience [2173]. psychiatric [505, 1366, 2707, 2219, 3182]. psychiatry [3133, 3041, 3174]. psycholinguistic [1812]. psychological [1855, 844, 3042, 2513, 1361, 2779, 1608]. psychologism [1882]. psychologists [112, 1211]. Psychology [1239, 2911, 3258, 554, 383, 1778, 829, 902, 294, 165, 3364, 1609, 1860, 1629, 953]. psychopathology [2313]. Psychophysical [2761]. Psychophysics [2549, 383]. Psychophysiological [546]. psychotherapy [2860]. Public [505, 2666, 50, 1794, 506, 3056, 504, 1402, 3341, 1254, 2936, 3009]. publication [2018]. publics [3056]. Publisher [2522, 3329, 2621]. punctuated [2076]. punish [2025]. puns [2866]. pure [158, 2867, 2204, 874]. purely [2938, 420, 909]. purism [1310]. purport [1021]. purpose [2978, 475]. purposes [2895]. purposiveness [3135]. pursuit [3140, 459, 3300, 595, 3087]. pursuit-worthiness [3300, 3087]. pushing [499]. put [546]. Putnam [2278]. Putting [1930, 504, 373, 2338]. Puzzle [2071, 1590, 470, 1728, 2801, 3143, 1438, 2027, 823, 1004, 2892, 2953, 1693, 1830, 1906, 995, 1606, 3117, 3402]. Puzzles [2077, 2158, 2143, 214]. puzzling [2006]. Pyrrhonist [2911].

QFT [2633]. qua} [1334]. quadratic [974]. quagmire [741]. Qualia [2566, 3131]. qualitative [691, 1950, 156, 866, 1168]. qualities [3225, 960, 316, 2080, 2647]. quality [2711, 3176, 1014]. Quantification [2611, 2061, 2134, 2467, 715, 1961, 870, 2968]. quantificational [2970]. quantified [2194, 2162]. quantitative [1073, 1917, 791, 2529]. Quantity [194]. quantization [1064]. Quantum [3033, 2314, 3257, 1654, 206, 1009, 2203, 1630, 414, 211, 1597, 949, 551, 666, 2301, 2456, 3154, 207, 966, 1639, 3304, 1417, 448, 1138, 1144, 1533, 1743, 2997, 155, 2040, 2180, 1148, 413, 1145, 1146, 2178, 102, 197, 45, 2079, 3169, 213, 886]. Quarc [1074]. quarrel [1992]. Quasi [537, 936, 2180, 2403, 114, 3188]. Quasi-canonical [537]. Quasi-Conventions [2403]. Quasi-realism [936, 114]. quasi-realist [3188]. Quasi-set [2180]. QUD [1544]. QUD-shifting [1544]. Queer [3411]. quest [3201]. question [63, 3397, 488]. question-begging [63]. question-sensitive [3397]. Questioning [2418, 683, 2277, 1402, 2124]. questioning-attitude [2277]. Questions [3071, 1816, 1716, 3395]. quibbles [741]. Quick [1434]. quid} [2754]. Quiddities [1995, 3336]. Quine [2086, 1744, 659]. quit [1381]. Quotative [3130]. quote [2716]. quote-tweetable [2716].

race [2318, 2673]. racial [125, 396, 3062]. racialised [2315]. racist [127]. Radical [384, 2769, 290, 1455, 3294, 304, 284, 2773, 480, 1911, 1189, 2011, 2590, 390, 1502, 152, 1532, 387, 1134, 738, 279, 278, 962, 289, 293]. radicalized [1330, 1329]. Radicalizing [303]. radically [305, 288]. radio [1715]. Ramsey [1089, 1944, 216, 1286]. random [679]. randomness [3084, 3236]. range [97, 3274, 2187]. Ranking [2717]. rapid [1564]. rather [2020, 2799]. Rational [120, 2878, 701, 2782, 1466, 2391, 2128, 125, 3135, 2590, 250, 1753, 2279, 1877, 8, 2049, 431, 567, 589, 855]. rationalism [2845, 377, 381, 2166]. Rationality [22, 987, 1701, 2897, 2712, 2878, 1363, 2851, 1388, 891, 3219, 1083, 745, 2399, 2272, 3208, 442, 453, 2122, 658, 1355, 2144, 959, 2439, 2474, 2419, 305, 1098, 2002, 1077, 934]. rationalization [3399]. Rationally [3300, 121]. Rawls [2944]. Rawlsian [2747]. Razor [1491, 3327]. Re [1947, 2580, 3088, 54, 300, 386, 1587, 356, 526]. re-appraisal [526]. re-embedding [1587]. Re-engineering [1947, 54]. re-examined [300]. re-examining [356]. Re-remembering [3088]. reach [937]. reaction [1286]. read [1962]. Reading [622, 687]. Real [779, 1897, 2897, 1724, 2807, 3358, 2497, 436, 284]. real-life [1724]. Realism [2660, 570, 2610, 718, 2320, 324, 1362, 3035, 3336, 3238, 936, 2847, 405, 2353, 2592, 1142, 2984, 771, 427, 3308, 3329, 437, 114, 3122, 1596, 1716, 2472, 3364, 870, 3129, 1723, 950, 3294, 1333, 3205, 1545, 2024, 769, 1123, 1779, 686, 941, 3093, 1112, 2733, 2793, 1806, 2284, 2918, 798, 2369, 2001, 2259, 2779, 1443, 2382, 3393, 232, 942]. Realist [2380, 3232, 2550, 137, 2985, 1143, 1560, 2012, 3188, 2259, 3211]. realistic [1175, 213]. realists [2550]. Reality [1750, 419, 418, 691, 2571, 1149, 2616, 849, 551]. realizability [2285, 1148]. realization [133]. realizations [1978, 669]. realizes [1708]. Realizing [2744]. really [2329, 653, 1480, 126, 806, 2167]. reappraisal [2140, 2601]. reason [2969, 905, 50, 2256, 77, 432, 1292, 1286, 2669, 1491]. Reasonable [456, 2344, 1206, 897]. reasonableness [1227, 3093]. reasoner [1796]. Reasoning [2298, 458, 1409, 1584, 2620, 3333, 959, 1176, 2755, 1302, 1849, 1157, 174, 3368, 2467, 2534, 2281, 2350, 2465, 2841, 466, 580, 2975, 926, 1464, 2118, 2688, 2595, 1268, 1917, 1406, 767, 1278, 246, 1508, 322, 1461, 1460, 1526, 62, 1619, 76, 3167, 1439, 2618, 1608, 645, 2833]. Reasons [1718, 763, 1381, 668, 1871, 2957, 1237, 816, 2688, 2027, 3011, 1327, 2544, 104, 1220, 2064, 1300, 1883, 1469, 2263]. Reassessing [712, 838, 3313]. Reassessment [2175]. rebuttal [1853]. rebutting [2032]. REC [1330, 1329, 280, 1126]. recalcitrant [2299]. recalibration [1401]. receiver [90]. reciprocal [1262, 1261]. reciprocity [974]. reclamation [2734]. recognition [2766, 2572]. Recognizing [1804]. Recombining [691]. recommender [1201]. reconceived [1922]. Reconceiving [305]. reconception [2550]. reconciliation [1528]. Reconciling [2978]. reconsideration [645]. reconsidered [1854, 2296, 3101, 2912]. reconsidering [2300, 116]. reconstitutor [2146]. reconstructed [2278]. reconstruction [2738]. record [562]. records [2950]. recruitment [1702]. recursive [2197]. red [2080, 2953, 839]. redress [1980]. Reducing [2292]. reduction [3183, 3353, 101, 347]. Reductionism [2802, 3033, 2665]. reductionist [2200, 3161, 145]. reductions [2121]. reductive [2121, 1289, 2102, 116, 1801]. redundancy [3376, 2005, 1734]. Redux [457]. Reference [2826, 3223, 1959, 2308, 1096, 2586, 2822, 419, 418, 2415, 3384, 3023, 2086, 3012, 3031, 47, 2795, 2867, 1434]. referential [2983, 2538]. Refined [85]. refining [3197, 559]. Refitting [1788]. reflect [1055, 1503]. Reflecting [393]. Reflection [596, 417, 118, 502, 2575]. reflections [433, 2286, 3185, 568, 962]. Reflective [939, 3195, 2876, 2794, 2749, 2824, 2815, 54, 1947, 2858, 2781, 2700, 2747, 2051, 2906, 3039, 2788]. reflective-equilibrium [1947]. Reflex [1720]. Reflexive [2983]. reflexive-referential [2983]. reflexivity [2983]. reform [3005]. reformulation [646]. reformulations [3048]. reframing [3361]. refusal [2491, 2522]. refusing [2147]. refute [1434]. refuted [778]. regained [3366, 244]. regard [1559]. regarding [791, 2227]. regeneration [1399]. regional [1102, 3369]. regress [1632, 808, 2789, 644, 172]. regresses [1305]. Regression [72, 3191]. regret [3149]. regular [3291]. regularities [992, 819]. Regularity [1188, 1680]. regulation [2986, 3168, 1483, 2092, 2506, 1317]. regulative [1541]. rehearsal [1367]. Reichenbach [1538, 2243, 1682, 1992, 1595, 2203, 1599, 2007, 1964, 1560, 1545, 1559, 1715, 2044, 1638, 1649, 2079]. Reichenbachian [1869]. Reinach [2730]. reinforcement [585, 1774, 2945]. reinvigoration [985]. Rejecting [1293, 2684, 2720]. rejection [1927]. rejoinder [2229]. related [2328, 2782, 1851]. relation [2197, 3346, 1419, 96, 1017, 2928, 995, 395, 2396]. Relational [3336, 3380, 1566, 1828, 411, 312, 2624, 2035, 1630, 2178, 3179, 3393]. relationist [1502, 2229]. relations [578, 1328, 531, 2234, 800, 991, 538, 2087, 2811, 409, 2146, 886]. relationship [3154, 2479, 1199, 1324, 2875, 882]. relationships [2328]. relative [338, 523, 737, 1020]. Relativism [950, 1710, 2009, 2344, 202, 2593, 3217, 1373, 1324]. relativist [492]. relativistic [1764]. relativity [1065, 1828, 1138, 2243, 1148, 1540, 2275, 2156, 1043, 3058, 1973, 948, 1386, 849]. Relativizing [1870]. Relevance [1184, 1054, 1739, 1290, 1794, 1557, 1556, 957, 3207, 432, 2273, 671, 1529]. Relevant [2014, 3144, 3411, 3255, 1184, 1152, 901, 1442, 915, 1818]. Reliabilism [1467, 610, 325]. Reliabilist [2851, 174, 2687]. Reliability [916, 548, 552, 2986, 2270, 1614, 2111, 2065, 555, 1362, 2939]. Reliable [550, 551]. reliably [1501]. Relief [2016]. religious [2028, 3026, 2263]. Relinquishing [3392]. Relocating [2558, 1054]. relocation [316]. reluctant [614]. Rem} [856, 2199]. remain [229, 168]. remarks [646]. Remembering [1280, 2112, 1969, 1330, 1329, 2826, 3223, 3088, 2670]. Remote [3015]. remotivating [2990]. removing [3334]. renaissance [1259]. renormalization [155]. Renouncing [2330]. repeat [1738]. repeatables [1995]. repeated [1868]. replace [1491]. Replacement [174]. replicability [2900]. replications [848]. Replies [1234, 3244, 2599, 88]. Reply [1565, 3386, 1948, 3155, 1617, 646, 381, 968]. report [1715]. reporting [3205]. reports [842]. represent [2751, 3028, 1175, 2057, 327, 1879]. Representation [225, 688, 1948, 2340, 2897, 834, 1151, 300, 2997, 1419, 2886, 2304, 2117, 744, 1473, 959, 332, 1222, 2024, 3123, 521, 2485, 1516, 1861, 2355, 2968, 2726, 291, 994]. Representational [3254, 1132, 1431, 1391, 1334, 1975, 478, 288]. Representationalism [2897, 599, 1672, 884, 2057, 1563, 718, 3106]. representationalist [1563, 687]. representations [1930, 2878, 1411, 2743, 1726, 12, 2186, 188, 528, 189, 1733, 523, 927, 2887, 1431, 1774, 1482, 1481, 2704, 1520, 2810, 164, 479, 190, 1397, 582, 3241, 2875, 94]. Representing [3207, 1280, 1334]. Reproducees [3171]. reproducers [3171]. reproduction [1050]. Republic [3119]. republican [2265]. repugnance [1303]. requests [3071]. Requiem [904]. require [2581]. requirement [2405, 1044]. requirements [226]. rescue [2313]. Rescuing [2448]. Research [3311, 2662, 883, 1956, 2889, 507, 501, 2328, 1133, 732, 616, 188, 2135, 504, 3184, 2315, 1807, 396, 1307, 3240, 1983, 2044, 3019, 3176, 2302, 1014, 1168, 3062]. resentment [1420]. resilient [553]. Resisting [3281]. Resolution [2737, 2049]. Resolutions [1381, 816]. resolve [39]. Resolving [3143, 2077, 1603, 2026]. Resonance [1134]. resource [2073]. respect [3056]. Respecting [2254]. responding [2032]. response [2300, 1675, 1569, 826, 2037, 1544, 2247, 878, 973, 2024, 1485, 3292, 1188, 1092, 1091, 2382, 3098, 88]. response-dependence [1569]. responses [3127, 1777]. responsibility [2160, 2591, 3143, 2612, 2325, 1620, 3361, 3065, 1004, 1185, 1858, 273, 508]. Responsible [501, 507, 504, 1482, 1481, 502, 404, 508]. responsiveness [1447]. rest [2824, 3396]. Restricted [2865, 2061]. result [18]. results [339, 1767, 978, 1507]. Resurrecting [2487]. retail [693]. rethink [2242]. Rethinking [418, 2889, 529, 952, 3246, 2604, 2110, 2146, 419]. retract [2910]. Retraction [2666]. retrieval [3278]. retrocausality [2201]. Retrodictive [3104]. retronasal [1354]. retrospect [1771]. retrospection [2493]. Retweeting [1051]. reuse [185, 186]. Revaluing [847]. revamping [1050]. revealing [734, 3323]. reveals [1516]. Revelation [1667]. revenge [820]. reversal [2560, 2333]. Reverse [36, 1849]. Reversing [1994]. reverts [609]. review [2248, 1066]. reviewers [3386]. Reviews [2785]. revised [3306]. revising [2707, 85, 2308]. revision [2860, 250, 926, 1323]. revisionary [58, 685]. revisited [48, 1543, 2384, 2410, 1556, 28, 9, 388, 1932, 1805, 640, 2210, 1557]. Revisiting [2713, 1682, 3083, 1126, 3010, 992, 819]. Reviving [1400]. revolution [3298, 2743]. revolutions [3025]. Reward [585, 1178]. rewarded [1474]. rewards [1774]. rhetic [2580]. rhyme [1292]. Richard [963]. rides [1623]. Right [96, 2614, 1223, 397, 1796, 365]. rightness [2702, 2724]. rights [50]. rigid [2710]. rigidity [422]. Rigor [2823, 2120]. rigorous [1268]. Rigour [1736]. riken [1971]. rise [462]. Risk [2081, 2694, 2988, 2194, 1184, 2990, 2103, 147, 2069, 2460, 219, 2305]. Risk-limited [2081]. riskiness [1677]. risks [2194, 3357, 2667, 2325, 2139]. risky [923]. rival [2789]. rivalry [666, 2969]. roadmap [1243]. roads [2201, 68]. Robbers [3333]. robots [2534, 1359]. Robust [1842, 810, 589, 1815]. robustly [3310]. robustness [2638, 1811, 526]. rock [1401]. Roger [2754]. role [1585, 1202, 577, 2889, 1334, 226, 2873, 3246, 633, 2559, 2915, 66, 1405, 1130, 1614, 3063, 3056, 2235, 1399, 1725, 1621, 2929, 2799, 1278, 2808, 434, 572, 3261, 3363, 2181, 582, 600]. roles [448, 3209, 1527, 3150]. roots [3052, 2396]. rotten [1344]. Rousseau [2031]. route [16, 1536, 519]. routes [1279]. RRI [504]. RRT [2983]. RT [1111]. Rubicon [1826]. rule [902, 926, 1562, 594, 107]. rule-following [926]. rules [911, 3188, 880, 1686]. RUM [3079]. ruminate [1738]. run [1687]. runabout [378]. runaway [143]. running [3327]. runs [3079]. Russell [818, 2634, 1545, 2839, 762]. Russellian [1961]. Russellians [272]. Russo [2260]. Ryle [3297]. Rylean [2118, 172].

S. [646]. saboteurs [364]. safe [2148, 2041, 203, 180, 900]. Safety [2916, 257, 2041, 641, 270, 254, 1256, 259, 2290, 1916, 2732, 166, 258]. said [2395, 3046, 1043]. Salience [1636, 1460, 2384, 90, 1461, 3274]. Salient [2992, 816]. Salmón [2966]. Same [1897, 425]. Same-tracking [1897]. Sameness [2461, 138]. Sample [994, 1831]. samples [1334]. sampling [1868]. satellites [527]. satisfaction [224]. satisficing [1815]. save [2503]. Saving [259]. saying [2340]. scaffolded [1201]. scaffolding [189]. scale [1828, 564, 1909, 1279]. scale-free [1279]. scale-relativity [1828]. scales [3120]. Scaling [287, 716]. Scaling-up [287]. scenarios [2929, 3047, 3170]. Scepticism [457, 121, 1082, 2916, 1230, 1191, 1205]. sceptics [3100, 3069]. schema [1323]. schemas [2249]. schemata [1176]. schemes [2096, 2013, 3254, 3369]. Schiffer [2966, 995]. Schiller [3385]. Schmidentity [1079, 1031]. scholarly [3202]. School [2312]. Schurz [88]. Science [2233, 1953, 2752, 2674, 1809, 2107, 1858, 2301, 3095, 1495, 3004, 3154, 3168, 2344, 2420, 1956, 2358, 3044, 1856, 507, 3366, 2367, 566, 2248, 2896, 2776, 1847, 290, 1133, 732, 75, 1568, 569, 430, 1711, 96, 1130, 1385, 440, 2664, 1765, 2622, 2778, 506, 1306, 2877, 406, 1444, 95, 3152, 2262, 2903, 695, 2673, 402, 1621, 504, 704, 2807, 728, 446, 3000, 1299, 2021, 1402, 1232, 570, 3278, 1319, 1999, 1125, 2331, 1990, 1129, 1132, 3352, 1917, 417, 1829, 1640, 2048, 3189, 2012, 1384, 791, 3022, 3229, 1136, 2565, 2675, 1453, 3343, 321, 572, 3406, 2844, 3009]. science [2085, 88, 567, 568, 608, 1967, 2646, 659, 2756, 2216, 1866, 110, 790, 2392, 1336, 2933, 1197, 751, 526, 2512, 2341, 284, 1039]. science-their [440]. sciences [1897, 1561, 2393, 529, 447, 3042, 2841, 1131, 2260, 1587, 285, 1444, 1252, 1800, 387, 2280, 944, 2583, 2100, 1378, 3330, 1014, 994]. Scientia [997]. Scientific [834, 1531, 592, 427, 2252, 606, 2809, 895, 2289, 3145, 556, 2675, 2443, 708, 3019, 2533, 2676, 3238, 2300, 1921, 883, 3298, 2847, 1793, 2731, 315, 2450, 2242, 3231, 159, 1657, 2465, 1786, 1133, 656, 1419, 2076, 1341, 426, 137, 3056, 1716, 181, 2597, 1851, 1999, 1273, 1560, 33, 1175, 3205, 1545, 1384, 560, 558, 3356, 1123, 1758, 3264, 654, 686, 3093, 2793, 571, 2485, 3259, 2756, 304, 2133, 2379, 1169, 2432, 1324, 912, 2726]. scientifically [2524]. scientists [2280]. scope [689, 2548, 858]. scoping [251]. Scoring [83, 1535]. Screams [461]. Screening [1538]. scrutability [381]. Scrutinizing [2894]. scrutiny [53]. Scylla [3275]. se} [1036, 2925, 106]. seamless [2363]. search [448, 2157, 1726, 3340, 1596, 2417]. second [1385]. secret [1642]. secret-of [1642]. section [2583]. secular [3403]. secure [615]. Seeing [2549, 2039, 733, 2702, 3194]. seeing-as [3194]. Seeking [3039, 2559]. seeming [1912]. Seemings [843]. seen [1117]. segment [1864]. Segmenting [3265]. Seiberg [3169]. selected [2706]. selection [3375, 2429, 2601, 2507, 2486, 2876, 1664, 1050, 858, 3062]. selectionism [1716]. selectionist [2012]. selective [1142]. selectivity [634]. Self [1263, 67, 1165, 3245, 823, 106, 2976, 2969, 2329, 3147, 1445, 831, 1265, 3357, 2505, 2506, 2805, 2984, 1047, 119, 2082, 2489, 432, 2709, 1601, 1355, 1971, 2849, 2872, 47, 2814, 1685, 637, 556, 3296, 2323, 2616, 2271, 661, 3389, 2945, 2153, 1749, 359, 2969]. Self-ascription [106]. self-awareness [1047, 2489, 1971]. self-censorship [3357]. Self-colocation [823]. self-consciousness [2984]. Self-control [1263, 1445, 2969]. self-correction [556]. self-coupling [1685]. self-deception [2082]. self-defeating [2329, 2805]. self-determination [119]. self-discovery [2814]. self-doubt [1265]. Self-equilibration [3245]. self-evaluation [359]. self-experience [1601]. Self-forming [67]. self-fulfilling [1355]. self-identity [2849, 661]. self-ignorance [831]. self-illness [3296]. self-improvement [2505]. self-indication [2976]. self-knowledge [2505, 1047, 432, 637, 2323, 2616, 1749]. self-location [2709]. self-modelling [2969]. self-reference [47]. self-regulation [2506]. self-stultification [3389]. self-subsisting [2153]. Self-supervision [1165]. self-trust [3147]. self-tuning [2945]. selfextensional [3355]. Selfless [1107]. Sellars [784, 3097, 1854, 1006, 114, 3127, 1690, 1450]. Sellarsian [3072]. selve [3296]. selves [2326]. Semantic [2538, 941, 3160, 1856, 3232, 598, 2073, 3224, 3292, 204, 561]. semantics [537, 1688, 3267, 2450, 1376, 1345, 617, 2073, 2137, 2147, 1306, 3055, 3397, 2445, 2735, 1550, 2199, 1054, 2871, 3191, 2992, 178, 21, 324, 2727]. semantics/pragmatics [21]. sender [90]. sender-receiver [90]. sensation [1854]. sensational [750]. Sense [1475, 960, 881, 1918, 639, 2790, 2722, 2904, 851, 390, 3038, 3337, 299, 2872, 2671, 2888]. sense-making [390, 299, 2888]. senses [2070, 1776, 1354]. sensible [316]. sensitive [489, 205, 2148, 1549, 3397, 2848, 3144, 968]. Sensitivity [2290, 3342, 470, 1493, 2908, 270]. Sensorimotor [471, 3369]. Sensory [1026, 2375, 1450]. sentience [2677]. sentiment [891]. sequence [932, 1637]. seriously [1448, 2260, 2532, 3056]. seriousness [3095]. Set [1219, 2772, 2665, 2180, 158, 38, 582, 2563]. Set-theoretic [1219, 158]. sets [2359, 1722, 37, 38, 1168]. setting [406]. settings [2666, 3209]. settled [1379]. severely [2148]. severity [3343]. sexist [127]. shadow [2877]. shall [2439]. shallow [2297]. Shannon [1294]. shape [3306, 2024, 1368]. shapes [2798]. shaping [2903]. share [2566]. Shared [2192, 3045, 788, 2623, 1210, 2272, 2796, 3334]. Sharing [1904]. sharp [859, 21]. Sharpening [2230]. shell [3048]. Sherlock [1039]. Shifting [3190, 1544]. shifts [1681]. shifty [2862]. Ships [2937]. shooting [2105]. Shopping [1996, 1088]. shortcomings [1668]. Should [1576, 167, 2673, 1605, 1717, 2780, 2422, 859, 122, 2896, 1711, 1043, 3320, 2951, 3138, 2669, 404, 2287, 2341, 291]. Show [791]. shows [1366]. side [2822, 332, 274, 2174, 2210, 132]. side-effect [274, 2174, 2210, 132]. sight [903]. signaling [2028, 1360, 1294, 3348, 1778, 2421, 90, 1106]. signalling [1012]. signalman [1069]. Signals [972]. significance [3378, 1422, 596, 1237, 2449, 328, 1785, 681, 1355, 1999, 1174, 1615, 2812, 2038, 2926, 1883, 3079, 2668]. signs [2471]. silence [3186]. silver [1931]. Similarity [1855, 984, 279, 1054, 1431, 1818]. Similarity-based [279].

Simon [2122]. Simple [243, 1969, 785, 719, 1995, 1730]. Simplicity [2420, 1008, 2362, 1720, 1380]. Simply [2304]. simulating [609, 2747]. Simulation [2709, 1679, 1183, 2433, 1331, 1761, 1187, 1209, 1679]. simulation-based [1761]. simulation-modeling [1679, 1183]. simulationist [2826]. simulations [1673, 2534, 2075, 3261, 571]. simultaneity [1429, 1656]. sin [617]. since [436]. Singer [1053]. Single [1010, 66, 2149]. Single-domain [1010]. Singular [2886, 2973, 3125, 1886, 1782]. singularities [3325]. Singularity [1777]. situate [96]. Situated [768, 2615, 1971, 605]. situating [305]. Situationism [770, 119, 610, 1675]. situationist [671]. situations [391]. size [2372, 1571]. skeptic [637, 3206]. skeptical [1987, 1824, 369, 2067, 3292]. Skepticism [449, 2200, 2699, 2329, 2030, 2264, 2017, 2037, 3127, 973, 3075, 1022, 366]. sketching [1316]. skewed [3340]. Skilful [118]. Skill [1432, 1445, 1898, 2614, 2239, 2617, 2447]. Skilled [1918, 1724, 1503, 2235, 287, 1359, 1532, 2702, 1456, 1609]. skillful [2991, 1364, 1896]. skills [1231, 759]. sleep [1456, 477]. Sleeping [2237, 690, 1053]. slice [489]. slice-sensitive [489]. slippery [2597]. slips [1231]. slot [412]. slow [624]. Slurs [1471, 3359, 2910, 2626, 880]. smallpox [397]. Smart [2752, 3237, 3001]. smarts [1171]. smell [2117, 1354]. Smellscapes [3021]. Smolin [3375]. Social [707, 1246, 2004, 33, 1702, 2565, 1668, 2583, 2935, 1542, 2613, 1179, 1897, 1561, 2393, 2318, 50, 881, 2517, 2776, 2234, 3214, 2171, 2260, 218, 2449, 81, 2336, 2137, 1549, 1216, 3310, 1587, 1030, 2118, 1989, 1252, 1800, 2763, 2589, 2673, 3053, 1319, 276, 2056, 2799, 1640, 2951, 767, 2006, 222, 3189, 1004, 2949, 1983, 1939, 2280, 2508, 1185, 2675, 2792, 2408, 2422, 2100, 1518, 2646, 828, 2859, 1378, 2940, 1169, 404, 2311, 73, 2888, 882, 994, 1247, 2056]. sociality [1709, 2914]. socially [1007, 404]. societies [2747]. society [507, 2674, 1321]. socio [187, 961, 276, 882]. socio-cognitive [961, 276, 882]. socio-cultural [187]. socioeconomic [1507]. sociological [3107]. Socrates [3226]. software [553, 2438]. software-based [2438]. solar [731]. solipsistic [2894]. solo [298]. Solution [1859, 2987, 922, 1230, 1389, 2632, 2687, 2143, 986, 1951, 1502, 808, 1916, 1004, 1693, 1108, 251, 121, 288]. solutions [28, 585, 2262, 3271, 1731, 738, 644]. solve [1632, 785, 2040, 272, 2000, 1424, 738, 438]. Solving [3106, 2239, 732, 210, 1758]. Some [1948, 2226, 683, 154, 646, 571, 2493, 125, 2911, 301, 531, 1574, 199, 339, 210, 3128, 2719]. Someone [2350, 1001]. Sophistry [1177]. sorites [2645, 1859]. sort [180]. sortal [429, 2497]. Sortals [2086]. sorts [1687]. Sosa [255, 2916, 258, 257, 262, 253, 260, 265]. sound [1879]. soundness [2250]. Sounds [2960]. source [2623, 205, 1614, 555, 551, 622]. sources [883, 2464, 3079]. Sourcing [1807]. Space [456, 3303, 2954, 3210, 104, 1795]. space- [2954]. spaces [1332, 1857, 2690]. Spacetime [3325, 1949, 1138, 1143, 1145, 1111, 1141, 3199, 3200, 913, 159, 1144, 2600, 3058, 1139, 2245]. Spatial [1149, 3008, 1568, 2604, 1021]. spatiotemporal [1146, 299]. Speak [1989]. speaker [2801, 3141]. speaker-meaning [3141]. speaking [218, 3314]. Special [2752, 393, 367, 230, 321, 278, 348, 140, 117, 449, 1065, 566, 53, 75, 467, 2275, 2156, 1043, 407, 417, 1784, 184, 1559, 13, 1513, 88, 849, 530, 1150, 488]. specialized [2289]. Species [425, 3050, 1835, 569, 427, 426, 1003, 1267, 3350]. species/genus [1835]. speciesism [635]. specific [1699, 1496, 356]. Specification [2468]. Specificity [2261]. spectrum [2371]. Speculative [3357]. Speech [1035, 3076, 3082, 1192, 2738, 2767, 1315, 1643, 2109, 1071, 1471]. speech-act [1071]. spin [3201, 1116]. spin-statistics [3201]. Spinozan [2299]. spirit [2285, 2203]. spoken [2478]. spontaneous [761]. Spot [756]. square [2523]. ST [2147]. stability [679, 1106, 757, 600, 2412]. stable [477]. stabs [3087]. stake [2296, 2446]. Stalnakerian [3270]. stance [1832, 2713]. stances [3266, 1324]. standard [2324, 3256, 3160, 463, 812, 1206, 264, 1951, 3026, 1513]. standards [2348, 1170, 1374]. Standing [1661, 957]. standpoint [3243]. standpoints [677]. state [1926, 3382, 2031, 1586, 3144]. state-based [3382]. state-sensitive [3144]. statements [784, 1440, 1995, 1801]. states [3017, 3053, 115, 3204, 1940, 228]. Statistical [1420, 2251, 2557, 1999, 2269, 1567, 1699, 2292, 10, 277, 155, 1249, 2167, 2438, 1868, 51]. Statistics [1761, 801, 3201, 1420, 1405, 2424, 3122, 1853, 2196]. Status [3182, 440, 1284, 2247, 1507]. stay [203]. Staying [3048, 3291]. steadfastness [809]. Stealing [364]. Stebbing [3385]. Steel [1215]. steering [357]. steps [1693]. stereotypical [1038]. stigma [3037]. stigmergic [3249]. stigmergy [2646]. still [1295]. stimulus [2492, 1187]. stimulus-control [2492]. stimulus-to-perception [1187]. stochastic [704]. Stochasticity [45, 221]. Stone [881]. Stop [2117, 2824, 291]. storey [1126]. stories [2978]. story [3408, 881, 2944, 1126, 1492]. straight [669]. Straightening [2756]. Strange [450, 201]. Stranger [201]. Strategic [2171, 1537, 2981, 1658, 1432, 66, 572]. strategies [2402, 1215, 3266, 2862, 732, 3289, 1990, 210, 759]. Strategy [595, 1837, 724, 58, 1173, 3129, 445, 1197]. stratification [3062]. Strawson [1966, 55]. stream [352]. Street [1171]. strength [1958, 1105]. Strengthening [1961, 885]. strict [2147]. strict/tolerant [2147]. strictly [540]. String [459]. Strong [3366, 161, 1553, 2479, 1849, 2771, 1613, 8, 855, 1044, 1708, 1890]. stronger [1424]. Structural [485, 2434, 1411, 1333, 927, 1048, 3336, 969, 1718, 771, 2985, 3294, 1788, 979, 2024, 709, 1818, 2671, 2259]. structural-causal [969]. Structuralism [1742, 1968, 2819, 2584, 2093, 1398, 3266, 1632, 2762, 2199, 856, 3240, 2953]. Structuralist [1791, 1887, 3294]. structurally [3234]. Structure [860, 2152, 523, 1245, 3318, 688, 2019, 2495, 3178, 3156, 2395, 67, 2522, 2491, 43, 42, 2127, 3346, 984, 2997, 3308, 3329, 2825, 2203, 1049, 407, 1178, 822, 814, 1983, 2518, 521, 976, 2704, 684, 103, 3186, 1260, 1443, 2689, 1064].

Structure-preserving [523]. Structured [2445, 153, 1365]. Structures [409, 578, 3073, 2257, 2209, 392, 1621, 652, 1175, 2153, 3015, 1449, 3230]. Stubborn [2469]. Stuck [1738]. students [504]. Studies [2107, 1226, 1950, 946, 26, 3062]. studio [1330, 1329]. Study [2266, 739, 3017, 3147, 3297, 3124, 463, 2552, 636, 6, 235, 2854, 2816, 2722, 1595, 504, 2642, 2331, 1273, 1889, 1917, 2811, 1983, 1496, 2514, 1008, 1187, 1209, 394, 2933, 626, 2895, 2139]. studying [2232]. stultification [3389]. style [2759]. styles [654]. Subject [356, 3328, 2544, 1358, 2392]. subject-dependency [3328]. Subject-specific [356]. Subjective [859, 590, 202, 2316, 950, 1933]. subjectivist [781]. subjects [2307]. subjunctive [3006, 2813]. subpersonal [1000]. subsisting [2153]. substance [411]. substances [413]. substantial [2505]. substantiality [3124]. substantivalism [1764]. substitute [3153]. substitution [2375]. substrates [1488]. substratum [1732]. substratum-trope [1732]. Substructural [520, 1150, 3271, 1161, 1153, 1159]. subversion [2883]. Success [3255, 3164, 2509, 2012]. successes [2734]. successful [2113]. successor [68]. Suffering [2728, 3210]. sufficiency [1046]. sufficient [3237]. Suffit [394]. suggest [516]. suicide [2988]. summativism [1848]. Super [868, 1199, 1764, 200]. Super-Humeanism [868, 1199, 200]. super-substantivalism [1764]. Superdeterminism [2140]. Supererogation [932]. superficially [1817]. superiority [320]. superposition [1484]. supersaturated [2284]. supersententialism [89]. superspace [2948]. Superstring [1777]. Supersymmetry [3169, 3201, 3405, 2742]. supervaluations [19]. supervenience [2294, 1371, 236, 340, 1025, 2942]. supervised [1604]. supervision [1165]. Supplementation [482]. support [191, 2286, 764, 641, 613, 270]. supporting [1925, 3161]. suppositional [1787, 590]. sure [1182]. surface [1792]. surprise [1070]. surprising [154]. surrealist [1511]. surrogates [2216]. Surrogative [2841, 2534]. survey [2967, 2876, 798, 2717, 1507]. surviving [2719]. suspended [249]. Suspending [700]. suspension [3117]. Suspicious [2022]. Swampman [2288, 1295, 361]. switching [2640]. syllables [1558]. syllogism [3250]. symbiosis [1262, 1261]. symbols [521, 522]. symmetric [966, 1328]. Symmetries [1653, 1208, 2655, 1177, 1546, 3221, 2164]. symmetry [761, 722, 800, 3254, 1441, 2560, 1851, 949]. symmetry-related [1851]. symposium [162]. symptom [3332]. symptoms [1366, 2219]. Synchronic [3350, 913]. Synergic [100]. synergies [3154]. Syntactic [2257, 2209, 598, 582]. syntax [2890]. synthese [367, 140, 2981, 2056, 2004, 184, 1245, 2512]. Synthesising [2686]. Synthetic [954, 2342, 1485, 3161, 2240, 2010, 1652]. syntheticity [2511]. System [1467, 1201, 1808, 2233, 731, 1784, 1628]. system-experience [1784]. systematic [3012, 1316, 57, 2923]. Systematicity [1336]. Systems [1618, 537, 1213, 861, 10, 1294, 2791, 1778, 474, 130, 1972, 954, 154, 704, 3347, 553, 3120, 1518, 3010, 1233, 277]. Szaszianism [2760]. table [593, 1019]. Tacit [806]. take [3404, 1534, 24, 2430, 3138, 762]. takes [2066, 2322]. Taking [1448, 2260, 2532, 3056]. tale [1720, 1765, 1578, 1126]. Talent [947]. talk [126, 2041, 1760, 3295, 2673, 2436, 2539]. Taming [378, 1742, 1847, 2748]. tampering [585]. Tap [3180]. Tap-dancing [3180]. TARDIS [2433]. Target [3274]. targets [470]. Tarskian [534, 531]. task [696, 1006, 1991]. tasks [2281]. taste [1454, 1635, 1634, 2267]. taxonomy [140, 737]. teach [1679, 1183, 3100, 3069, 2690, 526]. teaching [2044]. technical [2436, 1278, 3268]. technically [3242]. technological [1278]. Technologically [1201]. technologies [1866]. Technology [2216, 1030, 2661, 1539, 2980, 2432, 3190]. Technology-driven [2216, 2432]. teleological [2554, 1045, 3006, 2813]. Teleology [2416, 2459, 3187, 720, 1554, 2964, 2625, 2386, 3028, 680, 2446, 275]. teleology-neutrality [2386]. teleosemanticists [1295]. teleosemantics [826, 1277, 2288, 116]. teleparallel [3058]. telic [1825]. tells [601]. template [3231]. templates [2383, 2504, 2411]. Temporal [715, 3064, 46, 1780, 2757, 2578, 3304, 1902, 2571, 764, 1568, 2975, 2722, 2886, 2608, 3104, 2067, 2468, 588]. Temporalism [2912]. temporality [2624]. Temporalizing [990]. tendency [223]. Tennant [1155]. tense [2016, 1376, 2660, 2627, 948]. Tensed [2757, 764, 2975]. tenselessness [715]. tension [137]. teratogenicity [400, 399]. term [2597, 1815]. terms [1651, 926, 3064, 1915, 423, 2199, 2513, 424, 2867, 422]. territory [2849]. tessellations [1332]. test [1089, 2758, 656, 3031, 1767, 546, 596]. Testable [2241]. tested [2148]. Testimonial [2870, 3376, 697, 1920, 1441, 844]. testimony [1202, 1711, 1583, 181, 1232, 1287, 149, 3047, 3170, 2939]. testing [2540, 2424, 839, 1645, 78]. texts [2303, 2558]. texture [3091, 2120]. their [537, 2378, 1151, 996, 1488, 3239, 1531, 2667, 440, 1321, 1090, 310, 498, 2566, 2939, 282]. them [1301, 3043, 2910, 2280, 2109, 594]. Themes [1221]. themselves [1752, 3247]. theorem [3201, 3340, 1105, 1771]. theorems [3340, 332, 1604, 2427]. theoretic [3316, 50, 1926, 617, 514, 2073, 1219, 1306, 2278, 664, 158, 2332, 1226, 796, 333, 2523]. Theoretical [2358, 817, 1380, 1871, 1440, 1219, 1915, 1723, 2694, 2241, 317, 2145, 2320, 3167, 871]. Theories [3112, 1162, 3068, 2019, 542, 2028, 3200, 1803, 1028, 2022, 3354, 584, 466, 656, 3254, 2104, 2605, 2876, 44, 2617, 2806, 462, 1596, 1671, 293, 2849, 3031, 1383, 1382, 1123, 2948, 909, 2596, 941, 2110, 3330, 241, 1081, 233, 1159, 1697, 159]. theorists [1069]. theorize [112]. theorizing [1380, 330]. Theory [1076, 2162, 1214, 2238, 78, 2302, 1611, 666, 3033, 943, 1929, 433, 2329, 714, 2834, 542, 459, 2096, 2013, 1264, 1720, 1190, 2768, 3396, 2769, 42, 448, 2127, 2123, 3131, 1127, 837, 1902, 2927, 1752, 1904, 1310, 1480, 3359, 893, 892, 2962, 155, 3132, 2638, 535, 1309, 1992, 2086, 902, 455, 2200, 2828, 2797, 2515, 2611, 250, 1455, 2180, 262, 2983, 2657, 119, 965, 346, 963, 11, 199, 74, 1637, 2738, 1394, 1874, 776, 896, 2063, 420, 583, 1393, 1599, 1541, 1152, 1497, 1059, 2144, 1501, 2869, 845, 1788, 1537, 2981, 2236, 1194, 1332, 1045, 613, 587]. theory [45, 416, 1731, 2142, 1985, 3022, 553, 2867, 3123, 995, 2839, 3264, 1056, 1850, 906, 2193, 3379, 2064, 742, 2715, 840, 64, 810, 1762, 874, 787, 890, 2863, 403, 3169, 282, 849, 213, 1777, 582, 2563, 269, 2352, 1608, 3087, 3099, 2907, 1697, 2015, 2484, 43]. theory-change [403]. theory-heavy [2834, 2907]. Theory-ladenness [78, 943, 1731]. therapy [3245]. There [1957, 1494, 2189, 1491, 2456, 229, 168, 1371, 863, 2815, 2197, 512, 3395, 1878]. Thermal [413, 1346]. thermodynamic [1727]. thermodynamics [2292, 2213, 1385, 2048]. These [2224, 1045]. Theseus [2937]. thesis [3238, 3372, 2790, 2286, 2260, 1057, 3273, 2977, 329, 232, 2498, 2621]. thick [2848, 2363, 3311]. thin [3318, 2363, 3311]. thin-objects [3318]. thing [914, 594]. things [240, 124, 179, 3247, 375, 2750, 112, 888, 3268, 2347, 302, 151]. think [1941, 1792, 403, 365]. Thinking [1498, 1060, 1298, 558, 172, 920, 2055, 3260, 3136, 1121, 3036, 738, 654, 3249, 1434, 1748, 1407]. Third [32, 2599]. third-factor [2599]. thirders [1115]. Thomasson [2576]. Thought [2677, 2338, 364, 863, 1673, 2834, 1124, 1440, 2213, 2886, 2521, 1970, 903, 287, 3038, 2000, 1792, 669, 1782, 91, 751]. thoughts [1581, 1962]. threat [1244, 992, 819, 3065]. threatens [1254]. Threats [1509]. Three [2653, 3328, 3240, 2184, 2402, 800, 2157, 2327, 2830]. three-factor [2327]. three-level [2830]. threshold [2075, 2546]. thresholds [135, 3236]. Throwing [1033]. thrust [357]. Tichý [2468]. ticket [378, 938]. tile [922]. Till [229, 2703, 168]. Time [678, 3018, 241, 2016, 2623, 1280, 425, 1486, 1902, 242, 1802, 2624, 2168, 2612, 1656, 3063, 776, 1009, 2560, 2333, 2067, 2948, 2468, 1061, 2079, 1780, 436, 3015]. time-bias [2016]. time-lag [1656]. time-reversal [2560, 2333]. Timelessness [220, 1009]. times [2954, 1895, 2690]. timestamp [2727]. Tinbergen [1716]. titles [2803]. today [30]. together [302, 1861]. togetherness [3045]. token [2066, 2322]. tokens [3101, 3053]. Tolerance [423]. Tolerant [1157, 2147]. toleration [975]. tongue [2636]. tonk [1928, 544]. Too [2090, 914, 2304, 3101]. tool [2612, 3410, 1339, 551]. toolbox [846]. tools [1055, 2221, 2650, 2230, 2225, 882]. top [2386, 390, 1351]. top-down [2386, 1351]. topic [1789, 1299, 2021, 1690]. topic-modeling [1299, 2021]. Topical [1892, 2106, 2393, 3195, 3082, 2967, 1221, 2981, 2056, 2004, 1228, 1243, 2887, 1893, 1245, 2512]. topics [3144]. topological [2543, 1332, 332, 1024]. topologies [2079]. topology [2291, 1976, 1972, 2719]. Tortoise [2829]. totalities [2762]. touch [3177]. Tough [1815]. Toy [2475]. Tracers [1349]. Traces [3249, 2822]. Tracking [1567, 782, 3352, 811, 1897, 81, 396, 2432]. Tractability [2097, 1811, 846]. Tractatus [3109, 1795, 1530]. trade [2351]. trade-offs [2351]. tradeoff [3356]. trading [2646]. tradition [1927]. traditional [956, 2448]. traditions [2300, 535]. tragedy [2068]. training [2642, 1906]. trait [953]. Traits [829, 1169]. transcendental [2015, 2461, 431, 1113, 438]. transdisciplinary [2504]. Transfer [2046, 529, 2653, 3012, 2369]. transferability [2433]. transfinite [2693]. transformation [2493]. transformational [1291]. transformations [1528]. Transformative [2594, 3149, 3137, 2651, 1466, 2574, 2391, 2798]. transforming [1719, 2432]. transhumanist [2286, 3351]. transient [2341]. transition [2980]. Transitions [752]. transitive [675]. Transitivity [138, 2169]. translatability [2019, 1735]. translation [2021, 2162, 3392]. translations [545]. Transmission [2052, 3388, 146, 2200, 2170, 2466, 2263]. Transparency [884, 2034, 1979, 131, 1848, 974, 1910, 2778, 3301, 2114, 795]. transparent [3343, 2468]. travel [1486, 678, 241]. traveler [3018]. Traveling [2433]. travels [242]. traversal [3057]. treatise [1188]. Treatment [2095]. trees [1433, 1640]. trends [2018, 1789, 95, 2482]. trespassing [2187]. trial [2095, 394]. triangulation [2305]. trio [2372]. tripartite [1995]. triplet [1419]. trivial [1160, 153]. trivialism [1492]. Triviality [2061, 3222, 2989]. trolley [2051]. Trolls [609]. Trope [1390, 1732]. Tropes [1690, 1797]. trouble [1170, 1072]. Troubles [1347]. trousers [607]. True [1734, 2691, 639, 898, 2547, 2842, 1467, 246, 1221]. Truetemp [2412]. truism [311]. truly [1997]. trust [3147, 2306, 1576, 2778, 506, 2999, 2661, 181, 2490, 2796, 3341, 2936, 2678, 1185, 3009, 1209]. trust-based [2999]. Trusting [1793, 1114]. trustworthiness [505, 406, 1114, 2775]. trustworthy [1771]. Truth [312, 2321, 2215, 2176, 2714, 81, 2659, 931, 264, 310, 3109, 3259, 2428, 946, 3197, 2821, 1353, 2757, 311, 1712, 2009, 1186, 2947, 419, 418, 1666, 3372, 2674, 2640, 2994, 919, 308, 1166, 2762, 712, 3276, 1088, 3143, 2532, 843, 1162, 2147, 2509, 2556, 2034, 3340, 85, 830, 3136, 973, 3337, 595, 3150, 1228, 2177, 350, 1442, 813, 835, 2759, 1850, 309, 516, 324, 710, 1159, 3013, 968, 306, 3393, 307]. truth-constants [1442]. truth-evaluability [712]. truth-grounding [3197]. truth-maker [2147]. truth-makers [843]. truth-value [973]. Truthlikeness [1579, 84, 86, 1535, 2372, 1435, 1677]. Truthmaker [2158, 87, 2121, 3055, 2735]. Truthmakers [1025, 726]. Truthmaking [630, 318, 2377, 1834, 1025, 173, 729]. truths [1567, 16, 2975, 167, 1523, 158, 1892, 547, 726, 380, 3385]. tungsten [2225]. tuning [2956, 2020, 3248, 2945]. Turbulence [564]. turbulent [503]. Turing [1, 2361]. Turkish [2044]. turn [2847, 315, 2555, 2723, 1670, 2756]. Turning [2051, 954, 30]. Turri [968]. Twain [2439]. Tweet [2716]. tweetable [2716]. Twelve [88]. twenty [3187]. twenty-first [3187]. Twin [862, 2955, 2632, 2679, 371]. twins [642]. Twitter [2949]. Two [2201, 883, 68, 832, 1696, 2937, 1743, 535, 2394, 3134, 2802, 343, 208, 2047, 1027, 2408, 2042, 1077, 2227, 790, 1140, 1729, 384, 3401, 565, 16, 2720, 2684, 677, 3384, 2777, 1976, 958, 3002, 2214, 474, 1765, 1578, 2545, 388, 3127, 2903, 3060, 2597, 276, 2006, 1004, 1149, 567, 2396, 508, 1281, 2117, 1984, 1810]. two-dimensional [958, 2006, 1004]. two-dimensionalism [16]. two-standpoints [677]. twofold [1126]. Type [2077, 43, 42, 45, 848]. Types [2939, 140, 1957, 832, 549]. typicality [2161]. Typing [1583]. typology [2328].

u [2202]. ubiquity [944]. ultimate [3146, 2879]. ultrafast [299]. unam [3026]. unawareness [3073]. unbearable [1326, 2702]. uncertain [1745]. Uncertainty [1589, 3342, 2878, 2927, 3193, 658, 3333, 1048, 2005, 889, 644, 897, 1900, 2462]. uncompromising [685]. Unconscious [2651, 1737, 2905, 3233]. unconsciously [2039]. uncontrolled [3112]. uncritical [2844]. undecidability [1283]. undefeated [1467]. Underdetermination [1829, 625, 666, 1987, 2274, 1998, 474, 1887, 2333, 774, 3058]. underestimated [3269]. underlying [502]. undermine [802, 613]. Undermining [2032, 1499, 2846]. understand [3384, 1599, 1203]. Understanding [2303, 1124, 2240, 1657, 1786, 2502, 442, 1249, 1694, 769, 1714, 302, 2222, 2225, 2988, 939, 2128, 905, 894, 1651, 1198, 689, 159, 701, 1437, 3348, 807, 1094, 2728, 3118, 851, 1216, 1896, 354, 1836, 895, 2203, 3364, 1267, 1384, 3259, 2355, 395, 3019, 1573, 132, 751, 2831, 2784, 2570, 778]. understanding-assent [1573]. understanding-why [1437]. undetached [2718, 2649]. Undoing [124]. Unfelt [92]. Unfinished [3076]. Unification [1495, 3408, 1992, 1129, 1132, 850, 1923]. unified [1666, 3224, 2148, 455, 294, 1613, 2144, 854]. unify [1707, 2553]. unifying [592, 1132]. unintentional [2658]. unique [808]. uniqueness [2221]. unit [2146]. Unity [1598, 1410]. Univalent [334]. Universal [2046, 3250, 2979]. Universalism [2410, 491]. universality [491]. Universals [1490, 1884, 780, 2771, 1872, 1801]. universe [237, 37, 38]. University [2044]. unknowable [758]. unknown [1193]. unlike [307]. Unmanifested [1884]. unobserved [2743]. unobserved-or [2743]. Unpacking [3410]. Unrealistic [1175]. unreality [2319]. unreasonable [2111]. unreliable [557]. unreliably [1501]. unrestricted [2445, 1961]. Unrevisability [723]. Unsettledness [1895]. Unspeakable [2370]. untangling [2923]. untenability [2314]. untestable [78]. untimely [818]. untouchable [2308]. Untying [1488]. unusual [1509]. Unveiling [2693]. Unwanted [2466]. unwavering [2059]. Unzipping [2723]. update [2223]. updates [2767]. Updating [3360]. uploading [3034]. upon [2948]. uptake [1325]. use [1151, 951, 1930, 398, 1339, 478, 2225]. used [1627, 2346]. useful [3342, 2704]. usefulness [189]. uses [2659, 3012, 2513]. Using [562, 2006, 3324, 565, 2221, 3037, 2308, 1507]. utilities [1552]. utility [2540, 1497, 914, 1194]. utterances [2464].

Vacillating [1486]. vacuous [2759]. vague [1628]. vagueness [1847, 3373, 2247]. vale [1000]. Validity [47, 1158, 3049, 2611, 1587, 3299, 1163, 3271]. valley [2018]. valuable [2232, 898]. valuational [535]. Value [3388, 2778, 973, 2609, 3370, 2662, 2116, 1775, 1534, 2881, 2911, 3044, 2365, 1000, 2776, 1535, 1745, 2663, 764, 1166, 923, 1790, 2400, 1399, 1637, 1102, 1051, 3278, 265, 1228, 3341, 650, 2603, 2921, 1693, 813, 1830, 2977, 2814, 835, 3229, 3406, 3371, 2756, 64, 268, 1628, 757, 3067, 2760]. Value-aligned [3370]. Value-based [2609]. value-free [2662, 2881, 3044, 2776, 2977, 3229, 3406]. value-laden [2756, 2760]. Valued [1859, 536, 1637, 545, 538, 2054]. Values [2524, 2936, 2900, 859, 3044, 1448, 1076, 2292, 3116, 3388, 218, 2990, 2958, 3000, 1232, 2519, 2796, 791, 2756, 2432, 394]. valuing [2728, 2545]. Vanderschraaf [1904]. variability [2073, 1203, 3079]. Variable [1625, 1973, 2840]. variables [2086]. variants [310]. variation [2385, 1691, 133, 1396]. Variational [2879, 2587]. variations [1690]. varies [40]. Varieties [1002, 3162, 530, 508, 1281, 3214, 2170, 1687]. variety [2221, 1057, 1601]. various [1317, 2087]. varying [2144]. vats [3292]. vehicles [1236, 401, 1431]. veil [865]. Vendler [1728]. Verbal [1231, 65, 3141]. verdict [2251]. verification [1977, 2455]. Veritism [778, 2926]. veritistic [222]. versa [1385, 516]. version [1330, 1329]. versions [365]. versus [1097, 2623, 379, 1736, 1198, 2256, 227, 28, 653, 735, 2556, 525, 2283, 830, 1129, 46, 276, 2330, 2032, 1354, 1946]. Vertical [653, 2026]. Vertical-horizontal [2026]. very [1161]. via [2077]. Vice [2409, 2295, 1385, 1836, 2323, 516, 1181]. vices [2982, 1838, 2452, 2551]. vicious [1590]. victorious [2236]. Video [1747, 2880, 2690]. videogames [3051]. Videos [2345]. view [3197, 1074, 2454, 1856, 2624, 212, 749, 2149, 439, 1380, 524, 2196, 1037, 3161, 204, 1777, 288, 2461]. viewed [329]. views [1591, 2075, 1952, 278, 304]. violations [1058]. Virtual [2616, 3166, 2602, 546]. Virtue [1611, 3310, 1012, 2494, 610, 1751, 117, 3300, 1310, 770, 2200, 2421, 2687, 851, 1655, 119, 263, 2617, 681, 1226, 118, 1501, 149, 997, 431, 1181, 269, 3067]. virtue-as-skill [2617]. virtue-economic [3300]. virtue-theoretic [1226]. Virtues [1542, 2672, 3119, 2358, 3267, 660, 1219, 681, 1380, 759, 1706, 835, 1598, 3087, 359]. virtuous [1321, 1012, 900]. visibility [1444]. vision [467, 2388, 1804, 1368, 477, 2549, 479]. Visual [1258, 735, 750, 476, 471, 884, 474, 1033, 1368, 138, 684]. visualization [1401]. visually [469]. visuomotor [735, 479]. vital [3220]. Vividness [1304]. voice [1192, 1643]. Volterra [2653]. voluntarism [2713]. voluntary [2312]. voting [1612, 2927]. VR [1281]. vs [3404, 2434, 3297, 1886, 1992, 3326, 3363, 3030]. Vulnerability [1904, 2184, 1931]. vulnerable [1904].

Wales [238]. wandering [1511]. wanderings [1742]. want [1525]. wanted [2721, 2739]. War [751]. warnings [699]. warranted [3300, 1958, 2778]. Wars [1761]. Was [2773, 3298]. Wasserman [2654]. water [116]. Wave [1700, 210, 874]. Wave-functionalism [1700]. Wavefunction [1806]. waves [2885]. way [2101, 3239, 1427, 1428, 126, 353, 1523, 1680, 556, 1878, 2389]. ways [3384, 2547, 2842, 1339, 1303, 1960]. we-intentions [784]. we-perspective [1110]. Weak [482, 1849, 3397, 1152, 1112, 1553]. weakly [2350, 3126]. weakness [816, 2144]. wear [607]. weaving [1462]. Weber [3235]. Weighing [792, 1489, 1397]. weight [3013]. weighting [82]. Weinberg [446]. welfare [2889]. Well [506, 1681, 1756, 608]. well-being [1681, 1756]. Well-ordered [506, 608]. wellbeing [2106, 3001]. Western [117]. Weyl [922, 2520]. wh [1269]. wh-misidentification [1269]. Whales [737]. What-it’s-like [2436]. Where [2318, 1468, 398, 1833, 1301, 2629, 1364, 674, 2439, 151]. whether [596]. which [159, 3202]. while [3381, 203, 3241]. whistles [2109]. white [3283]. who [3384, 2938]. wholes [498]. why-interrogatives [793]. Wigner [3047, 3170]. Wikipedia [609]. wild [1092, 1091]. Wilfrid [1854, 1006]. Will [1423, 2437, 67, 677, 2506, 938, 868, 816, 2126, 3347, 2144, 3280, 1034, 3081, 1370, 3149, 2058]. Willful [3103]. Williamson [1787, 2260, 170, 2198, 2119, 2820, 880]. Williamsonian [1761]. willpower [2672]. wills [2760]. Wilson [1742, 2055]. window [3169]. wisdom [1648]. withholding [1173]. within [129, 3304, 1899, 466, 276, 3174, 1518, 659]. without [1429, 3112, 2340, 513, 2862, 2194, 159, 1078, 139, 699, 2176, 1338, 2605, 2886, 2509, 1970, 101, 2149, 491, 181, 797, 443, 1617, 1733, 2837, 2531, 293, 1292, 1050, 3045, 637, 718, 810, 2001, 1046, 1218, 1868, 3106, 543, 1156, 166, 3323, 1841]. withstand [3206]. witnessably [1268]. Witten [3169]. Wittgenstein [3362, 2773, 2961, 2520, 1647, 3226, 296, 3121, 1286, 755, 3148, 1795, 1962]. Wittgensteinian [3404, 3279]. woke [2613]. Women [596, 2088, 1767, 2085]. won’t [2759]. Word [739, 2206, 405, 1957, 614, 2753, 1325, 297, 2066, 2322]. Words [2209, 2071, 2340, 2636, 124, 2414, 2478, 2736, 845, 1756, 2006, 2319, 2708, 2130, 1575, 2257]. work [1371, 2274, 793, 2328, 3209, 791, 2759, 742]. working [1189, 3385]. works [2740]. World [415, 1793, 1297, 1296, 1017, 3090, 413, 1687, 210, 1580, 477, 2276, 642, 852, 1876, 71]. world-making [1017]. worldly [1417, 211]. worlds [3304, 771, 2040, 1620, 2188, 521, 1633, 522]. Worldview [1933]. worrisome [1295]. worry [1371, 1846, 2989, 2106]. worth [697]. worthiness [3300, 3087]. would [1564, 126, 3046]. writing [576]. wrong [1420, 2421, 189, 2599, 1136, 30, 871, 365, 2831, 2784]. Wrongfulness [1474]. Wronging [3139].

X [2949, 2470]. X-Phi [2470].

Yang [462]. young [1509]. you’re [365]. yours [1238]. YouTube [1201]. Yudja [194]. Yule [2046].

Zen [753]. Zeno [785]. zero [193]. zetetic [2723, 2812, 2917, 2668]. ZFC [2772]. Zika [400, 399]. zombie [1235]. zombies [2268]. zones [2646].

References

Copeland:1996:ATA

[1] B. Jack Copeland and Diane Proudfoot. On Alan Turing’s anticipation of connectionism. Synthese, 108(3):361–377, September 1996. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/BF00413694. See correction

[2361].

Finlay:2010:WOP

[2] Stephen Finlay. What ought probably means, and why you can’t detach it. Synthese, 177(1):67–89, November 2010. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-009-9640-7;   https://link.
springer.com/content/pdf/10.1007/s11229-009-9640-7.pdf. See

correction [1963].

Kim:2013:WN

[3] Joongol Kim. What are numbers? Synthese, 190(6):1099–1112, April 2013. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-011-9883-y;   https://link.springer.com/content/pdf/
10.1007/s11229-011-9883-y.pdf. See comments [11].

Cousins:2017:JLP

[4] Robert D. Cousins. The Jeffreys–Lindley paradox and discovery criteria in high energy physics. Synthese, 194(2):395–432, February 2017. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-014-0525-z;   https://link.springer.com/content/pdf/
10.1007/s11229-014-0525-z.pdf. See erratum [182].

Khudairi:2018:GCM

[5] Hasen Khudairi. Grounding, conceivability, and the mind-body problem. Synthese, 195(2):919–926, February 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-016-1254-2;   https://link.springer.
com/content/pdf/10.1007/s11229-016-1254-2.pdf. See corrections

[1937, 3085].

Fisher:2018:LAM

[6] A. R. J. Fisher. On Lewis against magic: a study of method in metaphysics. Synthese, 195(5):2335–2353, May 2018. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-015-
0679-3;   https://link.springer.com/content/pdf/10.1007/s11229-

015-0679-3.pdf. See correction [235].

Miller:2018:HEP

[7] Mark Miller and Andy Clark. Happily entangled: prediction, emotion, and the embodied mind. Synthese, 195(6):2559–2575, June 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-017-1399-7;   https://link.springer.com/content/pdf/
10.1007/s11229-017-1399-7.pdf. See correction [183].

Pedersen:2018:NRA

[8] Nikolaj Jang Lee Linding Pedersen. Non-rational action in the face of disagreement: an argument against (strong) non-conformism. Synthese, 195(7):2935–2966, July 2018. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-016-1086-0;   https://link.springer.
com/content/pdf/10.1007/s11229-016-1086-0.pdf. See correction

[855].

Gao:2019:MPR

[9] Shan Gao. The measurement problem revisited. Synthese, 196 (1):299–311, January 2019. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1476-y;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1476-y.pdf. See reply [646].

Feintzeig:2019:DDI

[10] Benjamin H. Feintzeig. Deduction and definability in infinite statistical systems. Synthese, 196(5):1831–1861, May 2019. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-017-
1497-6;   https://link.springer.com/content/pdf/10.1007/s11229-

017-1497-6.pdf. See correction [277].

Kim:2019:BCO

[11] Namjoong Kim. Bad company objection to Joongol Kim’s adverbial theory of numbers. Synthese, 196(8):3389–3407, August 2019. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-017-1602-x;   https://link.springer.com/content/pdf/
10.1007/s11229-017-1602-x.pdf. See [3, 74] and correction [74].

Hommen:2019:OCF

[12] David Hommen. Ontological commitments of frame-based knowledge representations. Synthese, 196(10):4155–4183, October 2019. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-017-
1649-8;   https://link.springer.com/content/pdf/10.1007/s11229-
017-1649-8.pdf. See [94].

Rossi:2020:NPM

[13] Carlo Rossi and Kyle Mitchell. New perspectives in metaontology: introduction to the special issue. Synthese, 197(1):1–5, January 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02408-
1;   https://link.springer.com/content/pdf/10.1007/s11229-019-
02408-1.pdf.

vanInwagen:2020:NC

[14] Peter van Inwagen. The neo-Carnapians. Synthese, 197(1):7–32, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-016-1110-4;   https://link.springer.com/content/pdf/
10.1007/s11229-016-1110-4.pdf.

Button:2020:DMO

[15] Tim Button. Deflationary metaphysics and ordinary language. Synthese, 197(1):33–57, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-016-1212-z;   https://link.springer.
com/content/pdf/10.1007/s11229-016-1212-z.pdf.

Biggs:2020:ATD

[16] Stephen Biggs and Jessica Wilson. Abductive two-dimensionalism: a new route to the a priori identification of necessary truths. Synthese, 197(1):59–93, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1444-6;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1444-6.pdf.

Schnieder:2020:GD

[17] Benjamin Schnieder. Grounding and dependence. Synthese, 197 (1):95–124, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1378-z;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1378-z.pdf.

Kelp:2020:CAA

[18] Christoph Kelp and Mona Simion. The C account of assertion: a negative result. Synthese, 197(1):125–137, January 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
1760-5;   https://link.springer.com/content/pdf/10.1007/s11229-
018-1760-5.pdf.

Oms:2020:MSF

[19] Sergi Oms. Minimalism, supervaluations and fixed points. Synthese, 197(1):139–153, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1831-7;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1831-7.pdf.

Berghofer:2020:TPC

[20] Philipp Berghofer. Towards a phenomenological conception of experiential justification. Synthese, 197(1):155–183, January 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
1744-5;   https://link.springer.com/content/pdf/10.1007/s11229-
018-1744-5.pdf.

Stotts:2020:TSS

[21] Megan Henricks Stotts. Toward a sharp semantics/pragmatics distinction. Synthese, 197(1):185–208, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1733-8;   https://link.
springer.com/content/pdf/10.1007/s11229-018-1733-8.pdf.

Klein:2020:RC

[22] Dominik Klein, Johannes Marx, and Simon Scheller. Rationality in context. Synthese, 197(1):209–232, January 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
1773-0;   https://link.springer.com/content/pdf/10.1007/s11229-
018-1773-0.pdf.

Andersen:2020:AGM

[23] Line Edslev Andersen. Acceptable gaps in mathematical proofs. Synthese, 197(1):233–247, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1778-8;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1778-8.pdf.

Jefferson:2020:WDI

[24] Anneli Jefferson. What does it take to be a brain disorder? Synthese, 197(1):249–262, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1784-x;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1784-x.pdf.

Read:2020:MD

[25] James Read and Thomas Møller-Nielsen. Motivating dualities. Synthese, 197(1):263–291, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1817-5;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1817-5.pdf.

Zhou:2020:CIS

[26] Liqian Zhou. Complementarity in information studies. Synthese, 197(1):293–310, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1786-8;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1786-8.pdf.

Moltmann:2020:EP

[27] Friederike Moltmann. Existence predicates. Synthese, 197(1):311–335, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1847-z;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1847-z.pdf.

Dietrich:2020:DPR

[28] Michael Dietrich and Phillip Honenberger. Duhem’s problem revisited: logical versus epistemic formulations and solutions. Synthese, 197(1):337–354, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1845-1;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1845-1.pdf.

Parisi:2020:AO

[29] Andrew Parisi. Atomic ontology. Synthese, 197(1):355–379, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1725-8;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1725-8.pdf.

Shan:2020:KWT

[30] Yafeng Shan. Kuhn’s “wrong turning” and legacy today. Synthese, 197(1):381–406, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1740-9;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1740-9.pdf.

Bruner:2020:BDD

[31] Justin P. Bruner. Bargaining and the dynamics of divisional norms. Synthese, 197(1):407–425, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1729-4;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1729-4.pdf.

Klenk:2020:TFE

[32] Michael Klenk. Third factor explanations and disagreement in metaethics. Synthese, 197(1):427–446, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1875-8;   https://link.
springer.com/content/pdf/10.1007/s11229-018-1875-8.pdf.

Niiniluoto:2020:SAS

[33] Ilkka Niiniluoto. Social aspects of scientific knowledge. Synthese, 197(1):447–468, January 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1868-7;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1868-7.pdf.

Antos:2020:I

[34] Carolin Antos, Neil Barton, Sy-David Friedman, Claudio Ternullo, and John Wigglesworth. Introduction. Synthese, 197(2):469–475, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-019-
02422-3;   https://link.springer.com/content/pdf/10.1007/s11229-
019-02422-3.pdf.

Warren:2020:MCM

[35] Jared Warren and Daniel Waxman. A metasemantic challenge for mathematical determinacy. Synthese, 197(2):477–495, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-016-1266-y;   https://link.springer.com/content/pdf/
10.1007/s11229-016-1266-y.pdf.

Sanders:2020:RF

[36] Sam Sanders. Reverse formalism 16. Synthese, 197(2):497–544, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-017-1322-2;   https://link.springer.com/content/pdf/
10.1007/s11229-017-1322-2.pdf.

Scambler:2020:IUS

[37] Chris Scambler. An indeterminate universe of sets. Synthese, 197 (2):545–573, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-016-1297-4;   https://link.springer.
com/content/pdf/10.1007/s11229-016-1297-4.pdf.

Soysal:2020:WUS

[38] Zeynep Soysal. Why is the universe of sets not a set? Synthese, 197(2):575–597, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1513-x;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1513-x.pdf.

Lingamneni:2020:CWR

[39] Shivaram Lingamneni. Can we resolve the continuum hypothesis? Synthese, 197(2):599–622, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-017-1648-9;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1648-9.pdf.

Barton:2020:MOH

[40] Neil Barton and Sy-David Friedman. Maximality and ontology: how axiom content varies across philosophical frameworks. Synthese, 197 (2):623–649, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1336-9;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1336-9.pdf.

Tsementzis:2020:MEH

[41] Dimitris Tsementzis. A meaning explanation for HoTT. Synthese, 197(2):651–680, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02052-1;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02052-1.pdf.

Corfield:2020:ESH

[42] David Corfield. Expressing ‘the structure of’ in homotopy type theory. Synthese, 197(2):681–700, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-017-1569-7;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1569-7.pdf. See correction

[43].

Corfield:2020:CES

[43] David Corfield. Correction to: Expressing ‘the structure of’ in homotopy type theory. Synthese, 197(2):701, February 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
1724-9;   https://link.springer.com/content/pdf/10.1007/s11229-
018-1724-9.pdf. See [42].

Johns:2020:ETO

[44] Richard Johns. Epistemic theories of objective chance. Synthese, 197(2):703–730, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1719-6;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1719-6.pdf.

Oldofredi:2020:SBT

[45] Andrea Oldofredi. Stochasticity and Bell-type quantum field theory. Synthese, 197(2):731–750, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1720-0;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1720-0.pdf.

Miller:2020:TPP

[46] Kristie Miller, Alex Holcombe, and Andrew James Latham. Temporal phenomenology: phenomenological illusion versus cognitive error. Synthese, 197(2):751–771, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1730-y;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1730-y.pdf.

Pailos:2020:VDS

[47] Federico Matias Pailos. Validity, dialetheism and self-reference. Synthese, 197(2):773–792, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1731-x;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1731-x.pdf.

Antill:2020:EFR

[48] Gregory Antill. Epistemic freedom revisited. Synthese, 197(2):793–815, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1735-6;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1735-6.pdf.

Barrett:2020:ME

[49] Jacob Barrett. Is Maximin egalitarian? Synthese, 197(2):817–837, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1736-5;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1736-5.pdf.

Chung:2020:DRS

[50] Hun Chung and Brian Kogelmann. Diversity and rights: a social choice-theoretic analysis of the possibility of public reason. Synthese, 197(2):839–865, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1737-4;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1737-4.pdf.

Wilhelm:2020:SAL

[51] Isaac Wilhelm. A statistical analysis of luck. Synthese, 197(2):867–885, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1745-4;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1745-4.pdf.

Hsiung:2020:WPD

[52] Ming Hsiung. What paradoxes depend on. Synthese, 197(2):887–913, February 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1748-1;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1748-1.pdf.

Eder:2020:PMU

[53] Anna-Maria A. Eder, Insa Lawler, and Raphael van Riel. Philosophical methods under scrutiny: introduction to the special issue philosophical methods. Synthese, 197(3):915–923, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02051-2;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02051-2.pdf.

Brun:2020:CRE

[54] Georg Brun. Conceptual re-engineering: from explication to reflective equilibrium. Synthese, 197(3):925–954, March 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-017-
1596-4;   https://link.springer.com/content/pdf/10.1007/s11229-
017-1596-4.pdf.

Pinder:2020:SCE

[55] Mark Pinder. On Strawson’s critique of explication as a method in philosophy. Synthese, 197(3):955–981, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-017-1614-6;   https://link.
springer.com/content/pdf/10.1007/s11229-017-1614-6.pdf.

Cordes:2020:CE

[56] Moritz Cordes. The constituents of an explication. Synthese, 197(3):983– 1010, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-017-1615-5;   https://link.springer.com/content/pdf/
10.1007/s11229-017-1615-5.pdf.

Novaes:2020:CEA

[57] Catarina Dutilh Novaes. Carnapian explication and ameliorative analysis: a systematic comparison. Synthese, 197(3):1011–1034, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1732-9;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1732-9.pdf.

Kitsik:2020:ESR

[58] Eve Kitsik. Explication as a strategy for revisionary philosophy. Synthese, 197(3):1035–1056, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1774-z;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1774-z.pdf.

Kusch:2020:GME

[59] Martin Kusch and Robin McKenna. The genealogical method in epistemology. Synthese, 197(3):1057–1076, March 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
1675-1;   https://link.springer.com/content/pdf/10.1007/s11229-
018-1675-1.pdf.

Egler:2020:PEU

[60] Miguel Egler and Lewis D. Ross. Philosophical expertise under the microscope. Synthese, 197(3):1077–1098, March 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
1757-0;   https://link.springer.com/content/pdf/10.1007/s11229-
018-1757-0.pdf.

Lutz:2020:APN

[61] Sebastian Lutz. Armchair philosophy naturalized. Synthese, 197 (3):1099–1125, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1653-z;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1653-z.pdf.

Sgaravatti:2020:ERC

[62] Daniele Sgaravatti. Experience and reasoning: challenging the a priori/a posteriori distinction. Synthese, 197(3):1127–1148, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1718-7;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1718-7.pdf.

Elgin:2020:QBA

[63] Samuel Z. Elgin. On question-begging and analytic content. Synthese, 197(3):1149–1163, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1669-4;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1669-4.pdf.

Steinberg:2020:PPF

[64] Alex Steinberg. Pleonastic propositions and the face value theory. Synthese, 197(3):1165–1180, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1742-7;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1742-7.pdf.

Balaguer:2020:WMD

[65] Mark Balaguer. Why metaphysical debates are not merely verbal (or how to have a non-verbal metaphysical debate). Synthese, 197(3):1181–1201, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-017-1375-2;   https://link.springer.com/content/pdf/
10.1007/s11229-017-1375-2.pdf.

Genot:2020:BAS

[66] Emmanuel J. Genot and Justine Jacot. The brain attics: the strategic role of memory in single and multi-agent inquiry. Synthese, 197(3):1203– 1224, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1743-6;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1743-6.pdf.

Campbell:2020:SFA

[67] Neil Campbell. Self-forming actions, contrastive explanations, and the structure of the will. Synthese, 197(3):1225–1240, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1749-0;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1749-0.pdf.

Buijsman:2020:TRS

[68] Stefan Buijsman. Two roads to the successor axiom. Synthese, 197 (3):1241–1261, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1752-5;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1752-5.pdf.

SantAnna:2020:HCM

[69] André Sant’Anna. The hybrid contents of memory. Synthese, 197 (3):1263–1290, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1753-4;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1753-4.pdf.

Berger:2020:IAA

[70] Jacob Berger. Implicit attitudes and awareness. Synthese, 197(3):1291– 1312, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1754-3;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1754-3.pdf.

Werner:2020:ECC

[71] Konrad Werner. Enactment and construction of the cognitive niche: toward an ontology of the mind-world connection. Synthese, 197(3):1313– 1341, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1756-1;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1756-1.pdf.

McCutcheon:2020:RMJ

[72] Randall G. McCutcheon. Regression to the mean and Judy Benjamin. Synthese, 197(3):1343–1355, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1761-4;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1761-4.pdf.

Wahlberg:2020:CPS

[73] Tobias Hansson Wahlberg. Causal powers and social ontology. Synthese, 197(3):1357–1377, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1763-2;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1763-2.pdf.

Kim:2020:CBC

[74] Namjoong Kim. Correction to: Bad company objection to Joongol Kim’s adverbial theory of numbers. Synthese, 197(3):1379, March 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1738-3;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1738-3.pdf. See [11].

Feldbacher-Escamilla:2020:ISI

[75] Christian J. Feldbacher-Escamilla, Alexander Gebharter, and Peter Brössel Markus Werning. Introduction to the special issue “Logical perspectives on science and cognition”. Synthese, 197(4):1381–1390, April 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-019-
02334-2;   https://link.springer.com/content/pdf/10.1007/s11229-
019-02334-2.pdf.

Spohn:2020:DNR

[76] Wolfgang Spohn. Defeasible normative reasoning. Synthese, 197(4):1391– 1428, April 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02083-2;   https://link.springer.com/content/pdf/
10.1007/s11229-019-02083-2.pdf.

Evans:2020:HWW

[77] Jonathan St. B. T. Evans and Shira Elqayam. How and why we reason from is to ought. Synthese, 197(4):1429–1446, April 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-02041-
4;   https://link.springer.com/content/pdf/10.1007/s11229-018-
02041-4.pdf.

Votsis:2020:TLT

[78] Ioannis Votsis. Theory-ladenness: testing the ‘untestable’. Synthese, 197(4):1447–1465, April 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01992-y;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01992-y.pdf.

Gebharter:2020:NPH

[79] Alexander Gebharter and Nina Retzlaff. A new proposal how to handle counterexamples to Markov causation à la Cartwright, or: fixing the chemical factory. Synthese, 197(4):1467–1486, April 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-02014-
7;   https://link.springer.com/content/pdf/10.1007/s11229-018-
02014-7.pdf.

Huttemann:2020:PPE

[80] Andreas Hüttemann. Processes, pre-emption and further problems. Synthese, 197(4):1487–1509, April 2020. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02058-9;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02058-9.pdf.

Hahn:2020:TTP

[81] Ulrike Hahn, Jens Ulrik Hansen, and Erik J. Olsson. Truth tracking performance of social networks: how connectivity and clustering can make groups less competent. Synthese, 197(4):1511–1541, April 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01936-
6;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01936-6.pdf.

Feldbacher-Escamilla:2020:OAE

[82] Christian J. Feldbacher-Escamilla. An optimality-argument for equal weighting. Synthese, 197(4):1543–1563, April 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02028-1;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02028-1.pdf.

Douven:2020:SC

[83] Igor Douven. Scoring in context. Synthese, 197(4):1565–1580, April 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1867-8;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1867-8.pdf.

Niiniluoto:2020:TON

[84] Ilkka Niiniluoto. Truthlikeness: old and new debates. Synthese, 197(4):1581–1599, April 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01975-z;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01975-z.pdf.

Kuipers:2020:RNT

[85] Theo A. F. Kuipers. Refined nomic truth approximation by revising models and postulates. Synthese, 197(4):1601–1625, April 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1755-2;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1755-2.pdf.

Cevolani:2020:PCA

[86] Gustavo Cevolani and Roberto Festa. A partial consequence account of truthlikeness. Synthese, 197(4):1627–1646, April 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01947-
3;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01947-3.pdf.

Brendel:2020:TMT

[87] Elke Brendel. Truthmaker maximalism and the truthmaker paradox. Synthese, 197(4):1647–1660, April 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01980-2;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01980-2.pdf.

Schurz:2020:TGP

[88] Gerhard Schurz. Twelve great papers: comments and replies. Response to a special issue on logical perspectives on science and cognition — the philosophy of Gerhard Schurz. Synthese, 197(4):1661–1695, April 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02329-
z;   https://link.springer.com/content/pdf/10.1007/s11229-019-
02329-z.pdf.

Sud:2020:PSP

[89] Rohan Sud. Plurivaluationism, supersententialism and the problem of the many languages. Synthese, 197(4):1697–1723, April 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1764-1;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1764-1.pdf.

LaCroix:2020:SSS

[90] Travis LaCroix. On salience and signaling in sender-receiver games: partial pooling, learning, and focal points. Synthese, 197(4):1725–1747, April 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1766-z;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1766-z.pdf.

Williams:2020:PCT

[91] Daniel Williams. Predictive coding and thought. Synthese, 197(4):1749– 1775, April 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1768-x;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1768-x.pdf.

Reuter:2020:UP

[92] Kevin Reuter and Justin Sytsma. Unfelt pain. Synthese, 197(4):1777– 1801, April 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1770-3;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1770-3.pdf.

Kovac:2020:CFC

[93] Srećko Kovac. On causality as the fundamental concept of Gödel’s philosophy. Synthese, 197(4):1803–1838, April 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1771-2;   https://link.
springer.com/content/pdf/10.1007/s11229-018-1771-2.pdf.

Hommen:2020:COC

[94] David Hommen. Correction to: Ontological commitments of framebased knowledge representations. Synthese, 197(4):1839–1840, April 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1772-1;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1772-1.pdf. See [12].

Kistler:2020:INT

[95] Max Kistler. Introduction: new trends in the metaphysics of science. Synthese, 197(5):1841–1846, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1799-3;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1799-3.pdf.

Guay:2020:RBH

[96] Alexandre Guay and Thomas Pradeu. Right out of the box: how to situate metaphysics of science in relation to other metaphysical approaches. Synthese, 197(5):1847–1866, May 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-017-
1576-8;   https://link.springer.com/content/pdf/10.1007/s11229-
017-1576-8.pdf.

Bryant:2020:KCC

[97] Amanda Bryant. Keep the chickens cooped: the epistemic inadequacy of free range metaphysics. Synthese, 197(5):1867–1887, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-017-1398-8;   https://link.springer.com/content/pdf/
10.1007/s11229-017-1398-8.pdf.

Esfeld:2020:PMO

[98] Michael Esfeld. A proposal for a minimalist ontology. Synthese, 197(5):1889–1905, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1426-8;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1426-8.pdf.

Woodward:2020:PML

[99] James Woodward. Physical modality, laws, and counterfactuals. Synthese, 197(5):1907–1929, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1400-5;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1400-5.pdf.

Martinez:2020:SK

[100] Manolo Martínez. Synergic kinds. Synthese, 197(5):1931–1946, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-017-1480-2;   https://link.springer.com/content/pdf/
10.1007/s11229-017-1480-2.pdf.

Khalidi:2020:NCR

[101] Muhammad Ali Khalidi. Neural correlates without reduction: the case of the critical period. Synthese, 197(5):1947–1959, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-017-1439-3;   https://link.springer.com/content/pdf/
10.1007/s11229-017-1439-3.pdf.

Norton:2020:IFM

[102] Joshua Norton. Incubating a future metaphysics: quantum gravity. Synthese, 197(5):1961–1982, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1473-1;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1473-1.pdf.

Statham:2020:NCC

[103] Georgie Statham. Normative commitments, causal structure, and policy disagreement. Synthese, 197(5):1983–2003, May 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
1775-y;   https://link.springer.com/content/pdf/10.1007/s11229-
018-1775-y.pdf.

Queloz:2020:HGC

[104] Matthieu Queloz. How genealogies can affect the space of reasons. Synthese, 197(5):2005–2027, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1777-9;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1777-9.pdf.

Schindler:2020:NHN

[105] Thomas Schindler. A note on Horwich’s notion of grounding. Synthese, 197(5):2029–2038, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1780-1;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1780-1.pdf.

Openshaw:2020:SAS

[106] James Openshaw. Self-ascription and the de se. Synthese, 197(5):2039– 2050, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1781-0;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1781-0.pdf.

Welch:2020:CCB

[107] John R. Welch. Credence for conclusions: a brief for Jeffrey’s rule. Synthese, 197(5):2051–2072, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1782-z;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1782-z.pdf.

Taylor:2020:ECI

[108] Henry Taylor. Emotions, concepts and the indeterminacy of natural kinds. Synthese, 197(5):2073–2093, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1783-y;   https://link.
springer.com/content/pdf/10.1007/s11229-018-1783-y.pdf.

Nutting:2020:BFA

[109] Eileen S. Nutting. Benacerraf, Field, and the agreement of mathematicians. Synthese, 197(5):2095–2110, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1785-9;   https://link.
springer.com/content/pdf/10.1007/s11229-018-1785-9.pdf.

Tambolo:2020:CNM

[110] Luca Tambolo. So close no matter how far: counterfactuals in history of science and the inevitability/contingency controversy. Synthese, 197(5):2111–2141, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1787-7;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1787-7.pdf.

Milojevic:2020:EMF

[111] Miljana Milojevic. Extended mind, functionalism and personal identity. Synthese, 197(5):2143–2170, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1797-5;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1797-5.pdf.

Lohr:2020:CCD

[112] Guido Löhr. Concepts and categorization: do philosophers and psychologists theorize about different things? Synthese, 197(5):2171–2191, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1798-4;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1798-4.pdf.

Wawer:2020:BAF

[113] Jacek Wawer and Alex Malpass. Back to the actual future. Synthese, 197(5):2193–2213, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1802-z;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1802-z.pdf.

Klemick:2020:SMQ

[114] Griffin Klemick. Sellars’ metaethical quasi-realism. Synthese, 197 (5):2215–2243, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1804-x;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1804-x.pdf.

Neufeld:2020:CWP

[115] Eleonore Neufeld. Can we perceive mental states? Synthese, 197(5):2245–2269, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1807-7;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1807-7.pdf.

Schulte:2020:WMC

[116] Peter Schulte. Why mental content is not like water: reconsidering the reductive claims of teleosemantics. Synthese, 197(5):2271–2290, May 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1808-6;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1808-6.pdf.

Carter:2020:ISI

[117] J. Adam Carter, Chienkuo Mi, S. Orestis Palermos, and Duncan Pritchard. Introduction to special issue: knowledge, virtue and action

— Eastern and Western perspectives. Synthese, 197(6):2291–2294, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-020-
02662-8;   https://link.springer.com/content/pdf/10.1007/s11229-
020-02662-8.pdf.

Mi:2020:SRM

[118] Chienkuo Mi and Shane Ryan. Skilful reflection as a master virtue. Synthese, 197(6):2295–2308, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-016-1192-z;   https://link.springer.
com/content/pdf/10.1007/s11229-016-1192-z.pdf.

Iizuka:2020:SVE

[119] Rie Iizuka. Situationism, virtue epistemology, and self-determination theory. Synthese, 197(6):2309–2332, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1750-7;   https://link.
springer.com/content/pdf/10.1007/s11229-018-1750-7.pdf.

Chuang:2020:RDH

[120] Christina Chuang. Rational devotion and human perfection. Synthese, 197(6):2333–2355, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1323-1;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1323-1.pdf.

Wang:2020:SCR

[121] Ju Wang. Scepticism, closure and rationally grounded knowledge: a new solution. Synthese, 197(6):2357–2374, June 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-017-
1337-8;   https://link.springer.com/content/pdf/10.1007/s11229-
017-1337-8.pdf.

Belkoniene:2020:WSW

[122] Miloud Belkoniene. What should we believe about the future? Synthese, 197(6):2375–2386, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1789-5;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1789-5.pdf.

Steglich-Petersen:2020:DLE

[123] Asbjørn Steglich-Petersen. Does luck exclude knowledge or certainty? Synthese, 197(6):2387–2397, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1790-z;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1790-z.pdf.

Caponetto:2020:UTW

[124] Laura Caponetto. Undoing things with words. Synthese, 197(6):2399– 2414, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1805-9;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1805-9.pdf.

Bolinger:2020:RIA

[125] Renée Jorgensen Bolinger. The rational impermissibility of accepting (some) racial generalizations. Synthese, 197(6):2415–2431, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1809-5;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1809-5.pdf.

Hansen:2020:NWR

[126] Nat Hansen. ”Nobody would really talk that way!”: the critical project in contemporary ordinary language philosophy. Synthese, 197(6):2433– 2464, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1812-x;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1812-x.pdf.

Silva:2020:BEI

[127] Paul Silva, Jr. A Bayesian explanation of the irrationality of sexist and racist beliefs involving generic content. Synthese, 197(6):2465–2487, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1813-9;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1813-9.pdf.

Roberts:2020:ALC

[128] Pendaran Roberts. Another look at color primitivism. Synthese, 197(6):2489–2506, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1821-9;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1821-9.pdf.

Chiffi:2020:AIW

[129] Daniele Chiffi and Ahti-Veikko Pietarinen. Abductive inference within a pragmatic framework. Synthese, 197(6):2507–2523, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1824-6;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1824-6.pdf.

Jaag:2020:MBS

[130] Siegfried Jaag and Christian Loew. Making best systems best for us. Synthese, 197(6):2525–2550, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1829-1;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1829-1.pdf.

Bozickovic:2020:AIT

[131] Vojislav Bozickovic. Anti-individualism and transparency. Synthese, 197(6):2551–2564, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1830-8;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1830-8.pdf.

Wilkenfeld:2020:ECD

[132] Daniel A. Wilkenfeld and Tania Lombrozo. Explanation classification depends on understanding: extending the epistemic side-effect effect. Synthese, 197(6):2565–2592, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1835-3;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1835-3.pdf.

Morris:2020:MRC

[133] Kevin Morris. Multiple realization and compositional variation. Synthese, 197(6):2593–2611, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1837-1;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1837-1.pdf.

Hu:2020:EAF

[134] Xingming Hu. The epistemic account of faultless disagreement. Synthese, 197(6):2613–2630, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1848-y;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1848-y.pdf.

Lutz:2020:BBP

[135] Matt Lutz. Background beliefs and plausibility thresholds: defending explanationist evidentialism. Synthese, 197(6):2631–2647, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1849-x;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1849-x.pdf.

Roeber:2020:HAP

[136] Blake Roeber. How to argue for pragmatic encroachment. Synthese, 197(6):2649–2664, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1850-4;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1850-4.pdf.

Ishida:2020:DTR

[137] Yoichi Ishida. In dialectical tension: realist and instrumentalist attitudes in scientific practice. Synthese, 197(6):2665–2694, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1855-z;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1855-z.pdf.

Skrzypulec:2020:TVS

[138] Błazej Skrzypulec. Transitivity of visual sameness. Synthese, 197 (6):2695–2719, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1858-9;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1858-9.pdf.

Donhauser:2020:IEM

[139] Justin Donhauser. Informative ecological models without ecological forces. Synthese, 197(6):2721–2743, June 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1859-8;   https://link.
springer.com/content/pdf/10.1007/s11229-018-1859-8.pdf.

Broncano-Berrocal:2020:TTE

[140] Fernando Broncano-Berrocal and Jesús Vega-Encabo. A taxonomy of types of epistemic dependence: introduction to the Synthese special issue on epistemic dependence. Synthese, 197(7):2745–2763, July 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02233-
6;   https://link.springer.com/content/pdf/10.1007/s11229-019-
02233-6.pdf.

McMyler:2020:MOO

[141] Benjamin McMyler. On not making up one’s own mind. Synthese, 197(7):2765–2781, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1563-0;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1563-0.pdf.

Goldberg:2020:EEE

[142] Sanford C. Goldberg. Epistemically engineered environments. Synthese, 197(7):2783–2802, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1413-0;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1413-0.pdf.

Nguyen:2020:CIR

[143] C. Thi Nguyen. Cognitive islands and runaway echo chambers: problems for epistemic dependence on experts. Synthese, 197(7):2803–2821, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1692-0;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1692-0.pdf.

Hallsson:2020:DDE

[144] Bjørn G. Hallsson and Klemens Kappel. Disagreement and the division of epistemic labor. Synthese, 197(7):2823–2847, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
1788-6;   https://link.springer.com/content/pdf/10.1007/s11229-
018-1788-6.pdf.

Simion:2020:HAR

[145] Mona Simion and Christoph Kelp. How to be an anti-reductionist. Synthese, 197(7):2849–2866, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1722-y;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1722-y.pdf.

Greco:2020:TKG

[146] John Greco. The transmission of knowledge and garbage. Synthese, 197(7):2867–2878, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02090-3;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02090-3.pdf.

Pritchard:2020:ARE

[147] Duncan Pritchard. Anti-risk epistemology and negative epistemic dependence. Synthese, 197(7):2879–2894, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-017-1586-6;   https://link.
springer.com/content/pdf/10.1007/s11229-017-1586-6.pdf.

Broncano-Berrocal:2020:EDC

[148] Fernando Broncano-Berrocal. Epistemic dependence and cognitive ability. Synthese, 197(7):2895–2912, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-017-1559-9;   https://link.
springer.com/content/pdf/10.1007/s11229-017-1559-9.pdf.

Page:2020:DVE

[149] Walker Page. Defending virtue epistemology: epistemic dependence in testimony and extended cognition. Synthese, 197(7):2913–2936, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-017-1518-5;   https://link.springer.com/content/pdf/
10.1007/s11229-017-1518-5.pdf.

Carter:2020:IAE

[150] J. Adam Carter. Intellectual autonomy, epistemic dependence and cognitive enhancement. Synthese, 197(7):2937–2961, July 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-017-
1549-y;   https://link.springer.com/content/pdf/10.1007/s11229-
017-1549-y.pdf.

Viebahn:2020:PEW

[151] Emanuel Viebahn. Presentism, eternalism and where things are located. Synthese, 197(7):2963–2974, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1816-6;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1816-6.pdf.

Mariscal:2020:LLO

[152] Carlos Mariscal and W. Ford Doolittle. Life and life only: a radical alternative to life definitionism. Synthese, 197(7):2975–2989, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1852-2;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1852-2.pdf.

Pickel:2020:SPT

[153] Bryan Pickel. Structured propositions and trivial composition. Synthese, 197(7):2991–3006, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1853-1;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1853-1.pdf.

Laraudogoitia:2020:SSI

[154] Jon Pérez Laraudogoitia. Some surprising instabilities in idealized dynamical systems. Synthese, 197(7):3007–3026, July 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
1861-1;   https://link.springer.com/content/pdf/10.1007/s11229-
018-1861-1.pdf.

Fraser:2020:DRG

[155] Doreen Fraser. The development of renormalization group methods for particle physics: Formal analogies between classical statistical mechanics and quantum field theory. Synthese, 197(7):3027–3063, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1862-0;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1862-0.pdf.

Simmons:2020:ICN

[156] Byron Simmons. Impure concepts and non-qualitative properties. Synthese, 197(7):3065–3086, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1871-z;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1871-z.pdf.

Mitchell:2020:NCC

[157] Jonathan Mitchell. On the non-conceptual content of affectiveevaluative experience. Synthese, 197(7):3087–3111, July 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
1872-y;   https://link.springer.com/content/pdf/10.1007/s11229-
018-1872-y.pdf.

Leitgeb:2020:WPM

[158] Hannes Leitgeb. Why pure mathematical truths are metaphysically necessary: a set-theoretic explanation. Synthese, 197(7):3113–3120, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1873-x;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1873-x.pdf.

DeHaro:2020:PBW

[159] Sebastian De Haro and Henk W. de Regt. A precipice below which lies absurdity? Theories without a spacetime and scientific understanding. Synthese, 197(7):3121–3149, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1874-9;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1874-9.pdf.

Carrara:2020:ICE

[160] Massimiliano Carrara and Ciro De Florio. Identity criteria: an epistemic path to conceptual grounding. Synthese, 197(7):3151–3169, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1878-5;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1878-5.pdf.

Perrine:2020:SID

[161] Timothy Perrine. Strong internalism, doxastic involuntarism, and the costs of compatibilism. Synthese, 197(7):3171–3191, July 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
1879-4;   https://link.springer.com/content/pdf/10.1007/s11229-
018-1879-4.pdf.

Watkins:2020:MSK

[162] Eric Watkins, Marcus Willaschek, and Clinton Tolley. Minisymposium on Kant and cognition. Synthese, 197(8):3193–3194, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-017-1623-5;   https://link.springer.com/content/pdf/
10.1007/s11229-017-1623-5.pdf.

Willaschek:2020:KCK

[163] Marcus Willaschek and Eric Watkins. Kant on cognition and knowledge. Synthese, 197(8):3195–3213, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-017-1624-4;   https://link.
springer.com/content/pdf/10.1007/s11229-017-1624-4.pdf.

Tolley:2020:KPC

[164] Clinton Tolley. Kant on the place of cognition in the progression of our representations. Synthese, 197(8):3215–3244, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-017-1625-3;   https://link.springer.com/content/pdf/
10.1007/s11229-017-1625-3.pdf.

Langford:2020:KJC

[165] Simon Langford. Knowledge judgements and cognitive psychology. Synthese, 197(8):3245–3259, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1880-y;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1880-y.pdf.

Zhao:2020:KS

[166] Haicheng Zhao. Knowledge without safety. Synthese, 197(8):3261–3278, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1881-x;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1881-x.pdf.

Greenberg:2020:SBA

[167] Alexander Greenberg. Should I believe all the truths? Synthese, 197(8):3279–3303, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1882-9;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1882-9.pdf.

Atkinson:2020:TLT

[168] David Atkinson and Jeanne Peijnenburg. “Till at last there remain nothing”. Synthese, 197(8):3305–3323, August 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
1883-8;   https://link.springer.com/content/pdf/10.1007/s11229-

018-1883-8.pdf. See correction [229].

Ridge:2020:NCE

[169] Michael Ridge. Normative certitude for expressivists. Synthese, 197 (8):3325–3347, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1884-7;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1884-7.pdf.

Immerman:2020:WCK

[170] Daniel Immerman. Williamson, closure, and KK. Synthese, 197 (8):3349–3373, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1890-9;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1890-9.pdf.

Unwin:2020:CIE

[171] Nicholas Unwin. Continuants, identity and essentialism. Synthese, 197(8):3375–3394, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1891-8;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1891-8.pdf.

Valaris:2020:TDR

[172] Markos Valaris. Thinking by doing: Rylean regress and the metaphysics of action. Synthese, 197(8):3395–3412, August 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
1893-6;   https://link.springer.com/content/pdf/10.1007/s11229-
018-1893-6.pdf.

Saenz:2020:AT

[173] Noël Blas Saenz. An account of truthmaking. Synthese, 197(8): 3413–3435, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1894-5;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1894-5.pdf.

Constantin:2020:RRR

[174] Jan Constantin. Replacement and reasoning: a reliabilist account of epistemic defeat. Synthese, 197(8):3437–3457, August 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01895-
y;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01895-y.pdf.

Fassio:2020:GAK

[175] Davide Fassio. On the generality argument for the knowledge norm. Synthese, 197(8):3459–3480, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01896-x;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01896-x.pdf.

Samaroo:2020:PEC

[176] Ryan Samaroo. The principle of equivalence as a criterion of identity. Synthese, 197(8):3481–3505, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01897-w;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01897-w.pdf.

Vukov:2020:HPE

[177] Joseph Vukov and Charles Lassiter. How to power encultured minds. Synthese, 197(8):3507–3534, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01899-8;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01899-8.pdf.

Skipper:2020:HSF

[178] Mattias Skipper and Jens Christian Bjerring. Hyperintensional semantics: a Fregean approach. Synthese, 197(8):3535–3558, August 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01900-
4;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01900-4.pdf.

Duncan:2020:KT

[179] Matt Duncan. Knowledge of things. Synthese, 197(8):3559–3592, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
01904-0;   https://link.springer.com/content/pdf/10.1007/s11229-
018-01904-0.pdf.

Whiting:2020:KJS

[180] Daniel Whiting. Knowledge, justification, and (a sort of) safe belief. Synthese, 197(8):3593–3609, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01905-z;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01905-z.pdf.

Leefmann:2020:KSE

[181] Jon Leefmann and Steffen Lesle. Knowledge from scientific expert testimony without epistemic trust. Synthese, 197(8):3611–3641, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01908-
w;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01908-w.pdf.

Cousins:2020:EJL

[182] Robert D. Cousins. Erratum to: The Jeffreys–Lindley paradox and discovery criteria in high energy physics. Synthese, 197(8):3643, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-015-0687-3;   https://link.springer.com/content/pdf/
10.1007/s11229-015-0687-3.pdf. See [4].

Miller:2020:CHE

[183] Mark Miller and Andy Clark. Correction to: Happily entangled: prediction, emotion, and the embodied mind. Synthese, 197(8):3645, August 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-019-
02260-3;   https://link.springer.com/content/pdf/10.1007/s11229-
019-02260-3.pdf. See [7].

Pantsar:2020:MCE

[184] Markus Pantsar. Mathematical cognition and enculturation: introduction to the Synthese special issue. Synthese, 197(9):3647–3655, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-019-
02478-1;   https://link.springer.com/content/pdf/10.1007/s11229-
019-02478-1.pdf.

Jones:2020:NNR

[185] Max Jones. Numerals and neural reuse. Synthese, 197(9):3657–3681, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-01922-y;   https://link.springer.com/content/pdf/
10.1007/s11229-018-01922-y.pdf. See correction [186].

Jones:2020:CNN

[186] Max Jones. Correction to: Numerals and neural reuse. Synthese, 197(9):3683, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01990-0;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01990-0.pdf.

See [185].

Fabry:2020:CEC

[187] Regina E. Fabry. The cerebral, extra-cerebral bodily, and sociocultural dimensions of enculturated arithmetical cognition. Synthese, 197(9):3685–3720, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02238-1;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02238-1.pdf.

Johansen:2020:MRM

[188] Mikkel W. Johansen and Morten Misfeldt. Material representations in mathematical research practice. Synthese, 197(9):3721–3741, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
02033-4;   https://link.springer.com/content/pdf/10.1007/s11229-
018-02033-4.pdf.

Larvor:2020:WSW

[189] Brendan Larvor. Why ‘scaffolding’ is the wrong metaphor: the cognitive usefulness of mathematical representations. Synthese, 197(9):3743–3756, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02039-y;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02039-y.pdf.

Vold:2020:EMC

[190] Karina Vold and Dirk Schlimm. Extended mathematical cognition: external representations with non-derived content. Synthese, 197(9):3757–3777, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02097-w;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02097-w.pdf.

Buijsman:2020:HNS

[191] Stefan Buijsman. How numerals support new cognitive capacities. Synthese, 197(9):3779–3796, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01989-7;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01989-7.pdf.

Pelland:2020:WNI

[192] Jean-Charles Pelland. What’s new: innovation and enculturation of arithmetical practices. Synthese, 197(9):3797–3822, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-02060-
1;   https://link.springer.com/content/pdf/10.1007/s11229-018-
02060-1.pdf.

Barton:2020:APP

[193] Neil Barton. Absence perception and the philosophy of zero. Synthese, 197(9):3823–3850, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02220-x;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02220-x.pdf.

Rothstein:2020:QEY

[194] Susan Rothstein and Suzi Lima. Quantity evaluations in Yudja: judgements, language and cultural practice. Synthese, 197(9):3851–3873, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-02016-5;   https://link.springer.com/content/pdf/
10.1007/s11229-018-02016-5.pdf.

Rittberg:2020:EIM

[195] Colin Jakob Rittberg, Fenner Stanley Tanswell, and Jean Paul Van Bendegem. Epistemic injustice in mathematics. Synthese, 197(9):3875–3904, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01981-1;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01981-1.pdf.

vanGessel:2020:AMI

[196] Thom van Gessel. Action models in inquisitive logic. Synthese, 197 (9):3905–3945, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1886-5;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1886-5.pdf.

Okon:2020:CBQ

[197] Elias Okon and Miguel Ángel Sebastián. A consciousness-based quantum objective collapse model. Synthese, 197(9):3947–3967, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1887-4;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1887-4.pdf.

Schiller:2020:PD

[198] Henry Ian Schiller. Phenomenal dispositions. Synthese, 197(9):3969– 3980, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.

1007/s11229-018-01909-9;   https://link.springer.com/content/
pdf/10.1007/s11229-018-01909-9.pdf.

Kim:2020:ATN

[199] Joongol Kim. The adverbial theory of numbers: some clarifications. Synthese, 197(9):3981–4000, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01911-1;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01911-1.pdf.

Matarese:2020:CSH

[200] Vera Matarese. A challenge for Super-Humeanism: the problem of immanent comparisons. Synthese, 197(9):4001–4020, September 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01914-
y;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01914-y.pdf.

Morris:2020:SSL

[201] Rick Morris. Stranger in a strange land: an optimal-environments account of evolutionary mismatch. Synthese, 197(9):4021–4046, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
01915-x;   https://link.springer.com/content/pdf/10.1007/s11229-
018-01915-x.pdf.

Guardo:2020:MRS

[202] Andrea Guardo. Meaning relativism and subjective idealism. Synthese, 197(9):4047–4064, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01917-9;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01917-9.pdf.

Hirvela:2020:HSS

[203] Jaakko Hirvelä. How to stay safe while extending the mind. Synthese, 197(9):4065–4081, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01920-0;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01920-0.pdf.

Shagrir:2020:DSV

[204] Oron Shagrir. In defense of the semantic view of computation. Synthese, 197(9):4083–4108, September 2020. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01921-z;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01921-z.pdf.

Constantin:2020:EAP

[205] Jan Constantin and Thomas Grundmann. Epistemic authority: preemption through source sensitive defeat. Synthese, 197(9):4109–4130, September 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-01923-x;   https://link.springer.com/content/pdf/
10.1007/s11229-018-01923-x.pdf.

Ismael:2020:QHN

[206] Jenann Ismael and Jonathan Schaffer. Quantum holism: nonseparability as common ground. Synthese, 197(10):4131–4160, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-016-1201-2;   https://link.springer.com/content/pdf/
10.1007/s11229-016-1201-2.pdf.

Bigaj:2020:QEC

[207] Tomasz Bigaj. On quantum entanglement, counterfactuals, causality and dispositions. Synthese, 197(10):4161–4185, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-016-
1304-9;   https://link.springer.com/content/pdf/10.1007/s11229-
016-1304-9.pdf.

Miller:2020:TNH

[208] Elizabeth Miller. Two notions of holism. Synthese, 197(10):4187–4206, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1710-2;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1710-2.pdf.

Dewar:2020:B

[209] Neil Dewar. La Bohume. Synthese, 197(10):4207–4225, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1800-1;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1800-1.pdf.

Ney:2020:FWW

[210] Alyssa Ney. Finding the world in the wave function: some strategies for solving the macro-object problem. Synthese, 197(10):4227–4249, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-017-1349-4;   https://link.springer.com/content/pdf/
10.1007/s11229-017-1349-4.pdf.

Torza:2020:QMI

[211] Alessandro Torza. Quantum metaphysical indeterminacy and worldly incompleteness. Synthese, 197(10):4251–4264, October 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-017-
1581-y;   https://link.springer.com/content/pdf/10.1007/s11229-
017-1581-y.pdf.

Healey:2020:PVM

[212] Richard Healey. A pragmatist view of the metaphysics of entanglement. Synthese, 197(10):4265–4302, October 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-016-
1204-z;   https://link.springer.com/content/pdf/10.1007/s11229-
016-1204-z.pdf.

Wallace:2020:LRP

[213] David Wallace. Lessons from realistic physics for the metaphysics of quantum theory. Synthese, 197(10):4303–4318, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
1706-y;   https://link.springer.com/content/pdf/10.1007/s11229-
018-1706-y.pdf.

Kantin:2020:WCN

[214] Holly Kantin. Why compositional nihilism dissolves puzzles. Synthese, 197(10):4319–4340, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01924-w;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01924-w.pdf.

Boge:2020:AAG

[215] Florian J. Boge. An argument against global no miracles arguments. Synthese, 197(10):4341–4363, October 2020. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01925-9;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01925-9.pdf.

See correction [970].

Liu:2020:RJD

[216] Yang Liu and Huw Price. Ramsey and Joyce on deliberation and prediction. Synthese, 197(10):4365–4386, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01926-8;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01926-8.pdf.

Stegmann:2020:PME

[217] Ulrich E. Stegmann. Physical models and embodied cognition. Synthese, 197(10):4387–4405, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01927-7;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01927-7.pdf.

Goldberg:2020:WWO

[218] Sanford C. Goldberg. What we owe each other, epistemologically speaking: ethico-political values in social epistemology. Synthese, 197(10):4407–4423, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01928-6;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01928-6.pdf.

Stefansson:2020:RAI

[219] H. Orri Stefánsson. Is risk aversion irrational? Examining the “fallacy” of large numbers. Synthese, 197(10):4425–4437, October 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01929-
5;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01929-5.pdf.

Cyr:2020:TF

[220] Taylor W. Cyr. Timelessness and freedom. Synthese, 197(10):4439–4453, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-01930-y;   https://link.springer.com/content/pdf/
10.1007/s11229-018-01930-y.pdf.

Carruthers:2020:MCS

[221] Peter Carruthers. Mechanisms for constrained stochasticity. Synthese, 197(10):4455–4473, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01933-9;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01933-9.pdf.

Olsson:2020:DPP

[222] Erik J. Olsson. A diachronic perspective on peer disagreement in veritistic social epistemology. Synthese, 197(10):4475–4493, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01935-
7;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01935-7.pdf.

Kauss:2020:CDT

[223] Dominik Kauss. Credence as doxastic tendency. Synthese, 197(10):4495– 4518, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-018-01938-4;   https://link.springer.com/content/
pdf/10.1007/s11229-018-01938-4.pdf.

Pezet:2020:FSP

[224] Robert E. Pezet. Finding satisfaction in presentism. Synthese, 197(10): 4519–4531, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-01940-w;   https://link.springer.com/content/
pdf/10.1007/s11229-018-01940-w.pdf.

Butlin:2020:RAC

[225] Patrick Butlin. Representation and the active consumer. Synthese, 197(10):4533–4550, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01941-9;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01941-9.pdf.

Daoust:2020:ERC

[226] Marc-Kevin Daoust. The explanatory role of consistency requirements. Synthese, 197(10):4551–4569, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01942-8;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01942-8.pdf.

Darrigol:2020:CPV

[227] Olivier Darrigol. Constitutive principles versus comprehensibility conditions in post-Kantian physics. Synthese, 197(10):4571–4616, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
01948-2;   https://link.springer.com/content/pdf/10.1007/s11229-
018-01948-2.pdf.

Vanello:2020:AMS

[228] Daniel Vanello. Affect, motivational states, and evaluative concepts. Synthese, 197(10):4617–4636, October 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02120-0;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02120-0.pdf.

Atkinson:2020:CTL

[229] David Atkinson and Jeanne Peijnenburg. Correction to: “Till at last there remain nothing”. Synthese, 197(10):4637, October 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01969-
x;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01969-x.pdf. See [168].

Marfori:2020:LDK

[230] Marianna Antonutti Marfori and Pierluigi Graziani. The legacy of D. K. Lewis: Introduction to the special issue. Synthese, 197(11):4639–4644, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02822-w;   https://
link.springer.com/content/pdf/10.1007/s11229-020-02822-w.pdf.

Divers:2020:OAL

[231] John Divers and Jade Fletcher. (Once again) Lewis on the analysis of modality. Synthese, 197(11):4645–4668, November 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
1698-7;   https://link.springer.com/content/pdf/10.1007/s11229-
018-1698-7.pdf.

Roca-Royes:2020:GMR

[232] Sonia Roca-Royes. Genuine Modal Realism, the Humean thesis and advanced modalizing. Synthese, 197(11):4669–4690, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1728-5;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1728-5.pdf.

Varzi:2020:CTE

[233] Achille C. Varzi. Counterpart theories for everyone. Synthese, 197(11):4691–4715, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02720-1;   https://
link.springer.com/content/pdf/10.1007/s11229-020-02720-1.pdf.

Nolan:2020:KML

[234] Daniel Nolan. It’s a kind of magic: Lewis, magic and properties. Synthese, 197(11):4717–4741, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-014-0565-4;   https://link.
springer.com/content/pdf/10.1007/s11229-014-0565-4.pdf.

Fisher:2020:CLA

[235] A. R. J. Fisher. Correction to: On Lewis against magic: a study of method in metaphysics. Synthese, 197(11):4743, November 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02323-
5;   https://link.springer.com/content/pdf/10.1007/s11229-019-
02323-5.pdf. See [6].

Borghini:2020:MMH

[236] Andrea Borghini and Giorgio Lando. Mereological monism and Humean supervenience. Synthese, 197(11):4745–4765, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-016-1048-6;   https://link.springer.com/content/pdf/
10.1007/s11229-016-1048-6.pdf.

Calosi:2020:EU

[237] Claudio Calosi. An elegant universe. Synthese, 197(11):4767–4782, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-015-0952-5;   https://link.springer.com/content/pdf/
10.1007/s11229-015-0952-5.pdf.

Maslen:2020:CAP

[238] Cei Maslen. Causation, absences, and the Prince of Wales. Synthese, 197 (11):4783–4794, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-015-0980-1;   https://link.springer.
com/content/pdf/10.1007/s11229-015-0980-1.pdf.

Hajek:2020:MD

[239] Alan Hájek. Minkish dispositions. Synthese, 197(11):4795–4811, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-015-1011-y;   https://link.springer.com/content/pdf/
10.1007/s11229-015-1011-y.pdf.

Bottani:2020:BBI

[240] Andrea C. Bottani. Bringing back intrinsics to enduring things. Synthese, 197(11):4813–4834, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-016-1157-2;   https://link.springer.
com/content/pdf/10.1007/s11229-016-1157-2.pdf.

Torrengo:2020:TTC

[241] Giuliano Torrengo. Time travel and coincidence-free local dynamical theories. Synthese, 197(11):4835–4846, November 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-017-
1433-9;   https://link.springer.com/content/pdf/10.1007/s11229-
017-1433-9.pdf.

Fano:2020:PIL

[242] Vincenzo Fano and Giovanni Macchia. A physical interpretation of Lewis’ discrepancy between personal and external time in time travels. Synthese, 197(11):4847–4866, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-016-1104-2;   https://link.springer.
com/content/pdf/10.1007/s11229-016-1104-2.pdf.

Collins:2020:SB

[243] John Collins. Simple belief. Synthese, 197(11):4867–4885, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1746-3;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1746-3.pdf.

Piazza:2020:PRN

[244] Mario Piazza and Nevia Dolcini. Possibilities regained: neo-Lewisian contextualism and ordinary life. Synthese, 197(11):4887–4906, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-015-0963-2;   https://link.springer.com/content/pdf/
10.1007/s11229-015-0963-2.pdf.

Henderson:2020:EEE

[245] David Henderson and Terence Horgan. Evidentially embedded epistemic entitlement. Synthese, 197(11):4907–4926, November 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01952-
6;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01952-6.pdf.

Pauw:2020:DFP

[246] Sylvia Pauw. Deducing false propositions from true ideas: Nieuwentijt on mathematical reasoning. Synthese, 197(11):4927–4945, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01953-
5;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01953-5.pdf.

Palmira:2020:IDA

[247] Michele Palmira. Inquiry and the doxastic attitudes. Synthese, 197(11):4947–4973, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01955-3;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01955-3.pdf.

Ranalli:2020:DDH

[248] Chris Ranalli. Deep disagreement and hinge epistemology. Synthese, 197(11):4975–5007, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01956-2;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01956-2.pdf.

Masny:2020:FSJ

[249] Michal Masny. Friedman on suspended judgment. Synthese, 197(11):5009–5026, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01957-1;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01957-1.pdf.

Haas:2020:JTR

[250] Gordian Haas. JuDAS: a theory of rational belief revision. Synthese, 197(11):5027–5050, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01958-0;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01958-0.pdf.

Todd:2020:PFC

[251] Patrick Todd. The problem of future contingents: scoping out a solution. Synthese, 197(11):5051–5072, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01959-z;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01959-z.pdf.

Jackson:2020:BCE

[252] Elizabeth Jackson. Belief, credence, and evidence. Synthese, 197(11):5073–5092, November 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01965-1;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01965-1.pdf.

Kelp:2020:EES

[253] Christoph Kelp. The epistemology of Ernest Sosa: an introduction. Synthese, 197(12):5093–5100, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02897-5;   https://
link.springer.com/content/pdf/10.1007/s11229-020-02897-5.pdf.

Bernecker:2020:AGM

[254] Sven Bernecker. Against global method safety. Synthese, 197(12):5101– 5116, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-018-02008-5;   https://link.springer.com/content/
pdf/10.1007/s11229-018-02008-5.pdf.

Carter:2020:SKJ

[255] J. Adam Carter. Sosa on knowledge, judgment and guessing. Synthese, 197(12):5117–5136, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-016-1181-2;   https://link.springer.
com/content/pdf/10.1007/s11229-016-1181-2.pdf.

Chrisman:2020:PNH

[256] Matthew Chrisman. Performance normativity and here-and-now doxastic agency. Synthese, 197(12):5137–5145, December 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-017-
1641-3;   https://link.springer.com/content/pdf/10.1007/s11229-
017-1641-3.pdf.

Greco:2020:SS

[257] John Greco. Safety in Sosa. Synthese, 197(12):5147–5157, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1863-z;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1863-z.pdf. See correction [258].

Greco:2020:CSS

[258] John Greco. Correction to: Safety in Sosa. Synthese, 197(12):5159, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-01906-y;   https://link.springer.com/content/pdf/
10.1007/s11229-018-01906-y.pdf. See [257].

Grundmann:2020:SSC

[259] Thomas Grundmann. Saving safety from counterexamples. Synthese, 197(12):5161–5185, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1677-z;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1677-z.pdf.

Kelp:2020:HBN

[260] Christoph Kelp, Cameron Boult, Fernando Broncano-Berrocal, Paul Dimmock, Harmen Ghijsen, and Mona Simion. Hoops and barns: a new dilemma for Sosa. Synthese, 197(12):5187–5202, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-017-1461-5;   https://link.springer.com/content/pdf/
10.1007/s11229-017-1461-5.pdf.

Vargas:2020:CPP

[261] Miguel Ángel Fernández Vargas. The ‘Conditional Position Problem’ for epistemic externalism. Synthese, 197(12):5203–5224, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1765-0;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1765-0.pdf.

Horwich:2020:STK

[262] Paul Horwich. Sosa’s theory of knowledge. Synthese, 197(12):5225–5232, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1697-8;   https://link.springer.com/content/pdf/
10.1007/s11229-018-1697-8.pdf.

Kallestrup:2020:GVE

[263] Jesper Kallestrup. Group virtue epistemology. Synthese, 197(12): 5233–5251, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-016-1225-7;   https://link.springer.
com/content/pdf/10.1007/s11229-016-1225-7.pdf.

Littlejohn:2020:TKS

[264] Clayton Littlejohn. Truth, knowledge, and the standard of proof in criminal law. Synthese, 197(12):5253–5286, December 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-017-
1608-4;   https://link.springer.com/content/pdf/10.1007/s11229-
017-1608-4.pdf.

McGrath:2020:SEV

[265] Matthew McGrath. Sosa on epistemic value: a Kantian obstacle. Synthese, 197(12):5287–5300, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1717-8;   https://link.
springer.com/content/pdf/10.1007/s11229-018-1717-8.pdf.

Nolfi:2020:FBJ

[266] Kate Nolfi. Functional belief and judgmental belief. Synthese, 197 (12):5301–5317, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1510-0;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1510-0.pdf.

Soteriou:2020:DAJ

[267] Matthew Soteriou. Dreams, agency, and judgement. Synthese, 197 (12):5319–5334, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1496-7;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1496-7.pdf.

Sylvan:2020:CPE

[268] Kurt L. Sylvan. Can performance epistemology explain higher epistemic value? Synthese, 197(12):5335–5356, December 2020. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-017-
1577-7;   https://link.springer.com/content/pdf/10.1007/s11229-
017-1577-7.pdf.

Wedgwood:2020:IVT

[269] Ralph Wedgwood. The internalist virtue theory of knowledge. Synthese, 197(12):5357–5378, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1707-x;   https://link.springer.
com/content/pdf/10.1007/s11229-018-1707-x.pdf.

Zalabardo:2020:SSD

[270] José L. Zalabardo. Safety, sensitivity and differential support. Synthese, 197(12):5379–5388, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-017-1645-z;   https://link.springer.
com/content/pdf/10.1007/s11229-017-1645-z.pdf.

Thorn:2020:JAB

[271] Paul D. Thorn. The joint aggregation of beliefs and degrees of belief. Synthese, 197(12):5389–5409, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01966-0;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01966-0.pdf.

Hodgson:2020:RCS

[272] Thomas Hodgson. Russellians can solve the problem of empty names with nonsingular propositions. Synthese, 197(12):5411–5433, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01971-
3;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01971-3.pdf.

Ye:2020:HOD

[273] Ru Ye. Higher-order defeat and intellectual responsibility. Synthese, 197(12):5435–5455, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01972-2;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01972-2.pdf.

Paprzycka-Hausman:2020:KCE

[274] Katarzyna Paprzycka-Hausman. Knowledge of consequences: an explanation of the epistemic side-effect effect. Synthese, 197(12):5457–5490, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-01973-1;   https://link.springer.com/content/pdf/
10.1007/s11229-018-01973-1.pdf.

Rose:2020:FTD

[275] David Rose, Jonathan Schaffer, and Kevin Tobia. Folk teleology drives persistence judgments. Synthese, 197(12):5491–5509, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01974-
0;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01974-0.pdf.

Mirski:2020:ABV

[276] Robert Mirski and Arkadiusz Gut. Action-based versus cognitivist perspectives on socio-cognitive development: culture, language and social experience within the two paradigms. Synthese, 197(12):5511–5537, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01976-y;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01976-y.pdf.

Feintzeig:2020:CDD

[277] Benjamin H. Feintzeig. Correction to: Deduction and definability in infinite statistical systems. Synthese, 197(12):5539–5540, December 2020. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01968-
y;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01968-y.pdf. See [10].

Silva:2021:SIR

[278] Marcos Silva and Francicleber Ferreira. Special issue on radical views on cognition: introduction. Synthese, 198(S1):1–4, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02959-8.

Segundo-Ortin:2021:SBC

[279] Miguel Segundo-Ortin and Daniel D. Hutto. Similarity-based cognition: radical enactivism meets cognitive neuroscience. Synthese, 198 (S1):5–23, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02505-1.

Myin:2021:ICL

[280] Erik Myin. On the importance of correctly locating content: why and how REC can afford affordance perception. Synthese, 198(S1): 25–39, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02607-1.

Ramstead:2021:MIB

[281] Maxwell J. D. Ramstead, Michael D. Kirchhoff, and Karl J. Friston. Multiscale integration: beyond internalism and externalism. Synthese, 198 (S1):41–70, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02115-x.

Villalobos:2021:ATE

[282] Mario Villalobos and Simón Palacios. Autopoietic theory, enactivism, and their incommensurable marks of the cognitive. Synthese, 198(S1): 71–87, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02376-6.

Schlicht:2021:PEA

[283] Tobias Schlicht and Tobias Starzak. Prospects of enactivist approaches to intentionality and cognition. Synthese, 198(S1):89–113, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02361-z.

deOliveira:2021:REC

[284] Guilherme Sanches de Oliveira, Vicente Raja, and Anthony Chemero. Radical embodied cognitive science and “Real cognition”. Synthese, 198 (S1):115–136, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02475-4.

Keijzer:2021:DCC

[285] Fred Keijzer. Demarcating cognition: the cognitive life sciences. Synthese, 198(S1):137–157, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02797-8.

Buccella:2021:EPP

[286] Alessandra Buccella. Enactivism and the “problem” of perceptual presence. Synthese, 198(S1):159–173, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02704-1.

Kiverstein:2021:SSI

[287] Julian Kiverstein and Erik Rietveld. Scaling-up skilled intentionality to linguistic thought. Synthese, 198(S1):175–194, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02540-3.

vanEs:2021:EVC

[288] Thomas van Es. The embedded view, its critics, and a radically non-representational solution. Synthese, 198(S1):195–211, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02385-5.

Espejo-Serna:2021:ARE

[289] Juan Camilo Espejo-Serna. Against Radical Enactivism’s narrowmindedness about phenomenality. Synthese, 198(S1):213–228, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02261-2.

Epstein:2021:REC

[290] Eric Gordon Epstein. Radical embodied cognitive science and problems of intentionality. Synthese, 198(S1):229–266, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02330-6.

Zahnoun:2021:RHC

[291] Farid Zahnoun. On representation hungry cognition (and why we should stop feeding it). Synthese, 198(S1):267–284, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02277-8.

deCarvalho:2021:EAD

[292] Eros Moreira de Carvalho. An ecological approach to disjunctivism. Synthese, 198(S1):285–306, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02253-2.

Michaelian:2021:MCR

[293] Kourken Michaelian and André Sant’Anna. Memory without content? Radical enactivism and (post)causal theories of memory. Synthese, 198 (S1):307–335, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02119-7.

Heras-Escribano:2021:PEE

[294] Manuel Heras-Escribano. Pragmatism, enactivism, and ecological psychology: towards a unified approach to post-cognitivism. Synthese, 198 (S1):337–363, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02111-1.

Krueger:2021:EOM

[295] Joel Krueger. Enactivism, other minds, and mental disorders. Synthese, 198(S1):365–389, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02133-9.

Loughlin:2021:WCE

[296] Victor Loughlin. Wittgenstein’s challenge to enactivism. Synthese, 198 (S1):391–404, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02244-3.

Moyal-Sharrock:2021:DWG

[297] Danièle Moyal-Sharrock. From deed to word: gapless and kink-free enactivism. Synthese, 198(S1):405–425, January 2021. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02218-5.

Hoffding:2021:IES

[298] Simon Høffding and Glenda Satne. Interactive expertise in solo and joint musical performance. Synthese, 198(S1):427–445, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02339-x.

Mojica:2021:SES

[299] Laura Mojica and Tom Froese. On the spatiotemporal extensiveness of sense-making: ultrafast cognition and the historicity of normativity. Synthese, 198(S1):447–460, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02240-7.

Fanaya:2021:AEA

[300] Patrícia Fonseca Fanaya. Autopoietic enactivism: action and representation re-examined under Peirce’s light. Synthese, 198(S1):461–483, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02457-6.

Dreon:2021:FCD

[301] Roberta Dreon. Framing cognition: Dewey’s potential contributions to some enactivist issues. Synthese, 198(S1):485–506, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02212-x.

Satne:2021:UDT

[302] Glenda Satne. Understanding others by doing things together: an enactive account. Synthese, 198(S1):507–528, January 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02692-2.

Zahidi:2021:RNC

[303] Karim Zahidi. Radicalizing numerical cognition. Synthese, 198(S1): 529–545, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02956-x.

Steiner:2021:RVC

[304] Pierre Steiner. Radical views on cognition and the dynamics of scientific change. Synthese, 198(S1):547–569, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02268-9.

Rolla:2021:RRS

[305] Giovanni Rolla. Reconceiving rationality: situating rationality into radically enactive cognition. Synthese, 198(S1):571–590, January 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02362-
y.

Wyatt:2021:EIT

[306] Jeremy Wyatt. Editorial introduction to ‘truth: concept meets property’. Synthese, 198(S2):591–603, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02711-2.

Asay:2021:TCU

[307] Jamin Asay. Truth: a concept unlike any other. Synthese, 198(S2): 605–630, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-017-1661-z.

Eklund:2021:WDA

[308] Matti Eklund. What is deflationism about truth? Synthese, 198(S2): 631–645, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-017-1557-y.

Scharp:2021:CET

[309] Kevin Scharp. Conceptual engineering for truth: aletheic properties and new aletheic concepts. Synthese, 198(S2):647–688, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02491-4.

Moltmann:2021:TPT

[310] Friederike Moltmann. Truth predicates, truth bearers, and their variants. Synthese, 198(S2):689–716, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1814-8.

Barnard:2021:OTC

[311] Robert Barnard and Joseph Ulatowski. The objectivity of truth, a core truism? Synthese, 198(S2):717–733, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-017-1605-7.

Edwards:2021:TRP

[312] Douglas Edwards. Truth as a relational property. Synthese, 198(S2): 735–757, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-016-1277-8.

Cumpa:2021:EI

[313] Javier Cumpa. Editor’s introduction. Synthese, 198(S3):759–761, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02712-1.

Brogaard:2021:CII

[314] Berit Brogaard, Bartek Chomanski, and Dimitria Electra Gatzia. Consciousness and information integration. Synthese, 198(S3):763–792, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-020-02613-3.

Buonomo:2021:STM

[315] Valerio Buonomo. The scientific turn in metaphysics: a factualist approach. Synthese, 198(S3):793–807, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-017-1445-5.

Coates:2021:PNR

[316] Paul Coates. Perception naturalised: relocation and the sensible qualities. Synthese, 198(S3):809–829, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-017-1556-z.

Moreno:2021:LMC

[317] Luis Fernández Moreno. Language and metaphysics: the case of theoretical identities. Synthese, 198(S3):831–848, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-017-1498-5.

Heil:2021:TF

[318] John Heil. Truthmaking and fundamentality. Synthese, 198(S3):849– 860, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-016-1292-9.

Sinnott-Armstrong:2021:CMC

[319] Walter Sinnott-Armstrong. Contrastive mental causation. Synthese, 198(S3):861–883, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-019-02506-0.

Koons:2021:OES

[320] Robert C. Koons. The ontological and epistemological superiority of hylomorphism. Synthese, 198(S3):885–903, February 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-016-1295-6.

Salas:2021:SII

[321] Javier González de Prado Salas, Mauricio Suárez, and Jesús Zamora-Bonilla. Special issue: Inferentialism in philosophy of science and in epistemology — introduction. Synthese, 198(S4):905–907, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-02009-4.

Salas:2021:IDC

[322] Javier González de Prado Salas, Xavier de Donato Rodríguez, and Jesús Zamora Bonilla. Inferentialism, degrees of commitment, and ampliative reasoning. Synthese, 198(S4):909–927, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-017-1579-5.

Khalifa:2021:IEA

[323] Kareem Khalifa, Jared Millson, and Mark Risjord. Inference, explanation, and asymmetry. Synthese, 198(S4):929–953, February 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-1791-y.

Tebben:2021:RIS

[324] Nicholas Tebben. Realism, inferential semantics, and the truth norm. Synthese, 198(S4):955–973, February 2021. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-017-1516-7.

Zalabardo:2021:IKB

[325] José L. Zalabardo. Inferentialism and knowledge: Brandom’s arguments against reliabilism. Synthese, 198(S4):975–993, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-017-1506-9.

Marfori:2021:IMP

[326] Marianna Antonutti Marfori and Paula Quinon. Intensionality in mathematics: problems and prospects. Synthese, 198(S5):995–999, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03060-4.

Sarnecka:2021:LRE

[327] Barbara W. Sarnecka. Learning to represent exact numbers. Synthese, 198(S5):1001–1018, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-015-0854-6.

Heylen:2021:ESN

[328] Jan Heylen. The epistemic significance of numerals. Synthese, 198(S5): 1019–1045, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-014-0542-y.

Quinon:2021:CCT

[329] Paula Quinon. Can Church’s thesis be viewed as a Carnapian explication? Synthese, 198(S5):1047–1074, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02286-7.

Tucker:2021:PLT

[330] Dustin Tucker. Paradoxes and the limits of theorizing about propositional attitudes. Synthese, 198(S5):1075–1094, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-018-01902-2.

Sagi:2021:EL

[331] Gil Sagi. Extensionality and logicality. Synthese, 198(S5):1095–1119, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-017-1447-3.

Negri:2021:ISA

[332] Sara Negri. The intensional side of algebraic-topological representation theorems. Synthese, 198(S5):1121–1143, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-017-1331-1.

Tranchini:2021:PTH

[333] Luca Tranchini. Proof-theoretic harmony: towards an intensional account. Synthese, 198(S5):1145–1176, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-016-1200-3.

Angere:2021:IIU

[334] Staffan Angere. Identity and intensionality in univalent foundations and philosophy. Synthese, 198(S5):1177–1217, March 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-016-1301-z.

OConaill:2021:NFG

[335] Donnchadh Ó Conaill and Tuomas Tahko. New frontiers in ground, essence, and modality: introduction. Synthese, 198(S6):1219–1230, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-02067-8.

Rydehn:2021:GOD

[336] Henrik Rydéhn. Grounding and ontological dependence. Synthese, 198(S6):1231–1256, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1818-4.

Wallner:2021:GGE

[337] Michael Wallner. The ground of ground, essence, and explanation. Synthese, 198(S6):1257–1277, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1856-y.

Correia:2021:LRF

[338] Fabrice Correia. The logic of relative fundamentality. Synthese, 198(S6): 1279–1301, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1709-8.

Lubrano:2021:EGS

[339] Michele Lubrano. The emergence of ground: some limitative results. Synthese, 198(S6):1303–1315, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1739-2.

Chilovi:2021:GES

[340] Samuele Chilovi. Grounding entails supervenience. Synthese, 198(S6): 1317–1334, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1723-x.

Schnieder:2021:GC

[341] Benjamin Schnieder. On ground and consequence. Synthese, 198(S6): 1335–1363, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-02012-9.

Vetter:2021:PP

[342] Barbara Vetter. A plenitude of powers. Synthese, 198(S6):1365–1385, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1758-z.

Mallozzi:2021:TNM

[343] Antonella Mallozzi. Two notions of metaphysical modality. Synthese, 198(S6):1387–1408, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-1702-2.

Glazier:2021:DBE

[344] Martin Glazier. The difference between epistemic and metaphysical necessity. Synthese, 198(S6):1409–1424, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-017-1626-2.

Mackie:2021:PM

[345] Penelope Mackie. Persistence and modality. Synthese, 198(S6):1425– 1438, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-018-1776-x.

Jago:2021:EBT

[346] Mark Jago. Essential bundle theory and modality. Synthese, 198(S6): 1439–1454, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1819-3.

Wildman:2021:ARM

[347] Nathan Wildman. Against the reduction of modality to essence. Synthese, 198(S6):1455–1471, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-017-1667-6.

Borges:2021:ISI

[348] Rodrigo Borges. Introduction to the special issue ‘knowledge and justification: new perspectives’. Synthese, 198(S7):1473–1480, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02601-
7.

Elgin:2021:MPD

[349] Samuel Z. Elgin. Merely partial definition and the analysis of knowledge. Synthese, 198(S7):1481–1505, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1846-0.

Nado:2021:CET

[350] Jennifer Nado. Conceptual engineering, truth, and efficacy. Synthese, 198 (S7):1507–1527, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02096-x.

Toribio:2021:AIB

[351] Josefa Toribio. Accessibility, implicit bias, and epistemic justification. Synthese, 198(S7):1529–1547, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1795-7.

Jenkins:2021:LSC

[352] David Jenkins. Luminosity in the stream of consciousness. Synthese, 198 (S7):1549–1562, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1801-0.

Koziolek:2021:IWK

[353] Nicholas Koziolek. Inferring as a way of knowing. Synthese, 198(S7): 1563–1582, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-017-1632-4.

Kelp:2021:IKU

[354] Christoph Kelp. Inquiry, knowledge and understanding. Synthese, 198 (S7):1583–1593, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1803-y.

Pavese:2021:KHA

[355] Carlotta Pavese. Know-how, action, and luck. Synthese, 198(S7): 1595–1617, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1823-7.

Wallbridge:2021:SSI

[356] Kevin Wallbridge. Subject-specific intellectualism: re-examining know how and ability. Synthese, 198(S7):1619–1638, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-018-01944-6.

Kvart:2021:STP

[357] Igal Kvart. The steering thrust phenomenon in action-directed-pragmatics. Synthese, 198(S7):1639–1671, ???? 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02442-z.

Benton:2021:KHF

[358] Matthew A. Benton. Knowledge, hope, and fallibilism. Synthese, 198 (S7):1673–1689, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1794-8.

Wright:2021:EHV

[359] Sarah Wright. Epistemic harm and virtues of self-evaluation. Synthese, 198(S7):1691–1709, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-01993-x.

Pritchard:2021:IHE

[360] Duncan Pritchard. Intellectual humility and the epistemology of disagreement. Synthese, 198(S7):1711–1723, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02024-5.

Tolly:2021:SDP

[361] Jeffrey Tolly. Swampman: a dilemma for proper functionalism. Synthese, 198(S7):1725–1750, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1684-0.

Ghijsen:2021:PPF

[362] Harmen Ghijsen. Predictive processing and foundationalism about perception. Synthese, 198(S7):1751–1769, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1715-x.

Comesana:2021:EJD

[363] Juan Comesaña. Empirical justification and defeasibility. Synthese, 198 (S7):1771–1786, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1865-x.

Sorensen:2021:SHT

[364] Roy Sorensen. Stealing Harman’s thought: knowledge saboteurs and dogmatists. Synthese, 198(S7):1787–1799, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01945-5.

Williams:2021:OYT

[365] John N. Williams. Once you think you’re wrong, you must be right: new versions of the preface paradox. Synthese, 198(S7):1801–1825, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-01943-7.

Veber:2021:DKD

[366] Michael Veber. A different kind of dream-based skepticism. Synthese, 198 (S7):1827–1839, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-01910-2.

Mallozzi:2021:SIS

[367] Antonella Mallozzi. Special issue of synthese on new directions in the epistemology of modality: introduction. Synthese, 198(S8):1841–1859, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02358-8.

Clarke-Doane:2021:MAP

[368] Justin Clarke-Doane. Metaphysical and absolute possibility. Synthese, 198(S8):1861–1872, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-019-02093-0.

Priest:2021:MNS

[369] Graham Priest. Metaphysical necessity: a skeptical perspective. Synthese, 198(S8):1873–1885, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1885-6.

Wang:2021:EOM

[370] Jennifer Wang. The epistemological objection to modal primitivism. Synthese, 198(S8):1887–1898, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-01954-4.

Jackson:2021:DAN

[371] Frank Jackson. Deflationism about the necessary a posteriori and Twin Earth. Synthese, 198(S8):1899–1907, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1834-4.

Vaidya:2021:EMP

[372] Anand Jayprakash Vaidya and Michael Wallner. The epistemology of modality and the problem of modal epistemic friction. Synthese, 198 (S8):1909–1935, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1860-2.

Mallozzi:2021:PMM

[373] Antonella Mallozzi. Putting modal metaphysics first. Synthese, 198(S8): 1937–1956, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1828-2.

Kment:2021:EMK

[374] Boris Kment. Essence and modal knowledge. Synthese, 198(S8): 1957–1979, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-01903-1.

Jago:2021:KHT

[375] Mark Jago. Knowing how things might have been. Synthese, 198(S8): 1981–1999, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1869-6.

Hale:2021:EDA

[376] Bob Hale. Essence and definition by abstraction. Synthese, 198(S8): 2001–2017, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1726-7.

Goff:2021:EMR

[377] Philip Goff. Essentialist modal rationalism. Synthese, 198(S8):2019– 2027, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02109-9.

Berto:2021:TRI

[378] Francesco Berto. Taming the runabout imagination ticket. Synthese, 198 (S8):2029–2043, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1751-6.

Biggs:2021:AVC

[379] Stephen Biggs and Jessica Wilson. Abduction versus conceiving in modal epistemology. Synthese, 198(S8):2045–2076, ???? 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02117-9.

Thomasson:2021:HCW

[380] Amie L. Thomasson. How can we come to know metaphysical modal truths? Synthese, 198(S8):2077–2106, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1841-5.

Rabin:2021:MRD

[381] Gabriel Oak Rabin. Modal rationalism and the demonstrative reply to the scrutability argument against physicalism. Synthese, 198(S8): 2107–2134, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02174-0.

Isaac:2021:I

[382] Alistair M. C. Isaac and Dave Ward. Introduction. Synthese, 198(S9): 2135–2151, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02391-7.

Feest:2021:GPF

[383] Uljana Feest. Gestalt psychology, frontloading phenomenology, and psychophysics. Synthese, 198(S9):2153–2173, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02211-y.

Baggs:2021:RET

[384] Edward Baggs and Anthony Chemero. Radical embodiment in two directions. Synthese, 198(S9):2175–2190, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02020-9.

Albertazzi:2021:EPW

[385] Liliana Albertazzi. Experimental phenomenology: What it is and what it is not. Synthese, 198(S9):2191–2212, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02209-6.

Hatfield:2021:GGR

[386] Gary Hatfield. Gibson and Gestalt: (re)presentation, processing, and construction. Synthese, 198(S9):2213–2241, ???? 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02380-w.

Muller:2021:MPR

[387] Robin M. Muller. Merleau–Ponty and the radical sciences of mind. Synthese, 198(S9):2243–2277, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-02015-6.

Kiverstein:2021:FLA

[388] Julian Kiverstein, Ludger van Dijk, and Erik Rietveld. The field and landscape of affordances: Koffka’s two environments revisited. Synthese, 198(S9):2279–2296, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02123-x.

Sheredos:2021:MPI

[389] Benjamin Sheredos. Merleau-Ponty’s implicit critique of the new mechanists. Synthese, 198(S9):2297–2321, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02006-7.

Kee:2021:PNA

[390] Hayden Kee. Phenomenology and naturalism in autopoietic and radical enactivism: exploring sense-making and continuity from the top down. Synthese, 198(S9):2323–2343, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1851-3.

Crippen:2021:AAS

[391] Matthew Crippen. Aesthetics and action: situations, emotional perception and the Kuleshov effect. Synthese, 198(S9):2345–2363, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02110-2.

Jackson:2021:GSM

[392] Sarah Pawlett Jackson. Gestalt structures in multi-person intersubjectivity. Synthese, 198(S9):2365–2382, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02001-y.

Kaplan:2021:RWP

[393] Jonathan Michael Kaplan and Sean A. Valles. Reflecting on what philosophy of epidemiology is and does, as the field comes into its own: Introduction to the special issue on philosophy of epidemiology. Synthese, 198 (S10):2383–2392, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02252-3.

Varghese:2021:IPN

[394] Joby Varghese. Influence and prioritization of non-epistemic values in clinical trial designs: a study of Ebola ça Suffit trial. Synthese, 198(S10): 2393–2409, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-01912-0.

Taylor:2021:CEL

[395] Peter J. Taylor. Critical epidemiological literacy: understanding ideas better when placed in relation to alternatives. Synthese, 198(S10): 2411–2438, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-01960-6.

Meissner:2021:HDC

[396] Shelbi Nahwilet Meissner. How does the consideration of indigenous identities in the US complicate conversations about tracking folk racial categories in epidemiologic research? Synthese, 198(S10):2439–2462, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-02019-2.

Grune-Yanoff:2021:CRM

[397] Till Grüne-Yanoff. Choosing the right model for policy decisionmaking: the case of smallpox epidemiology. Synthese, 198(S10):2463– 2484, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-018-1827-3.

Leonelli:2021:WHE

[398] Sabina Leonelli and Niccolò Tempini. Where health and environment meet: the use of invariant parameters in big data analysis. Synthese, 198 (S10):2485–2504, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1844-2.

Williamson:2021:ETZ

[399] Jon Williamson. Establishing the teratogenicity of Zika and evaluating causal criteria. Synthese, 198(S10):2505–2518, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:/

/link.springer.com/article/10.1007/s11229-018-1866-9. See cor-

rection [400].

Williamson:2021:CET

[400] Jon Williamson. Correction to: Establishing the teratogenicity of Zika and evaluating causal criteria. Synthese, 198(S10):2519, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1889-2. See [399].

McClimans:2021:FPE

[401] Leah M. McClimans. First person epidemiological measures: vehicles for patient centered care. Synthese, 198(S10):2521–2537, ???? 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02094-
z.

Lee:2021:EES

[402] Keekok Lee. Epidemiology is ecosystem science. Synthese, 198(S10): 2539–2567, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02129-5.

Tulodziecki:2021:HTA

[403] Dana Tulodziecki. How (not) to think about theory-change in epidemiology. Synthese, 198(S10):2569–2588, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02135-7.

Valles:2021:PSR

[404] Sean A. Valles. A pluralistic and socially responsible philosophy of epidemiology field should actively engage with social determinants of health and health disparities. Synthese, 198(S10):2589–2611, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02161-5.

Broadbent:2021:CWP

[405] Alex Broadbent. The C-word, the P-word, and realism in epidemiology. Synthese, 198(S10):2613–2628, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02169-x.

Jukola:2021:CIA

[406] Saana Jukola. Commercial interests, agenda setting, and the epistemic trustworthiness of nutrition science. Synthese, 198(S10):2629– 2646, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02228-3.

Marmodoro:2021:ISI

[407] Anna Marmodoro and Michele Paolini Paoletti. Introduction to the special issue on form, structure and hylomorphism. Synthese, 198(S11): 2647–2656, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02441-0.

Heil:2021:HWL

[408] John Heil. Hylomorphism: what’s not to like? Synthese, 198(S11): 2657–2670, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1792-x.

Paoletti:2021:SR

[409] Michele Paolini Paoletti. Structures as relations. Synthese, 198(S11): 2671–2690, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-01918-8.

Skrzypek:2021:PAH

[410] Jeremy W. Skrzypek. From potency to act: hyloenergeism. Synthese, 198 (S11):2691–2716, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02089-w.

Dumsday:2021:CRS

[411] Travis Dumsday. Can a relational substance ontology be hylomorphic? Synthese, 198(S11):2717–2734, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02173-1.

Sattig:2021:PSG

[412] Thomas Sattig. Part, slot, ground: foundations for neo-Aristotelian mereology. Synthese, 198(S11):2735–2749, ???? 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02141-9.

Koons:2021:TSN

[413] Robert C. Koons. Thermal substances: a neo-Aristotelian ontology of the quantum world. Synthese, 198(S11):2751–2772, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02318-2.

Morganti:2021:QEH

[414] Matteo Morganti. Quantum entanglement: a hylomorphic account. Synthese, 198(S11):2773–2793, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02280-z.

Brungardt:2021:WEF

[415] John G. Brungardt. World enough and form: why cosmology needs hylomorphism. Synthese, 198(S11):2795–2827, ???? 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02112-0.

Owen:2021:CCP

[416] Matthew Owen. Circumnavigating the causal pairing problem with hylomorphism and the integrated information theory of consciousness. Synthese, 198(S11):2829–2851, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02403-6.

Moreno:2021:RNK

[417] Luis Fernández Moreno. Reflection on natural kinds. Introduction to the special issue on natural kinds: language, science, and metaphysics. Synthese, 198(S12):2853–2862, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02325-3.

Boyd:2021:RNK

[418] Richard Boyd. Rethinking natural kinds, reference and truth: towards more correspondence with reality, not less. Synthese, 198(S12): 2863–2903, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02138-4. See correction [419].

Boyd:2021:CRN

[419] Richard Boyd. Correction to: Rethinking natural kinds, reference and truth: towards more correspondence with reality, not less. Synthese, 198(S12):2905, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02322-6. See [418].

Lemeire:2021:NPE

[420] Olivier Lemeire. No purely epistemic theory can account for the naturalness of kinds. Synthese, 198(S12):2907–2925, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-018-1806-8.

Allen:2021:KBB

[421] Sophie R. Allen. Kinds behaving badly: intentional action and interactive kinds. Synthese, 198(S12):2927–2956, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1870-0.

Schwartz:2021:ARN

[422] Stephen P. Schwartz. Against rigidity for natural kind terms. Synthese, 198(S12):2957–2971, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1857-x.

Marti:2021:TFA

[423] Genoveva Martí and Lorena Ramírez-Ludeña. Tolerance, flexibility and the application of kind terms. Synthese, 198(S12):2973–2986, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-01898-9.

Parent:2021:DAK

[424] T. Parent. A dilemma about kinds and kind terms. Synthese, 198(S12): 2987–3006, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02278-7.

Casetta:2021:SST

[425] Elena Casetta and Davide Vecchi. Species are, at the same time, kinds and individuals: a causal argument based on an empirical approach to species identity. Synthese, 198(S12):3007–3025, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02199-5.

Hazelwood:2021:SCS

[426] Caleb C. Hazelwood. The species category as a scientific kind. Synthese, 198(S12):3027–3040, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-02025-4.

Godman:2021:SRH

[427] Marion Godman. Scientific realism with historical essences: the case of species. Synthese, 198(S12):3041–3057, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02034-3.

vanRooij:2021:NKD

[428] Robert van Rooij and Katrin Schulz. Natural kinds and dispositions: a causal analysis. Synthese, 198(S12):3059–3084, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02184-y.

Freund:2021:PSC

[429] Max A. Freund. Predication and sortal concepts. Synthese, 198(S12): 3085–3106, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-02030-7.

Gava:2021:IKP

[430] Gabriele Gava. Introduction: Kant’s philosophical method and contemporary perspectives in epistemology, metaphysics, philosophy of mind, philosophy of science and ethics. Synthese, 198(S13):3107–3111, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-03022-2.

Schafer:2021:KVE

[431] Karl Schafer. A Kantian virtue epistemology: rational capacities and transcendental arguments. Synthese, 198(S13):3113–3136, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-02005-8.

Land:2021:EAS

[432] Thomas Land. Epistemic agency and the self-knowledge of reason: on the contemporary relevance of Kant’s method of faculty analysis. Synthese, 198(S13):3137–3154, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-02027-2.

Audi:2021:MRK

[433] Robert Audi. Methodological reflections on Kant’s ethical theory. Synthese, 198(S13):3155–3170, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-01977-x.

Rosefeldt:2021:KER

[434] Tobias Rosefeldt. Kant on the epistemic role of the imagination. Synthese, 198(S13):3171–3192, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02100-4.

Leech:2021:FMJ

[435] Jessica Leech. The function of modal judgment and the Kantian gap. Synthese, 198(S13):3193–3212, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1810-z.

Watkins:2021:WRP

[436] Eric Watkins. What real progress has metaphysics made since the time of Kant? Kant and the metaphysics of grounding. Synthese, 198(S13): 3213–3229, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02180-2.

Heidemann:2021:KFR

[437] Dietmar Heidemann. Kant and the forms of realism. Synthese, 198(S13): 3231–3252, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02502-4. See correction [942].

Stephenson:2021:HSK

[438] Andrew Stephenson. How to solve the knowability paradox with transcendental epistemology. Synthese, 198(S13):3253–3278, ???? 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-1822-8.

Massimi:2021:PVK

[439] Michela Massimi. Points of view: Kant on perspectival knowledge. Synthese, 198(S13):3279–3296, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1876-7.

Herfeld:2021:IFP

[440] Catherine Herfeld and Milena Ivanova. Introduction: first principles in science-their status and justification. Synthese, 198(S14):3297–3308, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02801-1.

Hoover:2021:FPF

[441] Kevin D. Hoover. First principles, fallibilism, and economics. Synthese, 198(S14):3309–3327, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-02021-8.

Herfeld:2021:URP

[442] Catherine Herfeld. Understanding the rationality principle in economics as a functional a priori principle. Synthese, 198(S14):3329–3358, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02730-z.

Linsbichler:2021:AEE

[443] Alexander Linsbichler. Austrian economics without extreme apriorism: construing the fundamental axiom of praxeology as analytic. Synthese, 198(S14):3359–3390, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-019-02150-8.

Hendry:2021:EFP

[444] Robin Findlay Hendry. Elements and (first) principles in chemistry. Synthese, 198(S14):3391–3411, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02312-8.

Okasha:2021:SEE

[445] Samir Okasha. The strategy of endogenization in evolutionary biology. Synthese, 198(S14):3413–3435, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1832-6.

Luchetti:2021:CES

[446] Michele Luchetti. Constitutive elements in science beyond physics: the case of the Hardy-Weinberg principle. Synthese, 198(S14):3437–3461, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-1833-5.

Colombo:2021:FPL

[447] Matteo Colombo and Cory Wright. First principles in the life sciences: the free-energy principle, organicism, and mechanism. Synthese, 198 (S14):3463–3488, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-01932-w.

Crowther:2021:DCR

[448] Karen Crowther. Defining a crisis: the roles of principles in the search for a theory of quantum gravity. Synthese, 198(S14):3489–3516, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-01970-4.

Carter:2021:ISI

[449] J. Adam Carter, Guillaume Dechauffour, and Grégoire Lefftz. Introduction to special issue: Skepticism and epistemic angst. Synthese, 198 (S15):3517–3519, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03241-1.

Coliva:2021:SBP

[450] Annalisa Coliva. Strange bedfellows: on Pritchard’s disjunctivist hinge epistemology. Synthese, 198(S15):3521–3532, ???? 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-018-02046-z.

Zhang:2021:CDH

[451] Xiaoxing Zhang. Closure, deduction and hinge commitments. Synthese, 198(S15):3533–3551, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1679-x. See correction [2444].

Simion:2021:ENC

[452] Mona Simion, Johanna Schnurr, and Emma Gordon. Epistemic norms, closure, and No-Belief hinge epistemology. Synthese, 198(S15):3553– 3564, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02165-1.

Jope:2021:CCR

[453] Matt Jope. Closure, credence and rationality: a problem for non-belief hinge epistemology. Synthese, 198(S15):3565–3575, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02153-5.

Neta:2021:EAH

[454] Ram Neta. An evidentialist account of hinges. Synthese, 198(S15): 3577–3591, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-02061-0.

Greco:2021:HEP

[455] John Greco. Hinge epistemology and the prospects for a unified theory of knowledge. Synthese, 198(S15):3593–3607, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02225-6.

Vorms:2021:SRD

[456] Marion Vorms and Ulrike Hahn. In the space of reasonable doubt. Synthese, 198(S15):3609–3633, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02488-z.

Pritchard:2021:SEA

[457] Duncan Pritchard. Scepticism and epistemic Angst, redux. Synthese, 198 (S15):3635–3664, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02504-2.

Eva:2021:RP

[458] Benjamin Eva and Stephan Hartmann. Reasoning in physics. Synthese, 198(S16):3665–3669, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02756-3.

Cabrera:2021:STN

[459] Frank Cabrera. String theory, non-empirical theory assessment, and the context of pursuit. Synthese, 198(S16):3671–3699, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-018-01987-9.

Crowther:2021:WWC

[460] Karen Crowther, Niels S. Linnemann, and Christian Wüthrich. What we cannot learn from analogue experiments. Synthese, 198(S16):3701– 3726, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02190-0.

Hossenfelder:2021:SEF

[461] Sabine Hossenfelder. Screams for explanation: finetuning and naturalness in the foundations of physics. Synthese, 198(S16):3727–3745, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02377-5.

Koberinski:2021:MDR

[462] Adam Koberinski. Mathematical developments in the rise of Yang–Mills gauge theories. Synthese, 198(S16):3747–3777, ???? 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-018-02070-z.

Chall:2021:BSM

[463] Cristin Chall, Martin King, and Michael Stöltzner. From a boson to the standard model Higgs: a case study in confirmation and model dynamics. Synthese, 198(S16):3779–3811, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02216-7.

Curiel:2021:FCN

[464] Erik Curiel. Framework confirmation by Newtonian abduction. Synthese, 198(S16):3813–3851, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02400-9.

Norton:2021:EIW

[465] John D. Norton. Eternal inflation: when probabilities fail. Synthese, 198 (S16):3853–3875, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1734-7.

Fletcher:2021:CRW

[466] Samuel C. Fletcher. Counterfactual reasoning within physical theories. Synthese, 198(S16):3877–3898, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02085-0.

Ferretti:2021:BVA

[467] Gabriele Ferretti and Silvano Zipoli Caiani. Between vision and action: introduction to the special issue. Synthese, 198(S17):3899–3911, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02518-w.

Briscoe:2021:BAN

[468] Robert Eamon Briscoe. Bodily awareness and novel multisensory features. Synthese, 198(S17):3913–3941, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02156-2.

Brogaard:2021:IBV

[469] Berit Brogaard. Implicit biases in visually guided action. Synthese, 198 (S17):3943–3967, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02588-1.

Butterfill:2021:GTD

[470] Stephen A. Butterfill. Goals and targets: a developmental puzzle about sensitivity to others’ actions. Synthese, 198(S17):3969–3990, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02214-9.

Cavedon-Taylor:2021:SEV

[471] Dan Cavedon-Taylor. Sensorimotor expectations and the visual field. Synthese, 198(S17):3991–4006, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01946-4.

Cosentino:2021:AA

[472] Erica Cosentino. Artifacts and affordances. Synthese, 198(S17):4007– 4026, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02297-4.

deVignemont:2021:PPA

[473] Frédérique de Vignemont. Peripersonal perception in action. Synthese, 198(S17):4027–4044, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-01962-4.

Grunbaum:2021:TVS

[474] Thor Grünbaum. The two visual systems hypothesis and contrastive underdetermination. Synthese, 198(S17):4045–4068, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-018-01984-y.

Nanay:2021:PAP

[475] Bence Nanay. Perception is not all-purpose. Synthese, 198(S17): 4069–4080, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-01937-5.

Raleigh:2021:VAA

[476] Thomas Raleigh. Visual acquaintance, action and the explanatory gap. Synthese, 198(S17):4081–4106, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01907-x.

Rosen:2021:DSW

[477] Melanie Rosen. Dreaming of a stable world: vision and action in sleep. Synthese, 198(S17):4107–4142, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02149-1.

Shepherd:2021:IAG

[478] Joshua Shepherd. Intelligent action guidance and the use of mixed representational formats. Synthese, 198(S17):4143–4162, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-018-1892-7.

Toribio:2021:VRC

[479] Josefa Toribio. Are visuomotor representations cognitively penetrable? Biasing action-guiding vision. Synthese, 198(S17):4163–4181, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-1854-0.

Caiani:2021:IBA

[480] Silvano Zipoli Caiani. Intensional biases in affordance perception: an explanatory issue for radical enactivism. Synthese, 198(S17):4183–4203, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-02049-w.

Carrara:2021:MI

[481] Massimiliano Carrara and Giorgio Lando. Mereology and identity. Synthese, 198(S18):4205–4227, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02592-5.

Cotnoir:2021:WSA

[482] A. J. Cotnoir. Is weak supplementation analytic? Synthese, 198(S18): 4229–4245, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-02066-9.

Guillon:2021:CP

[483] Jean-Baptiste Guillon. Coincidence as parthood. Synthese, 198(S18): 4247–4276, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02105-z.

Maffezioli:2021:IM

[484] Paolo Maffezioli and Achille C. Varzi. Intuitionistic mereology. Synthese, 198(S18):4277–4302, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-02035-2.

Azzano:2021:SPM

[485] Lorenzo Azzano. Structural properties, mereology, and modal magic. Synthese, 198(S18):4303–4329, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02040-5.

Wallace:2021:PP

[486] Meg Wallace. The polysemy of ‘part’. Synthese, 198(S18):4331–4354, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02088-x.

Smid:2021:EFP

[487] Jeroen Smid. Extensionality for fusions and pluralities. Synthese, 198 (S18):4355–4374, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-02003-w.

Spencer:2021:EDS

[488] Joshua Spencer. On the explanatory demands of the Special Composition Question. Synthese, 198(S18):4375–4388, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02147-3.

Bricker:2021:CIL

[489] Phillip Bricker. Composition as identity, Leibniz’s Law, and slicesensitive emergent properties. Synthese, 198(S18):4389–4409, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02106-
y.

Carrara:2021:CCI

[490] Massimiliano Carrara and Giorgio Lando. Contingent composition as identity. Synthese, 198(S18):4411–4440, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01934-8.

Lechthaler:2021:NUG

[491] Manuel Lechthaler. No universalism without gunk? Composition as identity and the universality of identity. Synthese, 198(S18):4441–4452, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02122-y.

Lipman:2021:RAM

[492] Martin A. Lipman. On relativist approaches to many-one identity. Synthese, 198(S18):4453–4465, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-01916-w.

Yi:2021:CI

[493] Byeong uk Yi. Is composition identity? Synthese, 198(S18):4467– 4501, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-018-02000-z.

Calosi:2021:PI

[494] Claudio Calosi. Is parthood identity? Synthese, 198(S18):4503–4517, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-02057-w.

Loss:2021:ACI

[495] Roberto Loss. On atomic composition as identity. Synthese, 198(S18): 4519–4542, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02295-6. See correction [496].

Loss:2021:CAC

[496] Roberto Loss. Correction to: On atomic composition as identity. Synthese, 198(S18):4543, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02617-z. See [495].

Botti:2021:CAM

[497] Martina Botti. Composition as analysis: the meta-ontological origins (and future) of composition as identity. Synthese, 198(S18):4545–4570, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02396-2.

Payton:2021:HIW

[498] Jonathan D. Payton. How to identify wholes with their parts. Synthese, 198(S18):4571–4593, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02191-z.

Bohn:2021:CIP

[499] Einar Duenger Bohn. Composition as identity: pushing forward. Synthese, 198(S18):4595–4607, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02193-x.

Woods:2021:MO

[500] Evan T. Woods. Many, but one. Synthese, 198(S18):4609–4626, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02162-4.

Carrier:2021:RRI

[501] Martin Carrier and Gürol Irzik. Responsible research and innovation: coming to grips with an ambitious concept. Synthese, 198(S19): 4627–4633, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02319-1.

Timmermans:2021:CHR

[502] Job Timmermans and Vincent Blok. A critical hermeneutic reflection on the paradigm-level assumptions underlying responsible innovation. Synthese, 198(S19):4635–4666, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-1839-z.

vonSchomberg:2021:TAI

[503] Lucien von Schomberg and Vincent Blok. The turbulent age of innovation. Synthese, 198(S19):4667–4683, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01950-8.

Limson:2021:PRR

[504] Janice Limson. Putting responsible research and innovation into practice: a case study for biotechnology research, exploring impacts and RRI learning outcomes of public engagement for science students. Synthese, 198 (S19):4685–4710, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-02063-y.

Bueter:2021:PET

[505] Anke Bueter. Public epistemic trustworthiness and the integration of patients in psychiatric classification. Synthese, 198(S19):4711–4729, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-01913-z.

Irzik:2021:WOS

[506] Gürol Irzik and Faik Kurtulmus. Well-ordered science and public trust in science. Synthese, 198(S19):4731–4748, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02022-7.

Carrier:2021:HCS

[507] Martin Carrier. How to conceive of science for the benefit of society: prospects of responsible research and innovation. Synthese, 198(S19): 4749–4768, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02254-1.

vandePoel:2021:VRT

[508] Ibo van de Poel and Martin Sand. Varieties of responsibility: two problems of responsible innovation. Synthese, 198(S19):4769–4787, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-01951-7.

Caret:2021:PPL

[509] Colin R. Caret and Teresa Kouri Kissel. Pluralistic perspectives on logic: an introduction. Synthese, 198(S20):4789–4800, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02525-x.

Kissel:2021:MNL

[510] Teresa Kouri Kissel. Metalinguistic negotiation and logical pluralism. Synthese, 198(S20):4801–4812, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02264-z.

Wyatt:2021:ALG

[511] Nicole Wyatt and Gillman Payette. Against logical generalism. Synthese, 198(S20):4813–4830, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-02073-w.

Golan:2021:TNM

[512] Rea Golan. Is there a neutral metalanguage? Synthese, 198(S20): 4831–4858, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-02078-5.

Blake-Turner:2021:LPN

[513] Christopher Blake-Turner and Gillian Russell. Logical pluralism without the normativity. Synthese, 198(S20):4859–4877, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-018-01939-3.

Ferrari:2021:PTP

[514] Filippo Ferrari and Eugenio Orlandelli. Proof-theoretic pluralism. Synthese, 198(S20):4879–4903, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02217-6.

Finn:2021:LLP

[515] Suki Finn. Limiting logical pluralism. Synthese, 198(S20):4905–4923, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02134-8.

Shapiro:2021:DLP

[516] Stewart Shapiro and Michael Lynch. Does logical pluralism imply, or suggest, truth pluralism, or vice versa? Synthese, 198(S20):4925–4936, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02175-z.

Priest:2021:NMP

[517] Graham Priest. A note on mathematical pluralism and logical pluralism. Synthese, 198(S20):4937–4946, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02292-9.

Caret:2021:WLP

[518] Colin R. Caret. Why logical pluralism? Synthese, 198(S20):4947– 4968, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02132-w.

French:2021:DRL

[519] Rohan French. A dialogical route to logical pluralism. Synthese, 198 (S20):4969–4989, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02404-5.

Barrio:2021:SLP

[520] Eduardo Alejandro Barrio, Federico Pailos, and Damian Szmuc. Substructural logics, pluralism and collapse. Synthese, 198(S20):4991–5007, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-01963-3.

SanPedro:2021:SPR

[521] Iñaki San Pedro. From structure preserving representation to making worlds with symbols. Synthese, 198(S21):5009–5013, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03290-6.

Teller:2021:MWS

[522] Paul Teller. Making worlds with symbols. Synthese, 198(S21):5015– 5036, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-018-1811-y.

Mormann:2021:SPR

[523] Thomas Mormann. Structure-preserving representations, constitution, and the relative a priori. Synthese, 198(S21):5037–5060, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-1759-y.

Odenbaugh:2021:MMM

[524] Jay Odenbaugh. Models, models, models: a deflationary view. Synthese, 198(S21):5061–5076, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-017-1665-8.

Knuuttila:2021:IEE

[525] Tarja Knuuttila. Imagination extended and embedded: artifactual versus fictional accounts of models. Synthese, 198(S21):5077–5097, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/

s11229-017-1545-2. See correction [2283].

Winsberg:2021:WDR

[526] Eric Winsberg. What does robustness teach us in climate science: a reappraisal. Synthese, 198(S21):5099–5122, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01997-7.

Wilson:2021:CMS

[527] Katia Wilson. The case of the missing satellites. Synthese, 198(S21): 5123–5143, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-017-1509-6.

Keyser:2021:EEC

[528] Vadim Keyser. Experimental effects and causal representations. Synthese, 198(S21):5145–5176, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-017-1633-3.

Chandrasekharan:2021:RCH

[529] Sanjay Chandrasekharan and Nancy J. Nersessian. Rethinking correspondence: how the process of constructing models leads to discoveries and transfer in the bioengineering sciences. Synthese, 198(S21): 5177–5206, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-017-1463-3.

Wansing:2021:VEI

[530] Heinrich Wansing and Marco Ruffino. Varieties of entailment: introduction to the special issue. Synthese, 198(S22):5207–5211, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02778-
x.

Drobyshevich:2021:TCR

[531] Sergey Drobyshevich. Tarskian consequence relations bilaterally: some familiar notions. Synthese, 198(S22):5213–5240, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02267-w.

Shramko:2021:NEI

[532] Yaroslav Shramko and Heinrich Wansing. The nature of entailment: an informational approach. Synthese, 198(S22):5241–5261, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/

s11229-019-02474-5. See correction [533].

Shramko:2021:CNE

[533] Yaroslav Shramko and Heinrich Wansing. Correction to: The nature of entailment: an informational approach. Synthese, 198(S22): 5263–5264, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02640-0. See [532].

Barrio:2021:MIL

[534] Eduardo Alejandro Barrio, Federico Pailos, and Damian Szmuc. (Meta)inferential levels of entailment beyond the Tarskian paradigm. Synthese, 198(S22):5265–5289, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02411-6.

French:2021:TTA

[535] Rohan French and David Ripley. Two traditions in abstract valuational model theory. Synthese, 198(S22):5291–5313, ???? 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02447-8.

Chemla:2021:MVC

[536] Emmanuel Chemla and Paul Egré. From many-valued consequence to many-valued connectives. Synthese, 198(S22):5315–5352, ???? 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02344-
0.

Avron:2021:QCS

[537] Arnon Avron. Quasi-canonical systems and their semantics. Synthese, 198(S22):5353–5371, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-018-02045-0.

Marcelino:2021:AND

[538] Sérgio Marcelino and Carlos Caleiro. Axiomatizing non-deterministic many-valued generalized consequence relations. Synthese, 198(S22): 5373–5390, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02142-8.

Pimentel:2021:ENE

[539] Elaine Pimentel, Luiz Carlos Pereira, and Valeria de Paiva. An ecumenical notion of entailment. Synthese, 198(S22):5391–5413, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02226-
5.

Verdee:2021:OIM

[540] Peter Verdée. Obtaining infinitely many degrees of inconsistency by adding a strictly paraconsistent negation to classical logic. Synthese, 198 (S22):5415–5449, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02638-8.

Rodrigues:2021:MEP

[541] Abilio Rodrigues, Juliana Bueno-Soler, and Walter Carnielli. Measuring evidence: a probabilistic approach to an extension of Belnap–Dunn logic. Synthese, 198(S22):5451–5480, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02571-w.

Blasio:2021:WLT

[542] Carolina Blasio, Carlos Caleiro, and João Marcos. What is a logical theory? On theories containing assertions and denials. Synthese, 198 (S22):5481–5504, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02183-z.

Zalabardo:2021:LM

[543] José L. Zalabardo. Logic without metaphysics. Synthese, 198(S22): 5505–5532, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02124-w.

Ramirez-Camara:2021:KNI

[544] Elisángela Ramírez-Cámara and Luis Estrada-González. Knot is not that nasty (but it is hardier than tonk). Synthese, 198(S22):5533–5554, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02498-x.

Kubyshkina:2021:CTF

[545] Ekaterina Kubyshkina. Conservative translations of four-valued logics in modal logic. Synthese, 198(S22):5555–5571, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02139-3.

Rudnicki:2021:PAL

[546] Konrad Rudnicki and Piotr Lukowski. Psychophysiological approach to the Liar paradox: Jean Buridan’s virtual entailment principle put to the test. Synthese, 198(S22):5573–5592, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02107-x.

Ruffino:2021:CPT

[547] Marco Ruffino. Contingent a priori truths and performatives. Synthese, 198(S22):5593–5613, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02762-5.

Bonzio:2021:RI

[548] Stefano Bonzio, Jürgen Landes, and Barbara Osimani. Reliability: an introduction. Synthese, 198(S23):5615–5624, ???? 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02725-w.

Tolly:2021:KEM

[549] Jeffrey Tolly. Knowledge, evidence, and multiple process types. Synthese, 198(S23):5625–5652, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02146-4.

Dunn:2021:RGB

[550] Jeffrey Dunn. Reliable group belief. Synthese, 198(S23):5653–5677, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-02075-8.

deRonde:2021:QPR

[551] C. de Ronde, H. Freytes, and G. Sergioli. Quantum probability: a reliable tool for an agent or a reliable source of reality? Synthese, 198(S23): 5679–5699, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02177-x.

Lalumera:2021:RMI

[552] Elisabetta Lalumera, Stefano Fanti, and Giovanni Boniolo. Reliability of molecular imaging diagnostics. Synthese, 198(S23):5701–5717, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02419-y.

Primiero:2021:TCP

[553] Giuseppe Primiero, Franco Raimondi, and Taolue Chen. A theory of change for prioritised resilient and evolvable software systems. Synthese, 198(S23):5719–5744, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02305-7.

Dutilh:2021:FYF

[554] Gilles Dutilh, Alexandra Sarafoglou, and Eric-Jan Wagenmakers. Flexible yet fair: blinding analyses in experimental psychology. Synthese, 198 (S23):5745–5772, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02456-7.

Merdes:2021:FMS

[555] Christoph Merdes, Momme von Sydow, and Ulrike Hahn. Formal models of source reliability. Synthese, 198(S23):5773–5801, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02595-2.

Romero:2021:SSC

[556] Felipe Romero and Jan Sprenger. Scientific self-correction: the Bayesian way. Synthese, 198(S23):5803–5823, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02697-x.

Holman:2021:EOI

[557] Bennett Holman. An ethical obligation to ignore the unreliable. Synthese, 198(S23):5825–5848, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02483-4.

Portides:2021:TAI

[558] Demetris Portides. Thinking about idealization, abstraction and scientific models: an introduction. Synthese, 198(S24):5849–5853, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03212-6.

Levy:2021:IAR

[559] Arnon Levy. Idealization and abstraction: refining the distinction. Synthese, 198(S24):5855–5872, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1721-z.

Portides:2021:IAS

[560] Demetris Portides. Idealization and abstraction in scientific modeling. Synthese, 198(S24):5873–5895, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01919-7.

French:2021:ICI

[561] Steven French. Identity conditions, idealisations and isomorphisms: a defence of the Semantic Approach. Synthese, 198(S24):5897–5917, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-017-1564-z.

Bokulich:2021:UMC

[562] Alisa Bokulich. Using models to correct data: paleodiversity and the fossil record. Synthese, 198(S24):5919–5940, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-018-1820-x.

Nguyen:2021:MOL

[563] James Nguyen and Roman Frigg. Mathematics is not the only language in the book of nature. Synthese, 198(S24):5941–5962, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-017-1526-5.

Morrison:2021:TEM

[564] Margaret Morrison. Turbulence, emergence and multi-scale modelling. Synthese, 198(S24):5963–5985, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-1825-5.

Bechtel:2021:EFF

[565] William Bechtel. Explaining features of fine-grained phenomena using abstract analyses of phenomena and mechanisms: two examples from chronobiology. Synthese, 198(S24):5987–6009, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-017-1469-x.

Dellsen:2021:DSI

[566] Finnur Dellsén and Maria Baghramian. Disagreement in science: introduction to the special issue. Synthese, 198(S25):6011–6021, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02767-0.

Seidel:2021:KTA

[567] Markus Seidel. Kuhn’s two accounts of rational disagreement in science: an interpretation and critique. Synthese, 198(S25):6023–6051, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02113-z.

Shaw:2021:FMD

[568] Jamie Shaw. Feyerabend and manufactured disagreement: reflections on expertise, consensus, and science policy. Synthese, 198(S25):6053– 6084, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-020-02538-x.

Frank:2021:DDI

[569] David M. Frank. Disagreement or denialism? “Invasive species denialism” and ethical disagreement in science. Synthese, 198(S25): 6085–6113, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02259-w.

Massimi:2021:RPD

[570] Michela Massimi. Realism, perspectivism, and disagreement in science. Synthese, 198(S25):6115–6141, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02500-6.

Seselja:2021:SLS

[571] Dunja Seselja. Some lessons from simulations of scientific disagreements. Synthese, 198(S25):6143–6158, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02182-0.

Santana:2021:LAD

[572] Carlos Santana. Let’s not agree to disagree: the role of strategic disagreement in science. Synthese, 198(S25):6159–6177, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02202-z.

Weatherall:2021:EEF

[573] James Owen Weatherall and Cailin O’Connor. Endogenous epistemic factionalization. Synthese, 198(S25):6179–6200, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02675-3.

Anzola:2021:DDB

[574] David Anzola. Disagreement in discipline-building processes. Synthese, 198(S25):6201–6224, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-019-02438-9.

Lowe:2021:EMC

[575] Benedikt Löwe, Ursula Martin, and Alison Pease. Enabling mathematical cultures: introduction. Synthese, 198(S26):6225–6231, ???? 2021.

CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02858-
y.

Andersen:2021:MWM

[576] Line Edslev Andersen, Mikkel Willum Johansen, and Henrik Kragh Sørensen. Mathematicians writing for mathematicians. Synthese, 198 (S26):6233–6250, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02145-5.

Ashton:2021:ARM

[577] Zoe Ashton. Audience role in mathematical proof development. Synthese, 198(S26):6251–6275, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02619-x.

Barany:2021:ARB

[578] Michael J. Barany. Abstract relations: bibliography and the infrastructures of modern mathematics. Synthese, 198(S26):6277–6290, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02683-3.

Johnson:2021:IFP

[579] Timothy Johnson. The influence of financial practice in developing mathematical probability. Synthese, 198(S26):6291–6331, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02636-w.

Friedman:2021:MRF

[580] Michael Friedman and Colin Jakob Rittberg. The material reasoning of folding paper. Synthese, 198(S26):6333–6367, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02131-x.

Inglis:2021:FEM

[581] Matthew Inglis and Juan Pablo Mejía-Ramos. Functional explanation in mathematics. Synthese, 198(S26):6369–6392, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02234-5.

Weber:2021:RSR

[582] Keith Weber. The role of syntactic representations in set theory. Synthese, 198(S26):6393–6412, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02157-1.

Liu:2021:EDT

[583] Yang Liu, Stephan Hartmann, and Huw Price. Editorial to “Decision theory and the future of AI”. Synthese, 198(S27):6413–6414, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03316-z.

Easwaran:2021:CNP

[584] Kenny Easwaran. A classification of Newcomb problems and decision theories. Synthese, 198(S27):6415–6434, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02272-z.

Everitt:2021:RTP

[585] Tom Everitt, Marcus Hutter, and Victoria Krakovna. Reward tampering problems and solutions in reinforcement learning: a causal influence diagram perspective. Synthese, 198(S27):6435–6467, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03141-4.

Kipper:2021:IID

[586] Jens Kipper. Intuition, intelligence, data compression. Synthese, 198 (S27):6469–6489, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02118-8.

Oesterheld:2021:ADA

[587] Caspar Oesterheld. Approval-directed agency and the decision theory of Newcomb-like problems. Synthese, 198(S27):6491–6504, ???? 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02148-
2.

Stern:2021:CCT

[588] Reuben Stern. Causal concepts and temporal ordering. Synthese, 198 (S27):6505–6527, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02235-4.

Zaffalon:2021:DFR

[589] Marco Zaffalon and Enrique Miranda. Desirability foundations of robust rational decision making. Synthese, 198(S27):6529–6570, ???? 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-02010-
x.

Zhang:2021:SCN

[590] Jiji Zhang, Teddy Seidenfeld, and Hailin Liu. Subjective causal networks and indeterminate suppositional credences. Synthese, 198(S27): 6571–6597, ???? 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02512-2.

Pallavicini:2021:PGB

[591] Josefine Pallavicini, Bjørn Hallsson, and Klemens Kappel. Polarization in groups of Bayesian agents. Synthese, 198(1):1–55, January 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01978-
w;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01978-w.pdf.

Dieveney:2021:SEU

[592] Patrick Dieveney. Scientific explanation, unifying mathematics, and indispensability arguments (a new Cicada MES). Synthese, 198(1):57–77, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-01979-9;   https://link.springer.com/content/pdf/
10.1007/s11229-018-01979-9.pdf.

Ross:2021:CEP

[593] Lauren N. Ross. Causal explanation and the periodic table. Synthese, 198(1):79–103, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01982-0;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01982-0.pdf.

Tallant:2021:OTR

[594] Jonathan Tallant and Sam Baron. It’s one thing to rule them all and another thing to bind them. Synthese, 198(1):105–115, January 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-01983-
z;   https://link.springer.com/content/pdf/10.1007/s11229-018-
01983-z.pdf.

Merdes:2021:SPT

[595] Christoph Merdes. Strategy and the pursuit of truth. Synthese, 198(1):117–138, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01985-x;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01985-x.pdf.

Easton:2021:WPP

[596] Christina Easton. Women and ‘the philosophical personality’: evaluating whether gender differences in the Cognitive Reflection Test have significance for explaining the gender gap in philosophy. Synthese, 198(1):139–167, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01986-w;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01986-w.pdf.

Darnell:2021:HPA

[597] Eamon Darnell and Aaron Thomas-Bolduc. Is Hume’s Principle analytic? Synthese, 198(1):169–185, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01988-8;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01988-8.pdf.

Demey:2021:ADS

[598] Lorenz Demey. Aristotelian diagrams for semantic and syntactic consequence. Synthese, 198(1):187–207, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01994-w;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01994-w.pdf.

deOliveira:2021:RDE

[599] Guilherme Sanches de Oliveira. Representationalism is a dead end. Synthese, 198(1):209–235, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01995-9;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01995-9.pdf.

Woodward:2021:EAR

[600] James Woodward. Explanatory autonomy: the role of proportionality, stability, and conditional irrelevance. Synthese, 198(1):237–265, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
01998-6;   https://link.springer.com/content/pdf/10.1007/s11229-
018-01998-6.pdf.

Zhou:2021:WPA

[601] Ziqian Zhou. What the progressive aspect tells us about processes. Synthese, 198(1):267–293, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01999-5;   https://
link.springer.com/content/pdf/10.1007/s11229-018-01999-5.pdf.

Nguyen:2021:FN

[602] Anthony Nguyen. A functional naturalism. Synthese, 198(1):295–313, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-02002-x;   https://link.springer.com/content/pdf/
10.1007/s11229-018-02002-x.pdf.

Reiland:2021:EMP

[603] Indrek Reiland. On experiencing moral properties. Synthese, 198(1):315– 325, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-018-02004-9;   https://link.springer.com/content/
pdf/10.1007/s11229-018-02004-9.pdf.

Koch:2021:ECC

[604] Steffen Koch. The externalist challenge to conceptual engineering. Synthese, 198(1):327–348, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02007-6;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02007-6.pdf.

vanDijk:2021:SA

[605] Ludger van Dijk and Erik Rietveld. Situated anticipation. Synthese, 198(1):349–371, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02013-8;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02013-8.pdf.

John:2021:SD

[606] Stephen John. Scientific deceit. Synthese, 198(1):373–394, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-02017-
4;   https://link.springer.com/content/pdf/10.1007/s11229-018-
02017-4.pdf.

Duncan:2021:DDD

[607] Michael Duncan, Kristie Miller, and James Norton. Ditching determination and dependence: or, how to wear the crazy trousers. Synthese, 198(1):395–418, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02023-6;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02023-6.pdf.

Shaw:2021:FWO

[608] Jamie Shaw. Feyerabend’s well-ordered science: how an anarchist distributes funds. Synthese, 198(1):419–449, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-02026-
3;   https://link.springer.com/content/pdf/10.1007/s11229-018-
02026-3.pdf.

Lageard:2021:TBR

[609] Valentin Lageard and Cédric Paternotte. Trolls, bans and reverts: simulating Wikipedia. Synthese, 198(1):451–470, January 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-02029-
0;   https://link.springer.com/content/pdf/10.1007/s11229-018-
02029-0.pdf.

Bland:2021:CBS

[610] Steven Bland. Cognitive bias, situationism, and virtue reliabilism. Synthese, 198(1):471–490, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02031-6;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02031-6.pdf.

Kosterec:2021:EEP

[611] Milos Kosterec. On the essence of empty properties. Synthese, 198(1):491–507, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02036-1;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02036-1.pdf.

Park:2021:CCC

[612] Ilho Park. Correcting credences with chances. Synthese, 198(1):509–536, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-02037-0;   https://link.springer.com/content/pdf/
10.1007/s11229-018-02037-0.pdf.

Northcott:2021:PEC

[613] Robert Northcott. Pre-emption cases may support, not undermine, the counterfactual theory of causation. Synthese, 198(1):537–555, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
02038-z;   https://link.springer.com/content/pdf/10.1007/s11229-
018-02038-z.pdf.

DiPaolo:2021:WRC

[614] Joshua DiPaolo. The word of a reluctant convert. Synthese, 198(1):557– 582, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-018-02042-3;   https://link.springer.com/content/
pdf/10.1007/s11229-018-02042-3.pdf.

Boyce:2021:WIB

[615] Kenneth Boyce. Why inference to the best explanation doesn’t secure empirical grounds for mathematical Platonism. Synthese, 198(1):583–595, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-02043-2;   https://link.springer.com/content/pdf/
10.1007/s11229-018-02043-2.pdf.

Hauswald:2021:EEC

[616] Rico Hauswald. The epistemic effects of close entanglements between research fields and activist movements. Synthese, 198(1):597–614, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-018-
02047-y;   https://link.springer.com/content/pdf/10.1007/s11229-
018-02047-y.pdf.

Dicher:2021:OSP

[617] Bogdan Dicher and Francesco Paoli. The original sin of proof-theoretic semantics. Synthese, 198(1):615–640, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02048-x;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02048-x.pdf.

Fernandez:2021:OMI

[618] Jordi Fernández. Observer memory and immunity to error through misidentification. Synthese, 198(1):641–660, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-018-02050-
3;   https://link.springer.com/content/pdf/10.1007/s11229-018-
02050-3.pdf.

Banick:2021:HAA

[619] Kyle Banick. How to be an adverbialist about phenomenal intentionality. Synthese, 198(1):661–686, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02053-0;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02053-0.pdf.

Bellucci:2021:ADP

[620] Francesco Bellucci. Analysis and decomposition in Peirce. Synthese, 198(1):687–706, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02054-z;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02054-z.pdf.

Fisher:2021:CBV

[621] David Fisher. A conflict between van Inwagen’s metaontology and his ontology. Synthese, 198(1):707–722, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02055-y;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02055-y.pdf.

vanWoudenberg:2021:RSK

[622] René van Woudenberg. Reading as a source of knowledge. Synthese, 198(1):723–742, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02056-x;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02056-x.pdf.

Coren:2021:ECB

[623] Daniel Coren. Epistemic conservatism and bare beliefs. Synthese, 198(1):743–756, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02059-8;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02059-8.pdf.

Turri:2021:LFS

[624] Angelo Turri and John Turri. Lying, fast and slow. Synthese, 198(1):757–775, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02062-z;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02062-z.pdf.

Pruss:2021:UIP

[625] Alexander R. Pruss. Underdetermination of infinitesimal probabilities. Synthese, 198(1):777–799, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02064-x;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02064-x.pdf.

Wallach:2021:HNE

[626] Efraim Wallach. Historiographic narratives and empirical evidence: a case study. Synthese, 198(1):801–821, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02065-w;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02065-w.pdf.

Deery:2021:FAN

[627] Oisín Deery. Free actions as a natural kind. Synthese, 198(1):823–843, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-02068-7;   https://link.springer.com/content/pdf/
10.1007/s11229-018-02068-7.pdf.

Kammerer:2021:ICE

[628] François Kammerer. The illusion of conscious experience. Synthese, 198(1):845–866, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02071-y;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02071-y.pdf.

Skrzypulec:2021:NMO

[629] Błazej Skrzypulec. The nonclassical mereology of olfactory experiences. Synthese, 198(1):867–886, January 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02072-x;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02072-x.pdf.

Fine:2021:TOG

[630] Kit Fine. Truthmaking and the is-ought gap. Synthese, 198(2):887–914, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-018-01996-8;   https://link.springer.com/content/pdf/
10.1007/s11229-018-01996-8.pdf.

Coates:2021:EIP

[631] Ashley Coates. Essence and the inference problem. Synthese, 198(2):915– 931, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-018-02074-9;   https://link.springer.com/content/
pdf/10.1007/s11229-018-02074-9.pdf.

Sedlar:2021:HLE

[632] Igor Sedlár. Hyperintensional logics for everyone. Synthese, 198(2):933– 956, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-018-02076-7;   https://link.springer.com/content/
pdf/10.1007/s11229-018-02076-7.pdf.

Finocchiaro:2021:IRM

[633] Peter Finocchiaro. Ideology and its role in metaphysics. Synthese, 198(2):957–983, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02077-6;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02077-6.pdf.

Naar:2021:PFL

[634] Hichem Naar. The possibility of fitting love: irreplaceability and selectivity. Synthese, 198(2):985–1010, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-02079-4;   https://
link.springer.com/content/pdf/10.1007/s11229-018-02079-4.pdf.

Jaquet:2021:DAA

[635] François Jaquet. A debunking argument against speciesism. Synthese, 198(2):1011–1027, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02080-5;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02080-5.pdf.

Fischer:2021:EOL

[636] Eugen Fischer, Paul E. Engelhardt, Joachim Horvath, and Hiroshi Ohtani. Experimental ordinary language philosophy: a cross-linguistic study of defeasible default inferences. Synthese, 198(2):1029–1070, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02081-4;   https://link.springer.com/content/pdf/
10.1007/s11229-019-02081-4.pdf.

Roche:2021:APH

[637] Michael Roche and William Roche. Authority without privilege: How to be a Dretskean conciliatory skeptic on self-knowledge. Synthese, 198(2):1071–1087, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02082-3;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02082-3.pdf.

Smyth:2021:MDN

[638] Nicholas Smyth. Moral disagreement and non-moral ignorance. Synthese, 198(2):1089–1108, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02084-1;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02084-1.pdf.

Dethier:2021:FTP

[639] Corey Dethier. Forces in a true and physical sense: from mathematical models to metaphysical conclusions. Synthese, 198(2):1109–1122, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02086-z;   https://link.springer.com/content/pdf/
10.1007/s11229-019-02086-z.pdf.

Schloder:2021:CKR

[640] Julian J. Schlöder. Counterfactual knowability revisited. Synthese, 198(2):1123–1137, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02087-y;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02087-y.pdf.

Neil:2021:SDD

[641] Charles Neil. Safety, domination, and differential support. Synthese, 198(2):1139–1152, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02091-2;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02091-2.pdf.

Satta:2021:ETM

[642] Mark Satta. Evil twins and the multiverse: distinguishing the world of difference between epistemic and physical possibility. Synthese, 198(2):1153–1160, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02092-1;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02092-1.pdf.

Lehet:2021:IED

[643] Ellen Lehet. Induction and explanatory definitions in mathematics. Synthese, 198(2):1161–1175, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02095-y;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02095-y.pdf.

Trammell:2021:FPS

[644] Philip Trammell. Fixed-point solutions to the regress problem in normative uncertainty. Synthese, 198(2):1177–1199, February 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02098-
9;   https://link.springer.com/content/pdf/10.1007/s11229-019-
02098-9.pdf.

Yarandi:2021:ECR

[645] Seyed Mohammad Yarandi. Externalism and critical reasoning: a reconsideration. Synthese, 198(2):1201–1216, February 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02099-
8;   https://link.springer.com/content/pdf/10.1007/s11229-019-
02099-8.pdf.

Oldofredi:2021:SRM

[646] Andrea Oldofredi. Some remarks on the mentalistic reformulation of the measurement problem: a reply to S. Gao. Synthese, 198(2):1217–1233, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02101-3;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02101-3.pdf.

See [9].

Locke:2021:CMN

[647] Theodore D. Locke. Counterpossibles for modal normativists. Synthese, 198(2):1235–1257, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02103-1;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02103-1.pdf.

Xu:2021:LKW

[648] Chao Xu, Yanjing Wang, and Thomas Studer. A logic of knowing why. Synthese, 198(2):1259–1285, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02104-0;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02104-0.pdf.

Steel:2021:IEE

[649] Daniel Steel, Sina Fazelpour, Bianca Crewe, and Kinley Gillette. Information elaboration and epistemic effects of diversity. Synthese, 198(2):1287–1307, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02108-w;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02108-w.pdf.

Morris:2021:DME

[650] Rebecca Lea Morris. Do mathematical explanations have instrumental value? Synthese, 198(2):1309–1328, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02114-y;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02114-y.pdf.

Stuart:2021:TDP

[651] Michael T. Stuart. Towards a dual process epistemology of imagination. Synthese, 198(2):1329–1350, February 2021. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02116-w;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02116-w.pdf.

Lutz:2021:GEA

[652] Sebastian Lutz. Generalizing empirical adequacy II: partial structures. Synthese, 198(2):1351–1380, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02121-z;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02121-z.pdf.

Donaldson:2021:VVH

[653] John Donaldson. Vertical versus horizontal: What is really at issue in the exclusion problem? Synthese, 198(2):1381–1396, February 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02126-
8;   https://link.springer.com/content/pdf/10.1007/s11229-019-
02126-8.pdf.

Rodrigues:2021:ASS

[654] Mariana Vitti Rodrigues and Claus Emmeche. Abduction and styles of scientific thinking. Synthese, 198(2):1397–1425, February 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02127-
7;   https://link.springer.com/content/pdf/10.1007/s11229-019-
02127-7.pdf.

Byrd:2021:WWC

[655] Nick Byrd. What we can (and can’t) infer about implicit bias from debiasing experiments. Synthese, 198(2):1427–1455, February 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02128-
6;   https://link.springer.com/content/pdf/10.1007/s11229-019-
02128-6.pdf.

Friedman:2021:ACT

[656] Daniel A. Friedman and Eirik Søvik. The ant colony as a test for scientific theories of consciousness. Synthese, 198(2):1457–1480, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02130-
y;   https://link.springer.com/content/pdf/10.1007/s11229-019-
02130-y.pdf.

Attie-Picker:2021:FCL

[657] Mario Attie-Picker. Is the folk concept of luck normative? Synthese, 198(2):1481–1515, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02137-5;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02137-5.pdf.

Kozyreva:2021:IUE

[658] Anastasia Kozyreva and Ralph Hertwig. The interpretation of uncertainty in ecological rationality. Synthese, 198(2):1517–1547, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL https://link.springer.com/article/10.1007/s11229-019-
02140-w;   https://link.springer.com/content/pdf/10.1007/s11229-
019-02140-w.pdf.

Smith:2021:QNN

[659] James Andrew Smith, Jr. Quine on naturalism, nominalism, and philosophy’s place within science. Synthese, 198(2):1549–1567, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02143-
7;   https://link.springer.com/content/pdf/10.1007/s11229-019-
02143-7.pdf.

Coren:2021:EEV

[660] Daniel Coren. Evaluating epistemic virtues. Synthese, 198(2):1569–1578, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02144-6;   https://link.springer.com/content/pdf/
10.1007/s11229-019-02144-6.pdf.

Spinelli:2021:SIE

[661] Nicola Spinelli. Is self-identity essential to objects? Synthese, 198(2):1579–1595, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02151-7;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02151-7.pdf.

Overgaard:2021:CBG

[662] Søren Overgaard and Alessandro Salice. Consciousness, belief, and the group mind hypothesis. Synthese, 198(2):1597–1621, February 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02152-
6;   https://link.springer.com/content/pdf/10.1007/s11229-019-
02152-6.pdf.

Khalaj:2021:KHP

[663] M. Hosein M. A. Khalaj. Knowledge-how and the problems of masking and finkishness. Synthese, 198(2):1623–1641, February 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02155-
3;   https://link.springer.com/content/pdf/10.1007/s11229-019-
02155-3.pdf.

Kramer:2021:GTE

[664] Stephan Krämer. Ground-theoretic equivalence. Synthese, 198(2):1643– 1683, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02154-4;   https://link.springer.com/content/
pdf/10.1007/s11229-019-02154-4.pdf.

Darby:2021:MDI

[665] George Darby and Martin Pickup. Modelling deep indeterminacy. Synthese, 198(2):1685–1710, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02158-0;   https://
link.springer.com/content/pdf/10.1007/s11229-019-02158-0.pdf.

Acuna:2021:CLI

[666] Pablo Acuña. Charting the landscape of interpretation, theory rivalry, and underdetermination in quantum mechanics. Synthese, 198(2):1711– 1740, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02159-z;   https://link.springer.com/content/
pdf/10.1007/s11229-019-02159-z.pdf.

Elzinga:2021:IKH

[667] Benjamin Elzinga. Intellectualizing know how. Synthese, 198(2):1741– 1760, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02160-6;   https://link.springer.com/content/
pdf/10.1007/s11229-019-02160-6.pdf.

Bondy:2021:ENI

[668] Patrick Bondy. The epistemic norm of inference and non-epistemic reasons for belief. Synthese, 198(2):1761–1781, February 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02163-
3;   https://link.springer.com/content/pdf/10.1007/s11229-019-
02163-3.pdf.

Saint-Germier:2021:GGS

[669] Pierre Saint-Germier. Getting Gettier straight: thought experiments, deviant realizations and default interpretations. Synthese, 198(2):1783– 1806, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02166-0;   https://link.springer.com/content/
pdf/10.1007/s11229-019-02166-0.pdf.

Schubbach:2021:JMP

[670] Arno Schubbach. Judging machines: philosophical aspects of deep learning. Synthese, 198(2):1807–1827, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02167-z.

Prychitko:2021:CSA

[671] Emily Prychitko. The causal situationist account of constitutive relevance. Synthese, 198(2):1829–1843, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02170-4.

Harris:2021:KHF

[672] Keith Harris. Knowledge-how and false belief. Synthese, 198(2):1845– 1861, February 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02172-2.

Grandjean:2021:HAB

[673] Vincent Grandjean. How is the asymmetry between the open future and the fixed past to be characterized? Synthese, 198(3):1863–1886, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02164-2.

Mandrigin:2021:WBA

[674] Alisa Mandrigin. The where of bodily awareness. Synthese, 198(3): 1887–1903, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02171-3.

Snyder:2021:HAT

[675] Eric Snyder, Richard Samuels, and Stewart Shaprio. Hale’s argument from transitive counting. Synthese, 198(3):1905–1933, March 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02178-
w.

Rescorla:2021:PFC

[676] Michael Rescorla. On the proper formulation of conditionalization. Synthese, 198(3):1935–1965, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02179-9.

Corrado:2021:FWF

[677] Michael Louis Corrado. Free will fallibilism and the “two-standpoints” account of freedom. Synthese, 198(3):1967–1982, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02181-1.

Fernandes:2021:TTC

[678] Alison Fernandes. Time travel and counterfactual asymmetry. Synthese, 198(3):1983–2001, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02186-w.

Deulofeu:2021:EBR

[679] Roger Deulofeu, Javier Suárez, and Alberto Pérez-Cervera. Explaining the behaviour of random ecological networks: the stability of the microbiome as a case of integrative pluralism. Synthese, 198(3): 2003–2025, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02187-9.

Hindriks:2021:FIE

[680] Frank Hindriks and Francesco Guala. The functions of institutions: etiology and teleology. Synthese, 198(3):2027–2043, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02188-8.

Kotsonis:2021:PCI

[681] Alkis Kotsonis. The platonic conception of intellectual virtues: its significance for virtue epistemology. Synthese, 198(3):2045–2060, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02189-7.

Landgrebe:2021:MAM

[682] Jobst Landgrebe and Barry Smith. Making AI meaningful again. Synthese, 198(3):2061–2081, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02192-y.

Hetherington:2021:SFK

[683] Stephen Hetherington. Some fallibilist knowledge: Questioning knowledge-attributions and open knowledge. Synthese, 198(3):2083– 2099, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02194-w.

Skrzypulec:2021:SAV

[684] Błazej Skrzypulec. The structure of audio-visual consciousness. Synthese, 198(3):2101–2127, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02195-9.

Kovacs:2021:HUR

[685] David Mark Kovacs. How to be an uncompromising revisionary ontologist. Synthese, 198(3):2129–2152, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02196-8.

Rowbottom:2021:MAA

[686] Darrell P. Rowbottom. A methodological argument against scientific realism. Synthese, 198(3):2153–2167, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02197-7.

Shahmoradi:2021:RRK

[687] Ayoob Shahmoradi. A representationalist reading of Kantian intuitions. Synthese, 198(3):2169–2191, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02198-6.

Alsmith:2021:BSB

[688] Adrian J. T. Alsmith. Bodily structure and body representation. Synthese, 198(3):2193–2222, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02200-1.

ChoGlueck:2021:BSO

[689] Christopher ChoGlueck. Broadening the scope of our understanding of mechanisms: lessons from the history of the morning-after pill. Synthese, 198(3):2223–2252, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02201-0.

Draper:2021:DIS

[690] Kaila Draper. Direct inference and the sleeping beauty problem. Synthese, 198(3):2253–2271, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02203-y.

Cowling:2021:RNQ

[691] Sam Cowling. Recombining non-qualitative reality. Synthese, 198(3): 2273–2295, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02204-x.

Matthews:2021:AGD

[692] Lucas J. Matthews and Eric Turkheimer. Across the great divide: pluralism and the hunt for missing heritability. Synthese, 198(3): 2297–2311, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02205-w.

Chen:2021:EIP

[693] Ruey-Lin Chen and Jonathon Hricko. Experimental individuation and philosophical retail arguments. Synthese, 198(3):2313–2332, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02207-

Morvarid:2021:NAI

[694] Mahmoud Morvarid. A new argument for the incompatibility of content externalism with justification internalism. Synthese, 198(3):2333– 2353, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02208-7.

Kovaka:2021:CCD

[695] Karen Kovaka. Climate change denial and beliefs about science. Synthese, 198(3):2355–2374, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02210-z.

Cupples:2021:MDH

[696] Laura M. Cupples. Measure development and the hermeneutic task. Synthese, 198(3):2375–2390, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02213-w.

Peet:2021:TW

[697] Andrew Peet. Testimonial worth. Synthese, 198(3):2391–2411, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02219-4.

Buckwalter:2021:EPC

[698] Wesley Buckwalter. Error possibility, contextualism, and bias. Synthese, 198(3):2413–2426, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02221-w.

Frigg:2021:MW

[699] Roman Frigg and James Nguyen. Mirrors without warnings. Synthese, 198(3):2427–2447, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-019-02222-9.

Raleigh:2021:SB

[700] Thomas Raleigh. Suspending is believing. Synthese, 198(3):2449–2474, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02223-8.

Dellsen:2021:RUT

[701] Finnur Dellsén. Rational understanding: toward a probabilistic epistemology of acceptability. Synthese, 198(3):2475–2494, March 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02224-
7.

Carter:2021:EA

[702] J. Adam Carter. Exercising abilities. Synthese, 198(3):2495–2509, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02227-4.

Martens:2021:DAN

[703] Niels C. M. Martens. The (un)detectability of absolute Newtonian masses. Synthese, 198(3):2511–2550, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02229-2.

List:2021:DSS

[704] Christian List and Marcus Pivato. Dynamic and stochastic systems as a framework for metaphysics and the philosophy of science. Synthese, 198 (3):2551–2612, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02231-8.

Sharadin:2021:EEI

[705] Nathaniel Sharadin. Ecumenical epistemic instrumentalism. Synthese, 198(3):2613–2639, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-019-02232-7.

Raikka:2021:DI

[706] Juha Räikkä. Distribution and ignorance. Synthese, 198(3):2641–2657, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02236-3.

Lohr:2021:SCC

[707] Guido Löhr. Social constructionism, concept acquisition and the mismatch problem. Synthese, 198(3):2659–2673, March 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02237-2.

Stegmann:2021:SMD

[708] Ulrich E. Stegmann. Scientific modelling with diagrams. Synthese, 198 (3):2675–2694, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02239-0.

Rosenblatt:2021:SCW

[709] Lucas Rosenblatt. On structural contraction and why it fails. Synthese, 198(3):2695–2720, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02241-6.

Warenski:2021:ENT

[710] Lisa Warenski. Epistemic norms: truth conducive enough. Synthese, 198 (3):2721–2741, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02242-5.

Korbak:2021:CEU

[711] Tomasz Korbak. Computational enactivism under the free energy principle. Synthese, 198(3):2743–2763, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02243-4.

Fisher:2021:RTE

[712] Sarah A. Fisher. Reassessing truth-evaluability in the Minimalism– Contextualism debate. Synthese, 198(3):2765–2782, March 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02245-
2.

McCutcheon:2021:HCE

[713] Randall G. McCutcheon. How to co-exist with nonexistent expectations. Synthese, 198(3):2783–2799, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02246-1.

Baravalle:2021:CET

[714] Lorenzo Baravalle. Cultural evolutionary theory as a theory of forces. Synthese, 198(3):2801–2820, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02247-0.

Graziani:2021:TOT

[715] Ernesto Graziani and Francesco Orilia. Temporal ontology: tenselessness and quantification. Synthese, 198(3):2821–2847, March 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02248-z.

Longino:2021:SSW

[716] Helen E. Longino. Scaling up; scaling down: What’s missing? Synthese, 198(4):2849–2863, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02249-y.

Nowak:2021:CDH

[717] Ethan Nowak. Complex demonstratives, hidden arguments, and presupposition. Synthese, 198(4):2865–2900, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02250-5.

Rydenfelt:2021:RR

[718] Henrik Rydenfelt. Realism without representationalism. Synthese, 198 (4):2901–2918, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02251-4.

Fuller:2021:MFS

[719] Jonathan Fuller. The myth and fallacy of simple extrapolation in medicine. Synthese, 198(4):2919–2939, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02255-0.

Peters:2021:TME

[720] Uwe Peters. Teleology and mentalizing in the explanation of action. Synthese, 198(4):2941–2957, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02256-z.

Predelli:2021:FD

[721] Stefano Predelli. Four Dthats. Synthese, 198(4):2959–2972, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02257-
y.

DeHaro:2021:SD

[722] Sebastian De Haro and Jeremy Butterfield. On symmetry and duality. Synthese, 198(4):2973–3013, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02258-x.

Hill:2021:U

[723] Christopher S. Hill. Unrevisability. Synthese, 198(4):3015–3031, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02262-1.

Heylen:2021:EIA

[724] Jan Heylen and Lars Arthur Tump. The enhanced indispensability argument, the circularity problem, and the interpretability strategy. Synthese, 198(4):3033–3045, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02263-0.

Kerr:2021:PK

[725] Alison Duncan Kerr. A plea for KR. Synthese, 198(4):3047–3071, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02265-y.

Schipper:2021:FTN

[726] Arthur Schipper. Fundamental truthmakers and non-fundamental truths. Synthese, 198(4):3073–3098, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02266-x.

Bronzo:2021:PCF

[727] Silver Bronzo. Propositional complexity and the Frege–Geach Point. Synthese, 198(4):3099–3130, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02270-1.

Lopez-Rubio:2021:DSM

[728] Ezequiel López-Rubio and Emanuele Ratti. Data science and molecular biology: prediction and mechanistic explanation. Synthese, 198(4): 3131–3156, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02271-0.

Simpson:2021:DT

[729] Matthew Simpson. Deflationism and truthmaking. Synthese, 198(4): 3157–3181, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02273-y.

Lougheed:2021:CEP

[730] Kirk Lougheed. Catherine Elgin on peerhood and the epistemic benefits of disagreement. Synthese, 198(4):3183–3202, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02274-x.

Campbell:2021:DSS

[731] Douglas Ian Campbell and Yi Yang. Does the solar system compute the laws of motion? Synthese, 198(4):3203–3220, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02275-w.

Fabry:2021:FLR

[732] Regina E. Fabry and Markus Pantsar. A fresh look at research strategies in computational cognitive science: The case of enculturated mathematical problem solving. Synthese, 198(4):3221–3263, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02276-9.

Matthen:2021:OSO

[733] Mohan Matthen. Objects, seeing, and object-seeing. Synthese, 198(4): 3265–3288, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02279-6.

Carbonell:2021:MRM

[734] Felipe Morales Carbonell. Mistakes as revealing and as manifestations of competence. Synthese, 198(4):3289–3308, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02281-y.

Ferretti:2021:VPV

[735] Gabriele Ferretti. Visual phenomenology versus visuomotor imagery: How can we be aware of action properties? Synthese, 198(4):3309– 3338, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02282-x.

Woods:2021:FGI

[736] John Woods. Four grades of ignorance-involvement and how they nourish the cognitive economy. Synthese, 198(4):3339–3368, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02283-w.

Taylor:2021:WFA

[737] Henry Taylor. Whales, fish and Alaskan bears: interest-relative taxonomy and kind pluralism in biology. Synthese, 198(4):3369–3387, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02284-9.

Reydon:2021:RSP

[738] Thomas A. C. Reydon. On radical solutions in the philosophy of biology: What does “individuals thinking” actually solve? Synthese, 198 (4):3389–3411, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02285-8.

Abrams:2021:WCM

[739] Lowell Abrams and Landon D. C. Elkind. Word choice in mathematical practice: a case study in polyhedra. Synthese, 198(4):3413–3441, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02287-6.

Vooys:2021:MMC

[740] Sarah Vooys and David G. Dick. Money and mental contents. Synthese, 198(4):3443–3458, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02288-5.

Krabbe:2021:QQD

[741] Erik C. W. Krabbe and Jan Albert van Laar. In the quagmire of quibbles: a dialectical exploration. Synthese, 198(4):3459–3476, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02289-4.

Siscoe:2021:NWT

[742] Robert Weston Siscoe. No work for a theory of epistemic dispositions. Synthese, 198(4):3477–3498, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02291-w.

McShane:2021:AEF

[743] Katie McShane. Against etiological function accounts of interests. Synthese, 198(4):3499–3517, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02293-8.

Mitchell:2021:ARA

[744] Jonathan Mitchell. Affective representation and affective attitudes. Synthese, 198(4):3519–3546, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02294-7.

Gigerenzer:2021:ARE

[745] Gerd Gigerenzer. Axiomatic rationality and ecological rationality. Synthese, 198(4):3547–3564, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02296-5.

Carrara:2021:WDK

[746] Massimiliano Carrara, Daniele Chiffi, Ciro De Florio, and Ahti-Veikko Pietarinen. We don’t know we don’t know: asserting ignorance. Synthese, 198(4):3565–3580, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02300-y.

Carr:2021:MP

[747] Jennifer Rose Carr. A modesty proposal. Synthese, 198(4):3581–3601, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02301-x.

Placek:2021:PPF

[748] Tomasz Placek. Past, present and future modally introduced. Synthese, 198(4):3603–3624, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02302-w.

Hiernaux:2021:HEP

[749] Quentin Hiernaux. History and epistemology of plant behaviour: a pluralistic view? Synthese, 198(4):3625–3650, April 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02303-9.

Munro:2021:VBS

[750] Daniel Munro. Visual and bodily sensational perception: an epistemic asymmetry. Synthese, 198(4):3651–3674, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02304-8.

Wiltsche:2021:FWU

[751] Harald A. Wiltsche. The forever war: understanding, science fiction, and thought experiments. Synthese, 198(4):3675–3698, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02306-6.

Bourrat:2021:TEF

[752] Pierrick Bourrat. Transitions in evolution: a formal analysis. Synthese, 198(4):3699–3731, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02307-5.

Elliott:2021:BAZ

[753] Edward Elliott. Betting against the Zen Monk: on preferences and partial belief. Synthese, 198(4):3733–3758, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02308-4.

Buttner:2021:HPP

[754] Kai Michael Büttner. Hume’s principle: a plea for austerity. Synthese, 198(4):3759–3781, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02309-3.

Persichetti:2021:LWG

[755] Alessio Persichetti. The later Wittgenstein’s guide to contradictions. Synthese, 198(4):3783–3799, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02310-w.

Redei:2021:HLB

[756] Miklós Rédei and Zalán Gyenis. Having a look at the Bayes Blind Spot. Synthese, 198(4):3801–3832, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02311-9.

Walker:2021:KFS

[757] Barnaby Walker. Knowledge first, stability and value. Synthese, 198 (4):3833–3854, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02313-7.

Wimmer:2021:KFB

[758] Simon Wimmer. Knowledge-first believing the unknowable. Synthese, 198(4):3855–3871, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-019-02314-6.

Parviainen:2021:NEC

[759] Jaana Parviainen and Lauri Lahikainen. Negative expertise in conditions of manufactured ignorance: epistemic strategies, virtues and skills. Synthese, 198(4):3873–3891, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02315-5.

Lange:2021:ACC

[760] Marc Lange. Asymmetry as a challenge to counterfactual accounts of non-causal explanation. Synthese, 198(4):3893–3918, April 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02317-
3.

Borrelli:2021:BSA

[761] Arianna Borrelli. Between symmetry and asymmetry: spontaneous symmetry breaking as narrative knowing. Synthese, 198(4):3919–3948, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02320-8.

Woods:2021:WDF

[762] John Woods. What did Frege take Russell to have proved? Synthese, 198 (4):3949–3977, April 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02324-4.

Alexander:2021:FRB

[763] David Alexander. Fallible reasons on behalf of fallibilism. Synthese, 198 (5):3979–3998, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-017-1452-6.

Fernandes:2021:DTA

[764] Alison Fernandes. Does the temporal asymmetry of value support a tensed metaphysics? Synthese, 198(5):3999–4016, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02269-8.

Lewis:2021:ACP

[765] Peter J. Lewis and Don Fallis. Accuracy, conditionalization, and probabilism. Synthese, 198(5):4017–4033, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02298-3.

Kindermann:2021:KE

[766] Dirk Kindermann. Knowledge embedded. Synthese, 198(5):4035–4055, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02326-2.

OMadagain:2021:JAM

[767] Cathal O’Madagain and Michael Tomasello. Joint attention to mental content and the social origin of reasoning. Synthese, 198(5):4057– 4078, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02327-1.

Arfini:2021:SID

[768] Selene Arfini. Situated ignorance: the distribution and extension of ignorance in cognitive niches. Synthese, 198(5):4079–4095, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02328-0.

Rice:2021:UR

[769] Collin Rice. Understanding realism. Synthese, 198(5):4097–4121, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02331-5.

Frierson:2021:SIV

[770] Patrick Frierson. Situationism and intellectual virtue: a Montessori perspective. Synthese, 198(5):4123–4144, May 2021. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02332-4.

Glick:2021:PSR

[771] David Glick. Pluralist structural realism: The best of both worlds? Synthese, 198(5):4145–4166, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02333-3.

DAgostino:2021:GKD

[772] Fred D’Agostino. Growth of knowledge: dual institutionalization of disciplines and brokerage. Synthese, 198(5):4167–4190, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02335-1.

Calosi:2021:DLI

[773] Claudio Calosi. Determinables, location, and indeterminacy. Synthese, 198(5):4191–4204, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02336-0.

Mohammadian:2021:ACD

[774] Mousa Mohammadian. Abduction − the context of discovery + underdetermination = inference to the best explanation. Synthese, 198 (5):4205–4228, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02337-z.

Steinberg:2021:HPL

[775] Alex Steinberg. How to properly lose direction. Synthese, 198(5): 4229–4250, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02338-y.

Latham:2021:ONT

[776] Andrew J. Latham, Kristie Miller, and James Norton. Is our naïve theory of time dynamical? Synthese, 198(5):4251–4271, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02340-4.

Baron:2021:NNM

[777] Samuel Baron, Mark Colyvan, and Michael Rubin. Non-naturalistic moral explanation. Synthese, 198(5):4273–4294, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02341-3.

Nawar:2021:VRU

[778] Tamer Nawar. Veritism refuted? Understanding, idealization, and the facts. Synthese, 198(5):4295–4313, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02342-2.

Sune:2021:RPI

[779] Abel Suñé and Manolo Martínez. Real patterns and indispensability. Synthese, 198(5):4315–4330, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02343-1.

Costa:2021:AAA

[780] Damiano Costa. An argument against Aristotelian universals. Synthese, 198(5):4331–4338, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02345-z.

Gallow:2021:SGD

[781] J. Dmitri Gallow. A subjectivist’s guide to deterministic chance. Synthese, 198(5):4339–4372, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02346-y.

Doulatova:2021:TIP

[782] Maria Doulatova. Tracking intentionalism and the phenomenology of mental effort. Synthese, 198(5):4373–4389, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02347-x.

Heyes:2021:MGN

[783] Cecilia Heyes. Is morality a gadget? Nature, nurture and culture in moral development. Synthese, 198(5):4391–4414, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02348-w.

Dach:2021:SWI

[784] Stefanie Dach. Sellars, we-intentions and ought-statements. Synthese, 198(5):4415–4439, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-019-02349-9.

Chen:2021:DSI

[785] Lu Chen. Do simple infinitesimal parts solve Zeno’s paradox of measure? Synthese, 198(5):4441–4456, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02350-2.

Green:2021:BDM

[786] E. J. Green. Binding and differentiation in multisensory object perception. Synthese, 198(5):4457–4491, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02351-1.

Tate:2021:PPT

[787] Alex James Miller Tate. A predictive processing theory of motivation. Synthese, 198(5):4493–4521, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02354-y.

Roy:2021:SIL

[788] Olivier Roy and Anne Schwenkenbecher. Shared intentions, loose groups, and pooled knowledge. Synthese, 198(5):4523–4541, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02355-x.

Malaterre:2021:BCD

[789] Christophe Malaterre and Jean-François Chartier. Beyond categorical definitions of life: a data-driven approach to assessing lifeness. Synthese, 198(5):4543–4572, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02356-w.

Taylor:2021:TKE

[790] Samuel D. Taylor. Two kinds of explanatory integration in cognitive science. Synthese, 198(5):4573–4601, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02357-9.

Plaisance:2021:SMN

[791] Kathryn S. Plaisance, Alexander V. Graham, and Jay Michaud. Show me the numbers: a quantitative portrait of the attitudes, experiences, and values of philosophers of science regarding broadly engaged work. Synthese, 198(5):4603–4633, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02359-7.

Atkinson:2021:WAD

[792] C. J. Atkinson. Weighing aims in doxastic deliberation. Synthese, 198 (5):4635–4650, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02360-0.

Cox:2021:HWI

[793] Ryan Cox. How why-interrogatives work. Synthese, 198(5):4651–4688, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02364-w.

Andrada:2021:MN

[794] Gloria Andrada. Mind the notebook. Synthese, 198(5):4689–4708, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02365-9.

Weksler:2021:TEN

[795] Assaf Weksler, Hilla Jacobson, and Zohar Z. Bronfman. The transparency of experience and the neuroscience of attention. Synthese, 198 (5):4709–4730, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02366-8.

Ranalli:2021:CII

[796] Christopher Ranalli and René van Woudenberg. Collective ignorance: an information theoretic account. Synthese, 198(5):4731–4750, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02367-7.

Leuenberger:2021:ELC

[797] Stephan Leuenberger and Martin Smith. Epistemic logic without closure. Synthese, 198(5):4751–4774, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02368-6.

Tahko:2021:SLR

[798] Tuomas E. Tahko. A survey of logical realism. Synthese, 198(5): 4775–4790, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02369-5.

Kirchhoff:2021:HDB

[799] Michael D. Kirchhoff and Julian Kiverstein. How to determine the boundaries of the mind: a Markov blanket proposal. Synthese, 198 (5):4791–4810, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02370-y.

Fletcher:2021:IAS

[800] Samuel C. Fletcher. An invitation to approximate symmetry, with three applications to intertheoretic relations. Synthese, 198(5):4811– 4831, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02371-x.

Bolinger:2021:DSD

[801] Renée Jorgensen Bolinger. Demographic statistics in defensive decisions. Synthese, 198(5):4833–4850, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02372-w.

Andreotta:2021:CDU

[802] Adam J. Andreotta. Confabulation does not undermine introspection for propositional attitudes. Synthese, 198(5):4851–4872, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02373-9.

Gebharter:2021:CBN

[803] Alexander Gebharter and Florian Fischer. A causal Bayes net analysis of dispositions. Synthese, 198(5):4873–4895, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02374-8.

Avnur:2021:JLN

[804] Yuval Avnur. Justification as a loaded notion. Synthese, 198(5): 4897–4916, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02375-7.

Won:2021:MCJ

[805] Chiwook Won. Mental causation as joint causation. Synthese, 198(5): 4917–4937, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02378-4.

Himmelreich:2021:ERT

[806] Johannes Himmelreich. Existence, really? Tacit disagreements about “existence” in disputes about group minds and corporate agents. Synthese, 198(5):4939–4953, May 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02379-3.

Gordon:2021:EUH

[807] Emma C. Gordon. Elgin on understanding: How does it involve know-how, endorsement and factivity? Synthese, 198(6):4955–4972, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02381-9.

Meester:2021:IER

[808] Ronald Meester and Timber Kerkvliet. The infinite epistemic regress problem has no unique solution. Synthese, 198(6):4973–4983, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02383-
7.

Kraay:2021:ECO

[809] Klaas J. Kraay. Elgin’s community-oriented steadfastness. Synthese, 198 (6):4985–5008, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02384-6.

Story:2021:JAR

[810] Daniel Story. Joint action without robust theory of mind. Synthese, 198 (6):5009–5026, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02386-4.

Troquard:2021:TMD

[811] Nicolas Troquard. Tracking and managing deemed abilities. Synthese, 198(6):5027–5045, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-019-02387-3.

Ciuni:2021:IIS

[812] Roberto Ciuni and Giovanni Tuzet. Inevitable ignorance as a standard for excusability: an epistemological analysis. Synthese, 198(6): 5047–5066, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02388-2.

Piazza:2021:VTN

[813] Tommaso Piazza. The value of truth and the normativity of evidence. Synthese, 198(6):5067–5088, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02389-1.

Noichl:2021:MSR

[814] Maximilian Noichl. Modeling the structure of recent philosophy. Synthese, 198(6):5089–5100, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02390-8.

Nebel:2021:DPA

[815] Jonathan Nebel. Doubting Pritchard’s account of hinge propositions. Synthese, 198(6):5101–5113, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02392-6.

Johnson:2021:RSR

[816] Christa M. Johnson. Resolutions, salient reasons, and weakness of will. Synthese, 198(6):5115–5138, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02393-5.

DeHaro:2021:TED

[817] Sebastian De Haro. Theoretical equivalence and duality. Synthese, 198 (6):5139–5177, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02394-4.

Rosenhagen:2021:NRH

[818] Raja Rosenhagen. Norwood Russell Hanson’s account of experience: an untimely defense. Synthese, 198(6):5179–5204, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02395-3.

Filomeno:2021:NAR

[819] Aldo Filomeno. Are non-accidental regularities a cosmic coincidence? Revisiting a central threat to Humean laws. Synthese, 198(6):5205– 5227, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02397-1. See correction [992].

Gasparri:2021:HMR

[820] Luca Gasparri and Michael Murez. Hearing meanings: the revenge of context. Synthese, 198(6):5229–5252, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02398-0.

Grimau:2021:DHL

[821] Berta Grimau. In defence of Higher-Level Plural Logic: drawing conclusions from natural language. Synthese, 198(6):5253–5280, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02399-z.

Mozersky:2021:NCN

[822] M. Joshua Mozersky. Nominalism, contingency, and natural structure. Synthese, 198(6):5281–5296, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02401-8.

Mooney:2021:SCC

[823] Justin Mooney. Self-colocation: a colocation puzzle for endurantists. Synthese, 198(6):5297–5309, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02402-7.

Papineau:2021:DK

[824] David Papineau. The disvalue of knowledge. Synthese, 198(6):5311– 5332, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02405-4. See correction [825].

Papineau:2021:CDK

[825] David Papineau. Correction to: The disvalue of knowledge. Synthese, 198(6):5333, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02434-z. See [824].

Hundertmark:2021:EHP

[826] Fabian Hundertmark. Explaining how to perceive the new: causalinformational teleosemantics and productive response functions. Synthese, 198(6):5335–5350, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02406-3.

Holguin:2021:LK

[827] Ben Holguín. Lying and knowing. Synthese, 198(6):5351–5371, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02407-2.

Singer:2021:DFF

[828] Daniel J. Singer, Aaron Bramson, Bennett Holman, Karen Kovaka, Jiin Jung, and William J. Berger. Don’t forget forgetting: the social epistemic importance of how we forget. Synthese, 198(6):5373– 5394, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02409-0.

Ghijsen:2021:TBD

[829] Harmen Ghijsen. Traits, beliefs and dispositions in a pluralistic folk psychology. Synthese, 198(6):5395–5413, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02412-5.

LeBihan:2021:PTV

[830] Soazig Le Bihan. Partial truth versus felicitous falsehoods. Synthese, 198 (6):5415–5436, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02413-4.

Bortolotti:2021:CBC

[831] Lisa Bortolotti and Ema Sullivan-Bissett. Is choice blindness a case of self-ignorance? Synthese, 198(6):5437–5454, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02414-3.

Cote-Bouchard:2021:TTE

[832] Charles Côté-Bouchard. Two types of epistemic instrumentalism. Synthese, 198(6):5455–5475, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02415-2.

Singh:2021:EDM

[833] Keshav Singh. Evidentialism doesn’t make an exception for belief. Synthese, 198(6):5477–5494, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02416-1.

Boesch:2021:SRD

[834] Brandon Boesch. Scientific representation and dissimilarity. Synthese, 198(6):5495–5513, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-019-02417-0.

Pritchard:2021:IVE

[835] Duncan Pritchard. Intellectual virtues and the epistemic value of truth. Synthese, 198(6):5515–5528, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02418-z.

Zubcic:2021:INI

[836] Marko-Luka Zubcić. Ignorance, norms and instrumental pluralism: Hayekian institutional epistemology. Synthese, 198(6):5529–5545, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02420-5.

Eder:2021:FOM

[837] Günther Eder. Frege and the origins of model theory in nineteenth century geometry. Synthese, 198(6):5547–5575, June 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02421-4.

Sperry-Taylor:2021:REE

[838] Ashton T. Sperry-Taylor. Reassessing equilibrium explanations: When are they causal explanations? Synthese, 198(6):5577–5598, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02423-

Mohseni:2021:EMD

[839] Aydin Mohseni, Cailin O’Connor, and Hannah Rubin. On the emergence of minority disadvantage: testing the cultural Red King hypothesis. Synthese, 198(6):5599–5621, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02424-1.

Solomon:2021:CDT

[840] Toby Charles Penhallurick Solomon. Causal decision theory’s predetermination problem. Synthese, 198(6):5623–5654, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02425-0.

Wirling:2021:IDH

[841] Ylwa Sjölin Wirling. An integrative design? How liberalised modal empiricism fails the integration challenge. Synthese, 198(6):5655–5673, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02426-z.

Antonsen:2021:FR

[842] Paal Fjeldvig Antonsen. Fictional reports. Synthese, 198(6):5675–5688, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02427-y.

Harrison:2021:STM

[843] Eilidh Harrison. Seemings, truth-makers, and epistemic justification. Synthese, 198(6):5689–5708, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02428-x.

Diaz:2021:YJB

[844] Rodrigo Díaz and Manuel Almagro. You are just being emotional! Testimonial injustice and folk-psychological attributions. Synthese, 198(6): 5709–5730, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02429-w.

Miller:2021:BTW

[845] J. T. M. Miller. A bundle theory of words. Synthese, 198(6):5731– 5748, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02430-3.

Rich:2021:NTA

[846] Patricia Rich, Mark Blokpoel, Maria Otworowska, Marieke Sweers, Todd Wareham, and Iris van Rooij. Naturalism, tractability and the adaptive toolbox. Synthese, 198(6):5749–5784, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02431-2.

Alksnis:2021:RBG

[847] Nikolai Alksnis and Jack Reynolds. Revaluing the behaviorist ghost in enactivism and embodied cognition. Synthese, 198(6):5785–5807, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02432-1.

Rubin:2021:WTT

[848] Mark Rubin. What type of type I error? Contrasting the Neyman– Pearson and Fisherian approaches in the context of exact and direct replications. Synthese, 198(6):5809–5834, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02433-0.

Wahlberg:2021:CIR

[849] Tobias Hansson Wahlberg. The creation of institutional reality, special theory of relativity, and mere Cambridge change. Synthese, 198 (6):5835–5860, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02435-y.

Salimkhani:2021:EUP

[850] Kian Salimkhani. Explaining unification in physics internally. Synthese, 198(6):5861–5882, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02436-x.

Hu:2021:WSU

[851] Xingming Hu. In what sense is understanding an intellectual virtue? Synthese, 198(6):5883–5895, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02437-w.

Shumener:2021:HW

[852] Erica Shumener. Humeans are out of this world. Synthese, 198(6): 5897–5916, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02439-8.

Kubyshkina:2021:LFI

[853] Ekaterina Kubyshkina and Mattia Petrolo. A logic for factive ignorance. Synthese, 198(6):5917–5928, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02440-1.

Soe:2021:UAI

[854] Sille Obelitz Søe. A unified account of information, misinformation, and disinformation. Synthese, 198(6):5929–5949, June 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02444-x.

Pedersen:2021:CNR

[855] Nikolaj Jang Lee Linding Pedersen. Correction to: Non-rational action in the face of disagreement: an argument against (strong) nonconformism. Synthese, 198(6):5951, June 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-018-01964-2. See [8].

Murphy:2021:EAA

[856] T. G. Murphy. An encoding approach to Ante Rem structuralism. Synthese, 198(7):5953–5976, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02185-x.

Motchoulski:2021:ADD

[857] Alexander Motchoulski. Adjudicating distributive disagreement. Synthese, 198(7):5977–6008, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02443-y.

Witteveen:2021:EDA

[858] Joeri Witteveen. Evolutionary debunking arguments and the explanatory scope of natural selection. Synthese, 198(7):6009–6024, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02446-9.

Asper:2021:SVS

[859] Jon Marc Asper. Subjective values should be sharp. Synthese, 198(7): 6025–6043, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02448-7.

Fritz:2021:SPA

[860] Peter Fritz. Structure by proxy, with an application to grounding. Synthese, 198(7):6045–6063, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02450-z.

Fazekas:2021:DSA

[861] Peter Fazekas, Balázs Gyenis, and Gergely Kertész. A dynamical systems approach to causation. Synthese, 198(7):6065–6087, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02451-y.

Almang:2021:TEP

[862] Jan Almäng. Twin Earth and perceptual content. Synthese, 198(7): 6089–6109, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02452-x.

Barkasi:2021:TEC

[863] Michael Barkasi. Are there epistemic conditions necessary for demonstrative thought? Synthese, 198(7):6111–6138, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02453-w.

Tee:2021:MGM

[864] Sim-Hui Tee. Mechanisms and generative material models. Synthese, 198 (7):6139–6157, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02454-9.

Stefansson:2021:AAB

[865] H. Orri Stefánsson. Ambiguity aversion behind the veil of ignorance. Synthese, 198(7):6159–6182, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02455-8.

Simmons:2021:FNQ

[866] Byron Simmons. Fundamental non-qualitative properties. Synthese, 198 (7):6183–6206, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02458-5.

Witek:2021:IAF

[867] Maciej Witek. Illocution and accommodation in the functioning of presumptions. Synthese, 198(7):6207–6244, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02459-4.

Esfeld:2021:SHF

[868] Michael Esfeld. Super-Humeanism and free will. Synthese, 198(7): 6245–6258, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02460-x.

Dragos:2021:EAG

[869] Chris Dragos. Epistemic autonomy and group knowledge. Synthese, 198 (7):6259–6279, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02461-w.

Marschall:2021:EOQ

[870] Benjamin Marschall. Easy Ontology, quantification, and realism. Synthese, 198(7):6281–6295, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02463-8.

Werner:2021:CCT

[871] Konrad Werner. Cognitive confinement: theoretical considerations on the construction of a cognitive niche, and on how it can go wrong. Synthese, 198(7):6297–6328, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02464-7.

Lassiter:2021:ABC

[872] Charles Lassiter. Arational belief convergence. Synthese, 198(7): 6329–6350, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02465-6.

Lavelle:2021:ICM

[873] Jane Suilin Lavelle. The impact of culture on mindreading. Synthese, 198 (7):6351–6374, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02466-5.

Tappenden:2021:ETP

[874] Paul Tappenden. Everettian theory as pure wave mechanics plus a nocollapse probability postulate. Synthese, 198(7):6375–6402, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02467-
4.

Fisher:2021:CNM

[875] David Fisher, Hao Hong, and Timothy Perrine. A challenge to the new metaphysics: deRosset, priority, and explanation. Synthese, 198 (7):6403–6433, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02468-3.

Zahle:2021:LLM

[876] Julie Zahle. Limits to levels in the methodological individualism-holism debate. Synthese, 198(7):6435–6454, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02469-2.

Fabry:2021:GIP

[877] Regina E. Fabry. Getting it: A predictive processing approach to irony comprehension. Synthese, 198(7):6455–6489, July 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02470-9.

Mays:2021:IPR

[878] Chris Mays. Ignorance as a productive response to epistemic perturbations. Synthese, 198(7):6491–6507, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02471-8.

Aas:2021:PE

[879] Sean Aas. Prosthetic embodiment. Synthese, 198(7):6509–6532, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02472-7.

Valtonen:2021:GIR

[880] Pasi Valtonen. Generic inferential rules for slurs: Dummett and Williamson on ethnic pejoratives. Synthese, 198(7):6533–6551, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-019-02473-
6.

Currie:2021:ASS

[881] Greg Currie and Xuanqi Zhu. Aesthetic sense and social cognition: a story from the Early Stone Age. Synthese, 198(7):6553–6572, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02476-3.

Zawidzki:2021:NPR

[882] Tadeusz W. Zawidzki. A new perspective on the relationship between metacognition and social cognition: metacognitive concepts as sociocognitive tools. Synthese, 198(7):6573–6596, July 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02477-2.

Bedessem:2021:TCS

[883] Baptiste Bedessem. Two conceptions of the sources of conservatism in scientific research. Synthese, 198(7):6597–6614, July 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02479-0.

Gert:2021:TRV

[884] Joshua Gert. Transparency, representationalism, and visual noise. Synthese, 198(7):6615–6629, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02480-7.

Rellihan:2021:SEA

[885] Matthew Rellihan. Strengthening the exclusion argument. Synthese, 198 (7):6631–6659, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02481-6.

deRonde:2021:NOD

[886] C. de Ronde and C. Massri. A new objective definition of quantum entanglement as potential coding of intensive and effective relations. Synthese, 198(7):6661–6688, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02482-5.

Bourrat:2021:HCI

[887] Pierrick Bourrat. Heritability, causal influence and locality. Synthese, 198 (7):6689–6715, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02484-3.

Nolfi:2021:ENA

[888] Kate Nolfi. Epistemic norms, all things considered. Synthese, 198(7): 6717–6737, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02485-2.

Staffel:2021:NUP

[889] Julia Staffel. Normative uncertainty and probabilistic moral knowledge. Synthese, 198(7):6739–6765, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02486-1.

Thoma:2021:JAN

[890] Johanna Thoma. Judgementalism about normative decision theory. Synthese, 198(7):6767–6787, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02487-0.

Ervas:2021:MIS

[891] Francesca Ervas. Metaphor, ignorance and the sentiment of (ir)rationality. Synthese, 198(7):6789–6813, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02489-y.

Fiebich:2021:DPT

[892] Anika Fiebich. In defense of pluralist theory. Synthese, 198(7):6815– 6834, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-019-02490-5. See correction [893].

Fiebich:2021:CDP

[893] Anika Fiebich. Correction to: In defense of pluralist theory. Synthese, 198(7):6835, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02552-z. See [892].

Boyd:2021:GU

[894] Kenneth Boyd. Group understanding. Synthese, 198(7):6837–6858, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02492-3.

Lawler:2021:SUF

[895] Insa Lawler. Scientific understanding and felicitous legitimate falsehoods. Synthese, 198(7):6859–6887, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02495-0.

Lazar:2021:DDT

[896] Seth Lazar and Peter A. Graham. Deontological decision theory and lesser-evil options. Synthese, 198(7):6889–6916, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02496-z.

Vazard:2021:RDA

[897] Juliette Vazard. (Un)reasonable doubt as affective experience: obsessivecompulsive disorder, epistemic anxiety and the feeling of uncertainty. Synthese, 198(7):6917–6934, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02497-y.

Hutchinson:2021:WCW

[898] Jim Hutchinson. Why can’t what is true be valuable? Synthese, 198 (7):6935–6954, July 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02499-w.

Shupe:2021:PD

[899] Eli Shupe. Perceiving deviance. Synthese, 198(8):6955–6967, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02493-2.

Zhao:2021:BVT

[900] Haicheng Zhao. Better virtuous than safe. Synthese, 198(8):6969– 6991, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02501-5.

Ogaard:2021:NBCa

[901] Tore Fjetland Øgaard. Non-Boolean classical relevant logics I. Synthese, 198(8):6993–7024, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02507-z.

Goldhaber:2021:DPP

[902] Charles Goldhaber. Does perceptual psychology rule out disjunctivism in the theory of perception? Synthese, 198(8):7025–7047, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02508-y.

Ivanov:2021:PST

[903] Ivan V. Ivanov. Properties in sight and in thought. Synthese, 198(8): 7049–7071, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02509-x.

Dicher:2021:RLN

[904] Bogdan Dicher. Requiem for logical nihilism, or: Logical nihilism annihilated. Synthese, 198(8):7073–7096, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02510-4.

Boisserie-Lacroix:2021:ERU

[905] Celine Boisserie-Lacroix and Marco Inchingolo. Empathy for a reason? From understanding agency to phenomenal insight. Synthese, 198(8): 7097–7118, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02511-3.

Rossi:2021:PTM

[906] Mauro Rossi. A perceptual theory of moods. Synthese, 198(8):7119– 7147, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02513-1.

Faller:2021:HGP

[907] August Faller. How groups persist. Synthese, 198(8):7149–7163, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02514-0.

McWilliams:2021:EBP

[908] Emily C. McWilliams. Evidentialism and belief polarization. Synthese, 198(8):7165–7196, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-019-02515-z.

Roche:2021:DDH

[909] William Roche and Elliott Sober. Disjunction and distality: the hard problem for purely probabilistic causal theories of mental content. Synthese, 198(8):7197–7230, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02516-y.

Bricker:2021:NPN

[910] Adam Michael Bricker. Neural phase: a new problem for the modal account of epistemic luck. Synthese, 198(8):7231–7248, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02517-x.

Brincus:2021:OER

[911] Constantin C. Brîncus. Are the open-ended rules for negation categorical? Synthese, 198(8):7249–7256, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02519-9.

Weatherall:2021:CSN

[912] James Owen Weatherall and Cailin O’Connor. Conformity in scientific networks. Synthese, 198(8):7257–7278, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02520-2.

Crowther:2021:BBS

[913] Karen Crowther. As below, so before: ‘synchronic’ and ‘diachronic’ conceptions of spacetime emergence. Synthese, 198(8):7279–7307, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02521-1.

Meacham:2021:TMG

[914] Christopher J. G. Meacham. Too much of a good thing: decision-making in cases with infinitely many utility contributions. Synthese, 198(8):

7309–7349, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02522-0.

Poggiolesi:2021:GPR

[915] Francesca Poggiolesi. Grounding principles for (relevant) implication. Synthese, 198(8):7351–7376, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02523-z.

Avigad:2021:RMI

[916] Jeremy Avigad. Reliability of mathematical inference. Synthese, 198(8): 7377–7399, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02524-y.

Khoo:2021:DAC

[917] Justin Khoo. Disjunctive antecedent conditionals. Synthese, 198(8): 7401–7430, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-018-1877-6.

Brenner:2021:MI

[918] Andrew Brenner. Mereology and ideology. Synthese, 198(8):7431– 7448, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02527-9.

Eder:2021:NCT

[919] Anna-Maria A. Eder. No commitment to the truth. Synthese, 198(8): 7449–7472, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02528-8.

Dellantonio:2021:IMC

[920] Sara Dellantonio and Luigi Pastore. Ignorance, misconceptions and critical thinking. Synthese, 198(8):7473–7501, August 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02529-7.

Magnani:2021:CDI

[921] Lorenzo Magnani. Computational domestication of ignorant entities. Synthese, 198(8):7503–7532, August 2021. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02530-5.

Chen:2021:ILD

[922] Lu Chen. Intrinsic local distances: a mixed solution to Weyl’s tile argument. Synthese, 198(8):7533–7552, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02531-4.

Goldschmidt:2021:IVR

[923] Zeev Goldschmidt and Ittay Nissan-Rozen. The intrinsic value of risky prospects. Synthese, 198(8):7553–7575, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02532-3.

Datteri:2021:IB

[924] Edoardo Datteri. Interactive biorobotics. Synthese, 198(8):7577–7595, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02533-2.

Castro:2021:JAC

[925] Víctor Fernández Castro and Elisabeth Pacherie. Joint actions, commitments and the need to belong. Synthese, 198(8):7597–7626, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02535-0.

Hedoin:2021:PRR

[926] Cyril Hédoin. Practical reasoning, rule-following and belief revision: an account in terms of Jeffrey’s rule. Synthese, 198(8):7627–7645, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02536-z.

Nirshberg:2021:SIR

[927] Gregory Nirshberg and Lawrence Shapiro. Structural and indicator representations: a difference in degree, not kind. Synthese, 198(8): 7647–7664, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02537-y.

Havstad:2021:CBC

[928] Joyce C. Havstad. Complexity begets crosscutting, dooms hierarchy (another paper on natural kinds). Synthese, 198(8):7665–7696, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02539-w.

Schippers:2021:LC

[929] Michael Schippers and Mark Siebel. The logic of coherence. Synthese, 198 (8):7697–7714, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02542-1.

Bovey:2021:EMI

[930] Gaétan Bovey. Essence, modality, and intrinsicality. Synthese, 198(8): 7715–7737, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02545-y.

Kyle:2021:TI

[931] Brent G. Kyle. Truth and ignorance. Synthese, 198(8):7739–7762, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02546-x.

Bales:2021:SS

[932] Adam Bales and Claire Benn. Supererogation and sequence. Synthese, 198(8):7763–7780, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02547-w.

Dolega:2021:FPB

[933] Krzysztof Dołega and Joe E. Dewhurst. Fame in the predictive brain: a deflationary approach to explaining consciousness in the prediction error minimization framework. Synthese, 198(8):7781–7806, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02548-9.

Williams:2021:MIR

[934] Daniel Williams. Motivated ignorance, rationality, and democratic politics. Synthese, 198(8):7807–7827, August 2021. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02549-8.

Stokke:2021:FNI

[935] Andreas Stokke. Fictional names and individual concepts. Synthese, 198 (8):7829–7859, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02550-1.

Bjorkholm:2021:QRN

[936] Stina Björkholm, Krister Bykvist, and Jonas Olson. Quasi-realism and normative certitude. Synthese, 198(8):7861–7869, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02553-y.

Smortchkova:2021:AER

[937] Joulia Smortchkova. After-effects and the reach of perceptual content. Synthese, 198(8):7871–7890, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02554-x.

Engel:2021:LKI

[938] Mylan Engel, Jr. Lotteries, knowledge, and inconsistent belief: why you know your ticket will lose. Synthese, 198(8):7891–7921, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02555-w.

Baumberger:2021:REU

[939] Christoph Baumberger and Georg Brun. Reflective equilibrium and understanding. Synthese, 198(8):7923–7947, August 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02556-9.

Peels:2021:EI

[940] Rik Peels and Duncan Pritchard. Educating for ignorance. Synthese, 198 (8):7949–7963, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02544-z.

Ruyant:2021:SRS

[941] Quentin Ruyant. Semantic realism in the semantic conception of theories. Synthese, 198(8):7965–7983, August 2021. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02557-8.

Heidemann:2021:CKF

[942] Dietmar Heidemann. Correction to: Kant and the forms of realism. Synthese, 198(8):7985, August 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02534-1. See [437].

Aktunc:2021:PTL

[943] M. Emrah Aktunc. Productive theory-ladenness in fMRI. Synthese, 198(9):7987–8003, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02125-9.

Rigato:2021:OUF

[944] Joana Rigato, Scott M. Rennie, and Zachary F. Mainen. The overlooked ubiquity of first-person experience in the cognitive sciences. Synthese, 198(9):8005–8041, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02136-6.

Huang:2021:CMM

[945] Linus Ta-Lun Huang. Can massive modularity explain human intelligence? Information control problem and implications for cognitive architecture. Synthese, 198(9):8043–8072, September 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02215-8.

Turri:2021:TFJ

[946] John Turri. Truth, fallibility, and justification: new studies in the norms of assertion. Synthese, 198(9):8073–8084, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02558-7.

Robb:2021:TD

[947] Catherine M. Robb. Talent dispositionalism. Synthese, 198(9):8085– 8102, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02559-6.

Read:2021:GTA

[948] James Read and Emily Qureshi-Hurst. Getting tense about relativity. Synthese, 198(9):8103–8125, September 2021. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02560-z.

Valente:2021:QSB

[949] Giovanni Valente. Quantum symmetry breaking and physical inequivalence: the case of ferromagnetism. Synthese, 198(9):8127–8148, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02561-y.

Merlo:2021:RRS

[950] Giovanni Merlo and Giulia Pravato. Relativism, realism, and subjective facts. Synthese, 198(9):8149–8165, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02562-x.

Brocker:2021:JEU

[951] Karen Brøcker. Justifying the evidential use of linguistic intuitions. Synthese, 198(9):8167–8189, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02563-w.

Chellappoo:2021:RPB

[952] Azita Chellappoo. Rethinking prestige bias. Synthese, 198(9):8191– 8212, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02565-8.

Westra:2021:FPP

[953] Evan Westra. Folk personality psychology: mindreading and mindshaping in trait attribution. Synthese, 198(9):8213–8232, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02566-7.

Knuuttila:2021:SFT

[954] Tarja Knuuttila and Rami Koskinen. Synthetic fictions: turning imagined biological systems into concrete ones. Synthese, 198(9):8233– 8250, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02567-6.

Cobreros:2021:CGN

[955] Pablo Cobreros, Elio La Rosa, and Luca Tranchini. (I can’t get no) antisatisfaction. Synthese, 198(9):8251–8265, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02570-x.

Dinges:2021:KNT

[956] Alexander Dinges. Knowledge and non-traditional factors: prospects for doxastic accounts. Synthese, 198(9):8267–8288, September 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02572-
9.

Gressis:2021:BNE

[957] Robert Gressis. Broad and narrow epistemic standing: its relevance to the epistemology of disagreement. Synthese, 198(9):8289–8306, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02573-8.

Fusco:2021:TDL

[958] Melissa Fusco. A two-dimensional logic for diagonalization and the a priori. Synthese, 198(9):8307–8322, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02574-7.

Munroe:2021:RRR

[959] Wade Munroe. Reasoning, rationality, and representation. Synthese, 198(9):8323–8345, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02575-6.

Coates:2021:MSP

[960] Ashley Coates. Making sense of powerful qualities. Synthese, 198(9): 8347–8363, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02576-5.

Fenici:2021:OMH

[961] Marco Fenici and Tadeusz Wiesław Zawidzki. The origins of mindreading: how interpretive socio-cognitive practices get off the ground. Synthese, 198(9):8365–8387, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02577-4.

vanDijk:2021:DPC

[962] Ludger van Dijk and Julian Kiverstein. Direct perception in context: radical empiricist reflections on the medium. Synthese, 198(9):8389– 8411, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02578-3.

Joyce:2021:CDC

[963] James M. Joyce. Conditional desirability: comments on Richard Bradley’s decision theory with a human face. Synthese, 198(9):8413– 8431, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02579-2.

Dawson:2021:HP

[964] Patrick Dawson. Hard presentism. Synthese, 198(9):8433–8461, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02580-9.

Ilgaz:2021:CCT

[965] Hande Ilgaz and Jedediah W. P. Allen. (Co-)Constructing a theory of mind: From language or through language? Synthese, 198(9):8463– 8484, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02581-8.

Bigaj:2021:DSL

[966] Tomasz Bigaj. On discernibility in symmetric languages: the case of quantum particles. Synthese, 198(9):8485–8502, September 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02582-
7.

Danon:2021:CA

[967] Laura Danón. The content of aliefs. Synthese, 198(9):8503–8520, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-020-02583-6.

Weissglass:2021:BET

[968] D. E. Weissglass. Is belief evaluation truth sensitive? A reply to Turri. Synthese, 198(9):8521–8532, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02584-5.

Buskell:2021:CNI

[969] Andrew Buskell. Cognitive novelties, informational form, and structuralcausal explanations. Synthese, 198(9):8533–8553, September 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02585-
4.

Boge:2021:CAA

[970] Florian J. Boge. Correction to: An argument against global no miracles arguments. Synthese, 198(9):8555, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02586-3. See [215].

Sterelny:2021:DCC

[971] Kim Sterelny. Demography and cultural complexity. Synthese, 198(9): 8557–8580, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02587-2.

Artiga:2021:SMC

[972] Marc Artiga. Signals are minimal causes. Synthese, 198(9):8581– 8599, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02589-0.

McCormick:2021:VBT

[973] Miriam Schleifer McCormick. Value beyond truth-value: a practical response to skepticism. Synthese, 198(9):8601–8619, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02590-7.

DAlessandro:2021:PQR

[974] William D’Alessandro. Proving quadratic reciprocity: explanation, disagreement, transparency and depth. Synthese, 198(9):8621–8664, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-020-02591-6.

Martinez-Ordaz:2021:IBI

[975] María del Rosario Martínez-Ordaz. The ignorance behind inconsistency toleration. Synthese, 198(9):8665–8686, September 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02593-4.

Santos:2021:ISE

[976] Gil Santos. Integrated-structure emergence and its mechanistic explanation. Synthese, 198(9):8687–8711, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02594-3.

Carnino:2021:HLC

[977] Pablo Carnino. Humean laws, circularity, and contrastivity. Synthese, 198(9):8713–8731, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02596-1.

Tamminga:2021:ERD

[978] Allard Tamminga, Hein Duijf, and Frederik Van De Putte. Expressivity results for deontic logics of collective agency. Synthese, 198(9):8733– 8753, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02597-0.

Paoletti:2021:ESP

[979] Michele Paolini Paoletti. Emergence and structural properties. Synthese, 198(9):8755–8778, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02599-y.

Heathcote:2021:MI

[980] Adrian Heathcote. Multiplicity and indiscernibility. Synthese, 198(9): 8779–8808, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02600-8.

Battich:2021:JAP

[981] Lucas Battich and Bart Geurts. Joint attention and perceptual experience. Synthese, 198(9):8809–8822, September 2021. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02602-6.

Ridge:2021:IG

[982] Michael Ridge. Individuating games. Synthese, 198(9):8823–8850, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-020-02603-5.

Egeland:2021:AOA

[983] Jonathan Egeland. Against overconfidence: arguing for the accessibility of memorial justification. Synthese, 198(9):8851–8871, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02604-4.

Fletcher:2021:SSD

[984] Samuel C. Fletcher. Similarity structure and diachronic emergence. Synthese, 198(9):8873–8900, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02605-3.

Roberts:2021:PPP

[985] Paul Roberts. Presumptuous or pluralistic presumptions of innocence? Methodological diagnosis towards conceptual reinvigoration. Synthese, 198(9):8901–8932, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02606-2.

Kovacs:2021:OSC

[986] David Mark Kovacs. The oldest solution to the circularity problem for Humeanism about the laws of nature. Synthese, 198(9):8933– 8953, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02608-0.

Stovall:2021:RAO

[987] Preston Stovall. Rationality, autonomy, and obedience to linguistic norms. Synthese, 198(9):8955–8980, September 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02609-z.

Mahtani:2021:AGD

[988] Anna Mahtani. Awareness growth and dispositional attitudes. Synthese, 198(9):8981–8997, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02611-5.

Conix:2021:ANK

[989] Stijn Conix and Pei-Shan Chi. Against natural kind eliminativism. Synthese, 198(9):8999–9020, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02614-2.

vanDijk:2021:TOC

[990] Ludger van Dijk. Temporalizing ontology: a case for pragmatic emergence. Synthese, 198(9):9021–9034, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02615-1.

Leuridan:2021:DCC

[991] Bert Leuridan and Thomas Lodewyckx. Diachronic causal constitutive relations. Synthese, 198(9):9035–9065, September 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02616-0.

Filomeno:2021:CNA

[992] Aldo Filomeno. Correction to: Are non-accidental regularities a cosmic coincidence? Revisiting a central threat to Humean laws. Synthese, 198 (9):9067, September 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02564-9. See [819].

Martin:2021:ILE

[993] Ben Martin. Identifying logical evidence. Synthese, 198(10):9069– 9095, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02618-y.

Zhao:2021:SRS

[994] Kino Zhao. Sample representation in the social sciences. Synthese, 198 (10):9097–9115, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02621-3.

Rinner:2021:MRT

[995] Stefan Rinner. The multiple relation theory and Schiffer’s puzzle. Synthese, 198(10):9117–9137, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02623-1.

Castroviejo:2021:MPL

[996] Elena Castroviejo, Katherine Fraser, and Agustín Vicente. More on pejorative language: insults that go beyond their extension. Synthese, 198 (10):9139–9164, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02624-0.

Parvizian:2021:SDC

[997] Saja Parvizian. Scientia, diachronic certainty, and virtue. Synthese, 198(10):9165–9192, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02626-y.

Loss:2021:CIP

[998] Roberto Loss. Composition, identity and plural ontology. Synthese, 198(10):9193–9210, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02628-w.

Watson:2021:EGF

[999] David S. Watson and Luciano Floridi. The explanation game: a formal framework for interpretable machine learning. Synthese, 198(10): 9211–9242, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02629-9.

Carter:2021:EVS

[1000] J. Adam Carter and Robert D. Rupert. Epistemic value in the subpersonal vale. Synthese, 198(10):9243–9272, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02631-1.

Cepollaro:2021:WIO

[1001] Bianca Cepollaro, Filippo Domaneschi, and Isidora Stojanovic. When is it ok to call someone a jerk? An experimental investigation of expressives. Synthese, 198(10):9273–9292, October 2021. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02633-z.

Buckley:2021:VEI

[1002] Daniel Buckley. Varieties of epistemic instrumentalism. Synthese, 198 (10):9293–9313, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02634-y.

Lean:2021:ISN

[1003] Christopher Hunter Lean. Invasive species and natural function in ecology. Synthese, 198(10):9315–9333, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02635-x.

Osborne:2021:SSP

[1004] Robert Carry Osborne. A social solution to the puzzle of doxastic responsibility: a two-dimensional account of responsibility for belief. Synthese, 198(10):9335–9356, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02637-9.

Gentry:2021:MPW

[1005] Brittany A. Gentry. Measuring the present: What is the duration of ‘now’? Synthese, 198(10):9357–9371, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02644-w.

Hicks:2021:WST

[1006] Michael R. Hicks. Wilfrid Sellars and the task of philosophy. Synthese, 198(10):9373–9400, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02646-8.

Gagne-Julien:2021:TSC

[1007] Anne-Marie Gagné-Julien. Towards a socially constructed and objective concept of mental disorder. Synthese, 198(10):9401–9426, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02647-7.

Terzian:2021:SWC

[1008] Giulia Terzian and María Inés Corbalán. Simplicity of what? A case study from generative linguistics. Synthese, 198(10):9427–9452, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02649-5.

Latham:2021:QGT

[1009] Andrew J. Latham and Kristie Miller. Quantum gravity, timelessness, and the folk concept of time. Synthese, 198(10):9453–9478, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02650-y.

Rami:2021:SDF

[1010] Dolf Rami. Single-domain free logic and the problem of compositionality. Synthese, 198(10):9479–9523, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02651-x.

Michaelian:2021:EMI

[1011] Kourken Michaelian. Episodic memory is not immune to error through misidentification: against Fernández. Synthese, 198(10):9525–9543, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02652-w.

Levy:2021:VSV

[1012] Neil Levy. Virtue signalling is virtuous. Synthese, 198(10):9545– 9562, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02653-9.

Cassell:2021:BC

[1013] Lisa Cassell. Bayesian coherentism. Synthese, 198(10):9563–9590, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02656-6.

Yan:2021:IQC

[1014] Karen Yan, Meng-Li Tsai, and Tsung-Ren Huang. Improving the quality of case-based research in the philosophy of contemporary sciences. Synthese, 198(10):9591–9610, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02657-5.

Djordjevic:2021:WCC

[1015] C. M. Djordjevic. When clarity and consistency conflicts with empirical adequacy: conceptual engineering, anthropology, and Evans-pritchard’s ethnography. Synthese, 198(10):9611–9637, October 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02666-4.

Khani:2021:IJD

[1016] Ali Hossein Khani. Interpretationism and judgement-dependence. Synthese, 198(10):9639–9659, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02670-8.

Jaksland:2021:EWM

[1017] Rasmus Jaksland. Entanglement as the world-making relation: distance from entanglement. Synthese, 198(10):9661–9693, October 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02671-
7.

Irmak:2021:PCA

[1018] Nurbay Irmak. The problem of creation and abstract artifacts. Synthese, 198(10):9695–9708, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02672-6.

Scerri:2021:CEC

[1019] Eric R. Scerri. Causation, electronic configurations and the periodic table. Synthese, 198(10):9709–9720, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02674-4.

Werner:2021:GBM

[1020] Jonas Werner. A grounding-based measure of relative fundamentality. Synthese, 198(10):9721–9737, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02676-2.

Rorot:2021:ESP

[1021] Wiktor Rorot. Explaining “spatial purport of perception”: a predictive processing approach. Synthese, 198(10):9739–9762, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02678-0.

Silins:2021:EDL

[1022] Nicholas Silins. The evil demon in the lab: skepticism, introspection, and introspection of introspection. Synthese, 198(10):9763–9785, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02680-6.

Langsam:2021:MIN

[1023] Harold Langsam. McDowell’s infallibilism and the nature of knowledge. Synthese, 198(10):9787–9801, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02682-4.

Ross:2021:DTC

[1024] Lauren N. Ross. Distinguishing topological and causal explanation. Synthese, 198(10):9803–9820, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02685-1.

De:2021:TTS

[1025] Michael De. Truthmakers or truthmaking supervenience? Synthese, 198(10):9821–9839, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02688-y.

Wadle:2021:SMN

[1026] Douglas C. Wadle. Sensory modalities and novel features of perceptual experiences. Synthese, 198(10):9841–9872, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02689-x.

Roberts:2021:TMA

[1027] Alexander Roberts. Two morals about a modal paradox. Synthese, 198 (10):9873–9896, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02693-1.

Dentith:2021:DCT

[1028] M. R. X. Dentith. Debunking conspiracy theories. Synthese, 198(10): 9897–9911, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02694-0.

Fleisher:2021:HEC

[1029] Will Fleisher. How to endorse conciliationism. Synthese, 198(10): 9913–9939, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02695-z.

Johnson:2021:ABI

[1030] Gabbrielle M. Johnson. Algorithmic bias: on the implicit biases of social technology. Synthese, 198(10):9941–9961, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02696-y.

Fraissler:2021:SI

[1031] Hannes Fraissler. Schmidentity and informativity. Synthese, 198(10): 9963–9989, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02698-w. See correction [1079].

Pettigrew:2021:LIL

[1032] Richard Pettigrew. Logical ignorance and logical learning. Synthese, 198(10):9991–10020, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02699-9.

Lopez-Rubio:2021:TLB

[1033] Ezequiel López-Rubio. Throwing light on black boxes: emergence of visual categories from deep learning. Synthese, 198(10):10021–10041, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02700-5.

Palmer:2021:FWC

[1034] David Palmer. Free will and control: a noncausal approach. Synthese, 198(10):10043–10062, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02701-4.

Ruffino:2021:SAM

[1035] Marco Ruffino, Luca San Mauro, and Giorgio Venturi. Speech acts in mathematics. Synthese, 198(10):10063–10087, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02702-3.

Verdejo:2021:PSI

[1036] Víctor M. Verdejo. Perspectives on de se immunity. Synthese, 198(10): 10089–10107, October 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03122-7.

Pezlar:2021:PVA

[1037] Ivo Pezlar. The placeholder view of assumptions and the Curry–Howard correspondence. Synthese, 198(11):10109–10125, November 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02706-
z.

Fischer:2021:ISI

[1038] Eugen Fischer, Paul E. Engelhardt, and Justin Sytsma. Inappropriate stereotypical inferences? An adversarial collaboration in experimental ordinary language philosophy. Synthese, 198(11):10127–10168, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-020-02708-x.

Genot:2021:HLS

[1039] Emmanuel J. Genot. The Holmesian logician: Sherlock Holmes’ “Science of Deduction and Analysis” and the logic of discovery. Synthese, 198(11):10169–10186, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02709-w.

Zanetti:2021:GAA

[1040] Luca Zanetti. Grounding and auto-abstraction. Synthese, 198(11):10187– 10205, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02710-3.

Thyssen:2021:DF

[1041] Pieter Thyssen and Sylvia Wenmackers. Degrees of freedom. Synthese, 198(11):10207–10235, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02714-z.

Horden:2021:GP

[1042] John Horden and Dan López de Sa. Groups as pluralities. Synthese, 198(11):10237–10271, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02715-y.

Kugler:2021:WBS

[1043] Peter Kügler. What Bergson should have said about special relativity. Synthese, 198(11):10273–10288, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02716-x.

Rudy-Hiller:2021:DSP

[1044] Fernando Rudy-Hiller. In defense of a strong persistence requirement on intention. Synthese, 198(11):10289–10312, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02719-8.

Murray:2021:CGI

[1045] Samuel Murray and Peter Finocchiaro. These confabulations are guaranteed to improve your marriage! Toward a teleological theory of confabulation. Synthese, 198(11):10313–10339, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02721-0.

Vaassen:2021:CEC

[1046] Bram Vaassen. Causal exclusion without causal sufficiency. Synthese, 198(11):10341–10353, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02723-y.

Giananti:2021:KHK

[1047] Andrea Giananti. I know how I know: perception, self-awareness, selfknowledge. Synthese, 198(11):10355–10375, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02726-9.

Pacchetti:2021:SUT

[1048] Marina Baldissera Pacchetti. Structural uncertainty through the lens of model building. Synthese, 198(11):10377–10393, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02727-8.

Malmgren:2021:GAA

[1049] Anna-Sara Malmgren. Goodness, availability, and argument structure. Synthese, 198(11):10395–10427, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02728-7.

Papale:2021:EMN

[1050] François Papale. Evolution by means of natural selection without reproduction: revamping Lewontin’s account. Synthese, 198(11):10429– 10455, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02729-6.

Marsili:2021:RLE

[1051] Neri Marsili. Retweeting: its linguistic and epistemic value. Synthese, 198(11):10457–10483, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02731-y.

Brauer:2021:DCN

[1052] Ethan Brauer. The dependence of computability on numerical notations. Synthese, 198(11):10485–10511, November 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02732-x.

Kim:2021:LAS

[1053] Namjoong Kim. Limitation of analogy in the Sleeping Beauty debate: the case of Singer’s argument. Synthese, 198(11):10513–10528, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-020-02733-w.

Nichols:2021:RFR

[1054] Cory Nichols. Relevance first: relocating similarity in counterfactual semantics. Synthese, 198(11):10529–10564, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02735-8.

Amilburu:2021:DLE

[1055] Alba Amilburu, Álvaro Moreno, and Kepa Ruiz-Mirazo. Definitions of life as epistemic tools that reflect and foster the advance of biological knowledge. Synthese, 198(11):10565–10585, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02736-7.

Romano:2021:MFB

[1056] Davide Romano. Multi-field and Bohm’s theory. Synthese, 198(11): 10587–10609, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02737-6.

Landes:2021:VET

[1057] Jürgen Landes. The variety of evidence thesis and its independence of degrees of independence. Synthese, 198(11):10611–10641, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02738-5.

Gibbs:2021:CLV

[1058] Cameron Gibbs. Counterfactuals and laws with violations. Synthese, 198(11):10643–10659, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02740-x.

Melchior:2021:MTD

[1059] Guido Melchior. A modal theory of discrimination. Synthese, 198 (11):10661–10684, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02747-4.

Meloni:2021:TEG

[1060] Maurizio Meloni and Jack Reynolds. Thinking embodiment with genetics: epigenetics and postgenomic biology in embodied cognition and enactivism. Synthese, 198(11):10685–10708, November 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02748-3. See cor-

rection [1407].

Shardlow:2021:EPB

[1061] Jack Shardlow, Ruth Lee, and Alison S. Fernandes. Exploring people’s beliefs about the experience of time. Synthese, 198(11):10709– 10731, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02749-2.

Ferretti:2021:AIM

[1062] Gabriele Ferretti. Anti-intellectualist motor knowledge. Synthese, 198 (11):10733–10763, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02750-9.

Aranyosi:2021:PMP

[1063] István Aranyosi. Preteriception: memory as past-perception. Synthese, 198(11):10765–10792, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02751-8.

Yaghmaie:2021:DQA

[1064] Aboutorab Yaghmaie. Deformation quantization as an appropriate guide to ontic structure. Synthese, 198(11):10793–10815, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02752-7.

Coffey:2021:OPM

[1065] Kevin Coffey. On the ontology of particle mass and energy in special relativity. Synthese, 198(11):10817–10846, November 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02754-5.

Velasco:2021:AEP

[1066] Pablo Fernandez Velasco and Slawa Loev. Affective experience in the predictive mind: a review and new integrative account. Synthese, 198 (11):10847–10882, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02755-4.

Himelright:2021:PPP

[1067] Jack Himelright. Paraphrasing away properties with pluriverse counterfactuals. Synthese, 198(11):10883–10902, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02757-2.

Banick:2021:MAP

[1068] Kyle Banick. A modal analysis of phenomenal intentionality: horizonality and object-directed phenomenal presence. Synthese, 198(11):10903– 10922, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02758-1.

Hansen:2021:SAG

[1069] Casper Storm Hansen. The signalman against the glut and gap theorists. Synthese, 198(11):10923–10937, November 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02759-0.

Shogenji:2021:PPS

[1070] Tomoji Shogenji. Probability and proximity in surprise. Synthese, 198 (11):10939–10957, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02761-6.

Schwartz:2021:ESA

[1071] Jeremy Schwartz and Christopher Hom. Embedding speech-act propositions. Synthese, 198(11):10959–10977, November 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02763-4.

Longenecker:2021:CPM

[1072] Michael Tze-Sung Longenecker. Is consequentialist perdurantism in moral trouble? Synthese, 198(11):10979–10990, November 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02764-
3.

Bonino:2021:LAP

[1073] Guido Bonino, Paolo Maffezioli, and Paolo Tripodi. Logic in analytic philosophy: a quantitative analysis. Synthese, 198(11):10991– 11028, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02770-5.

Ben-Yami:2021:BFN

[1074] Hanoch Ben-Yami. The Barcan formulas and necessary existence: the view from Quarc. Synthese, 198(11):11029–11064, November 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02771-

Cohen:2021:PPN

[1075] Michael Cohen. The problem of perception and the no-miracles principle. Synthese, 198(11):11065–11080, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02772-3.

Dotan:2021:TCN

[1076] Ravit Dotan. Theory choice, non-epistemic values, and machine learning. Synthese, 198(11):11081–11101, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02773-2.

Sridharan:2021:TFR

[1077] Vishnu Sridharan. Two faces of rationality. Synthese, 198(11):11103– 11124, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02774-1.

Deutsch:2021:CAC

[1078] Max Deutsch. Conceptual analysis without concepts. Synthese, 198 (11):11125–11157, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02775-0.

Fraissler:2021:CSI

[1079] Hannes Fraissler. Correction to: Schmidentity and informativity. Synthese, 198(11):11159, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02818-6. See [1031].

Amaya:2021:H

[1080] Santiago Amaya. Out of habit. Synthese, 198(12):11161–11185, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02780-3.

Tugby:2021:GTP

[1081] Matthew Tugby. Grounding theories of powers. Synthese, 198(12):11187– 11216, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02781-2.

Boucher:2021:EDA

[1082] Sandy C. Boucher. Evolutionary debunking arguments, commonsense and scepticism. Synthese, 198(12):11217–11239, December 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02782-
1.

Galeazzi:2021:ERD

[1083] Paolo Galeazzi and Alessandro Galeazzi. The ecological rationality of decision criteria. Synthese, 198(12):11241–11264, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02785-y.

Pruss:2021:ADB

[1084] Alexander R. Pruss. Avoiding Dutch Books despite inconsistent credences. Synthese, 198(12):11265–11289, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02786-x.

Nolfi:2021:EFF

[1085] Kate Nolfi. Epistemically flawless false beliefs. Synthese, 198(12):11291– 11309, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02787-w.

Lam:2021:BPA

[1086] Derek Lam. Being pessimistic about the objective present. Synthese, 198(12):11311–11326, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02789-8.

Latham:2021:FBA

[1087] Andrew J. Latham, Kristie Miller, and Christian Tarsney. Future bias in action: does the past matter more when you can affect it? Synthese, 198(12):11327–11349, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02791-0.

Gamester:2021:STP

[1088] Will Gamester. Shopping for truth pluralism. Synthese, 198(12):11351– 11377, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02792-z.

Andreas:2021:ADC

[1089] Holger Andreas. An analysis of disposition concepts by the Ramsey test. Synthese, 198(12):11379–11397, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02794-x.

Kostochka:2021:WMC

[1090] Tatyana A. Kostochka. Why moods change: their appropriateness and connection to beliefs. Synthese, 198(12):11399–11420, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02795-w.

Taggart:2021:RPW

[1091] Christopher P. Taggart. A response to the problem of wild coincidences. Synthese, 198(12):11421–11435, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02800-2. See correction

[1092].

Taggart:2021:CRP

[1092] Christopher P. Taggart. Correction to: A response to the problem of wild coincidences. Synthese, 198(12):11437, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02832-8. See

[1091].

Glennan:2021:CC

[1093] Stuart Glennan. Corporeal composition. Synthese, 198(12):11439– 11462, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02805-x.

Grodniewicz:2021:PLU

[1094] J. P. Grodniewicz. The process of linguistic understanding. Synthese, 198(12):11463–11481, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02807-9.

Kuan:2021:BLC

[1095] Ko-Hung Kuan. Beyond linear conciliation. Synthese, 198(12):11483– 11504, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02809-7.

Taylor:2021:RNF

[1096] David E. Taylor. Reference for neo-Fregeans. Synthese, 198(12):11505– 11536, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02811-z.

Arenhart:2021:EAP

[1097] Jonas Rafael Becker Arenhart. The evidence approach to paraconsistency versus the paraconsistent approach to evidence. Synthese, 198(12):11537– 11559, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02813-x.

Skipper:2021:DRD

[1098] Mattias Skipper. Does rationality demand higher-order certainty? Synthese, 198(12):11561–11585, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02814-w.

Pollock:2021:CIC

[1099] Joey Pollock. Content internalism and conceptual engineering. Synthese, 198(12):11587–11605, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02815-9.

Baras:2021:HCN

[1100] Dan Baras. How can necessary facts call for explanation? Synthese, 198(12):11607–11624, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02817-7.

Kim:2021:HFB

[1101] Junyeol Kim. The horizontal in Frege’s Begriffsschrift. Synthese, 198 (12):11625–11644, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02820-y.

Lloyd:2021:ADA

[1102] Elisabeth A. Lloyd, Melissa Bukovsky, and Linda O. Mearns. An analysis of the disagreement about added value by regional climate models. Synthese, 198(12):11645–11672, December 2021. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02821-x.

Horvath:2021:KN

[1103] Joachim Horvath and Jennifer Nado. Knowledge and normality. Synthese, 198(12):11673–11694, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02823-9.

Jacoby:2021:DIP

[1104] Franklin Jacoby. Data identity and perspectivism. Synthese, 198(12): 11695–11711, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02824-8.

Nielsen:2021:SFC

[1105] Michael Nielsen. The strength of de Finetti’s coherence theorem. (French) []. Synthese, 198(12):11713–11724, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02825-7.

Muhlenbernd:2021:ESA

[1106] Roland Mühlenbernd. Evolutionary stability of ambiguity in context signaling games. Synthese, 198(12):11725–11753, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02826-6.

Tebben:2021:SAK

[1107] Nicholas Tebben. Selfless assertions and the Knowledge Norm. Synthese, 198(12):11755–11774, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02827-5.

Stenwall:2021:GPS

[1108] Robin Stenwall. A grounding physicalist solution to the causal exclusion problem. Synthese, 198(12):11775–11795, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02829-3.

Smart:2021:DDD

[1109] J. A. Smart. Disbelief is a distinct doxastic attitude. Synthese, 198 (12):11797–11813, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02830-w.

Crone:2021:FWP

[1110] Katja Crone. Foundations of a we-perspective. Synthese, 198(12):11815– 11832, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02834-6.

Bain:2021:RFD

[1111] Jonathan Bain. The RT formula and its discontents: spacetime and entanglement. Synthese, 198(12):11833–11860, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02836-4.

Sass:2021:OWE

[1112] Reuben Sass. An ontology of weak entity realism for HPC kinds. Synthese, 198(12):11861–11880, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02838-2.

Shafiei:2021:KHT

[1113] Mohammad Shafiei and Ahmad Ali Akbar Mesgari. On Kant and Husserl on transcendental logic. Synthese, 198(12):11881–11896, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02839-1.

Pace:2021:TOI

[1114] Michael Pace. Trusting in order to inspire trustworthiness. Synthese, 198(12):11897–11923, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02840-8.

Korzukhin:2021:DBC

[1115] Theodore Korzukhin. A Dutch book for CDT thirders. Synthese, 198 (12):11925–11941, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02841-7.

Sebens:2021:PFM

[1116] Charles T. Sebens. Particles, fields, and the measurement of electron spin. Synthese, 198(12):11943–11975, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02843-5.

Westfall:2021:OMN

[1117] Mason Westfall. Other minds are neither seen nor inferred. Synthese, 198(12):11977–11997, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02844-4.

Peters:2021:AEC

[1118] Uwe Peters. An argument for egalitarian confirmation bias and against political diversity in academia. Synthese, 198(12):11999–12019, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02846-2.

Miracchi:2021:CIE

[1119] Lisa Miracchi. A case for integrative epistemology. Synthese, 198(12): 12021–12039, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02848-0.

Bokros:2021:DME

[1120] Sofia Ellinor Bokros. A deference model of epistemic authority. Synthese, 198(12):12041–12069, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02849-z.

McBrayer:2021:EET

[1121] Justin P. McBrayer. Everybody else is thinking it, so why can’t we? Synthese, 198(12):12071–12087, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02850-6.

Wronski:2021:OCP

[1122] Leszek Wroński. Objective consequentialism and the plurality of chances. Synthese, 198(12):12089–12105, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02851-5.

Rivat:2021:ETI

[1123] Sébastien Rivat. Effective theories and infinite idealizations: a challenge for scientific realism. Synthese, 198(12):12107–12136, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02852-4.

Carter:2021:UCT

[1124] J. Adam Carter, Emma C. Gordon, and J. P. Grodniewicz. Understanding a communicated thought. Synthese, 198(12):12137–12151, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02854-2.

McLoone:2021:CCS

[1125] Brian McLoone. Calculus and counterpossibles in science. Synthese, 198(12):12153–12174, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02855-1.

Myin:2021:TTO

[1126] Erik Myin and Jasper C. van den Herik. A twofold tale of one mind: revisiting REC’s multi-storey story. Synthese, 198(12):12175– 12193, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02857-z.

Dougherty:2021:NIT

[1127] John Dougherty. The non-ideal theory of the Aharonov–Bohm effect. Synthese, 198(12):12195–12221, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02859-x.

Ehring:2021:FAH

[1128] Douglas Ehring. Fission and anticipating having an experience. Synthese, 198(12):12223–12234, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02860-4.

Milkowski:2021:ECS

[1129] Marcin Miłkowski and Mateusz Hohol. Explanations in cognitive science: unification versus pluralism. Synthese, 199(S1):1–17, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02777-y.

Harbecke:2021:MRM

[1130] Jens Harbecke. The methodological role of mechanistic-computational models in cognitive science. Synthese, 199(S1):19–41, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02568-
5.

Elber-Dorozko:2021:ICM

[1131] Lotem Elber-Dorozko and Oron Shagrir. Integrating computation into the mechanistic hierarchy in the cognitive and neural sciences. Synthese, 199(S1):43–66, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02230-9.

Milkowski:2021:RUC

[1132] Marcin Miłkowski and Przemysław Nowakowski. Representational unification in cognitive science: Is embodied cognition a unifying perspective? Synthese, 199(S1):67–88, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02445-w.

Erdin:2021:ACM

[1133] Haydar Oguz Erdin. Appraisal of certain methodologies in cognitive science based on Lakatos’s methodology of scientific research programmes. Synthese, 199(S1):89–112, November 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02612-4.

Raja:2021:RRE

[1134] Vicente Raja. Resonance and radical embodiment. Synthese, 199(S1): 113–141, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02610-6.

Carls-Diamante:2021:MYM

[1135] Sidney Carls-Diamante. Make up your mind: octopus cognition and hybrid explanations. Synthese, 199(S1):143–158, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02102-2.

Ritchie:2021:WWM

[1136] J. Brendan Ritchie. What’s wrong with the minimal conception of innateness in cognitive science? Synthese, 199(S1):159–176, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02543-0.

Dupre:2021:LEE

[1137] Gabe Dupre. Linguistics and the explanatory economy. Synthese, 199 (S1):177–219, November 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-019-02290-x.

Crowther:2021:SFG

[1138] Karen Crowther, Niels S. Linnemann, and Christian Wüthrich. Spacetime functionalism in general relativity and quantum gravity. Synthese, 199(S2):221–227, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02722-z.

Read:2021:LIF

[1139] James Read and Tushar Menon. The limitations of inertial frame spacetime functionalism. Synthese, 199(S2):229–251, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02299-2.

Weatherall:2021:TDD

[1140] James Owen Weatherall. Two dogmas of dynamicism. Synthese, 199(S2): 253–275, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02880-0.

Baker:2021:KIS

[1141] David John Baker. Knox’s inertial spacetime functionalism (and a better alternative). Synthese, 199(S2):277–298, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02598-z.

Duerr:2021:AFG

[1142] Patrick M. Duerr. Against ‘functional gravitational energy’: a critical note on functionalism, selective realism, and geometric objects and gravitational energy. Synthese, 199(S2):299–333, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-019-02503-3.

Lam:2021:SFR

[1143] Vincent Lam and Christian Wüthrich. Spacetime functionalism from a realist perspective. Synthese, 199(S2):335–353, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02642-y.

Esfeld:2021:ADS

[1144] Michael Esfeld. Against the disappearance of spacetime in quantum gravity. Synthese, 199(S2):355–369, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02168-y.

LeBihan:2021:SEQ

[1145] Baptiste Le Bihan. Spacetime emergence in quantum gravity: functionalism and the hard problem. Synthese, 199(S2):371–393, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-019-02449-6.

Linnemann:2021:ECS

[1146] Niels Linnemann. On the empirical coherence and the spatiotemporal gap problem in quantum gravity: and why functionalism does not (have to) help. Synthese, 199(S2):395–412, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02659-3.

Baron:2021:EID

[1147] Sam Baron. Empirical incoherence and double functionalism. Synthese, 199(S2):413–439, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-019-02462-9.

Jaksland:2021:MRG

[1148] Rasmus Jaksland. The multiple realizability of general relativity in quantum gravity. Synthese, 199(S2):441–467, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-019-02382-8.

Saad:2021:SES

[1149] Bradford Saad. Spatial experience, spatial reality, and two paths to primitivism. Synthese, 199(S2):469–491, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-019-02410-7.

Zardini:2021:SAP

[1150] Elia Zardini. Substructural approaches to paradox: an introduction to the special issue. Synthese, 199(S3):493–525, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03415-x.

Brady:2021:UDT

[1151] Ross T. Brady. The use of definitions and their logical representation in paradox derivation. Synthese, 199(S3):527–546, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-017-1362-7.

Mares:2021:IIW

[1152] Edwin Mares. An informational interpretation of weak relevant logic and relevant property theory. Synthese, 199(S3):547–569, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-017-1524-7.

TerresVillalonga:2021:SAE

[1153] Pilar Terrés Villalonga. A substructural analysis of embedded conditionals. Synthese, 199(S3):571–595, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02922-7.

Tennant:2021:NCE

[1154] Neil Tennant. Normalizability, cut eliminability and paradox. Synthese, 199(S3):597–616, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-016-1119-8.

Schroeder-Heister:2021:HEC

[1155] Peter Schroeder-Heister and Luca Tranchini. How to Ekman a Crabbé-Tennant. Synthese, 199(S3):617–639, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-018-02018-3.

Zardini:2021:CB

[1156] Elia Zardini. Closed without boundaries. Synthese, 199(S3):641– 679, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02352-0.

Cobreros:2021:TRN

[1157] Pablo Cobreros, Paul Egré, Dave Ripley, and Robert van Rooij. Tolerant reasoning: nontransitive or nonmonotonic? Synthese, 199(S3): 681–705, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-017-1584-8.

Barrio:2021:CNV

[1158] Eduardo Barrio, Lucas Rosenblatt, and Diego Tajer. Capturing naive validity in the Cut-free approach. Synthese, 199(S3):707–723, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-016-1199-5.

Weber:2021:CTS

[1159] Zach Weber. On closure and truth in substructural theories of truth. Synthese, 199(S3):725–739, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-016-1226-6.

Cintula:2021:MCT

[1160] Petr Cintula and Francesco Paoli. Is multiset consequence trivial? Synthese, 199(S3):741–765, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-016-1209-7.

Shapiro:2021:VIS

[1161] Lionel Shapiro. The very idea of a substructural approach to paradox. Synthese, 199(S3):767–786, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-016-1230-x.

Hjortland:2021:TTM

[1162] Ole Thomassen Hjortland. Theories of truth and the maxim of minimal mutilation. Synthese, 199(S3):787–818, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-017-1612-8.

Murzi:2021:NV

[1163] Julien Murzi and Lorenzo Rossi. Naïve validity. Synthese, 199(S3): 819–841, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-017-1541-6.

Jacob:2021:MG

[1164] Pierre Jacob and Thom Scott-Phillips. Is mindreading a gadget? Synthese, 199(1–2):1–27, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02620-4.

Hohwy:2021:SSN

[1165] Jakob Hohwy. Self-supervision, normativity and the free energy principle. Synthese, 199(1–2):29–53, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02622-2.

Ferrari:2021:APV

[1166] Filippo Ferrari. Alethic pluralism and the value of truth. Synthese, 199 (1–2):55–79, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02625-z.

Turner:2021:PCL

[1167] Cameron Rouse Turner and Lachlan Douglas Walmsley. Preparedness in cultural learning. Synthese, 199(1–2):81–100, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02627-x.

Zahle:2021:ODS

[1168] Julie Zahle. Objective data sets in qualitative research. Synthese, 199(1– 2):101–117, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02630-2.

Tommasi:2021:ATS

[1169] Marco Tommasi, Paolo Petricca, and Claudia Casadio. Attitudes towards scientific knowledge: social dispositions and personality traits. Synthese, 199(1–2):119–139, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02632-0.

King:2021:TSP

[1170] Zoë A. Johnson King. The trouble with standards of proof. Synthese, 199(1–2):141–159, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02639-7.

Curry:2021:SS

[1171] Devin Sanchez Curry. Street smarts. Synthese, 199(1–2):161–180, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02641-z.

Hajek:2021:CC

[1172] Alan Hájek. Contra counterfactism. Synthese, 199(1–2):181–210, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02643-x.

Koskinen:2021:OCW

[1173] Inkeri Koskinen. Objectivity in contexts: withholding epistemic judgement as a strategy for mitigating collective bias. Synthese, 199(1–2): 211–225, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02645-9.

Newton:2021:ESM

[1174] Lilith Newton. The epistemic significance of modal factors. Synthese, 199(1–2):227–248, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02648-6.

Papayannopoulos:2021:UMR

[1175] Philippos Papayannopoulos. Unrealistic models for realistic computations: how idealisations help represent mathematical structures and found scientific computing. Synthese, 199(1–2):249–283, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02654-8.

Berio:2021:CES

[1176] Leda Berio. Culturally embedded schemata for false belief reasoning. Synthese, 199(1–2):285–314, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02655-7.

Martens:2021:SAS

[1177] Niels C. M. Martens and James Read. Sophistry about symmetries? Synthese, 199(1–2):315–344, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02658-4.

Morris:2021:IGR

[1178] Rebecca Lea Morris. Intellectual generosity and the reward structure of mathematics. Synthese, 199(1–2):345–367, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02660-w.

Aydinonat:2021:ALF

[1179] N. Emrah Aydinonat, Samuli Reijula, and Petri Ylikoski. Argumentative landscapes: the function of models in social epistemology. Synthese, 199(1–2):369–395, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02661-9.

Lossau:2021:KC

[1180] Tammo Lossau. Knowledge and cancelability. Synthese, 199(1–2):397– 405, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02663-7.

Tanswell:2021:MPE

[1181] Fenner Stanley Tanswell and Ian James Kidd. Mathematical practice and epistemic virtue and vice. Synthese, 199(1–2):407–426, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02664-6.

Steele:2021:HIY

[1182] Katie Steele. How to be imprecise and yet immune to sure loss. Synthese, 199(1–2):427–444, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02665-5.

Boge:2021:PHN

[1183] Florian J. Boge and Christian Zeitnitz. Polycratic hierarchies and networks: what simulation-modeling at the LHC can teach us about the epistemology of simulation. Synthese, 199(1–2):445–480, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/

s11229-020-02667-3. See correction [1679].

Gardiner:2021:RRH

[1184] Georgi Gardiner. Relevance and risk: How the relevant alternatives framework models the epistemology of risk. Synthese, 199(1–2):481– 511, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02668-2.

Rolin:2021:OTS

[1185] Kristina H. Rolin. Objectivity, trust and social responsibility. Synthese, 199(1–2):513–533, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02669-1.

Bjelde:2021:AT

[1186] Joseph Bjelde. Anything but the truth. Synthese, 199(1–2):535– 549, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02673-5.

Thorn:2021:SPC

[1187] Paul D. Thorn. The stimulus-to-perception connection: a simulation study in the epistemology of perception. Synthese, 199(1–2):551– 578, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02677-1.

Rocknak:2021:RCH

[1188] Stefanie Rocknak. Regularity and certainty in Hume’s treatise: a Humean response to Husserl. Synthese, 199(1–2):579–600, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02679-z.

Fano:2021:WHL

[1189] Vincenzo Fano and Pierluigi Graziani. A working hypothesis for the logic of radical ignorance. Synthese, 199(1–2):601–616, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02681-5.

Cohon:2021:HPT

[1190] Rachel Cohon. Hume’s practice theory of promises and its dissimilar descendants. Synthese, 199(1–2):617–635, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02684-2.

Hajek:2021:CSA

[1191] Alan Hájek. Counterfactual scepticism and antecedent-contextualism. Synthese, 199(1–2):637–659, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02686-0.

Drozdzowicz:2021:BBV

[1192] Anna Drozdzowicz. Bringing back the voice: on the auditory objects of speech perception. Synthese, 199(1–2):661–687, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02687-z.

Arfini:2021:IKU

[1193] Selene Arfini and Lorenzo Magnani. Introduction: knowing the unknown. Synthese, 199(1–2):689–693, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02690-4.

Mongin:2021:EUT

[1194] Philippe Mongin and Jean Baccelli. Expected utility theory, Jeffrey’s decision theory, and the paradoxes. Synthese, 199(1–2):695–713, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02691-3.

Pritchard:2021:ALE

[1195] Duncan Pritchard. Anti-luck epistemology and pragmatic encroachment. Synthese, 199(1–2):715–729, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02703-2.

McKenna:2021:AAC

[1196] Robin McKenna and Michael Hannon. Assertion, action, and context. Synthese, 199(1–2):731–743, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02705-0.

Walmsley:2021:SMB

[1197] Lachlan Douglas Walmsley. The strategy of model building in climate science. Synthese, 199(1–2):745–765, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02707-y.

Chirimuuta:2021:PVU

[1198] M. Chirimuuta. Prediction versus understanding in computationally enhanced neuroscience. Synthese, 199(1–2):767–790, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02713-0.

Matarese:2021:SHP

[1199] Vera Matarese. Super-Humeanism and physics: a merry relationship? Synthese, 199(1–2):791–813, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02717-w.

Boehm:2021:HPE

[1200] Miren Boehm. Hume’s “projectivism” explained. Synthese, 199(1–2): 815–833, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02718-9.

Alfano:2021:TSA

[1201] Mark Alfano, Amir Ebrahimi Fard, J. Adam Carter, Peter Clutton, and Colin Klein. Technologically scaffolded atypical cognition: the case of YouTube’s recommender system. Synthese, 199(1–2):835–858, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02724-x.

Andersen:2021:RTM

[1202] Line Edslev Andersen, Hanne Andersen, and Henrik Kragh Sørensen. The role of testimony in mathematics. Synthese, 199(1–2):859–870, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02734-9.

Poirier:2021:NHB

[1203] Pierre Poirier and Luc Faucher. A New Hope: a better ICM to understand human cognitive architectural variability. Synthese, 199(1–2): 871–903, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02739-4.

Segal:2021:HME

[1204] Aaron Segal. Humeanisms: metaphysical and epistemological. Synthese, 199(1–2):905–925, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02741-w.

Sandgren:2021:LCS

[1205] Alexander Sandgren and Katie Steele. Levelling counterfactual scepticism. Synthese, 199(1–2):927–947, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02742-9.

Jellema:2021:RDS

[1206] Hylke Jellema. The reasonable doubt standard as inference to the best explanation. Synthese, 199(1–2):949–973, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02743-8.

Blome-Tillmann:2021:IPE

[1207] Michael Blome-Tillmann. Impurism, pragmatic encroachment, and the Argument from Principles. Synthese, 199(1–2):975–982, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02744-7.

Dardashti:2021:ESA

[1208] Radin Dardashti, Mathias Frisch, and Giovanni Valente. Editorial: Symmetries and asymmetries in physics. Synthese, 199(1–2):983–989, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/

s11229-020-02745-6. See correction [1653].

Trpin:2021:LML

[1209] Borut Trpin, Anna Dobrosovestnova, and Sebastian J. Götzendorfer. Lying, more or less: a computer simulation study of graded lies and trust dynamics. Synthese, 199(1–2):991–1018, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02746-5.

Cochrane:2021:CSC

[1210] Tom Cochrane. A case of shared consciousness. Synthese, 199(1–2): 1019–1037, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02753-6.

McGeer:2021:EFP

[1211] Victoria McGeer. Enculturating folk psychologists. Synthese, 199(1–2): 1039–1063, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02760-7.

Loewer:2021:PDA

[1212] Barry Loewer. The package deal account of laws and properties (PDA). Synthese, 199(1–2):1065–1089, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02765-2.

Barack:2021:MKD

[1213] David L. Barack. Mental kinematics: dynamics and mechanics of neurocognitive systems. Synthese, 199(1–2):1091–1123, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02766-1.

Rubio-Fernandez:2021:PMM

[1214] Paula Rubio-Fernandez. Pragmatic markers: the missing link between language and theory of mind. Synthese, 199(1–2):1125–1158, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02768-z.

Batterman:2021:SBM

[1215] Robert W. Batterman and Sara Green. Steel and bone: mesoscale modeling and middle-out strategies in physics and biology. Synthese, 199(1–2): 1159–1184, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02769-y.

Jager:2021:SFU

[1216] Christoph Jäger and Federica Isabella Malfatti. The social fabric of understanding: equilibrium, authority, and epistemic empathy. Synthese, 199(1–2):1185–1205, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02776-z.

Crasnow:2021:COM

[1217] Sharon Crasnow. Coherence objectivity and measurement: the example of democracy. Synthese, 199(1–2):1207–1229, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02779-w.

Vekony:2021:IAK

[1218] Romy Vekony, Alfred Mele, and David Rose. Intentional action without knowledge. Synthese, 199(1–2):1231–1243, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02783-0.

Heron:2021:STJ

[1219] John Heron. Set-theoretic justification and the theoretical virtues. Synthese, 199(1–2):1245–1267, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02784-z.

Radcliffe:2021:HEA

[1220] Elizabeth S. Radcliffe. A Humean explanation of acting on normative reasons. Synthese, 199(1–2):1269–1292, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02788-9.

Malfatti:2021:ITC

[1221] Federica Isabella Malfatti. Introduction to the topical collection “True enough? Themes from Elgin”. Synthese, 199(1–2):1293–1305, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02790-1.

Poldrack:2021:PR

[1222] Russell A. Poldrack. The physics of representation. Synthese, 199(1–2): 1307–1325, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02793-y.

Fridland:2021:IJF

[1223] Ellen Fridland and Matt Stichter. It just feels right: an account of expert intuition. Synthese, 199(1–2):1327–1346, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02796-9.

Woensdregt:2021:CMC

[1224] Marieke Woensdregt, Chris Cummins, and Kenny Smith. A computational model of the cultural co-evolution of language and mindreading. Synthese, 199(1–2):1347–1385, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02798-7.

Fritz:2021:BCM

[1225] James Fritz and Elizabeth Jackson. Belief, credence, and moral encroachment. Synthese, 199(1–2):1387–1408, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02799-6.

Martin:2021:PVT

[1226] James V. Martin. Prolegomena to virtue-theoretic studies in the philosophy of mathematics. Synthese, 199(1–2):1409–1434, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02802-0.

Lawlor:2021:KR

[1227] Krista Lawlor. Knowledge and reasonableness. Synthese, 199(1–2):1435– 1451, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02803-z.

Moretti:2021:VTI

[1228] Luca Moretti, Peter Hartl, and Akos Gyarmathy. The value of truth: introduction to the topical collection. Synthese, 199(1–2):1453–1460, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-020-02804-y.

Paakkonen:2021:HIM

[1229] Juho Pääkkönen and Petri Ylikoski. Humanistic interpretation and machine learning. Synthese, 199(1–2):1461–1497, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02806-w.

Dohrn:2021:NMS

[1230] Daniel Dohrn. The non-maximality-solution to counterfactual scepticism. Synthese, 199(1–2):1499–1520, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02808-8.

Monteleone:2021:VSI

[1231] John M. Monteleone. Verbal slips and the intentionality of skills. Synthese, 199(1–2):1521–1537, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02810-0.

Margitay:2021:ETV

[1232] Tihamér Margitay. Epistemology of testimony and values in science. Synthese, 199(1–2):1539–1553, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02812-y.

vanEck:2021:MIS

[1233] Dingmar van Eck and Cory Wright. Mechanist idealisation in systems biology. Synthese, 199(1–2):1555–1575, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02816-8.

Elgin:2021:R

[1234] Catherine Z. Elgin. Replies. Synthese, 199(1–2):1577–1597, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02819-5.

Mohammadian:2021:ICC

[1235] Mousa Mohammadian. If consciousness causes collapse, the zombie argument fails. Synthese, 199(1–2):1599–1615, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02828-4.

Burnston:2021:CVC

[1236] Daniel C. Burnston. Contents, vehicles, and complex data analysis in neuroscience. Synthese, 199(1–2):1617–1639, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02831-9.

Goldberg:2021:ESP

[1237] Sanford C. Goldberg. On the epistemic significance of practical reasons to inquire. Synthese, 199(1–2):1641–1658, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02833-7.

Worsdale:2021:MOB

[1238] Rosie Worsdale and Jack Wright. My objectivity is better than yours: contextualising debates about gender inequality. Synthese, 199(1–2): 1659–1683, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02835-5.

Andrews:2021:IFP

[1239] Kristin Andrews, Shannon Spaulding, and Evan Westra. Introduction to Folk Psychology: Pluralistic Approaches. Synthese, 199(1–2):1685– 1700, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02837-3.

Dohrn:2021:HME

[1240] Daniel Dohrn. A Humean modal epistemology. Synthese, 199(1–2): 1701–1725, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02842-6.

Kotevska:2021:MIT

[1241] Laura Kotevska. Moral improvement through mathematics: Antoine Arnauld and Pierre Nicole’s Nouveaux éléments de géométrie. Synthese, 199(1–2):1727–1749, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02845-3.

Moore:2021:CEM

[1242] Richard Moore. The cultural evolution of mind-modelling. Synthese, 199(1–2):1751–1776, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02853-3.

Muszynski:2021:REP

[1243] Eric Muszynski and Christophe Malaterre. A roadmap to explanatory pluralism: introduction to the topical collection The Biology of Behaviour. Synthese, 199(1–2):1777–1789, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02856-0.

Beck:2021:DAP

[1244] Ori Beck. The direct argument is a prima facie threat to compatibilism. Synthese, 199(1–2):1791–1817, December 2021. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02861-3.

Vernazzani:2021:SPO

[1245] Alfredo Vernazzani, Błazej Skrzypulec, and Tobias Schlicht. Structure of perceptual objects: introduction to the Synthese topical collection. Synthese, 199(1–2):1819–1830, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02862-2.

Miyazono:2021:SEC

[1246] Kengo Miyazono and Alessandro Salice. Social epistemological conception of delusion. Synthese, 199(1–2):1831–1851, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02863-1. See

correction [1247].

Miyazono:2021:CSE

[1247] Kengo Miyazono and Alessandro Salice. Correction to: Social epistemological conception of delusion. Synthese, 199(1–2):1853–1854, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03036-4. See [1246].

Dings:2021:MA

[1248] Roy Dings. Meaningful affordances. Synthese, 199(1–2):1855–1875, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02864-0.

Jebeile:2021:UCC

[1249] Julie Jebeile, Vincent Lam, and Tim Räz. Understanding climate change with statistical downscaling and machine learning. Synthese, 199(1–2): 1877–1897, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02865-z.

Zavaleta:2021:DL

[1250] Martín Abreu Zavaleta. Disagreement lost. Synthese, 199(1–2):1899– 1932, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02866-y. See correction [1251].

Zavaleta:2021:CDL

[1251] Martín Abreu Zavaleta. Correction to: Disagreement lost. Synthese, 199(1–2):1933–1934, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02921-8. See [1250].

Kincaid:2021:LEI

[1252] Harold Kincaid. Live empirical issues in debates over objectivity in the social sciences. Synthese, 199(1–2):1935–1954, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02867-x.

Koch:2021:EWC

[1253] Steffen Koch. Engineering what? On concepts in conceptual engineering. Synthese, 199(1–2):1955–1975, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02868-w.

OLeary:2021:MMM

[1254] Diane O’Leary. Medicine’s metaphysical morass: how confusion about dualism threatens public health. Synthese, 199(1–2):1977–2005, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02869-9.

Bursten:2021:CCA

[1255] Julia R. S. Bursten. Classifying and characterizing active materials. Synthese, 199(1–2):2007–2026, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02870-2.

Dietz:2021:EES

[1256] Christina H. Dietz. Emotions, evidence, and safety. Synthese, 199(1–2): 2027–2050, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02871-1.

Vetter:2021:ED

[1257] Barbara Vetter. Explanatory dispositionalism. Synthese, 199(1–2):2051– 2075, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02872-0.

Ferretti:2021:VAP

[1258] Gabriele Ferretti and Francesco Marchi. Visual attention in pictorial perception. Synthese, 199(1–2):2077–2101, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02873-z.

Favela:2021:DRN

[1259] Luis H. Favela. The dynamical renaissance in neuroscience. Synthese, 199(1–2):2103–2127, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02874-y.

Viola:2021:BPB

[1260] Marco Viola. Beyond the Platonic Brain: facing the challenge of individual differences in function-structure mapping. Synthese, 199(1–2): 2129–2155, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02875-x.

Sims:2021:HCB

[1261] Matthew Sims. How to count biological minds: symbiosis, the free energy principle, and reciprocal multiscale integration. Synthese, 199(1–2): 2157–2179, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02876-w. See correction [1262].

Sims:2021:CHC

[1262] Matthew Sims. Correction to: How to count biological minds: symbiosis, the free energy principle, and reciprocal multiscale integration. Synthese, 199(1–2):2181, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02945-0. See [1261].

Altehenger:2021:SCS

[1263] Hannah Altehenger. Self-control and the self. Synthese, 199(1–2):2183– 2198, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02877-9.

Cea:2021:IIT

[1264] Ignacio Cea. Integrated information theory of consciousness is a functionalist emergentism. Synthese, 199(1–2):2199–2224, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02878-8.

Bradley:2021:BSD

[1265] Darren Bradley. Bayesianism and self-doubt. Synthese, 199(1–2):2225– 2243, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02879-7.

Turri:2021:EOF

[1266] John Turri. Evaluating objections to a factive norm of belief. Synthese, 199(1–2):2245–2250, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02881-z.

Monso:2021:DCU

[1267] Susana Monsó and Antonio J. Osuna-Mascaró. Death is common, so is understanding it: the concept of death in other species. Synthese, 199(1–2):2251–2275, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02882-y.

Livingston:2021:PRW

[1268] Eric Livingston. Practical reasoning and the witnessably rigorous proof. Synthese, 199(1–2):2277–2291, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02883-x.

McGlynn:2021:IWM

[1269] Aidan McGlynn. Immunity to wh-misidentification. Synthese, 199(1–2): 2293–2313, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02885-9.

Wakil:2021:OOP

[1270] Samantha Wakil. Objectivity and orgasm: the perils of imprecise definitions. Synthese, 199(1–2):2315–2333, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02886-8.

Mokriski:2021:CMP

[1271] David Mokriski. Conciliatory metaontology, permissive ontology, and nature’s joints. Synthese, 199(1–2):2335–2351, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02887-7.

Holmes:2021:CDM

[1272] Travis Holmes. Cognitive dynamical models as minimal models. Synthese, 199(1–2):2353–2373, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02888-6.

Mizrahi:2021:CSP

[1273] Moti Mizrahi. Conceptions of scientific progress in scientific practice: an empirical study. Synthese, 199(1–2):2375–2394, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02889-5.

Vlerick:2021:EHA

[1274] Michael Vlerick. Explaining human altruism. Synthese, 199(1–2):2395– 2413, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02890-y.

Vandenburgh:2021:CLT

[1275] Jonathan Vandenburgh. Conditional learning through causal models. Synthese, 199(1–2):2415–2437, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02891-x.

Mason:2021:EDF

[1276] Cathy Mason. The epistemic demands of friendship: friendship as inherently knowledge-involving. Synthese, 199(1–2):2439–2455, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02892-w.

Millikan:2021:NT

[1277] Ruth Garrett Millikan. Neuroscience and teleosemantics. Synthese, 199(1–2):2457–2465, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02893-9.

Pain:2021:MGM

[1278] Ross Pain and Rachael L. Brown. Mind the gap: a more evolutionarily plausible role for technical reasoning in cumulative technological culture. Synthese, 199(1–2):2467–2489, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02894-8.

Tozzi:2021:IDR

[1279] Arturo Tozzi and James F. Peters. Information-devoid routes for scalefree neurodynamics. Synthese, 199(1–2):2491–2504, December 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02895-
7.

Boyle:2021:RER

[1280] Alexandria Boyle. Remembering events and representing time. Synthese, 199(1–2):2505–2524, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02896-6.

Barkasi:2021:DWW

[1281] Michael Barkasi. Does what we dream feel present? Two varieties of presence and implications for measuring presence in VR. Synthese, 199(1–2): 2525–2551, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02898-4.

Gottlieb:2021:WNL

[1282] Joseph Gottlieb and Ali Rezaei. When nothing looks blue. Synthese, 199(1–2):2553–2561, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02899-3. See correction [1678].

Newhard:2021:AUA

[1283] Jay Newhard. Alethic undecidability and alethic indeterminacy. Synthese, 199(1–2):2563–2574, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02900-z.

Kappes:2021:ESE

[1284] Yannic Kappes. Explanation by status as empty-base explanation. Synthese, 199(1–2):2575–2595, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02901-y.

Hireche:2021:CNM

[1285] Salim Hirèche. Causal necessitarianism and the monotonicity objection. Synthese, 199(1–2):2597–2627, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02902-x.

Nakano:2021:RRA

[1286] Anderson Nakano. On Ramsey’s reason to amend Principia Mathematica’s logicism and Wittgenstein’s reaction. Synthese, 199(1–2):2629– 2646, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02903-w.

OBrien:2021:HET

[1287] Dan O’Brien. Humeanism and the epistemology of testimony. Synthese, 199(1–2):2647–2669, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02905-8.

Giannini:2021:NPD

[1288] Giacomo Giannini. New powers for dispositionalism. Synthese, 199(1–2): 2671–2700, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02906-7.

Klar:2021:WND

[1289] Philipp Klar. What is neurophilosophy: Do we need a non-reductive form? Synthese, 199(1–2):2701–2725, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02907-6.

Bilat:2021:CRM

[1290] Andrzej Biłat. The correctness and relevance of the modal ontological argument. Synthese, 199(1–2):2727–2743, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02908-5.

Humphreys:2021:ITE

[1291] Paul Humphreys. Invariances in transformational emergence. Synthese, 199(1–2):2745–2756, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02909-4.

Michel:2021:CBR

[1292] Matthias Michel and Megan A. K. Peters. Confirmation bias without rhyme or reason. Synthese, 199(1–2):2757–2772, December 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02910-
x.

Baysan:2021:RE

[1293] Umut Baysan. Rejecting epiphobia. Synthese, 199(1–2):2773–2791, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02911-w.

Figdor:2021:SFC

[1294] Carrie Figdor. Shannon + Friston = Content: Intentionality in predictive signaling systems. Synthese, 199(1–2):2793–2816, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02912-9.

Kim:2021:EMW

[1295] Dongwoo Kim. Explanation and modality: on why the Swampman is still worrisome to teleosemanticists. Synthese, 199(1–2):2817–2839, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02913-8.

Brown:2021:FMP

[1296] Christopher Devlin Brown. Fundamental mentality in a physical world. Synthese, 199(1–2):2841–2860, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02914-7. See correction

[1297].

Brown:2021:CFM

[1297] Christopher Devlin Brown. Correction to: Fundamental mentality in a physical world. Synthese, 199(1–2):2861, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02941-4. See

[1296].

Openshaw:2021:TAM

[1298] James Openshaw. Thinking about many. Synthese, 199(1–2):2863– 2882, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02904-9.

Malaterre:2021:EJE

[1299] Christophe Malaterre, Francis Lareau, Davide Pulizzotto, and Jonathan St-Onge. Eight journals over eight decades: a computational topicmodeling approach to contemporary philosophy of science. Synthese, 199(1–2):2883–2923, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02915-6.

Silva:2021:PRW

[1300] Paul Silva, Jr. Possessing reasons: why the awareness-first approach is better than the knowledge-first approach. Synthese, 199(1–2):2925– 2947, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02916-5.

Azzano:2021:DCW

[1301] Lorenzo Azzano. Dispositionality, categoricity, and where to find them. Synthese, 199(1–2):2949–2976, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02917-4.

Bradley:2021:OCR

[1302] Darren Bradley. Ought-contextualism and reasoning. Synthese, 199(1–2): 2977–2999, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02918-3.

Talbot:2021:ERF

[1303] Brian Talbot. Epistemic repugnance four ways. Synthese, 199(1–2): 3001–3022, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02919-2.

Tooming:2021:VNK

[1304] Uku Tooming and Kengo Miyazono. Vividness as a natural kind. Synthese, 199(1–2):3023–3043, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02920-9.

Atkinson:2021:PFB

[1305] David Atkinson and Jeanne Peijnenburg. Probability functions, belief functions and infinite regresses. Synthese, 199(1–2):3045–3059, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02923-6.

Jackson:2021:MTS

[1306] Brendan Balcerak Jackson. Model-theoretic semantics as model-based science. Synthese, 199(1–2):3061–3081, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02924-5.

Morgan:2021:EIR

[1307] Daniel Morgan. The essential indexical research program. Synthese, 199(1–2):3083–3100, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02925-4.

Hong:2021:MPM

[1308] Hao Hong. Many problems of the many. Synthese, 199(1–2):3101– 3116, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02926-3.

Gijsbers:2021:HJT

[1309] Victor Gijsbers. How (not) to judge a theory of causation. Synthese, 199(1–2):3117–3135, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02927-2.

Ernest:2021:MEP

[1310] Paul Ernest. Mathematics, ethics and purism: an application of MacIntyre’s virtue theory. Synthese, 199(1–2):3137–3167, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02928-1.

Xhignesse:2021:IFC

[1311] Michel-Antoine Xhignesse. Imagining fictional contradictions. Synthese, 199(1–2):3169–3188, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02929-0.

Arcangeli:2021:CNI

[1312] Margherita Arcangeli. The conceptual nature of imaginative content. Synthese, 199(1–2):3189–3205, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02930-7.

Priest:2021:LAN

[1313] Graham Priest. Logical abductivism and non-deductive inference. Synthese, 199(1–2):3207–3217, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02932-5.

Nguyen:2021:DFE

[1314] James Nguyen. Do fictions explain? Synthese, 199(1–2):3219–3244, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-020-02931-6.

Marsili:2021:LSA

[1315] Neri Marsili. Lying, speech acts, and commitment. Synthese, 199(1–2): 3245–3269, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02933-4.

Kuby:2021:DBE

[1316] Daniel Kuby. Decision-based epistemology: sketching a systematic framework of Feyerabend’s metaphilosophy. Synthese, 199(1–2):3271– 3299, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02934-3.

Henderson:2021:PNE

[1317] David Henderson. The place of non-epistemic matters in epistemology: norms and regulation in various communities. Synthese, 199(1–2):3301– 3323, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02935-2.

Schwartz:2021:PIL

[1318] Avshalom M. Schwartz. Political imagination and its limits. Synthese, 199(1–2):3325–3343, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02936-1.

May:2021:BSN

[1319] Joshua May. Bias in science: natural and social. Synthese, 199(1–2): 3345–3366, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02937-0.

Brouwer:2021:HCI

[1320] Thomas Brouwer, Roberta Ferrario, and Daniele Porello. Hybrid collective intentionality. Synthese, 199(1–2):3367–3403, December 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02938-
z.

Kant:2021:MVP

[1321] Deborah Kant and Deniz Sarikaya. Mathematizing as a virtuous practice: different narratives and their consequences for mathematics education and society. Synthese, 199(1–2):3405–3429, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02939-y.

Kilov:2021:BEW

[1322] Daniel Kilov. The brittleness of expertise and why it matters. Synthese, 199(1–2):3431–3455, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02940-5.

Roussos:2021:EDB

[1323] Joe Roussos. Expert deference as a belief revision schema. Synthese, 199(1–2):3457–3484, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02942-3.

Veigl:2021:NCR

[1324] Sophie Juliane Veigl. Notes on a complicated relationship: scientific pluralism, epistemic relativism, and stances. Synthese, 199(1–2):3485– 3503, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02943-2.

McDonald:2021:YWA

[1325] Lucy McDonald. Your word against mine: the power of uptake. Synthese, 199(1–2):3505–3526, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02944-1.

Raab:2021:UCE

[1326] Jonas Raab. The unbearable circularity of easy ontology. Synthese, 199(1–2):3527–3556, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02946-z.

Muller:2021:PEN

[1327] Jean Moritz Müller. Perceptualism and the epistemology of normative reasons. Synthese, 199(1–2):3557–3586, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02947-y.

Donnelly:2021:EDA

[1328] Maureen Donnelly. Explaining the differential application of nonsymmetric relations. Synthese, 199(1–2):3587–3610, December 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02948-
x.

Carmona:2021:PRM

[1329] Carla Carmona. Practices of remembering a movement in the dance studio: evidence for (a radicalized version of) the REC framework in the domain of memory. Synthese, 199(1–2):3611–3643, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/

s11229-020-02949-w. See correction [1330].

Carmona:2021:CPR

[1330] Carla Carmona. Correction to: Practices of remembering a movement in the dance studio: evidence for (a radicalized version of) the REC framework in the domain of memory. Synthese, 199(1–2): 3645, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-03021-3. See [1329].

Mayo-Wilson:2021:CPS

[1331] Conor Mayo-Wilson and Kevin J. S. Zollman. The computational philosophy: simulation as a core philosophical method. Synthese, 199(1–2): 3647–3673, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02950-3.

Mormann:2021:PPT

[1332] Thomas Mormann. Prototypes, poles, and tessellations: towards a topological theory of conceptual spaces. Synthese, 199(1–2):3675–3710, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02951-2.

Nefdt:2021:SRG

[1333] Ryan M. Nefdt. Structural realism and generative linguistics. Synthese, 199(1–2):3711–3737, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02952-1.

DalMagro:2021:RRE

[1334] Tamires Dal Magro and Matheus Valente. On the representational role of Euclidean diagrams: representing qua samples. Synthese, 199(1–2): 3739–3760, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02953-0.

Wheeler:2021:IFC

[1335] Gregory Wheeler and Fabio G. Cozman. On the imprecision of full conditional probabilities. Synthese, 199(1–2):3761–3782, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02954-z.

Varga:2021:MSS

[1336] Somogy Varga. Medicine as science. Systematicity and demarcation. Synthese, 199(1–2):3783–3804, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02955-y.

Aguilera:2021:HIM

[1337] Mariela Aguilera. Heterogeneous inferences with maps. Synthese, 199 (1–2):3805–3824, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02957-w.

Glauer:2021:NFP

[1338] Ramiro Glauer and Frauke Hildebrandt. No facts without perspectives. Synthese, 199(1–2):3825–3851, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02958-9.

Osiurak:2021:FWM

[1339] François Osiurak and Giovanni Federico. Four ways of (mis-)conceiving embodiment in tool use. Synthese, 199(1–2):3853–3879, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02960-1.

Brown:2021:AEP

[1340] Deborah J. Brown and Brian Key. Is absence of evidence of pain ever evidence of absence? Synthese, 199(1–2):3881–3902, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02961-0.

Grune-Yanoff:2021:JMC

[1341] Till Grüne-Yanoff. Justifying method choice: a heuristic-instrumentalist account of scientific methodology. Synthese, 199(1–2):3903–3921, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02962-z.

Mariani:2021:IPO

[1342] Cristian Mariani and Giuliano Torrengo. The indeterminate present and the open future. Synthese, 199(1–2):3923–3944, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02963-y.

Lai:2021:AEA

[1343] Changsheng Lai. Against epistemic absolutism. Synthese, 199(1–2): 3945–3967, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02964-x.

Tallant:2021:RCP

[1344] Jonathan Tallant and David Ingram. The rotten core of presentism. Synthese, 199(1–2):3969–3991, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02965-w.

Deng:2021:ICC

[1345] Duen-Min Deng and Kok Yong Lee. Indicative and counterfactual conditionals: a causal-modeling semantics. Synthese, 199(1–2):3993– 4014, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02966-9.

Anta:2021:CIT

[1346] Javier Anta. Can informational thermal physics explain the approach to equilibrium? Synthese, 199(1–2):4015–4038, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02967-8.

Raatikainen:2021:TCP

[1347] Panu Raatikainen. Troubles with the Canberra Plan. Synthese, 199(1–2): 4039–4060, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02968-7.

Montuschi:2021:FCO

[1348] Eleonora Montuschi. Finding a context for objectivity. Synthese, 199 (1–2):4061–4076, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02969-6.

Ross:2021:TNC

[1349] Lauren N. Ross. Tracers in neuroscience: Causation, constraints, and connectivity. Synthese, 199(1–2):4077–4095, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02970-z.

Broughton:2021:CF

[1350] Gabriel L. Broughton. Carnapian frameworks. Synthese, 199(1–2):4097– 4126, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02971-y.

Walsh:2021:PPT

[1351] Elena Walsh. Pathological prediction: a top-down cause of organic disease. Synthese, 199(1–2):4127–4150, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02972-x.

Ketland:2021:FAM

[1352] Jeffrey Ketland. Foundations of applied mathematics I. Synthese, 199 (1–2):4151–4193, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02973-w.

Atkinson:2021:PTA

[1353] David Atkinson and Jeanne Peijnenburg. Probabilistic truth approximation and fixed points. Synthese, 199(1–2):4195–4216, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02975-8.

Wilson:2021:ISS

[1354] Keith A. Wilson. Individuating the senses of ‘smell’: orthonasal versus retronasal olfaction. Synthese, 199(1–2):4217–4242, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02976-7.

Marxen:2021:ERS

[1355] Chad Marxen. On the epistemic rationality and significance of selffulfilling beliefs. Synthese, 199(1–2):4243–4260, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02977-6.

Krempel:2021:CEL

[1356] Raquel Krempel. Is color experience linguistically penetrable? Synthese, 199(1–2):4261–4285, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02978-5.

Meynell:2021:GA

[1357] Letitia Meynell and Andrew Lopez. Gendering animals. Synthese, 199 (1–2):4287–4311, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-02979-4.

Showler:2021:WSN

[1358] Paul D. G. Showler. Why subject naturalists need pragmatic genealogy. Synthese, 199(1–2):4313–4335, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02980-x.

Mingon:2021:WRC

[1359] McArthur Mingon and John Sutton. Why robots can’t haka: skilled performance and embodied knowledge in the Maori haka. Synthese, 199(1–2):4337–4365, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02981-w.

Cochran:2021:HSC

[1360] Calvin T. Cochran and Jeffrey A. Barrett. How signaling conventions are established. Synthese, 199(1–2):4367–4391, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02982-9.

Rudnicki:2021:PAP

[1361] Konrad Rudnicki. The psychological aspects of paraconsistency. Synthese, 199(1–2):4393–4414, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02983-8.

Topey:2021:RRE

[1362] Brett Topey. Realism, reliability, and epistemic possibility: on modally interpreting the Benacerraf–Field challenge. Synthese, 199(1–2):4415– 4436, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02984-7.

DeWitt:2021:DDA

[1363] Jason A. DeWitt. Dogramaci’s deflationism about rationality. Synthese, 199(1–2):4437–4455, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02985-6.

Hipolito:2021:ESP

[1364] Inês Hipólito, Manuel Baltieri, Karl Friston, and Maxwell J. D. Ramstead. Embodied skillful performance: where the action is. Synthese, 199(1–2):4457–4481, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02986-5.

Jespersen:2021:FAE

[1365] Bjørn Jespersen. First among equals: co-hyperintensionality for structured propositions. Synthese, 199(1–2):4483–4497, December 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-02987-
4.

Fellowes:2021:HAS

[1366] Sam Fellowes. How autism shows that symptoms, like psychiatric diagnoses, are ’constructed’: methodological and epistemic consequences. Synthese, 199(1–2):4499–4522, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02988-3.

Gallagher:2021:PPR

[1367] Shaun Gallagher and Zuzanna Rucińska. Prospecting performance: rehearsal and the nature of imagination. Synthese, 199(1–2):4523– 4541, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02989-2.

Roberts:2021:OVH

[1368] Alice Roberts. Objectification and vision: how images shape our early visual processes. Synthese, 199(1–2):4543–4560, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02990-9.

Baron-Schmitt:2021:CG

[1369] Nathaniel Baron-Schmitt. Contingent grounding. Synthese, 199(1–2): 4561–4580, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02991-8.

Rimkevicius:2021:MII

[1370] Paulius Rimkevicius. Mental imagery and the illusion of conscious will. Synthese, 199(1–2):4581–4600, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02992-7.

Baetu:2021:HIA

[1371] Tudor M. Baetu. How interventionist accounts of causation work in experimental practice and why there is no need to worry about supervenience. Synthese, 199(1–2):4601–4620, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-02993-6.

Lee:2021:OI

[1372] Junhyo Lee. “Ought” and intensionality. Synthese, 199(1–2):4621– 4643, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-02994-5.

Piedrahita:2021:CHE

[1373] Oscar A. Piedrahita. Can hinge epistemology close the door on epistemic relativism? Synthese, 199(1–2):4645–4671, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-02995-4.

Kompa:2021:EEN

[1374] Nikola Anna Kompa. Epistemic evaluation and the need for ‘impure’ epistemic standards. Synthese, 199(1–2):4673–4693, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02996-3.

Huggett:2021:MPN

[1375] Nick Huggett, Fedele Lizzi, and Tushar Menon. Missing the point in noncommutative geometry. Synthese, 199(1–2):4695–4728, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02998-1.

DeFlorio:2021:AFT

[1376] Ciro De Florio and Aldo Frigerio. Assertions and future tense semantics. Synthese, 199(1–2):4729–4747, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-02999-0.

Indrzejczak:2021:NAE

[1377] Andrzej Indrzejczak. A novel approach to equality. Synthese, 199(1–2): 4749–4774, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-03000-8.

Strohmaier:2021:ONN

[1378] David Strohmaier. Ontology, neural networks, and the social sciences. Synthese, 199(1–2):4775–4794, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-03002-6.

Xu:2021:ELO

[1379] Kevin Xu. Everything is learnable, once it is settled. Synthese, 199(1–2): 4795–4817, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-03003-5.

Mohammadian:2021:TVT

[1380] Mousa Mohammadian. Theoretical virtues and theorizing in physics: against the instrumentalist view of simplicity. Synthese, 199(1–2):4819– 4828, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-03004-4.

Bales:2021:RPR

[1381] Adam Bales and Toby Handfield. Resolutions provide reasons or: “how the Cookie Monster quit cookies”. Synthese, 199(1–2):4829– 4840, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-03005-3.

Pupa:2021:CTD

[1382] Francesco Pupa. Coalescent theories and divergent paraphrases: definites, non-extensional contexts, and familiarity. Synthese, 199(1–2): 4841–4862, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-03006-2. See correction [1383].

Pupa:2021:CCT

[1383] Francesco Pupa. Correction to: Coalescent theories and divergent paraphrases: definites, non-extensional contexts, and familiarity. Synthese, 199(1–2):4863–4864, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03091-x. See [1382].

Plaisance:2021:PIU

[1384] Kathryn S. Plaisance, Jay Michaud, and John McLevey. Pathways of influence: understanding the impact of philosophy of science in scientific domains. Synthese, 199(1–2):4865–4896, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-03007-1.

Hemmo:2021:CSL

[1385] Meir Hemmo and Orly Shenker. A challenge to the second law of thermodynamics from cognitive science and vice versa. Synthese, 199(1–2): 4897–4927, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-03008-0.

Vassallo:2021:DGR

[1386] Antonio Vassallo. Does general relativity highlight necessary connections in nature? Synthese, 199(1–2):4929–4951, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-020-03009-z.

Dorge:2021:NEI

[1387] Friedrich Christoph Dörge and Matthias Holweger. The non-existence of institutional facts. Synthese, 199(1–2):4953–4974, December 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-03010-
6.

Eder:2021:ECE

[1388] Anna-Maria A. Eder. Explicating the concept of epistemic rationality. Synthese, 199(1–2):4975–5000, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-03011-5.

Freitag:2021:ESM

[1389] Wolfgang Freitag and Nadja-Mira Yolcu. An expressivist solution to Moorean paradoxes. Synthese, 199(1–2):5001–5024, December 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-020-03012-
4.

Rennick:2021:TAF

[1390] Stephanie Rennick. Trope analysis and folk intuitions. Synthese, 199 (1–2):5025–5043, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-020-03013-3.

Brown:2021:PR

[1391] Thomas D. Brown. Propositions are not representational. Synthese, 199(1–2):5045–5060, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-03014-2.

Voltolini:2021:BMB

[1392] Alberto Voltolini. Beliefs, make-beliefs, and making believe that beliefs are not make-beliefs. Synthese, 199(1–2):5061–5078, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-03015-1.

Livanios:2021:CIT

[1393] Vassilis Livanios. Challenging the identity theory of properties. Synthese, 199(1–2):5079–5105, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-03016-0.

Kriegel:2021:COA

[1394] Uriah Kriegel. The concreteness of objects: an argument against mereological bundle theory. Synthese, 199(1–2):5107–5124, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-03017-z.

Li:2021:NPI

[1395] Zhanglyu Li and Frank Zenker. Newcomb’s problem isn’t a choice dilemma. Synthese, 199(1–2):5125–5143, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-020-03018-y.

Morrison:2021:PVI

[1396] John Morrison. Perceptual variation and ignorance. Synthese, 199(1–2): 5145–5173, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-03019-x.

Waller:2021:WDB

[1397] Robyn Repko Waller. Weighing in on decisions in the brain: neural representations of pre-awareness practical intention. Synthese, 199(1–2): 5175–5203, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-020-03020-4.

Bianchi:2021:GOS

[1398] Silvia Bianchi and Joaquim Giannotti. Grounding ontic structuralism. Synthese, 199(1–2):5205–5223, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-020-03001-7.

Jones:2021:DRD

[1399] Elis Jones. Distinguishing regeneration from degradation in coral ecosystems: the role of value. Synthese, 199(1–2):5225–5253, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03023-9.

Hubert:2021:RF

[1400] Mario Hubert. Reviving frequentism. Synthese, 199(1–2):5255–5284, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-021-03024-8.

Rucinska:2021:EPR

[1401] Zuzanna Rucińska. Enactive planning in rock climbing: recalibration, visualization and nested affordances. Synthese, 199(1–2):5285– 5310, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03025-7.

Malecka:2021:KBP

[1402] Magdalena Małecka. Knowledge, behaviour, and policy: questioning the epistemic presuppositions of applying behavioural science in public policymaking. Synthese, 199(1–2):5311–5338, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03026-6.

Elgin:2021:KCU

[1403] Samuel Z. Elgin. Knowledge is closed under analytic content. Synthese, 199(1–2):5339–5353, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03027-5.

Demeter:2021:FGH

[1404] Tamás Demeter. Fodor’s guide to the Humean mind. Synthese, 199(1–2): 5355–5375, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03028-4.

Grosholz:2021:PPM

[1405] Emily R. Grosholz. On the plains and prairies of Minnesota: The role of mathematical statistics in biological explanation. Synthese, 199(1–2): 5377–5393, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03029-3.

Mueller:2021:KNA

[1406] Andy Mueller. The knowledge norm of apt practical reasoning. Synthese, 199(1–2):5395–5414, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03030-w.

Meloni:2021:CTE

[1407] Maurizio Meloni and Jack Reynolds. Correction to: Thinking embodiment with genetics: epigenetics and postgenomic biology in embodied cognition and enactivism. Synthese, 199(1–2):5415–5416, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-020-02997-2. See [1060].

Mallozzi:2021:MBH

[1408] Antonella Mallozzi. In memoriam: Bob Hale. Synthese, 199(3–4): 5417, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-019-02363-x.

Godden:2021:RGA

[1409] David Godden and John Grey. Reasoning by grounded analogy. Synthese, 199(3–4):5419–5453, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-020-02974-9.

Peebles:2021:PUC

[1410] Graham Peebles. Phenomenology and the unity of consciousness. Synthese, 199(3–4):5455–5477, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03031-9.

Facchin:2021:SRD

[1411] Marco Facchin. Structural representations do not meet the job description challenge. Synthese, 199(3–4):5479–5508, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03032-8.

Cohen:2021:IKD

[1412] Michael Cohen. Inexact knowledge and dynamic introspection. Synthese, 199(3–4):5509–5531, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03033-7.

Williams:2021:NPP

[1413] Neil E. Williams. (A new paradigm for) the problem of the many. Synthese, 199(3–4):5533–5550, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03034-6.

Raleigh:2021:NAE

[1414] Thomas Raleigh. A new anti-expertise dilemma. Synthese, 199(3–4): 5551–5569, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03035-5.

Rittberg:2021:IHM

[1415] Colin Jakob Rittberg. Intellectual humility in mathematics. Synthese, 199(3–4):5571–5601, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03037-3.

Mankowitz:2021:HMD

[1416] Poppy Mankowitz. How to have a metalinguistic dispute. Synthese, 199(3–4):5603–5622, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03038-2.

Corti:2021:YAQ

[1417] Alberto Corti. Yet again, quantum indeterminacy is not worldly indecision. Synthese, 199(3–4):5623–5643, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03039-1.

Giberman:2021:DGD

[1418] Daniel Giberman. Deep gunk and deep junk. Synthese, 199(3–4):5645– 5667, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03040-8.

Gaspard:2021:FLT

[1419] Jeoffrey Gaspard. “A familiar logical triplet”: on Peirce’s grammar of representation and its relation to scientific inquiry. Synthese, 199(3–4): 5669–5686, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03041-7.

Enoch:2021:SRW

[1420] David Enoch and Levi Spectre. Statistical resentment, or: what’s wrong with acting, blaming, and believing on the basis of statistics alone. Synthese, 199(3–4):5687–5718, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03042-6.

Ingelstrom:2021:CHM

[1421] Mats Ingelström and Willem van der Deijl. Can happiness measures be calibrated? Synthese, 199(3–4):5719–5746, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03043-5.

Berghofer:2021:EEE

[1422] Philipp Berghofer. Evaluative experiences: the epistemological significance of moral phenomenology. Synthese, 199(3–4):5747–5768, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03044-4.

Lebens:2021:WGJ

[1423] Samuel Lebens. Will I get a job? Contextualism, belief, and faith. Synthese, 199(3–4):5769–5790, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03045-3.

Pantsar:2021:BIC

[1424] Markus Pantsar. Bootstrapping of integer concepts: the stronger deviant-interpretation challenge (and how to solve it). Synthese, 199 (3–4):5791–5814, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-021-03046-2.

Charbonneau:2021:FGP

[1425] Mathieu Charbonneau and Pierrick Bourrat. Fidelity and the grain problem in cultural evolution. Synthese, 199(3–4):5815–5836, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03047-1.

vanElswyk:2021:HIN

[1426] Peter van Elswyk and Yasha Sapir. Hedging and the ignorance norm on inquiry. Synthese, 199(3–4):5837–5859, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03048-0.

Evershed:2021:AWL

[1427] J. W. Evershed. Another way logic might be normative. Synthese, 199(3–4):5861–5881, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03049-z. See correction [1428].

Evershed:2021:CAW

[1428] J. W. Evershed. Correction to: Another way logic might be normative. Synthese, 199(3–4):5883, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03308-z. See [1427].

Amiriara:2021:PNE

[1429] Hassan Amiriara. On the possibility of non-eternalism without absolute simultaneity. Synthese, 199(3–4):5885–5898, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03050-6.

Lubacz:2021:EPI

[1430] Józef Lubacz. Epistemic and poietic intentional processes. Synthese, 199(3–4):5899–5915, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03051-5.

Roskies:2021:RSA

[1431] Adina L. Roskies. Representational similarity analysis in neuroimaging: proxy vehicles and provisional representations. Synthese, 199(3–4):

5917–5935, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03052-4.

Fridland:2021:SSC

[1432] Ellen Fridland. Skill and strategic control. Synthese, 199(3–4):5937– 5964, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03053-3.

Correia:2021:FGT

[1433] Fabrice Correia. Fundamentality from grounding trees. Synthese, 199 (3–4):5965–5994, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-021-03054-2.

Sussman:2021:QTH

[1434] Nathaniel F. Sussman. Quick thinking: how Einstein did (and did not) refute the ether frame of reference. Synthese, 199(3–4):5995– 6008, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03056-0.

Schurz:2021:PTC

[1435] Gerhard Schurz. Probabilistic truthlikeness, content elements, and metainductive probability optimization. Synthese, 199(3–4):6009–6037, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03057-z.

Duguid:2021:LHE

[1436] Callum Duguid. Lawful Humean explanations are not circular. Synthese, 199(3–4):6039–6059, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03058-y.

Egler:2021:WUW

[1437] Miguel Egler. Why understanding-why is contrastive. Synthese, 199 (3–4):6061–6083, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-021-03059-x.

Jaag:2021:PAL

[1438] Siegfried Jaag. A puzzle about laws and explanation. Synthese, 199(3–4): 6085–6102, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03061-3.

Wilde:2021:MRP

[1439] Michael Wilde. Mechanistic reasoning and the problem of masking. Synthese, 199(3–4):6103–6118, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03062-2.

ElSkaf:2021:PTS

[1440] Rawad El Skaf. Probing theoretical statements with thought experiments. Synthese, 199(3–4):6119–6147, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03063-1.

Jope:2021:SPT

[1441] Matthew Jope. The symmetry problem for testimonial conservatism. Synthese, 199(3–4):6149–6167, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03064-0.

Ogaard:2021:NBCb

[1442] Tore Fjetland Øgaard. Non-Boolean classical relevant logics II: Classicality through truth-constants. Synthese, 199(3–4):6169–6201, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03065-z.

Whittle:2021:MAR

[1443] Bruno Whittle. Mathematical anti-realism and explanatory structure. Synthese, 199(3–4):6203–6217, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03066-y.

Khelfaoui:2021:VPS

[1444] Mahdi Khelfaoui, Yves Gingras, and Thomas Pradeu. The visibility of philosophy of science in the sciences, 1980-2018. Synthese, 199(3–4): 6219–6249, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03067-x.

Bermudez:2021:SSC

[1445] Juan Pablo Bermúdez. The skill of self-control. Synthese, 199(3–4): 6251–6273, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03068-w.

Fox:2021:NC

[1446] Philip Fox. Naturalizing the contributory. Synthese, 199(3–4):6275– 6298, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03069-9.

Flores:2021:DER

[1447] Carolina Flores. Delusional evidence-responsiveness. Synthese, 199(3–4): 6299–6330, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03070-2.

Bykvist:2021:TVS

[1448] Krister Bykvist. Taking values seriously. Synthese, 199(3–4):6331– 6356, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03071-1.

Wilhelm:2021:CSM

[1449] Isaac Wilhelm. Comparing the structures of mathematical objects. Synthese, 199(3–4):6357–6369, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03072-0.

Williams:2021:HGS

[1450] Heath Williams. Is Husserl guilty of Sellars’ myth of the sensory given. Synthese, 199(3–4):6371–6389, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03073-z.

Faulkner:2021:PA

[1451] Paul Faulkner. The presumption of assurance. Synthese, 199(3–4):6391– 6406, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03074-y.

Beni:2021:CAM

[1452] Majid D. Beni. A critical analysis of Markovian monism. Synthese, 199(3–4):6407–6427, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03075-x.

Rossberg:2021:LSS

[1453] Marcus Rossberg and Stewart Shapiro. Logic and science: science and logic. Synthese, 199(3–4):6429–6454, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03076-w.

Kneer:2021:PPT

[1454] Markus Kneer. Predicates of personal taste: empirical data. Synthese, 199(3–4):6455–6471, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03077-9.

Hattiangadi:2021:RID

[1455] Anandi Hattiangadi and H. Orri Stefánsson. Radical interpretation and decision theory. Synthese, 199(3–4):6473–6494, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03078-8.

Rosen:2021:CDM

[1456] Melanie G. Rosen. I could do that in my sleep: skilled performance in dreams. Synthese, 199(3–4):6495–6522, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03079-7.

Eder:2021:FIO

[1457] Günther Eder. Frege on intuition and objecthood in projective geometry. Synthese, 199(3–4):6523–6561, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03080-0.

Lee:2021:IC

[1458] Wooram Lee. The independence of (in)coherence. Synthese, 199(3–4): 6563–6584, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03081-z.

Huttegger:2021:BCD

[1459] Simon M. Huttegger and Gerard J. Rothfus. Bradley conditionals and dynamic choice. Synthese, 199(3–4):6585–6599, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03082-y.

Schonherr:2021:SRC

[1460] Julius Schönherr. Salience reasoning in coordination games. Synthese, 199(3–4):6601–6620, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03083-x. See correction [1461].

Schonherr:2021:CSR

[1461] Julius Schönherr. Correction to: Salience reasoning in coordination games. Synthese, 199(3–4):6621, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03147-y. See [1460].

vanEs:2021:BPO

[1462] Thomas van Es and Michael D. Kirchhoff. Between pebbles and organisms: weaving autonomy into the Markov blanket. Synthese, 199(3–4): 6623–6644, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03084-w.

Antoniou:2021:PAO

[1463] Antonis Antoniou. A pragmatic approach to the ontology of models. Synthese, 199(3–4):6645–6664, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03085-9.

Henning:2021:EMN

[1464] Tim Henning. An epistemic modal norm of practical reasoning. Synthese, 199(3–4):6665–6686, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03086-8.

Lewinski:2021:IP

[1465] Marcin Lewiński. Illocutionary pluralism. Synthese, 199(3–4):6687– 6714, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03087-7.

Khan:2021:RPT

[1466] Saira Khan. Rational preference in transformative experiences. Synthese, 199(3–4):6715–6732, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03089-5.

Palermos:2021:SRB

[1467] Spyridon Orestis Palermos. System reliabilism and basic beliefs: defeasible, undefeated and likely to be true. Synthese, 199(3–4):6733– 6759, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03090-y.

Elliott:2021:WC

[1468] Katrina Elliott. Where are the chances? Synthese, 199(3–4):6761– 6783, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03092-w.

Wolf:2021:DRD

[1469] Aaron Wolf. Do reasons drain away? Synthese, 199(3–4):6785– 6802, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03093-9.

Greely:2021:EFM

[1470] Nathaniel Greely. Epistemic feelings, metacognition, and the Lima problem. Synthese, 199(3–4):6803–6825, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03094-8.

Stillman:2021:SBS

[1471] Richard P. Stillman. Slurs as ballistic speech. Synthese, 199(3–4):6827– 6843, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03095-7.

dAgostini:2021:CP

[1472] Franca d’Agostini. Conjunctive paraconsistency. Synthese, 199(3–4): 6845–6874, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03096-6.

Mollo:2021:WGC

[1473] Dimitri Coelho Mollo. Why go for a computation-based approach to cognitive representation. Synthese, 199(3–4):6875–6895, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03097-5.

OBrien:2021:WR

[1474] David O’Brien and Ben Schwan. Wrongfulness rewarded? Synthese, 199(3–4):6897–6916, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03098-4.

Pickel:2021:FCS

[1475] Bryan Pickel. The functional composition of sense. Synthese, 199(3–4): 6917–6942, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03099-3.

Fuqua:2021:EM

[1476] Jonathan Fuqua. Ethical Mooreanism. Synthese, 199(3–4):6943–6965, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-021-03100-z.

Harris:2021:CIH

[1477] Keith Raymond Harris. Collective intellectual humility and arrogance. Synthese, 199(3–4):6967–6979, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03101-y.

McGrew:2021:CCC

[1478] Lydia McGrew. Confirmation, coincidence, and contradiction. Synthese, 199(3–4):6981–7002, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03102-x.

Sarkia:2021:MIA

[1479] Matti Sarkia. Modeling intentional agency: a neo-gricean framework. Synthese, 199(3–4):7003–7030, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03103-w.

Fassio:2021:DWR

[1480] Davide Fassio and Jie Gao. Do we really need a knowledge-based decision theory? Synthese, 199(3–4):7031–7059, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03104-9.

Sebastian:2021:FPR

[1481] Miguel Ángel Sebastián. First-person representations and responsible agency in AI. Synthese, 199(3–4):7061–7079, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03105-8. See

correction [1482].

Sebastian:2021:CFP

[1482] Miguel Ángel Sebastián. Correction to: First-person representations and responsible agency in AI. Synthese, 199(3–4):7081, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03136-1. See [1481].

Chasid:2021:IIR

[1483] Alon Chasid. Imaginative immersion, regulation, and doxastic mediation. Synthese, 199(3–4):7083–7106, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03055-1.

Skokowski:2021:OS

[1484] Paul Skokowski. Observing a superposition. Synthese, 199(3–4):7107– 7129, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03106-7.

Qu:2021:SPJ

[1485] Hsueh Qu. Synthetic a priori judgments and Kant’s response to Hume on induction. Synthese, 199(3–4):7131–7157, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03107-6.

Effingham:2021:VTM

[1486] Nikk Effingham. Vacillating time: a metaphysics for time travel and Geachianism. Synthese, 199(3–4):7159–7180, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:/

/link.springer.com/article/10.1007/s11229-021-03108-5. See re-

ply [1565].

Yip:2021:EHL

[1487] Brandon Yip. Emotion as high-level perception. Synthese, 199(3–4): 7181–7201, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03109-4.

Cavedon-Taylor:2021:UKI

[1488] Dan Cavedon-Taylor. Untying the knot: imagination, perception and their neural substrates. Synthese, 199(3–4):7203–7230, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03110-x.

Peters:2021:WCE

[1489] Uwe Peters and Nikolaj Nottelmann. Weighing the costs: the epistemic dilemma of no-platforming. Synthese, 199(3–4):7231–7253, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03111-w.

Imaguire:2021:CAU

[1490] Guido Imaguire. On the coherence of Aristotelian universals. Synthese, 199(3–4):7255–7263, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03112-9.

Thunder:2021:TNR

[1491] Simon Thunder. There is no reason to replace the razor with the laser. Synthese, 199(3–4):7265–7282, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03113-8.

Sendlak:2021:CSP

[1492] Maciej Sendłak. Counterpossibles, story prefix and trivialism. Synthese, 199(3–4):7283–7301, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03114-7.

Kauss:2021:CSP

[1493] Dominik Kauss. Context-sensitivity and the preface paradox for credence. Synthese, 199(3–4):7303–7320, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03115-6.

Fiocco:2021:TNI

[1494] M. Oreste Fiocco. There is nothing to identity. Synthese, 199(3–4): 7321–7337, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03117-4.

Baron:2021:UME

[1495] Sam Baron. Unification and mathematical explanation in science. Synthese, 199(3–4):7339–7363, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03118-3.

Rostworowski:2021:CSL

[1496] Wojciech Rostworowski, Natalia Pietrulewicz, and Marcin Bedkowski. Conditionals and specific links — an experimental study. Synthese, 199(3–4):7365–7399, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03119-2.

Marxen:2021:EUT

[1497] Chad Marxen. Epistemic utility theory’s difficult future. Synthese, 199(3–4):7401–7421, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03120-9.

Cumby:2021:TBI

[1498] Jill Cumby. Thinking beyond imagining. Synthese, 199(3–4):7423– 7435, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03121-8.

delRio:2021:WUE

[1499] Andrew del Rio. Why undermining evolutionary debunkers is not enough. Synthese, 199(3–4):7437–7452, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03123-6.

Anaya:2021:PPM

[1500] Alfonso Anaya. Perceptual presentation and the myth of the given. Synthese, 199(3–4):7453–7476, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03124-5.

Michaelian:2021:IPR

[1501] Kourken Michaelian. Imagining the past reliably and unreliably: towards a virtue theory of memory. Synthese, 199(3–4):7477–7507, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03125-4.

Marchesi:2021:RRS

[1502] Andrea Marchesi. A radical relationist solution to the problem of intentional inexistence. Synthese, 199(3–4):7509–7534, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03126-3.

Buttingsrud:2021:BSP

[1503] Camille Buttingsrud. Bodies in skilled performance: how dancers reflect through the living body. Synthese, 199(3–4):7535–7554, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03127-2.

DePoe:2021:JA

[1504] John M. DePoe. Justification by acquaintance. Synthese, 199(3–4): 7555–7573, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03128-1.

Spaulding:2021:BB

[1505] Shannon Spaulding. Beliefs and biases. Synthese, 199(3–4):7575– 7594, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03129-0.

Simion:2021:BP

[1506] Mona Simion. Blame as performance. Synthese, 199(3–4):7595–7614, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-021-03130-7.

deBruin:2021:KAS

[1507] Boudewijn de Bruin. Knowledge attribution, socioeconomic status, and education: new results using the Great British Class Survey. Synthese, 199(3–4):7615–7657, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03131-6.

Pavese:2021:PCP

[1508] Carlotta Pavese. Practical concepts and productive reasoning. Synthese, 199(3–4):7659–7688, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03132-5.

Houlders:2021:TEA

[1509] Joseph W. Houlders, Lisa Bortolotti, and Matthew R. Broome. Threats to epistemic agency in young people with unusual experiences and beliefs. Synthese, 199(3–4):7689–7704, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03133-4.

Bellucci:2021:AEG

[1510] Francesco Bellucci and Ahti-Veikko Pietarinen. An analysis of Existential Graphs — part 2: Beta. Synthese, 199(3–4):7705–7726, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03134-3.

Arango-Munoz:2021:IMW

[1511] Santiago Arango-Muñoz and Juan Pablo Bermúdez. Intentional mindwandering as intentional omission: the surrealist method. Synthese, 199(3–4):7727–7748, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03135-2.

Read:2021:LMW

[1512] James Read and Baptiste Le Bihan. The landscape and the multiverse: What’s the problem? Synthese, 199(3–4):7749–7771, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03137-0.

Sartenaer:2021:NSA

[1513] Olivier Sartenaer and Umut Baysan. Non-standard approaches to emergence: introduction to the special issue. Synthese, 199(3–4):7773– 7776, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03138-z.

Kovaka:2021:FAF

[1514] Karen Kovaka. Fighting about frequency. Synthese, 199(3–4):7777– 7797, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03139-y. See correction [1515].

Kovaka:2021:CFA

[1515] Karen Kovaka. Correction to: Fighting about frequency. Synthese, 199 (3–4):7799, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03194-5. See [1514].

Sims:2021:MOW

[1516] Matt Sims and Giovanni Pezzulo. Modelling ourselves: what the free energy principle reveals about our implicit notions of representation. Synthese, 199(3–4):7801–7833, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03140-5.

Coninx:2021:PFA

[1517] Sabrina Coninx and Peter Stilwell. Pain and the field of affordances: an enactive approach to acute and chronic pain. Synthese, 199(3–4): 7835–7863, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03142-3.

Silver:2021:EWS

[1518] Kenneth Silver. Emergence within social systems. Synthese, 199(3–4): 7865–7887, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03143-2.

Curry:2021:HBL

[1519] Devin Sanchez Curry. How beliefs are like colors. Synthese, 199(3–4): 7889–7918, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03144-1.

Springle:2021:AII

[1520] Alison Ann Springle and Justin Humphreys. Anti-intellectualism, instructive representations, and the intentional action argument. Synthese, 199(3–4):7919–7955, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03145-0.

Alexander:2021:AGD

[1521] J. McKenzie Alexander and Julia Morley. Accounting for groups: the dynamics of intragroup deliberation. Synthese, 199(3–4):7957–7980, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03146-z.

Sassarini:2021:NGB

[1522] Elisabetta Sassarini. No ground to bridge the gap. Synthese, 199(3–4): 7981–7999, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03149-w.

Kuipers:2021:APD

[1523] Theo A. F. Kuipers. Approaching probabilistic and deterministic nomic truths in an inductive probabilistic way. Synthese, 199(3–4):8001– 8028, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03150-3.

Marshall:2021:CFG

[1524] Dan Marshall. Causation and fact granularity. Synthese, 199(3–4):8029– 8045, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03151-2.

Payton:2021:CIN

[1525] Jonathan D. Payton. Composition as identity, now with all the pluralities you could want. Synthese, 199(3–4):8047–8068, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03152-1.

Schulz:2021:PRD

[1526] Moritz Schulz. Practical reasoning and degrees of outright belief. Synthese, 199(3–4):8069–8090, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03153-0.

Gozzano:2021:PRD

[1527] Simone Gozzano. Phenomenal roles: a dispositional account of bodily pain. Synthese, 199(3–4):8091–8112, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03154-z.

Guo:2021:NPR

[1528] Hongyu Guo. A new paradox and the reconciliation of Lorentz and Galilean transformations. Synthese, 199(3–4):8113–8142, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03155-y.

Rucinska:2021:MIE

[1529] Zuzanna Rucińska and Shaun Gallagher. Making imagination even more embodied: imagination, constraint and epistemic relevance. Synthese, 199(3–4):8143–8170, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03156-x.

Welty:2021:IIL

[1530] Ivan Welty. Incompatibility, inconsistency, and logical analysis in Tractatus Logico–Philosophicus. Synthese, 199(3–4):8171–8186, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03157-w.

Dang:2021:SCN

[1531] Haixin Dang and Liam Kofi Bright. Scientific conclusions need not be accurate, justified, or believed by their authors. Synthese, 199(3–4): 8187–8203, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03158-9.

Mojica:2021:ANI

[1532] Laura Mojica and Melina Gastelum Vargas. The affective and normative intentionality of skilled performance: a radical embodied approach. Synthese, 199(3–4):8205–8230, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03159-8.

Faye:2021:BBQ

[1533] Jan Faye and Rasmus Jaksland. Barad, Bohr, and quantum mechanics. Synthese, 199(3–4):8231–8255, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03160-1.

Benson:2021:EVD

[1534] Jonathan Benson. The epistemic value of deliberative democracy: how far can diversity take us? Synthese, 199(3–4):8257–8279, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03161-0.

Douven:2021:STV

[1535] Igor Douven. Scoring, truthlikeness, and value. Synthese, 199(3–4): 8281–8298, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03162-z.

Correia:2021:KRG

[1536] Fabrice Correia. A kind route from grounding to fundamentality. Synthese, 199(3–4):8299–8315, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03163-y.

Moehler:2021:SJC

[1537] Michael Moehler. Strategic justice, conventionalism, and bargaining theory. Synthese, 199(3–4):8317–8334, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03164-x.

Atkinson:2021:SGR

[1538] David Atkinson and Jeanne Peijnenburg. Screening off generalized: Reichenbach’s legacy. Synthese, 199(3–4):8335–8354, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03165-w.

Pritchard:2021:ECA

[1539] Duncan Pritchard, Andrea R. English, and John Ravenscroft. Extended cognition, assistive technology and education. Synthese, 199(3–4):8355– 8377, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03166-9.

Janssen-Lauret:2021:AEM

[1540] Frederique Janssen-Lauret. Anti-essentialism, modal relativity, and alternative material-origin counterfactuals. Synthese, 199(3–4):8379– 8398, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03167-8.

Lusson:2021:RTB

[1541] Philippe A. Lusson. A regulative theory of basic intentional omissions. Synthese, 199(3–4):8399–8421, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03168-7.

Alfano:2021:VAD

[1542] Mark Alfano. Virtues for agents in directed social networks. Synthese, 199(3–4):8423–8442, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03169-6.

Barz:2021:DEA

[1543] Wolfgang Barz. The distinct existences argument revisited. Synthese, 199(3–4):8443–8463, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03170-z.

Keiser:2021:ALM

[1544] Jessica Keiser. The “All lives matter” response: QUD-shifting as epistemic injustice. Synthese, 199(3–4):8465–8483, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03171-y.

Pincock:2021:RRS

[1545] Christopher Pincock. Reichenbach, Russell and scientific realism. Synthese, 199(3–4):8485–8506, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03172-x.

Pruss:2021:NCP

[1546] Alexander R. Pruss. Non-classical probabilities invariant under symmetries. Synthese, 199(3–4):8507–8532, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03173-w. See correction

[2164].

Tarnovanu:2021:MHL

[1547] Horia Tarnovanu. Mapping higher-level causal efficacy. Synthese, 199 (3–4):8533–8554, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-021-03174-9.

Fritz:2021:FAP

[1548] James Fritz. Fitting anxiety and prudent anxiety. Synthese, 199(3–4): 8555–8578, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03175-8.

Jackson:2021:KAS

[1549] Alexander Jackson. Are knowledge ascriptions sensitive to social context? Synthese, 199(3–4):8579–8610, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03176-7.

Li:2021:NSL

[1550] Yanjun Li and Yanjing Wang. Neighborhood semantics for logic of knowing how. Synthese, 199(3–4):8611–8639, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03177-6.

Proietti:2021:DEL

[1551] Carlo Proietti and Antonio Yuste-Ginel. Dynamic epistemic logics for abstract argumentation. Synthese, 199(3–4):8641–8700, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03178-5.

Bartha:2021:MPP

[1552] Paul Bartha and C. Tyler DesRoches. Modeling the precautionary principle with lexical utilities. Synthese, 199(3–4):8701–8740, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03179-4.

Schulz:2021:SKW

[1553] Moritz Schulz. Strong knowledge, weak belief? Synthese, 199(3–4): 8741–8753, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03180-x.

Babcock:2021:ET

[1554] Gunnar Babcock and Daniel W. McShea. An externalist teleology. Synthese, 199(3–4):8755–8780, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03181-w.

Peregrin:2021:MAE

[1555] Jaroslav Peregrin and Vladimír Svoboda. Moderate anti-exceptionalism and earthborn logic. Synthese, 199(3–4):8781–8806, December 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-021-03182-
9.

Craver:2021:CRM

[1556] Carl F. Craver, Stuart Glennan, and Mark Povich. Constitutive relevance & mutual manipulability revisited. Synthese, 199(3–4):8807– 8828, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03183-8. See correction [1557].

Craver:2021:CCR

[1557] Carl F. Craver, Stuart Glennan, and Mark Povich. Correction to: Constitutive relevance & mutual manipulability revisited. Synthese, 199(3– 4):8829, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03248-8. See [1556].

Goodwyn:2021:BMH

[1558] Erik Goodwyn. Bodies and minds, heaps and syllables. Synthese, 199 (3–4):8831–8855, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-021-03184-7.

Redei:2021:MPS

[1559] Miklós Rédei and Zalán Gyenis. The Maxim of Probabilism, with special regard to Reichenbach. Synthese, 199(3–4):8857–8874, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03185-6.

Neuber:2021:RSR

[1560] Matthias Neuber. Reichenbach: scientific realist and logical empricist? Synthese, 199(3–4):8875–8897, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03186-5.

Beach:2021:EPE

[1561] Derek Beach. Evidential pluralism and evidence of mechanisms in the social sciences. Synthese, 199(3–4):8899–8919, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03187-4.

Merritt:2021:FRD

[1562] David Merritt. Feyerabend’s rule and dark matter. Synthese, 199(3–4): 8921–8942, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03188-3.

Rowbottom:2021:HCR

[1563] Darrell P. Rowbottom. How can representationalism accommodate degrees of belief? A dispositional representationalist proposal. Synthese, 199(3–4):8943–8964, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03189-2.

Brown:2021:CEA

[1564] Rachael L. Brown. Is cultural evolution always fast? Challenging the idea that cognitive gadgets would be capable of rapid and adaptive evolution. Synthese, 199(3–4):8965–8989, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03190-9.

Cumpa:2021:AMN

[1565] Javier Cumpa and Alexandre Declos. Against mereological nominalism: Reply to Effingham. Synthese, 199(3–4):8991–9011, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03191-8. See [1486].

Catala:2021:AEI

[1566] Amandine Catala, Luc Faucher, and Pierre Poirier. Autism, epistemic injustice, and epistemic disablement: a relational account of epistemic agency. Synthese, 199(3–4):9013–9039, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03192-7.

Baltag:2021:TPT

[1567] Alexandru Baltag, Soroush Rafiee Rad, and Sonja Smets. Tracking probabilistic truths: a logic for statistical learning. Synthese, 199(3–4): 9041–9087, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03193-6.

Fernandez-Prat:2021:PCS

[1568] Olga Fernández-Prat and Daniel Quesada. Philosophy and cognitive science on spatial and temporal experience. Synthese, 199(3–4):9089– 9108, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03195-4.

Hu:2021:ERD

[1569] Ivan Hu. Epistemicism and response-dependence. Synthese, 199(3–4): 9109–9131, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03196-3.

Taylor:2021:CCE

[1570] Samuel D. Taylor. Causation and cognition: an epistemic approach. Synthese, 199(3–4):9133–9160, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03197-2.

Paseau:2021:AEI

[1571] A. C. Paseau. Arithmetic, enumerative induction and size bias. Synthese, 199(3–4):9161–9184, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03198-1.

Wojtowicz:2021:DBA

[1572] Anna Wójtowicz and Krzysztof Wójtowicz. Dutch Book against Lewis. Synthese, 199(3–4):9185–9217, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03199-0.

Warren:2021:DUA

[1573] Jared Warren. Defending understanding-assent links. Synthese, 199(3–4): 9219–9236, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03200-w.

Gomez:2021:MEE

[1574] Javiera Perez Gomez. Moral encroachment and the epistemic impermissibility of (some) microaggressions. Synthese, 199(3–4):9237– 9256, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03201-9.

dosSantos:2021:EHO

[1575] César Frederico dos Santos. Enculturation and the historical origins of number words and concepts. Synthese, 199(3–4):9257–9287, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03202-8.

Duijf:2021:SOT

[1576] Hein Duijf. Should one trust experts? Synthese, 199(3–4):9289– 9312, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03203-7.

Juvshik:2021:AMD

[1577] Tim Juvshik. Artifacts and mind-dependence. Synthese, 199(3–4):9313– 9336, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03204-6.

Jacobs:2021:IET

[1578] Caspar Jacobs. Invariance or equivalence: a tale of two principles. Synthese, 199(3–4):9337–9357, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03205-5.

Garcia-Lapena:2021:TPL

[1579] Alfonso García-Lapeña. Truthlikeness for probabilistic laws. Synthese, 199(3–4):9359–9389, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03206-4.

Oddie:2021:PCA

[1580] Graham Oddie. Propositional and credal accuracy in an indeterministic world. Synthese, 199(3–4):9391–9410, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03207-3.

Stein:2021:IET

[1581] Nathanael Stein. Imagination, expectation, and “thoughts entangled in metaphors”. Synthese, 199(3–4):9411–9431, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03208-2.

Rodriguez:2021:HEP

[1582] Ángel García Rodríguez. How emotions are perceived. Synthese, 199 (3–4):9433–9461, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-021-03209-1.

Graham:2021:TT

[1583] Peter J. Graham. Typing testimony. Synthese, 199(3–4):9463–9477, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03210-8.

Jenkins:2021:RL

[1584] David Jenkins. Reasoning and its limits. Synthese, 199(3–4):9479– 9495, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03211-7.

Abbate:2021:RKC

[1585] Cheryl Abbate. On the role of knowers and corresponding epistemic role oughts. Synthese, 199(3–4):9497–9522, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03213-5.

Stradis:2021:MFA

[1586] Athamos Stradis. Memory, the fork asymmetry, and the initial state. Synthese, 199(3–4):9523–9547, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03214-4.

Jimenez-Buedo:2021:EPO

[1587] María Jiménez-Buedo and Federica Russo. Experimental practices and objectivity in the social sciences: re-embedding construct validity in the internal-external validity distinction. Synthese, 199(3–4):9549– 9579, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03215-3.

Najenson:2021:WWL

[1588] Jonathan Najenson. What have we learned about the engram? Synthese, 199(3–4):9581–9601, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03216-2.

Stewart:2021:UEF

[1589] Rush T. Stewart. Uncertainty, equality, fraternity. Synthese, 199(3–4): 9603–9619, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03217-1.

Bailey:2021:EVP

[1590] Olivia Bailey. Empathy with vicious perspectives? A puzzle about the moral limits of empathetic imagination. Synthese, 199(3–4):9621– 9647, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03219-z.

DaVee:2021:EPM

[1591] Dean Da Vee. An epistemological problem for minimalist views about composition. Synthese, 199(3–4):9649–9668, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03220-6.

Rich:2021:KKN

[1592] Patricia Rich. The key to the knowledge norm of action is ambiguity. Synthese, 199(3–4):9669–9698, December 2021. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03221-5.

Spafford:2021:EJE

[1593] Jesse Spafford. Explanation, justification, and egalitarianism. Synthese, 199(3–4):9699–9724, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03222-4.

Sagi:2021:LMD

[1594] Gil Sagi. Logic as a methodological discipline. Synthese, 199(3–4):9725– 9749, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03223-3.

Kocsis:2021:LNK

[1595] László Kocsis and Adam Tamas Tuboly. The liberation of nature and knowledge: a case study on Hans Reichenbach’s naturalism. Synthese, 199(3–4):9751–9784, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03224-2.

Lassiter:2021:SOT

[1596] Charles Lassiter and Joseph Vukov. In search of an ontology for 4E theories: from new mechanism to causal powers realism. Synthese, 199(3–4): 9785–9808, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03225-1.

Torza:2021:QM

[1597] Alessandro Torza. Quantum metametaphysics. Synthese, 199(3–4):9809– 9833, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03226-0.

Wilson:2021:UIV

[1598] Alan T. Wilson. Unity of the intellectual virtues. Synthese, 199(3–4): 9835–9854, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03227-z.

Luchetti:2021:CTE

[1599] Michele Luchetti. Coordination in theory extension: how Reichenbach can help us understand endogenization in evolutionary biology. Synthese, 199(3–4):9855–9880, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03228-y.

Groff:2021:CCP

[1600] Ruth Porter Groff. Conceptualizing causal powers: activity, capacity, essence, necessitation. Synthese, 199(3–4):9881–9896, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03229-x.

Lin:2021:VLS

[1601] Ying-Tung Lin and Vilius Dranseika. The variety and limits of selfexperience and identification in imagination. Synthese, 199(3–4):9897– 9926, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03230-4.

Lee:2021:BLL

[1602] Samuel Lee. Building low level causation out of high level causation. Synthese, 199(3–4):9927–9955, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03231-3.

Maziarz:2021:REC

[1603] Mariusz Maziarz. Resolving empirical controversies with mechanistic evidence. Synthese, 199(3–4):9957–9978, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03232-2.

Sterkenburg:2021:NFL

[1604] Tom F. Sterkenburg and Peter D. Grünwald. The no-free-lunch theorems of supervised learning. Synthese, 199(3–4):9979–10015, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03233-1.

Liggins:2021:SHO

[1605] David Liggins. Should a higher-order metaphysician believe in properties? Synthese, 199(3–4):10017–10037, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03234-0.

Verreault-Julien:2021:FIP

[1606] Philippe Verreault-Julien. Factive inferentialism and the puzzle of model-based explanation. Synthese, 199(3–4):10039–10057, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03235-z.

Smith:2021:PII

[1607] David Woodruff Smith. Phenomenal intentionality, inner awareness, and the given. Synthese, 199(3–4):10059–10076, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03236-y.

Winstanley:2021:PTR

[1608] M. A. Winstanley. A psychological theory of reasoning as logical evidence: a Piagetian perspective. Synthese, 199(3–4):10077–10108, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03237-x.

Segundo-Ortin:2021:NMN

[1609] Miguel Segundo-Ortin and Manuel Heras-Escribano. Neither mindful nor mindless, but minded: habits, ecological psychology, and skilled performance. Synthese, 199(3–4):10109–10133, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03238-w.

Werkhoven:2021:NKM

[1610] Sander Werkhoven. Natural kinds of mental disorder. Synthese, 199 (3–4):10135–10165, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03239-9.

Aberdein:2021:VTM

[1611] Andrew Aberdein, Colin Jakob Rittberg, and Fenner Stanley Tanswell. Virtue theory of mathematical practices: an introduction. Synthese, 199 (3–4):10167–10180, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03240-2.

Cato:2021:IPA

[1612] Susumu Cato, Eric Rémila, and Philippe Solal. Infinite-population approval voting: A proposal. Synthese, 199(3–4):10181–10209, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03242-0.

Hireche:2021:SAL

[1613] Salim Hirèche, Niels Linnemann, Robert Michels, and Lisa Vogt. The strong arm of the law: a unified account of necessary and contingent laws of nature. Synthese, 199(3–4):10211–10252, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03243-z.

Henderson:2021:RSR

[1614] Leah Henderson and Alexander Gebharter. The role of source reliability in belief polarisation. Synthese, 199(3–4):10253–10276, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03244-y.

Noordhof:2021:CSA

[1615] Paul Noordhof and Ema Sullivan-Bissett. The clinical significance of anomalous experience in the explanation of monothematic delusions. Synthese, 199(3–4):10277–10309, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03245-x.

Backes:2021:CGG

[1616] Marvin Backes. Can groups be genuine believers? The argument from interpretationism. Synthese, 199(3–4):10311–10329, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03246-w.

Lipski:2021:AEE

[1617] Jonas Lipski. Austrian economics without extreme apriorism: A critical reply. Synthese, 199(3–4):10331–10341, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03249-7.

Lehtinen:2021:EBG

[1618] Aki Lehtinen. The epistemic benefits of generalisation in modelling I: Systems and applicability. Synthese, 199(3–4):10343–10370, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03250-0.

Simion:2021:KR

[1619] Mona Simion. Knowledge and reasoning. Synthese, 199(3–4):10371– 10388, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03251-z.

Lampert:2021:HCW

[1620] Fabio Lampert and Pedro Merlussi. How (not) to construct worlds with responsibility. Synthese, 199(3–4):10389–10413, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03252-y.

Leng:2021:MSE

[1621] Mary Leng. Models, structures, and the explanatory role of mathematics in empirical science. Synthese, 199(3–4):10415–10440, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03253-x.

Dresow:2021:EAE

[1622] Max Dresow. Explaining the apocalypse: the end-permian mass extinction and the dynamics of explanation in geohistory. Synthese, 199 (3–4):10441–10474, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03254-w.

Carter:2021:FRM

[1623] Jessica Carter. “Free rides” in mathematics. Synthese, 199(3–4):10475– 10498, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03255-9.

Niiniluoto:2021:APL

[1624] Ilkka Niiniluoto. Approaching probabilistic laws. Synthese, 199(3–4): 10499–10519, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-021-03256-8.

Friend:2021:MVA

[1625] Toby Friend. Megarian variable actualism. Synthese, 199(3–4):10521– 10541, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03257-7.

OShea:2021:WMG

[1626] James R. O’Shea. What is the myth of the given? Synthese, 199 (3–4):10543–10567, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03258-6.

Andreoletti:2021:FAW

[1627] Giacomo Andreoletti and Giuseppe Spolaore. The future ain’t what it used to be. Synthese, 199(3–4):10569–10585, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03259-5.

Vogt:2021:VVI

[1628] Thomas Vogt. The value of vague ideas in the development of the periodic system of chemical elements. Synthese, 199(3–4):10587–10614, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-021-03260-y.

Vessonen:2021:CEO

[1629] Elina Vessonen. Conceptual engineering and operationalism in psychology. Synthese, 199(3–4):10615–10637, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03261-x.

Manero:2021:QPR

[1630] Jorge Manero. Quantum pointillism with relational identity. Synthese, 199(3–4):10639–10666, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03262-w.

Dupre:2021:CPP

[1631] John Dupré. Causally powerful processes. Synthese, 199(3–4):10667– 10683, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03263-9.

Busse:2021:ACS

[1632] Ralf Busse. Asymmetry cannot solve the circularity/regress problem of property structuralism. Synthese, 199(3–4):10685–10720, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03264-8.

Stern:2021:ICN

[1633] Reuben Stern. Interventionist counterfactuals and the nearness of worlds. Synthese, 199(3–4):10721–10737, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03265-7.

Wyatt:2021:NDM

[1634] Jeremy Wyatt. The nature of disagreement: matters of taste and environs. Synthese, 199(3–4):10739–10767, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03266-6. See cor-

rection [1635].

Wyatt:2021:CND

[1635] Jeremy Wyatt. Correction to: The nature of disagreement: matters of taste and environs. Synthese, 199(3–4):10769, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03357-4. See

[1634].

Corkum:2021:SME

[1636] Phil Corkum. Salience and metaphysical explanation. Synthese, 199 (3–4):10771–10792, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03267-5.

Knutsson:2021:MVL

[1637] Simon Knutsson. Many-valued logic and sequence arguments in value theory. Synthese, 199(3–4):10793–10825, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03268-4.

Schurz:2021:RBA

[1638] Gerhard Schurz. Reichenbach’s best alternative account to the problem of induction. Synthese, 199(3–4):10827–10838, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03269-3.

Cinti:2021:HLQ

[1639] Enrico Cinti and Marco Sanchioni. Humeanism in light of quantum gravity. Synthese, 199(3–4):10839–10863, December 2021. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03270-w.

Neto:2021:ISP

[1640] Celso Neto. From idealizations to social practices in science: the case of phylogenetic trees. Synthese, 199(3–4):10865–10884, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03271-9.

Brunet:2021:LC

[1641] T. D. P. Brunet. Local causation. Synthese, 199(3–4):10885–10908, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03272-8.

Strawson:2021:MSG

[1642] Galen Strawson. The mechanism-the secret-of the given. Synthese, 199 (3–4):10909–10928, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03273-7.

Patel:2021:SVC

[1643] Shivam Patel. From speech to voice: on the content of inner speech. Synthese, 199(3–4):10929–10952, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03274-6.

Iacona:2021:CF

[1644] Andrea Iacona and Samuele Iaquinto. Credible futures. Synthese, 199 (3–4):10953–10968, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03275-5.

Rubin:2021:WAA

[1645] Mark Rubin. When to adjust alpha during multiple testing: a consideration of disjunction, conjunction, and individual testing. Synthese, 199 (3–4):10969–11000, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03276-4.

Hall:2021:PPL

[1646] Geoffrey Hall. Phenomenal properties are luminous properties. Synthese, 199(3–4):11001–11022, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03277-3.

Fairhurst:2021:WDM

[1647] Jordi Fairhurst. Wittgenstein, deflationism and moral entities. Synthese, 199(3–4):11023–11050, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03278-2.

Sytsma:2021:MWC

[1648] Justin Sytsma, Ryan Muldoon, and Shaun Nichols. The meta-wisdom of crowds. Synthese, 199(3–4):11051–11074, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03279-1.

Seide:2021:PCB

[1649] Ansgar Seide. Points of convergence between logical empiricism and inductive metaphysics: Hans Reichenbach and Erich Becher in comparison. Synthese, 199(3–4):11075–11107, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03280-8.

Zanotti:2021:PBP

[1650] Giacomo Zanotti. Physicalism and the burden of parsimony. Synthese, 199(3–4):11109–11132, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03281-7.

Caamano-Alegre:2021:GHF

[1651] María Caamaño-Alegre. On glasses half full or half empty: understanding framing effects in terms of default implicatures. Synthese, 199 (3–4):11133–11159, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03282-6.

Zhao:2021:MBS

[1652] Xinkan Zhao. Metasemantics and Boydian synthetic moral naturalism. Synthese, 199(3–4):11161–11178, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03283-5.

Dardashti:2021:CES

[1653] Radin Dardashti, Mathias Frisch, and Giovanni Valente. Correction to: Editorial: Symmetries and asymmetries in physics. Synthese, 199 (3–4):11179–11180, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03284-4. See [1208].

Fletcher:2021:QIE

[1654] Samuel C. Fletcher and David E. Taylor. Quantum indeterminacy and the eigenstate-eigenvalue link. Synthese, 199(3–4):11181–11212, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03285-3.

Hundertmark:2021:MDV

[1655] Fabian Hundertmark and Steven Kindley. Making a difference in virtue epistemology. Synthese, 199(3–4):11213–11229, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03286-2.

Gu:2021:TLA

[1656] Zhiwei Gu. The time-lag argument and simultaneity. Synthese, 199 (3–4):11231–11248, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03287-1.

Dellsen:2021:USP

[1657] Finnur Dellsén. Understanding scientific progress: the noetic account. Synthese, 199(3–4):11249–11278, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03289-z.

Pawlowitsch:2021:SMB

[1658] Christina Pawlowitsch. Strategic manipulation in Bayesian dialogues. Synthese, 199(3–4):11279–11303, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03288-0.

Calosi:2021:GGG

[1659] Claudio Calosi. Gappy, glutty, glappy. Synthese, 199(3–4):11305– 11321, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03291-5.

Gonnerman:2021:KHP

[1660] Chad Gonnerman, Kaija Mortensen, and Jacob Robbins. KNOWING HOW as a philosophical hybrid. Synthese, 199(3–4):11323–11354, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03292-4.

Boult:2021:SEB

[1661] Cameron Boult. Standing to epistemically blame. Synthese, 199(3– 4):11355–11375, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-021-03293-3.

Pino:2021:GEC

[1662] Dani Pino. Group (epistemic) competence. Synthese, 199(3–4):11377– 11396, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03294-2.

Gladziejewski:2021:PJB

[1663] Paweł Gładziejewski. Perceptual justification in the Bayesian brain: a foundherentist account. Synthese, 199(3–4):11397–11421, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03295-1.

Lundgren:2021:EMD

[1664] Björn Lundgren. Ethical machine decisions and the input-selection problem. Synthese, 199(3–4):11423–11443, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03296-0.

Siscoe:2021:GPE

[1665] Robert Weston Siscoe. Grounding and a priori epistemology: challenges for conceptualism. Synthese, 199(3–4):11445–11463, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03297-z.

Cevolani:2021:ADP

[1666] Gustavo Cevolani and Roberto Festa. Approaching deterministic and probabilistic truth: a unified account. Synthese, 199(3–4):11465–11489, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-021-03298-y.

Liu:2021:RID

[1667] Michelle Liu. Revelation and the intuition of dualism. Synthese, 199 (3–4):11491–11515, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03299-x.

Rytila:2021:SCM

[1668] Jenni Rytilä. Social constructivism in mathematics? The promise and shortcomings of Julian Cole’s institutional account. Synthese, 199(3– 4):11517–11540, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-021-03300-7.

Frances:2021:MBA

[1669] Bryan Frances. Metaphysics, bullshit, and the analysis of philosophical problems. Synthese, 199(3–4):11541–11554, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03301-6.

Riggs:2021:DFT

[1670] Jared Riggs. Deflating the functional turn in conceptual engineering. Synthese, 199(3–4):11555–11586, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03302-5.

Lean:2021:BOM

[1671] Oliver M. Lean. Are bio-ontologies metaphysical theories? Synthese, 199(3–4):11587–11608, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03303-4.

Facchin:2021:PPA

[1672] Marco Facchin. Predictive processing and anti-representationalism. Synthese, 199(3–4):11609–11642, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03304-3.

Beisbart:2021:OTT

[1673] Claus Beisbart. Opacity thought through: on the intransparency of computer simulations. Synthese, 199(3–4):11643–11666, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03305-2.

Pincock:2021:DPL

[1674] Christopher Pincock. The derivation of Poiseuille’s law: heuristic and explanatory considerations. Synthese, 199(3–4):11667–11687, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03306-1.

Clemente:2021:DED

[1675] Noel L. Clemente. Degrees of epistemic dependence: an extension of Pritchard’s response to epistemic situationism. Synthese, 199(3–4): 11689–11705, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-021-03307-0.

Ramsey:2021:WEM

[1676] William M. Ramsey. What eliminative materialism isn’t. Synthese, 199 (3–4):11707–11728, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03309-y.

Vignero:2021:DRF

[1677] Leander Vignero and Sylvia Wenmackers. Degrees of riskiness, falsifiability, and truthlikeness. Synthese, 199(3–4):11729–11764, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03310-5.

Gottlieb:2021:CWN

[1678] Joseph Gottlieb and Ali Rezaei. Correction to: When nothing looks blue. Synthese, 199(3–4):11765–11766, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03218-0. See [1282].

Boge:2021:CPH

[1679] Florian J. Boge and Christian Zeitnitz. Correction to: Polycratic hierarchies and networks: what simulation-modeling at the LHC can teach us about the epistemology of simulation. Synthese, 199(3–4):11767– 11768, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03247-9. See [1183].

Pruss:2021:CWF

[1680] Alexander R. Pruss. A classical way forward for the regularity and normalization problems. Synthese, 199(5–6):11769–11792, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03311-4.

Mitchell:2021:ASM

[1681] Jonathan Mitchell. Affective shifts: mood, emotion and well-being. Synthese, 199(5–6):11793–11820, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03312-3.

Estrada-Gonzalez:2021:RRL

[1682] Luis Estrada-González and Fernando Cano-Jorge. Revisiting Reichenbach’s logic. Synthese, 199(5–6):11821–11845, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03313-2.

Munro:2021:WFI

[1683] Daniel Munro and Margot Strohminger. Are we free to imagine what we choose? Synthese, 199(5–6):11847–11864, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03314-1.

Turyn:2021:DM

[1684] Matthew Turyn. On dispositional masks. Synthese, 199(5–6):11865– 11886, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03315-0. See correction [2440].

Ritson:2021:CMM

[1685] Sophie Ritson. Creativity and modelling the measurement process of the Higgs self-coupling at the LHC and HL-LHC. Synthese, 199(5–6):11887– 11911, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03317-y.

Waszek:2021:CMC

[1686] David Waszek and Dirk Schlimm. Calculus as method or calculus as rules? Boole and Frege on the aims of a logical calculus. Synthese, 199 (5–6):11913–11943, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03318-x.

Mulder:2021:ARA

[1687] Jesse M. Mulder. ‘Animals run about the world in all sorts of paths’: varieties of indeterminism. Synthese, 199(5–6):11945–11961, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03319-w.

Behounek:2021:GSC

[1688] Libor Behounek and Ondrej Majer. A graded semantics for counterfactuals. Synthese, 199(5–6):11963–11994, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03320-3.

Szigeti:2021:EIJ

[1689] András Szigeti. Emotions as indeterminate justifiers. Synthese, 199 (5–6):11995–12017, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03321-2.

Nunziante:2021:TVT

[1690] Antonio M. Nunziante. Tropes variations: the topic of particulars beyond Sellars’s myth of the given. Synthese, 199(5–6):12019–12043, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03322-1.

Martina:2021:CVO

[1691] Giulia Martina. Contextual variation and objectivity in olfactory perception. Synthese, 199(5–6):12045–12071, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03323-0.

Gross:2021:BMC

[1692] Fridolin Gross. Between mechanical clocks and emergent flocks: complexities in biology. Synthese, 199(5–6):12073–12102, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03324-z.

Perrine:2021:PAE

[1693] Timothy Perrine. A puzzle about epistemic value and steps towards a solution. Synthese, 199(5–6):12103–12119, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03325-y.

Muller:2021:UC

[1694] Anselm Winfried Müller. Understanding causation. Synthese, 199(5– 6):12121–12153, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-021-03326-x.

Koon:2021:EED

[1695] Justis Koon. The epistemology of evolutionary debunking. Synthese, 199(5–6):12155–12176, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03327-w.

Crane:2021:TAN

[1696] Judith K. Crane. Two approaches to natural kinds. Synthese, 199 (5–6):12177–12198, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03328-9.

dosSantos:2021:ITC

[1697] César Frederico dos Santos. Intuitions, theory choice and the ameliorative character of logical theories. Synthese, 199(5–6):12199–12223, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03329-8.

Borghini:2021:DFC

[1698] Andrea Borghini, Nicola Piras, and Beatrice Serini. Defective food concepts. Synthese, 199(5–6):12225–12249, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03330-1.

DiBello:2021:WSE

[1699] Marcello Di Bello. When statistical evidence is not specific enough. Synthese, 199(5–6):12251–12269, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03331-0.

Allori:2021:WF

[1700] Valia Allori. Wave-functionalism. Synthese, 199(5–6):12271–12293, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03332-z.

vanBasshuysen:2021:RGI

[1701] Philippe van Basshuysen. Rationality in games and institutions. Synthese, 199(5–6):12295–12314, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03333-y.

Reijula:2021:SCM

[1702] Samuli Reijula. Social categories in the making: construction or recruitment? Synthese, 199(5–6):12315–12330, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03334-x.

Simpson:2021:GLC

[1703] James Simpson. Getting a little closure for closure. Synthese, 199 (5–6):12331–12361, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03335-w.

Gasparri:2021:PDP

[1704] Luca Gasparri. Proportional dispositional predicates. Synthese, 199 (5–6):12363–12383, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03336-9.

Davies:2021:CCI

[1705] Alex Davies. Communicating in contextual ignorance. Synthese, 199 (5–6):12385–12405, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03337-8.

Pettigrew:2021:PEV

[1706] Richard Pettigrew. On the pragmatic and epistemic virtues of inference to the best explanation. Synthese, 199(5–6):12407–12438, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03338-7.

Luque:2021:MPH

[1707] Victor J. Luque and Lorenzo Baravalle. The mirror of physics: on how the Price equation can unify evolutionary biology. Synthese, 199(5–6):12439– 12462, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03339-6.

vanHateren:2021:MRS

[1708] J. H. van Hateren. A mechanism that realizes strong emergence. Synthese, 199(5–6):12463–12483, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03340-z.

Borghi:2021:ACS

[1709] Anna M. Borghi. Affordances, context and sociality. Synthese, 199 (5–6):12485–12515, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-018-02044-1.

Berskyte:2021:JLE

[1710] Justina Berskyte. Jeremy is a … Expressive-relativism and expressives in predicative positions. Synthese, 199(5–6):12517–12539, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03341-y.

Goldberg:2021:WET

[1711] Sanford C. Goldberg. What epistemologists of testimony should learn from philosophers of science. Synthese, 199(5–6):12541–12559, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03342-x.

Barrio:2021:AET

[1712] Eduardo Alejandro Barrio, Federico Pailos, and Joaquín Toranzo Calderón. Anti-exceptionalism, truth and the BA-plan. Synthese, 199 (5–6):12561–12586, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03343-w.

Hesni:2021:GI

[1713] Samia Hesni. Generics as instructions. Synthese, 199(5–6):12587– 12602, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03344-9.

Sander:2021:UFN

[1714] Thorsten Sander. Understanding Frege’s notion of presupposition. Synthese, 199(5–6):12603–12624, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03346-7.

Richardson:2021:HRR

[1715] Alan W. Richardson. Hans Reichenbach, radio philosopher: a preliminary report. Synthese, 199(5–6):12625–12641, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03345-8.

Lee:2021:TFQ

[1716] Kok Yong Lee. Tinbergen’s four questions and the debate between scientific realism and selectionism. Synthese, 199(5–6):12643–12661, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03347-6.

Passmann:2021:SPP

[1717] Robert Passmann. Should pluralists be pluralists about pluralism? Synthese, 199(5–6):12663–12682, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03348-5.

Fritts:2021:REA

[1718] Megan Fritts. Reasons explanations (of actions) as structural explanations. Synthese, 199(5–6):12683–12704, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03349-4.

Bechtel:2021:ABM

[1719] William Bechtel and Andrew Bollhagen. Active biological mechanisms: transforming energy into motion in molecular motors. Synthese, 199 (5–6):12705–12729, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03350-x.

Chirimuuta:2021:RTC

[1720] M. Chirimuuta. Reflex theory, cautionary tale: misleading simplicity in early neuroscience. Synthese, 199(5–6):12731–12751, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03351-w.

Fresco:2021:IC

[1721] Nir Fresco, B. Jack Copeland, and Marty J. Wolf. The indeterminacy of computation. Synthese, 199(5–6):12753–12775, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03352-9.

Salmon:2021:PAS

[1722] Nathan Salmón. A paradox about sets of properties. Synthese, 199 (5–6):12777–12793, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03353-8.

McSweeney:2021:CCL

[1723] Michaela M. McSweeney. The cost of closure: logical realism, antiexceptionalism, and theoretical equivalence. Synthese, 199(5–6):12795– 12817, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03354-7.

Bruineberg:2021:MAR

[1724] Jelle Bruineberg, Ludovic Seifert, and Julian Kiverstein. Metastable attunement and real-life skilled behavior. Synthese, 199(5–6):12819– 12842, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03355-6.

Legault:2021:NNR

[1725] Mylène Legault, Jean-Nicolas Bourdon, and Pierre Poirier. From neurodiversity to neurodivergence: the role of epistemic and cognitive marginalization. Synthese, 199(5–6):12843–12868, December 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-021-03356-
5.

Gessell:2021:MPA

[1726] Bryce Gessell, Benjamin Geib, and Felipe De Brigard. Multivariate pattern analysis and the search for neural representations. Synthese, 199 (5–6):12869–12889, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03358-3.

teVrugt:2021:MTE

[1727] Michael te Vrugt. The mereology of thermodynamic equilibrium. Synthese, 199(5–6):12891–12921, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03359-2.

DAmbrosio:2021:VPA

[1728] Justin D’Ambrosio and Daniel Stoljar. Vendler’s puzzle about imagination. Synthese, 199(5–6):12923–12944, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03360-9.

Atkins:2021:PEG

[1729] Richard Kenneth Atkins. A Peircean examination of Gettier’s two cases. Synthese, 199(5–6):12945–12961, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03361-8.

Lewis:2021:KNA

[1730] Max Lewis. The knowledge norm of assertion: keep it simple. Synthese, 199(5–6):12963–12984, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03362-7.

Pantazakos:2021:PES

[1731] Themistoklis Pantazakos. Problems of empirical solutions to the theoryladenness of observation. Synthese, 199(5–6):12985–13007, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03363-6.

Puy:2021:EMS

[1732] Nemesio G. C. Puy. Eliminating Martin’s substratum-trope categorial ontology. Synthese, 199(5–6):13009–13033, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03364-5.

Lohr:2021:CEC

[1733] Guido Löhr. Commitment engineering: conceptual engineering without representations. Synthese, 199(5–6):13035–13052, December 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-021-03365-

Wiegmann:2021:TLM

[1734] Alex Wiegmann and Emanuel Viebahn. True lies and Moorean redundancy. Synthese, 199(5–6):13053–13066, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03366-3.

Schurz:2021:WCL

[1735] Gerhard Schurz. Why classical logic is privileged: justification of logics based on translatability. Synthese, 199(5–6):13067–13094, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03367-2.

Cartwright:2021:RVN

[1736] Nancy Cartwright. Rigour versus the need for evidential diversity. Synthese, 199(5–6):13095–13119, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03368-1.

Kind:2021:CIU

[1737] Amy Kind. Can imagination be unconscious? Synthese, 199(5–6):13121– 13141, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03369-0.

Russell:2021:SRW

[1738] Jodie Louise Russell. Stuck on repeat: Why do we continue to ruminate? Synthese, 199(5–6):13143–13162, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03370-7.

Behrens:2021:NNF

[1739] Singa Behrens. No normative free lunch: relevance and the autonomy of the normative domain. Synthese, 199(5–6):13163–13186, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03371-6.

Carmichael:2021:NPI

[1740] Chad Carmichael. A new problem for internalism. Synthese, 199(5– 6):13187–13199, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-021-03372-5.

LeBrun:2021:WEC

[1741] Alex LeBrun. What are empirical consequences? On dispensability and composite objects. Synthese, 199(5–6):13201–13223, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03373-4.

DeBenedetto:2021:TCW

[1742] Matteo De Benedetto. Taming conceptual wanderings: Wilson– structuralism. Synthese, 199(5–6):13225–13246, December 2021. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-021-03374-
3.

Fletcher:2021:TQL

[1743] Samuel C. Fletcher and David E. Taylor. Two quantum logics of indeterminacy. Synthese, 199(5–6):13247–13281, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03375-2.

Morris:2021:CQH

[1744] Sean Morris. Carnap, Quine, and the Humean condition. Synthese, 199 (5–6):13283–13312, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03376-1.

Dziurosz-Serafinowicz:2021:VCF

[1745] Patryk Dziurosz-Serafinowicz and Dominika Dziurosz-Serafinowicz. The value of cost-free uncertain evidence. Synthese, 199(5–6):13313–13343, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-021-03377-0.

Gisin:2021:IPI

[1746] Nicolas Gisin. Indeterminism in physics and intuitionistic mathematics. Synthese, 199(5–6):13345–13371, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03378-z.

Harris:2021:VDW

[1747] Keith Raymond Harris. Video on demand: what deepfakes do and how they harm. Synthese, 199(5–6):13373–13391, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03379-y.

Veit:2021:AT

[1748] Walter Veit. Agential thinking. Synthese, 199(5–6):13393–13419, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03380-5.

Winokur:2021:DFP

[1749] Benjamin Winokur. Davidson, first-person authority, and direct selfknowledge. Synthese, 199(5–6):13421–13440, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03381-4.

Gomori:2021:MER

[1750] Márton Gömöri and Gábor Hofer-Szabó. On the meaning of EPR’s reality criterion. Synthese, 199(5–6):13441–13469, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03382-3.

Cabrera:2021:EAI

[1751] Frank Cabrera. Is epistemic anxiety an intellectual virtue? Synthese, 199(5–6):13471–13495, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03383-2.

Engelhard:2021:DMT

[1752] Kristina Engelhard. Dispositions manifest themselves: an identity theory of properties. Synthese, 199(5–6):13497–13522, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-021-03385-0.

Leonard:2021:BRI

[1753] Nick Leonard. Belief and rational indeterminacy. Synthese, 199(5– 6):13523–13542, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-021-03386-z.

Kosterec:2021:MF

[1754] Milos Kosterec. On modality in fiction. Synthese, 199(5–6):13543– 13567, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03388-x.

Predelli:2021:DCC

[1755] Stefano Predelli. Disambiguation in conversation: the case of disambiguating parentheticals. Synthese, 199(5–6):13569–13582, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03389-w.

Munch-Jurisic:2021:LWA

[1756] Ditte Marie Munch-Jurisic. Lost for words: anxiety, well-being, and the costs of conceptual deprivation. Synthese, 199(5–6):13583–13600, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-021-03390-3.

Naar:2021:FE

[1757] Hichem Naar. The fittingness of emotions. Synthese, 199(5–6):13601– 13619, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03391-2.

Rizza:2021:MPS

[1758] Davide Rizza. Mathematical problem-solving in scientific practice. Synthese, 199(5–6):13621–13641, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03392-1.

Brenner:2021:FDE

[1759] Andrew Brenner. Four-dimensionalism, eternalism, and deprivationist accounts of the evil of death. Synthese, 199(5–6):13643–13660, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03393-0.

Kocurek:2021:LT

[1760] Alexander W. Kocurek. Logic talk. Synthese, 199(5–6):13661–13688, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-021-03394-z.

Radzvilas:2021:BSW

[1761] Mantas Radzvilas, William Peden, and Francesco De Pretis. A battle in the statistics wars: a simulation-based comparison of Bayesian, Frequentist and Williamsonian methodologies. Synthese, 199(5–6):13689– 13748, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03395-y.

Summa:2021:HFG

[1762] Michela Summa. How are fictions given? Conjoining the ‘artifactual theory’ and the ‘imaginary-object theory’. Synthese, 199(5–6):13749– 13769, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03396-x.

Ohlhorst:2021:EAL

[1763] Jakob Ohlhorst. Epistemic austerity: limits to entitlement. Synthese, 199(5–6):13771–13787, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03397-w.

Duerr:2021:GRC

[1764] Patrick M. Duerr and Claudio Calosi. The general-relativistic case for super-substantivalism. Synthese, 199(5–6):13789–13822, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03398-9.

Hirvonen:2021:MIS

[1765] Ilmari Hirvonen, Rami Koskinen, and Ilkka Pättiniemi. Modal inferences in science: a tale of two epistemologies. Synthese, 199(5–6):13823– 13843, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03399-8.

Loughrist:2021:DGP

[1766] Tim Loughrist. Defeaters and the generality problem. Synthese, 199 (5–6):13845–13860, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03400-4.

Reis-Dennis:2021:LWE

[1767] Samuel Reis-Dennis and Vida Yao. ‘I love women’: an explicit explanation of implicit bias test results. Synthese, 199(5–6):13861–13882, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-021-03401-3.

Buccella:2021:PPI

[1768] Alessandra Buccella. The problem of perceptual invariance. Synthese, 199(5–6):13883–13905, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03402-2.

Bailey:2021:DFF

[1769] Andrew M. Bailey and Amy Seymour. In defense of flip-flopping. Synthese, 199(5–6):13907–13924, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03403-1.

Chislenko:2021:HCB

[1770] Eugene Chislenko. How can belief be akratic? Synthese, 199(5–6):13925– 13948, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03404-0.

Spanos:2021:BGT

[1771] Aris Spanos. Bernoulli’s golden theorem in retrospect: error probabilities and trustworthy evidence. Synthese, 199(5–6):13949–13976, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03405-z.

Nanay:2021:IOE

[1772] Bence Nanay. Imagining one experience to be another. Synthese, 199 (5–6):13977–13991, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03406-y.

Fresco:2021:LAF

[1773] Nir Fresco. Long-arm functional individuation of computation. Synthese, 199(5–6):13993–14016, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03407-x.

Rusanen:2021:ACF

[1774] Anna-Mari Rusanen, Otto Lappi, Jesse Kuokkanen, and Jami Pekkanen. Action control, forward models and expected rewards: representations in reinforcement learning. Synthese, 199(5–6):14017–14033, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-021-03408-w.

Atiq:2021:AKV

[1775] Emad H. Atiq. Acquaintance, knowledge, and value. Synthese, 199 (5–6):14035–14062, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-021-03409-9.

Hornett:2021:PCD

[1776] William Hornett. Perceptual capacities, discrimination, and the senses. Synthese, 199(5–6):14063–14085, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03410-2.

Wang:2021:FDS

[1777] En Wang. Four dilemmas of the “Superstring theory” and new responses from the “Singularity theory” in the view of information ontology. Synthese, 199(5–6):14087–14095, December 2021. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-021-03411-1.

Funkhouser:2021:EPL

[1778] Eric Funkhouser. Evolutionary psychology, learning, and belief signaling: design for natural and artificial systems. Synthese, 199(5–6):14097– 14119, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-021-03412-0.

Roski:2021:DER

[1779] Stefan Roski. In defence of explanatory realism. Synthese, 199(5–6): 14121–14141, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-021-03413-z.

Tramacere:2022:TBD

[1780] Antonella Tramacere and Colin Allen. Temporal binding: digging into animal minds through time perception. Synthese, 200(1):1–24, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03456-w.

Bradley:2022:ADD

[1781] Richard Bradley. Attitudes, deliberation and decisions. Synthese, 200 (1):1–18, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03468-6.

Viejo:2022:CDS

[1782] José Manuel Viejo. Complex demonstratives, singular thought, and belief attributions. Synthese, 200(1):1–27, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03463-x.

Loew:2022:WFM

[1783] Christian Loew and Andreas Hüttemann. Are we free to make the laws? Synthese, 200(1):1–16, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03460-0.

Pan:2022:WSA

[1784] Shen Pan. What is so special about episodic memory: lessons from the system-experience distinction. Synthese, 200(1):1–26, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03500-9.

Hill:2022:LS

[1785] Jesse Hill. On luck and significance. Synthese, 200(1):1–18, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03533-0.

Emmerson:2022:USP

[1786] Nicholas Emmerson. Understanding and scientific progress: lessons from epistemology. Synthese, 200(1):1–18, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03501-8.

Berto:2022:WIS

[1787] Francesco Berto. Williamson on indicatives and suppositional heuristics. Synthese, 200(1):1–12, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03518-z.

Miracchi:2022:RMS

[1788] Lisa Miracchi and J. Adam Carter. Refitting the mirrors: on structural analogies in epistemology and action theory. Synthese, 200(1): 1–28, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03462-y.

Kinney:2022:DTT

[1789] David Kinney. Diachronic trends in the topic distributions of formal epistemology abstracts. Synthese, 200(1):1–34, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03466-8.

Hazlett:2022:VCK

[1790] Allan Hazlett. The value of common knowledge. Synthese, 200(1): 1–18, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03457-9.

Lorenzetti:2022:SAB

[1791] Lorenzo Lorenzetti. Structuralist approaches to Bohmian mechanics. Synthese, 200(1):1–15, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03489-1.

Rappe:2022:PMC

[1792] Sofiia Rappe. Predictive minds can think: addressing generality and surface compositionality of thought. Synthese, 200(1):1–22, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03502-7.

Boyd:2022:TSE

[1793] Kenneth Boyd. Trusting scientific experts in an online world. Synthese, 200(1):1–21, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03592-3.

Conix:2022:PRP

[1794] Stijn Conix, Olivier Lemeire, and Pei-Shan Chi. The public relevance of philosophy. Synthese, 200(1):1–28, February 2022. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03546-9.

Spinney:2022:LFL

[1795] Oliver Thomas Spinney. Logical form and logical space in Wittgenstein’s Tractatus. Synthese, 200(1):1–23, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03470-y.

Popa:2022:GCR

[1796] Elena Popa. Getting counterfactuals right: the perspective of the causal reasoner. Synthese, 200(1):1–18, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03492-6.

Giberman:2022:OT

[1797] Daniel Giberman. Ostrich tropes. Synthese, 200(1):1–25, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03494-4.

Hamri:2022:CF

[1798] Soufiane Hamri. Causal fundamentality. Synthese, 200(1):1–13, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03583-4.

Elgin:2022:DP

[1799] Catherine Z. Elgin. Disagreement in philosophy. Synthese, 200(1): 1–16, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03535-y.

Kincaid:2022:AES

[1800] Harold Kincaid and Julie Zahle. Are ABM explanations in the social sciences inevitably individualist? Synthese, 200(1):1–22, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03465-9.

White:2022:RAS

[1801] Ben White. A reductive analysis of statements about universals. Synthese, 200(1):1–21, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03573-6.

Farr:2022:CAT

[1802] Matt Farr. Conventionalism about time direction. Synthese, 200(1): 1–21, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03540-1.

Dahl:2022:FPP

[1803] Niklas Dahl. A fixed-point problem for theories of meaning. Synthese, 200(1):1–15, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03559-4.

Ritchie:2022:RWV

[1804] J. Brendan Ritchie. Recognizing why vision is inferential. Synthese, 200 (1):1–27, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03508-1.

Kulakova:2022:ORE

[1805] Eugenia Kulakova and Stefan Rinner. Oswald revisited: the effect of focus and context. Synthese, 200(1):1–15, February 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03478-4. See cor-

rection [1932].

Schroeren:2022:WRD

[1806] David Schroeren. Wavefunction realism does not ‘privilege position’. Synthese, 200(1):1–14, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03525-0.

Martin:2022:ICS

[1807] James V. Martin. Indeterminacy, coincidence, and “Sourcing Newness” in mathematical research. Synthese, 200(1):1–23, February 2022. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03493-
5.

Ardourel:2022:BMD

[1808] Vincent Ardourel. Brownian motion from a deterministic system of particles. Synthese, 200(1):1–15, February 2022. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03577-2.

Dethier:2022:SAC

[1809] Corey Dethier. Science, assertion, and the common ground. Synthese, 200(1):1–19, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03580-7.

Shapiro:2022:WLD

[1810] Lionel Shapiro. What is logical deflationism? Two non-metalinguistic conceptions of logic. Synthese, 200(1):1–28, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03547-8. See

correction [1984].

OLoughlin:2022:MRE

[1811] Ryan O’Loughlin and Dan Li. Model robustness in economics: the admissibility and evaluation of tractability assumptions. Synthese, 200 (1):1–23, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03608-y.

Fischer:2022:PLE

[1812] Eugen Fischer, Paul E. Engelhardt, and Aurélie Herbelot. Philosophers’ linguistic expertise: a psycholinguistic approach to the expertise objection against experimental philosophy. Synthese, 200(1):1–33, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03487-3.

Halak:2022:MEM

[1813] Jan Halák. Mathematics embodied: Merleau-Ponty on geometry and algebra as fields of motor enaction. Synthese, 200(1):1–28, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03526-z.

Nappo:2022:CA

[1814] Francesco Nappo. Confirmation by analogy. Synthese, 200(1):1–26, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-03545-w.

Mogensen:2022:TER

[1815] Andreas L. Mogensen and David Thorstad. Tough enough? Robust satisficing as a decision norm for long-term policy analysis. Synthese, 200 (1):1–26, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03566-5.

Habgood-Coote:2022:KMA

[1816] Joshua Habgood-Coote. Knowing more (about questions). Synthese, 200 (1):1–23, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03594-1.

Thorpe:2022:MSC

[1817] Joshua Rowan Thorpe. Merely superficially contingent a priori knowledge and the McKinsey paradox. Synthese, 200(1):1–15, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03604-2.

Schoonen:2022:PRS

[1818] Tom Schoonen. Possibility, relevant similarity, and structural knowledge. Synthese, 200(1):1–22, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03488-2.

Dorst:2022:LMP

[1819] Chris Dorst. Laws, melodies, and the paradox of predictability. Synthese, 200(1):1–21, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03595-0.

Melenovsky:2022:IC

[1820] C. M. Melenovsky. Interests from and in conventions. Synthese, 200 (1):1–21, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03589-y.

Heinzelmann:2022:DCC

[1821] Nora Heinzelmann and Stephan Hartmann. Deliberation and confidence change. Synthese, 200(1):1–13, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03584-3.

Curtis-Trudel:2022:DC

[1822] André Curtis-Trudel. The determinacy of computation. Synthese, 200 (1):1–28, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03568-3. See correction [2125].

Tegtmeyer:2022:OEC

[1823] Maximilian Tegtmeyer. Overcoming epistemic compositionalism by appreciating Kant’s insight. Synthese, 200(1):1–37, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03557-6.

Jaksland:2022:DPM

[1824] Rasmus Jaksland. Dependent philosophical majorities and the skeptical argument from disagreement. Synthese, 200(1):1–24, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03570-9.

Commandeur:2022:ATM

[1825] Leon Commandeur. Against telic monism in logic. Synthese, 200(1): 1–18, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03497-1.

Meneganzin:2022:BMI

[1826] Andra Meneganzin and Adrian Currie. Behavioural modernity, investigative disintegration & Rubicon expectation. Synthese, 200(1): 1–28, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03491-7.

Pietroiusti:2022:HDE

[1827] Giulio Pietroiusti. Having a disagreement: expression, persuasion and demand. Synthese, 200(1):1–12, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03509-0.

Cretu:2022:ASR

[1828] Ana-Maria Cretu. Authentication, scale-relativity, and relational kindhood. Synthese, 200(1):1–21, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03552-x.

Nelson:2022:UHF

[1829] Lynn Hankinson Nelson. Underdetermination, holism, and feminist philosophy of science. Synthese, 200(1):1–12, February 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03569-2.

Pitcovski:2022:CIV

[1830] Eli Pitcovski and Andrew Peet. Counterfactuals, indeterminacy, and value: a puzzle. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03464-w.

Dethier:2022:WEL

[1831] Corey Dethier. When is an ensemble like a sample? “Model-based” inferences in climate modeling. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03477-5.

Baker:2022:MS

[1832] Alan Baker. The mathematical stance. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03458-8.

Palermos:2022:CKW

[1833] Spyridon Orestis Palermos. Collaborative knowledge: Where the distributed and commitment models merge. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03459-7.

Sendlak:2022:TM

[1834] Maciej Sendłak. Truthmaking for meinongians. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-03541-0.

Delaney:2022:COC

[1835] Brent Delaney. Contraries, oppositions, and contradictions: a species/ genus account of Humean contrariety. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03473-9.

Kotsonis:2022:NUN

[1836] Alkis Kotsonis. A novel understanding of the nature of epistemic vice. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03519-y.

Basu:2022:CMS

[1837] Kaushik Basu. Conventions, morals and strategy: Greta’s dilemma and the incarceration game. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03475-7.

Bland:2022:DEV

[1838] Steven Bland. In defence of epistemic vices. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-03572-7.

Kaspers:2022:APP

[1839] Tom Kaspers. Alethic pluralism for pragmatists. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03558-5.

Fuller:2022:PCM

[1840] Jonathan Fuller. Preventive and curative medical interventions. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03579-0.

Zubcic:2022:DDE

[1841] Marko-Luka Zubcić. Disagreement without discovery and the epistemological argument for freedom from poverty. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-03490-8.

Latham:2022:RPP

[1842] Andrew J. Latham, Kristie Miller, Christian Tarsney, and Hannah Tierney. Robust passage phenomenology probably does not explain futurebias. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03514-3.

Demirtas:2022:CCD

[1843] Huzeyfe Demirtas. Causation comes in degrees. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03507-2.

Watson:2022:CCI

[1844] David S. Watson. Conceptual challenges for interpretable machine learning. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03485-5.

Turner:2022:NCO

[1845] Cody Turner. Neuromedia, cognitive offloading, and intellectual perseverance. Synthese, 200(1):??, February 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03472-w.

Maibom:2022:DWH

[1846] Heidi Lene Maibom. Don’t worry, be happy? Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03597-y.

Elek:2022:TVP

[1847] Gábor Elek and Eszter Babarczy. Taming vagueness: the philosophy of network science. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03622-0.

Byerly:2022:GIT

[1848] T. Ryan Byerly. Group intellectual transparency: a novel case for nonsummativism. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03617-x.

Calzavarini:2022:ARC

[1849] Fabrizio Calzavarini and Gustavo Cevolani. Abductive reasoning in cognitive neuroscience: weak and strong reverse inference. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03585-2.

Rosenblatt:2022:PKT

[1850] Lucas Rosenblatt and Camila Gallovich. Paradoxicality in Kripke’s theory of truth. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03625-x.

Luc:2022:ASP

[1851] Joanna Luc. Arguments from scientific practice in the debate about the physical equivalence of symmetry-related models. Synthese, 200 (2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-03618-w.

Werner:2022:ICE

[1852] Jonas Werner. Irreducibly collective existence and bottomless nihilism. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03623-z.

Muller:2022:LBA

[1853] Thomas Müller. “Let’s build an Anscombe box”: assessing Anscombe’s rebuttal of the statistics objection against indeterminism-based free agency. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03565-6.

Corti:2022:DMS

[1854] Luca Corti. Demystifying the myth of sensation: Wilfrid Sellars’ adverbialism reconsidered. Synthese, 200(2):??, April 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03615-z.

Bloch-Mullins:2022:SMH

[1855] Corinne L. Bloch-Mullins. Similarity in the making: how folk psychological concepts facilitate development of psychological concepts. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03600-6.

Calzavarini:2022:SVC

[1856] Fabrizio Calzavarini and Alfredo Paternoster. The semantic view of computation and the argument from the cognitive science practice. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03542-z.

Strossner:2022:CNC

[1857] Corina Strößner. Criteria for naturalness in conceptual spaces. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03610-4.

Sample:2022:SRP

[1858] Matthew Sample. Science, responsibility, and the philosophical imagination. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03612-2.

Akiba:2022:BMV

[1859] Ken Akiba. The Boolean many-valued solution to the Sorites Paradox. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03561-w.

Smith:2022:AIM

[1860] Ryan Smith, Maxwell J. D. Ramstead, and Alex Kiefer. Active inference models do not contradict folk psychology. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03480-w.

Sinigaglia:2022:MRA

[1861] Corrado Sinigaglia and Stephen A. Butterfill. Motor representation in acting together. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03539-8.

Brock:2022:PP

[1862] Stuart Brock and Joshua Glasgow. The paradox paradox. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03591-4.

Lehtinen:2022:EBG

[1863] Aki Lehtinen. The epistemic benefits of generalisation in modelling II: expressive power and abstraction. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03530-3.

Reese:2022:HCL

[1864] Brian Reese, Michael Vazquez, and Scott Weinstein. How can a line segment with extension be composed of extensionless points? Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03538-9.

DeToffoli:2022:WMD

[1865] Silvia De Toffoli. What are mathematical diagrams? Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03553-w.

Symons:2022:EID

[1866] John Symons and Ramón Alvarado. Epistemic injustice and data science technologies. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03631-z.

McGrath:2022:LAJ

[1867] Matthew McGrath. A limitation on agency in judgment. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03616-y.

Vos:2022:FSI

[1868] Paul Vos and Don Holbert. Frequentist statistical inference without repeated sampling. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03560-x.

Eberhardt:2022:CER

[1869] Frederick Eberhardt. A contemporary example of Reichenbachian coordination. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03571-8.

Touborg:2022:RPD

[1870] Caroline Torpe Touborg. Relativizing proportionality to a domain of events. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03596-z.

Brunero:2022:PRT

[1871] John Brunero. Practical reasons, theoretical reasons, and permissive and prohibitive balancing. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03628-8.

Giordani:2022:DAU

[1872] Alessandro Giordani and Eric Tremolanti. In defense of Aristotelian universals. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03614-0.

Koslow:2022:MCC

[1873] Allison Koslow. Meaning change and changing meaning. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03563-8.

Langland-Hassan:2022:PCT

[1874] Peter Langland-Hassan. Propping up the causal theory. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03635-9.

Heil:2022:KD

[1875] Roman Heil, Jakob Koscholke, Patricia Rich, and Moritz Schulz. Knowledge and decision. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03668-0.

Steeger:2022:OWP

[1876] Jer Steeger. One world is (probably) just as good as many. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03499-z.

Lissia:2022:CPR

[1877] Lina Maria Lissia. Cut-off points for the rational believer. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03510-7. See correction [2279].

Yetter-Chappell:2022:DAW

[1878] Helen Yetter-Chappell. Dualism all the way down: why there is no paradox of phenomenal judgment. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03654-6.

Soland:2022:DLR

[1879] Kim Soland. Does loudness represent sound intensity? Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03665-3.

Kind:2022:IEL

[1880] Amy Kind. Introduction: exploring the limits of imagination. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03669-z.

Tomasello:2022:WIL

[1881] Michael Tomasello. What is it like to be a chimpanzee? Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03574-5.

Rybarikova:2022:LCL

[1882] Zuzana Rybaríková. Łukasiewicz’s concept of logic and anti-psychologism. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03479-3.

Vahid:2022:NPR

[1883] Hamid Vahid. No practical reasons for belief: the epistemic significance of practical considerations. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03474-8.

Coates:2022:UPU

[1884] Ashley Coates. Unmanifested powers and universals. Synthese, 200 (2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-03476-6.

Jorem:2022:GBI

[1885] Sigurd Jorem. The good, the bad and the insignificant-assessing concept functions for conceptual engineering. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03548-7.

Garcia-Carpintero:2022:WDP

[1886] Manuel García-Carpintero and Michele Palmira. What do propositions explain? Inflationary vs. deflationary perspectives and the case of singular propositions. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03467-7.

Lee:2022:SAU

[1887] Chanwoo Lee. The structuralist approach to underdetermination. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03495-3.

Perrin:2022:EMF

[1888] Denis Perrin and André Sant’Anna. Episodic memory and the feeling of pastness: from intentionalism to metacognition. Synthese, 200 (2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-03567-4.

Mizrahi:2022:PGG

[1889] Moti Mizrahi and Michael Adam Dickinson. Philosophy’s gender gap and argumentative arena: an empirical study. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03587-0.

Rapetti:2022:CSA

[1890] Pablo A. Rapetti. A critique of strong anti-Archimedeanism: metaethics, conceptual jurisprudence, and legal disagreements. Synthese, 200(2): ??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-03624-y.

Aleksiev:2022:LHA

[1891] Damian Aleksiev. Lightweight and heavyweight anti-physicalism. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03506-3.

Niiniluoto:2022:APT

[1892] Ilkka Niiniluoto, Gustavo Cevolani, and Theo Kuipers. Approaching probabilistic truths: introduction to the topical collection. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03516-1.

Scheller:2022:CMP

[1893] Simon Scheller, Christoph Merdes, and Stephan Hartmann. Computational modeling in philosophy: introduction to a topical collection. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03481-9.

Filho:2022:DPM

[1894] Sérgio Farias de Souza Filho. A dual proposal of minimal conditions for intentionality. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03483-7.

Pickup:2022:UTC

[1895] Martin Pickup. Unsettledness in times of change. Synthese, 200(2): ??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-03607-z.

Jeuk:2022:PDH

[1896] Alexander Albert Jeuk. The pragmatist domestication of Heidegger: Dreyfus on ‘skillful’ understanding. Synthese, 200(2):??, April 2022. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03606-
0.

Bach:2022:STR

[1897] Theodore Bach. Same-tracking real kinds in the social sciences. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03521-4.

Blomkvist:2022:ISB

[1898] Andrea Blomkvist. Imagination as a skill: A Bayesian proposal. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03550-z.

Fjellstad:2022:ELD

[1899] Andreas Fjellstad. Expressing logical disagreement from within. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03667-1.

Brown:2022:BMU

[1900] Rebecca C. H. Brown, Mícheál de Barra, and Brian D. Earp. Broad Medical Uncertainty and the ethical obligation for openness. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03666-2. See correction [2462].

Kawamleh:2022:CCC

[1901] Suzanne Kawamleh. Confirming (climate) change: a dynamical account of model evaluation. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03659-1.

Eker:2022:TTT

[1902] Bahadir Eker. The A-theory of time, temporal passage, and comprehensiveness. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03532-1.

Vignolo:2022:MCL

[1903] Massimiliano Vignolo. Minimal contents, lying, and conventions of language. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03609-x.

Eriksson:2022:SVV

[1904] Lina Eriksson. Sharing with the vulnerable? The vulnerability objection and Vanderschraaf’s theory of justice as mutual advantage. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03599-w.

Fruge:2022:AN

[1905] Christopher Frugé. Artifactual normativity. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03621-1.

Ransom:2022:APP

[1906] Madeleine Ransom. Aesthetic perception and the puzzle of training. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03555-8.

Giustina:2022:IKA

[1907] Anna Giustina. Introspective knowledge by acquaintance. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03578-1.

Haueis:2022:DMM

[1908] Philipp Haueis. Descriptive multiscale modeling in data-driven neuroscience. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03551-y.

Sorensen:2022:LMI

[1909] Roy Sorensen. Lie for me: the intent to deceive fails to scale up. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03603-3.

Freitag:2022:TE

[1910] Wolfgang Freitag and Felix Bräuer. The transparency of expressivism. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03634-w.

Caiani:2022:III

[1911] Silvano Zipoli Caiani. Intelligence involves intensionality: An explanatory issue for radical enactivism (again). Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03527-y.

Kulp:2022:MIS

[1912] Christopher B. Kulp. Moral intuitions: seeming or believing? Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03461-z.

Aikin:2022:APA

[1913] Scott F. Aikin and John Casey. Argumentation and the problem of agreement. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03680-4.

Jacobs:2022:IIP

[1914] Caspar Jacobs. Invariance, intrinsicality and perspicuity. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03682-2.

Johannesson:2022:ITT

[1915] Eric Johannesson. On the indispensability of theoretical terms and entities. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03683-1.

Mortini:2022:NSS

[1916] Dario Mortini. A new solution to the safety dilemma. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03626-w.

Mizrahi:2022:PRA

[1917] Moti Mizrahi and Michael Adam Dickinson. Philosophical reasoning about science: a quantitative, digital study. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03670-6.

Deans:2022:SPC

[1918] Catherine Deans and Sarah Pini. Skilled performance in contact improvisation: the importance of interkinaesthetic sense of agency. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03629-7.

Allen:2022:NC

[1919] Sophie R. Allen and Javier Cumpa. The necessity of conceivability. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03534-z.

Croce:2022:TKF

[1920] Michel Croce. On testimonial knowledge and its functions. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03528-x.

Barrotta:2022:ISE

[1921] Pierluigi Barrotta and Roberto Gronda. Intelligence and scientific expertise. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03513-4.

deRidder:2022:GBR

[1922] Jeroen de Ridder. Group belief reconceived. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03633-x.

Tahko:2022:NKM

[1923] Tuomas E. Tahko. Natural kinds, mind-independence, and unification principles. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03661-7.

Wilson:2022:PDC

[1924] Mark Wilson. A plea for distinctions (counterfactuals and differential equations). Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03651-9.

Lange:2022:IBE

[1925] Marc Lange. Inference to the best explanation as supporting the expansion of mathematicians’ ontological commitments. Synthese, 200 (2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-03656-4.

Chung:2022:LSN

[1926] Hun Chung. Locke’s state of nature and its epistemic deficit: a gametheoretic analysis. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03582-5.

Martin:2022:AEA

[1927] Ben Martin and Ole Thomassen Hjortland. Anti-exceptionalism about logic as tradition rejection. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03653-7.

De:2022:KML

[1928] Michael De and Hitoshi Omori. Knot much like tonk. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03655-5.

Asif:2022:MTM

[1929] Nimra Asif. Minimal theory of mind — a Millikanian approach. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03662-6.

Cao:2022:PRU

[1930] Rosa Cao. Putting representations to use. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03522-3.

vonPlatz:2022:SBJ

[1931] Jeppe von Platz. The silver bullet: justice as mutual advantage and the vulnerability objection. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03575-4.

Kulakova:2022:COR

[1932] Eugenia Kulakova and Stefan Rinner. Correction to: Oswald revisited: the effect of focus and context. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03671-5. See

[1805].

Ooi:2022:WDS

[1933] Daryl Ooi. Worldview disagreement and subjective epistemic obligations. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03646-6.

Kurth:2022:IEM

[1934] Charlie Kurth. Inappropriate emotions, marginalization, and feeling better. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03619-9.

Vogt:2022:NDE

[1935] Lisa Vogt. Nominalist dispositional essentialism. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03588-z.

Holmes:2022:CEM

[1936] Travis Holmes. Cognitive extra-mathematical explanations. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03643-9.

Khudairi:2022:CGC

[1937] Hasen Khudairi. Correction to: Grounding, conceivability, and the mind– body problem. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03640-y. See [5].

Barrett:2022:LGE

[1938] Jeffrey A. Barrett and Jacob VanDrunen. Language games and the emergence of discourse. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03645-7.

Podosky:2022:CCE

[1939] Paul-Mikhail Catapang Podosky. Can conceptual engineering actually promote social justice? Synthese, 200(2):??, April 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03469-5.

Stout:2022:DAE

[1940] Rowland Stout. Dormant and active emotional states. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03652-8.

Andow:2022:HDP

[1941] James Andow. How do philosophers and nonphilosophers think about philosophy? And does personality make a difference? Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03639-5.

Siscoe:2022:CAK

[1942] Robert Weston Siscoe. Credal accuracy and knowledge. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03636-8.

Ackermans:2022:IFP

[1943] Lennart B. Ackermans. Infinite frequency principles of direct inference. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03642-w.

Gunther:2022:RC

[1944] Mario Günther and Caterina Sisti. Ramsey’s conditionals. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03586-1.

Steward:2022:WDC

[1945] Helen Steward. What does causality have to do with necessity? Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03482-8.

Kowalczyk:2022:JVJ

[1946] Kacper Kowalczyk. Johnston versus Johnston. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03593-2.

Brun:2022:REC

[1947] Georg Brun. Re-engineering contested concepts. A reflective-equilibrium approach. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03556-7.

Anderson:2022:SDA

[1948] Michael L. Anderson and Heather Champion. Some dilemmas for an account of neural representation: a reply to Poldrack. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03505-4.

Baron:2022:SFA

[1949] Sam Baron. Spacetime: function and approximation. Synthese, 200 (2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-03498-0.

Runhardt:2022:LEP

[1950] Rosa W. Runhardt. Limits to evidential pluralism: multi-method large-n qualitative analysis and the primacy of mechanistic studies. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03650-w.

Loewenstein:2022:ASS

[1951] Yael Loewenstein. Against the standard solution to the grandfather paradox. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03543-y.

Knoks:2022:CVH

[1952] Aleks Knoks. Conciliatory views, higher-order disagreements, and defeasible logic. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03503-6.

Blessenohl:2022:SAM

[1953] Simon Blessenohl and Deniz Sarikaya. Science advice: making credences accurate. Synthese, 200(2):??, April 2022. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03620-2.

Wilson:2022:WKP

[1954] Robert A. Wilson. Why kinship is progeneratively constrained: Extending anthropology. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03537-w.

Betz-Richman:2022:LHN

[1955] Noah Betz-Richman. Lying, hedging, and the norms of assertion. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03644-8.

Bonino:2022:WPS

[1956] Guido Bonino, Paolo Maffezioli, and Paolo Tripodi. When philosophy (of science) meets formal methods: a citation analysis of early approaches between research fields. Synthese, 200(2):??, April 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03484-6.

Carrara:2022:TOL

[1957] Massimiliano Carrara and Jeroen Smid. There is overlap in letters: a mereological definition of word types. Synthese, 200(2):??, April 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03664-4.

Fischer:2022:HAS

[1958] Martin Fischer. Is the HYPE about strength warranted? Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03673-3.

Rieppel:2022:RP

[1959] Michael Rieppel. Reference by proxy. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03681-
3.

Werner:2022:PPH

[1960] Konrad Werner. The problem of presentations: how it is that one object is perceptually given in multiple ways. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03486-4.

Luna:2022:SRA

[1961] Laureano Luna. Strengthening the Russellian argument against absolutely unrestricted quantification. Synthese, 200(3):??, June 2022. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03672-
4.

Wenzel:2022:CTR

[1962] Christian Helmut Wenzel. Can thoughts be read from the brain? Neuroscience contra Wittgenstein. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03504-5.

Finlay:2022:CWO

[1963] Stephen Finlay. Correction to: What ought probably means, and why you can’t detach it. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03605-1. See [2].

Muller:2022:EMR

[1964] Seán Mfundza Muller. Econometric methods and Reichenbach’s principle. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03684-0. See correction [2007].

Pritchard:2022:HCC

[1965] Duncan Pritchard. Hinge commitments and common knowledge. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03647-5.

Heyndels:2022:FLE

[1966] Sybren Heyndels. A forgotten logical expressivist: Strawson’s philosophy of logic and its challenges. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03564-7.

Sikorski:2022:FSC

[1967] Michał Sikorski. Is forensic science in crisis? Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03685-z.

Doherty:2022:SIP

[1968] F. T. Doherty and J. Dewhurst. Structuralism, indiscernibility, and physical computation. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03691-1.

Schwartz:2022:SR

[1969] Arieh Schwartz. Simple remembering. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03663-
5.

Irikefe:2022:ETE

[1970] Paul O. Irikefe. The epistemology of thought experiments without exceptionalist ingredients. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03690-2.

Miyahara:2022:SSA

[1971] Katsunori Miyahara and Miguel Segundo-Ortin. Situated self-awareness in expert performance: a situated normativity account of riken no ken. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03688-w.

Jha:2022:DME

[1972] Aditya Jha, Douglas Campbell, and Phillip L. Wilson. Not so distinctively mathematical explanations: topology and dynamical systems. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03697-9.

Parkkinen:2022:VRC

[1973] Veli-Pekka Parkkinen. Variable relativity of causation is good. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03676-0.

Tajer:2022:AEM

[1974] Diego Tajer. Anti-exceptionalism and methodological pluralism in logic. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03698-8.

Kieval:2022:MRM

[1975] Phillip Hintikka Kieval. Mapping representational mechanisms with deep neural networks. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03694-y.

Fletcher:2022:ITA

[1976] Samuel C. Fletcher and Nathan Lackey. The introduction of topology into analytic philosophy: two movements and a coda. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03689-9.

Elkind:2022:CVH

[1977] Landon D. C. Elkind. Computer verification for historians of philosophy. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03678-y. See correction [2455].

Brozek:2022:JLP

[1978] Anna Brozek. Jan Łukasiewicz’s program of the logicization of philosophy: its genesis, content and realizations. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03699-7.

Bochner:2022:ETD

[1979] Gregory Bochner. Externalism, transparency, and diagonal propositions. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03679-x.

Hull:2022:ER

[1980] George Hull. Epistemic redress. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03613-1.

Lyons:2022:CDC

[1981] Jack C. Lyons. Cognitive diversity and the contingency of evidence. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03660-8.

LoGuercio:2022:AOI

[1982] Nicolás Lo Guercio and Ramiro Caso. An account of overt intentional dogwhistling. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03511-6.

Petrovich:2022:ABN

[1983] Eugenio Petrovich. Acknowledgments-based networks for mapping the social structure of research fields. A case study on recent analytic philosophy. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03515-2.

Shapiro:2022:CWL

[1984] Lionel Shapiro. Correction to: What is logical deflationism? Two nonmetalinguistic conceptions of logic. Synthese, 200(3):??, June 2022. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03641-

x. See [1810].

Pfister:2022:TTA

[1985] Rolf Pfister. Towards a theory of abduction based on conditionals. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03581-6. See correction [2142].

Jaksland:2022:DMD

[1986] Rasmus Jaksland. Disagreement in metametaphysical dispute. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03598-x.

Baumann:2022:MUN

[1987] Marius Baumann. Moral underdetermination and a new skeptical challenge. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03529-w.

Biggs:2022:DAE

[1988] Stephen Biggs and Jessica M. Wilson. Does anti-exceptionalism about logic entail that logic is a posteriori? Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03657-3.

Kaspers:2022:SNI

[1989] Tom Kaspers, Jacob LiBrizzi, Duccio Calosi, and Yoichi Kobe. Speak no ill of the dead: the dead as a social group. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03512-5.

Merz:2022:OCB

[1990] Martina Merz and Helene Sorgner. Organizational complexity in big science: strategies and practices. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03649-3.

Jurgens:2022:FBT

[1991] Alan Jurgens. False-belief task know-how. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03630-0.

Giovanelli:2022:GVU

[1992] Marco Giovanelli. Geometrization vs. unification: the Reichenbach– Einstein quarrel about the Fernparallelismus field theory. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03531-2.

Carlson:2022:AEJ

[1993] Matthew Carlson. Anti-exceptionalism and the justification of basic logical principles. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03715-w.

Haze:2022:RLN

[1994] Tristan Grøtvedt Haze. Reversing logical nihilism. Synthese, 200(3): ??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03471-x.

Hennig:2022:QRT

[1995] Boris Hennig. Quiddities and repeatables: towards a tripartite analysis of simple predicative statements. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03627-9.

Contessa:2022:SE

[1996] Gabriele Contessa. Shopping for experts. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03590-5.

Dormandy:2022:LTE

[1997] Katherine Dormandy. Loving truly: An epistemic approach to the doxastic norms of love. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03562-9.

Drozdzowicz:2022:MIP

[1998] Anna Drozdzowicz. Making it precise-imprecision and underdetermination in linguistic communication. Synthese, 200(3):??, June 2022. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03544-
x.

Mayo:2022:SSC

[1999] Deborah G. Mayo and David Hand. Statistical significance and its critics: practicing damaging science, or damaging scientific practice? Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03692-0.

Meier:2022:CTE

[2000] Lukas J. Meier. Can thought experiments solve problems of personal identity? Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03637-7.

Tugby:2022:DRD

[2001] Matthew Tugby. Dispositional realism without dispositional essences. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03554-9.

Smithies:2022:DMA

[2002] Declan Smithies, Preston Lennon, and Richard Samuels. Delusions and madmen: against rationality constraints on belief. Synthese, 200(3): ??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03632-y.

Vazard:2022:EAD

[2003] Juliette Vazard. Everyday anxious doubt. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03711-0.

Moore:2022:SIS

[2004] Richard Moore and Rachael L. Brown. Social inheritance and the social mind: Introduction to the Synthese topical collection The Cultural Evolution of Human Social Cognition. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/

s11229-022-03496-2. See correction [2056].

Pedroso:2022:BBC

[2005] Makmiller Pedroso. Betting blind: coping with uncertainty through redundancy. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03717-8.

Oliphint:2022:UTD

[2006] Jared S. Oliphint. Using a two-dimensional model from social ontology to explain the puzzling metaphysical features of words. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03716-9.

Muller:2022:CEM

[2007] Seán Mfundza Muller. Correction to: Econometric methods and Reichenbach’s principle. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03729-4. See [1964].

Hoffmann:2022:EAA

[2008] Stephanie Hoffmann. An evaluation of the autopoietic account of interests. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03658-2.

Besson:2022:CTR

[2009] Corine Besson and Anandi Hattiangadi. Can truth relativism account for the indeterminacy of future contingents? Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03549-6.

Milliere:2022:DLS

[2010] Raphaël Millière. Deep learning and synthetic media. Synthese, 200 (3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-03739-2.

Giannotti:2022:BFR

[2011] Joaquim Giannotti. Brutalist fundamentalism: radical and moderate. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03705-y.

Park:2022:RSE

[2012] Seungbae Park. The realist and selectionist explanations for the success of science. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03703-0.

Catren:2022:GTS

[2013] Gabriel Catren and Fernando Cukierman. Grothendieck’s theory of schemes and the algebra–geometry duality. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/

s11229-022-03675-1. See correction [2096].

Tedder:2022:RPD

[2014] Andrew Tedder and Marta Bilková. Relevant propositional dynamic logic. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03732-9.

Chis-Ciure:2022:TDI

[2015] Robert Chis-Ciure. The transcendental deduction of Integrated Information Theory: connecting the axioms, postulates, and identity through categories. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03704-z.

Bacharach:2022:RTB

[2016] Julian Bacharach. Relief, time-bias, and the metaphysics of tense. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03707-w.

Dung:2022:DII

[2017] Leonard Dung. Does illusionism imply skepticism of animal consciousness? Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03710-1.

Johansen:2022:EVF

[2018] Mikkel Willum Johansen and Josefine Lomholt Pallavicini. Entering the valley of formalism: trends and changes in mathematicians’ publication practice — 1885 to 2015. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03741-8.

Barrett:2022:MTE

[2019] Thomas William Barrett and Hans Halvorson. Mutual translatability, equivalence, and the structure of theories. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03733-8.

Halpin:2022:FTH

[2020] John F. Halpin. Fine-tuning and Humean laws: fine-tuning as argument for a non-governing account of laws rather than for God or multiverse. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03712-z.

Malaterre:2022:EDC

[2021] Christophe Malaterre and Francis Lareau. The early days of contemporary philosophy of science: novel insights from machine translation and topic-modeling of non-parallel multilingual corpora. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03722-x.

Dentith:2022:SCT

[2022] M. R. X. Dentith. Suspicious conspiracy theories. Synthese, 200(3): ??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03602-4.

Arvan:2022:PAC

[2023] Marcus Arvan and Corey J. Maley. Panpsychism and AI consciousness. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03695-x.

Proszewska:2022:GSM

[2024] Agnieszka M. Proszewska. Goals shape means: a pluralist response to the problem of formal representation in ontic structural realism. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03706-x.

Rubin:2022:WIP

[2025] Hannah Rubin. When it pays to punish in the evolution of honesty and cooperation. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03743-6.

Kuokkanen:2022:VHD

[2026] Jesse Kuokkanen. Vertical-horizontal distinction in resolving the abstraction, hierarchy, and generality problems of the mechanistic account of physical computation. Synthese, 200(3):??, June 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03725-8.

Leary:2022:ERA

[2027] Stephanie Leary. Epistemic reasons for action: a puzzle for pragmatists. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03686-y.

Brusse:2022:ECR

[2028] Carl Brusse, Toby Handfield, and Kevin J. S. Zollman. Explaining costly religious practices: credibility enhancing displays and signaling theories. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03742-7.

Hale:2022:II

[2029] Benjamin Hale. Indeterminacy and impotence. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03718-7.

Coliva:2022:WPD

[2030] Annalisa Coliva and Louis Doulas. What philosophical disagreement and philosophical skepticism hinge on. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03735-6.

Juarez-Garcia:2022:EIR

[2031] Mario I. Juarez-Garcia and Alexander Schaefer. Exit & isolation: Rousseau’s state of nature. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03736-5.

Slome:2022:UVR

[2032] Emily Slome. Undermining versus rebutting: options for responding to evolutionary debunking arguments. Synthese, 200(3):??, June 2022. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03723-
w.

Mayr:2022:AIA

[2033] Erasmus Mayr. Anscombe and intentional agency incompatibilism (for human and animal agents). Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03523-2.

Kazemi:2022:TTN

[2034] Alireza Kazemi. Transparency and the truth norm of belief. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03724-9.

Harbin:2022:RCF

[2035] Ami Harbin. The relational calibration of fear. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03719-6.

DiBona:2022:HC

[2036] Elvira Di Bona. Hearing chimeras. Synthese, 200(3):??, June 2022. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03721-
y.

Immerman:2022:PRS

[2037] Daniel Immerman. The paradigm response to skepticism. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03727-6.

Robitzsch:2022:ESN

[2038] Andrea Robitzsch, Nikola Kompa, and Igal Kvart. The epistemic significance of non-epistemic factors: an introduction. Synthese, 200(3): ??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03611-3.

Zieba:2022:SCU

[2039] Paweł Jakub Zieba. Seeing colours unconsciously. Synthese, 200(3): ??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03687-x.

Hemmo:2022:PBP

[2040] Meir Hemmo and Orly Shenker. The preferred basis problem in the many-worlds interpretation of quantum mechanics: why decoherence does not solve it. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03713-y.

Himelright:2022:SFM

[2041] Jack Himelright. Safety first: making property talk safe for nominalists. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03714-x.

Smith:2022:TAA

[2042] Martin Smith. Two accounts of assertion. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03745-4.

Haziza:2022:AHD

[2043] Eliran Haziza. Assertion and the “How do you know?” challenge. Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03761-4.

Roure:2022:LET

[2044] Pascale Roure. Logical empiricism in Turkish exile: Hans Reichenbach’s research and teaching activities at Istanbul University (1933–1938). Synthese, 200(3):??, June 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03740-9.

Daoust:2022:IP

[2045] Marc-Kevin Daoust and David Montminy. Immodesty and permissivism. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03758-z.

Tieleman:2022:MTU

[2046] Sebastiaan Tieleman. Model transfer and universal patterns: Lessons from the Yule process. Synthese, 200(4):??, August 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03737-4.

Onofri:2022:TPC

[2047] Andrea Onofri. Two pictures of communication: from content identity to coordination. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03757-0.

Norton:2022:HAH

[2048] John D. Norton. How analogy helped create the new science of thermodynamics. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03708-9.

Popa:2022:RRD

[2049] Eugen Octav Popa. On the rational resolution of (deep) disagreements. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03753-4.

Bruckner:2022:IIP

[2050] Michael Bruckner. Internal identity is (partly) dispositional identity. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03747-2.

Rechnitzer:2022:TTR

[2051] Tanja Rechnitzer. Turning the trolley with reflective equilibrium. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03762-3.

Sosa:2022:JGT

[2052] Ernest Sosa. John Greco’s The Transmission of Knowledge. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03746-3.

Loev:2022:AAI

[2053] Slawa Loev. Affectivism about intuitions. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03749-0.

Schmidt:2022:PLM

[2054] Jon Alan Schmidt. Peirce and Łukasiewicz on modal and multi-valued logics. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03755-2.

Longworth:2022:CWK

[2055] Guy Longworth and Simon Wimmer. Cook Wilson on knowledge and forms of thinking. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03748-1.

Moore:2022:CSI

[2056] Richard Moore and Rachael L. Brown. Correction to: Social inheritance and the social mind: Introduction to the Synthese topical collection The Cultural Evolution of Human Social Cognition. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03764-1. See [2004].

Rolla:2022:DBR

[2057] Giovanni Rolla. Do babies represent? On a failed argument for representationalism. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03728-5.

vanMiltenburg:2022:CDF

[2058] Niels van Miltenburg. Causality, determination and free will: towards an Anscombean account of free action. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03700-3.

Jonsson:2022:IMA

[2059] Martin L. Jönsson and Jakob Bergman. Improving misrepresentations amid unwavering misrepresenters. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03744-5.

Kersten:2022:NMC

[2060] Luke Kersten. A new mark of the cognitive? Predictive processing and extended cognition. Synthese, 200(4):??, August 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03674-2.

Charlow:2022:TLR

[2061] Nate Charlow. Triviality and the logic of restricted quantification. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03730-x.

Lewis-Martin:2022:WKG

[2062] Jimmy Lewis-Martin. What kinds of groups are group agents? Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03766-z.

Lelyk:2022:TIC

[2063] Mateusz Lełyk and Carlo Nicolai. A theory of implicit commitment. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03601-5.

Shaver:2022:MET

[2064] Robert Shaver. Moral error theory and hypothetical reasons. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03759-y.

Liberatore:2022:BMA

[2065] Paolo Liberatore. Belief merging in absence of reliability information. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03750-7.

Munroe:2022:WIT

[2066] Wade Munroe. What it takes to make a word (token). Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03751-6. See correction [2322].

Reinikainen:2022:MTT

[2067] Jaakko Reinikainen. Meaning in time: on temporal externalism and Kripkenstein’s skeptical challenge. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03756-1.

LaCroix:2022:TAC

[2068] Travis LaCroix and Aydin Mohseni. The tragedy of the AI commons. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03763-2.

Pritchard:2022:DMA

[2069] Duncan Pritchard. In defence of the modal account of legal risk. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03693-z.

Forbes:2022:ES

[2070] Graeme A. Forbes and Nathan Wildman. Enduring senses. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03770-3.

Tarnowski:2022:WKP

[2071] Maciej Tarnowski and Maciej Głowacki. Words on Kripke’s puzzle. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03769-w.

Mahant:2022:MBM

[2072] Nikhil Mahant. The metaphysical burden of Millianism. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03754-3.

Francez:2022:PTS

[2073] Nissim Francez. Proof-theoretic semantics as a resource for expressing semantic variability. Synthese, 200(4):??, August 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03780-1.

Pesonen:2022:ACE

[2074] Renne Pesonen. Argumentation, cognition, and the epistemic benefits of cognitive diversity. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03786-9.

Huang:2022:NCT

[2075] Alice C. W. Huang. A normative comparison of threshold views through computer simulations. Synthese, 200(4):??, August 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03784-x.

Grim:2022:PES

[2076] Patrick Grim, Frank Seidl, Calum McNamara, Isabell N. Astor, and Caroline Diaso. The punctuated equilibrium of scientific change: a Bayesian network model. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03720-z.

Kvart:2022:RBT

[2077] Igal Kvart. Resolving bank-type puzzles via action-directed pragmatics. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03648-4.

Wojtowicz:2022:MFM

[2078] Krzysztof Wójtowicz and Bartłomiej Skowron. A metaphysical foundation for mathematical philosophy. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03760-
5.

Stoltzner:2022:PTF

[2079] Michael Stöltzner. From probabilistic topologies to Feynman diagrams: Hans Reichenbach on time, genidentity, and quantum physics. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03726-7.

Coleman:2022:FRO

[2080] Sam Coleman. Fred’s red: on the objectivity and physicality of mental qualities. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03772-1.

Axtell:2022:RLI

[2081] Guy Axtell. Risk-limited indulgent permissivism. Synthese, 200(4): ??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-03781-0.

Jurjako:2022:CPP

[2082] Marko Jurjako. Can predictive processing explain self-deception? Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03797-6.

Cimino:2022:HHI

[2083] Andrea Cimino. Husserl, hallucination, and intentionality. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03777-w.

Tang:2022:EMG

[2084] Refeng Tang. Exorcising the myth of the given: the idea of doxasticism. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03778-9.

Schroeder:2022:DSC

[2085] S. Andrew Schroeder. Diversifying science: comparing the benefits of citizen science with the benefits of bringing more women into science. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03774-z.

Glauer:2022:SBV

[2086] Ramiro Glauer and Frauke Hildebrandt. Sortals, bodies, and variables. A critique of Quine’s theory of reference. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03767-y.

Newman:2022:COO

[2087] Andrew Newman. A causal ontology of objects, causal relations, and various kinds of action. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03752-5.

Knoll:2022:NWM

[2088] Viktoria Knoll. Negotiating “women”: metalinguistic negotiations across languages. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03779-8.

Schonherr:2022:DJT

[2089] Julius Schönherr. Doxastic justification through dispositions to cause. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03800-0.

Horne:2022:TMC

[2090] Alex Horne. Too many cooks. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03799-4.

Pelczar:2022:CPC

[2091] Michael Pelczar. The case for panpsychism: a critical assessment. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03775-y.

Doulatova:2022:MER

[2092] Maria Doulatova. Mindreading, emotion-regulation, and oppression. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03805-9.

Bertran-San-Millan:2022:PSB

[2093] Joan Bertran-San-Millán. Peano’s structuralism and the birth of formal languages. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03782-z.

Herrmann:2022:PEA

[2094] Daniel A. Herrmann. Prediction with expert advice applied to the problem of prediction with expert advice. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03809-5.

Tresker:2022:TEG

[2095] Steven Tresker. Treatment effectiveness, generalizability, and the explanatory/pragmatic-trial distinction. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03517-0.

Catren:2022:CGT

[2096] Gabriel Catren and Fernando Cukierman. Correction: Grothendieck’s theory of schemes and the algebra–geometry duality. Synthese, 200 (4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03789-6. See [2013].

Wilhelm:2022:TL

[2097] Isaac Wilhelm. Tractability and laws. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03638-6.

Ohman:2022:IGD

[2098] Carl Öhman. The identification game: deepfakes and the epistemic limits of identity. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03798-5.

Carr:2022:CDN

[2099] Vanessa Carr. Causality, determination and necessitation in free human action. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03806-8.

Shan:2022:EME

[2100] Yafeng Shan and Jon Williamson. Evidential monism, evidential pluralism, or evidential contextualism? An introduction to evidential diversity in the social sciences. Synthese, 200(4):??, August 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03801-z.

Andersson:2022:CCP

[2101] Henrik Andersson and Anders Herlitz. Classifying comparability problems in a way that matters. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03795-8.

Sass:2022:HMO

[2102] Reuben Sass. How to minimize ontological commitments: a groundingreductive approach. Synthese, 200(4):??, August 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03768-x.

Newton:2022:EAE

[2103] Lilith Newton. Epistemic anxiety and epistemic risk. Synthese, 200 (4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03788-7.

Gough:2022:MTM

[2104] Joe Gough. The many theories of mind: eliminativism and pluralism in context. Synthese, 200(4):??, August 2022. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03804-w.

Leuschner:2022:ELD

[2105] Anna Leuschner and Manuela Fernandez Pinto. Exploring the limits of dissent: the case of shooting bias. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03783-y.

Vazard:2022:AAI

[2106] Juliette Vazard and Charlie Kurth. Apprehending anxiety: an introduction to the topical collection on worry and wellbeing. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03794-9.

Pence:2022:EID

[2107] Charles H. Pence and Luca Rivelli. Editorial introduction to “Digital Studies of Digital Science”. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03802-y.

Diggin:2022:EE

[2108] Steven Diggin. Ethical evidence. Synthese, 200(4):??, August 2022. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03808-
6.

Quaranto:2022:DWC

[2109] Anne Quaranto. Dog whistles, covertly coded speech, and the practices that enable them. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03791-y.

Shields:2022:RCT

[2110] Matthew Shields. Rethinking conspiracy theories. Synthese, 200(4): ??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-03811-x.

Larvor:2022:URM

[2111] Brendan Philip Larvor. On the unreasonable reliability of mathematical inference. Synthese, 200(4):??, August 2022. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03812-w.

Lai:2022:RKK

[2112] Changsheng Lai. Remembering is not a kind of knowing. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03814-8.

Pinder:2022:HAP

[2113] Mark Pinder. Is Haslanger’s ameliorative project a successful conceptual engineering project? Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03803-x.

Smart:2022:PTE

[2114] Paul R. Smart, Gloria Andrada, and Robert W. Clowes. Phenomenal transparency and the extended mind. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03824-6.

Janda:2022:HMB

[2115] Pavel Janda. How much are bold Bayesians favoured? Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03825-5.

Arroyo:2022:EVM

[2116] Raoni Wohnrath Arroyo and Jonas R. Becker Arenhart. The epistemic value of metaphysics. Synthese, 200(4):??, August 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03833-5.

Masciari:2022:SSW

[2117] Christopher F. Masciari. Stop and smell the what? Two kinds of olfactory representation. Synthese, 200(4):??, August 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03836-2.

Kalis:2022:NSA

[2118] Annemarie Kalis. Normativity in social accounts of reasoning: a Rylean approach. Synthese, 200(4):??, August 2022. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03796-7.

Khalaj:2022:KHF

[2119] M. Hosein M. A. Khalaj. Know-how-first anti-intellectualism: Williamson against Williamson. Synthese, 200(4):??, August 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03823-7. See cor-

rection [2198].

Zayton:2022:OTR

[2120] Benjamin Zayton. Open texture, rigor, and proof. Synthese, 200 (4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03842-4.

Ioannou:2022:CRT

[2121] Savvas Ioannou. Conceptual reductions, truthmaker reductive explanations, and ontological reductions. Synthese, 200(4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03831-7.

Khalil:2022:IIH

[2122] Elias L. Khalil. The information inelasticity of habits: Kahneman’s bounded rationality or Simon’s procedural rationality? Synthese, 200 (4):??, August 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03815-7.

DeBaerdemaeker:2022:MME

[2123] Siska De Baerdemaeker and Richard Dawid. MOND and meta-empirical theory assessment. Synthese, 200(5):??, October 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03830-8.

Teague:2022:PCQ

[2124] Richard Teague. The problem of closure and questioning attitudes. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03826-4.

Curtis-Trudel:2022:CDC

[2125] André Curtis-Trudel. Correction to: The determinacy of computation. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03835-3. See [1822].

Kieval:2022:PBP

[2126] Phillip Hintikka Kieval. Permission to believe is not permission to believe at will. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03845-1.

Dalbey:2022:ICP

[2127] Bryce Dalbey and Bradford Saad. Internal constraints for phenomenal externalists: a structure matching theory. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03829-1.

Belkoniene:2022:RDU

[2128] Miloud Belkoniene. The rational dimension of understanding. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03839-z.

Skibra:2022:OA

[2129] Daniel Skibra. Ought and agency. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03819-3.

Tasker:2022:CEA

[2130] Nick Tasker. The case for eliminativism about words. Synthese, 200 (5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03846-0.

Piekarski:2022:MCP

[2131] Michał Piekarski. Motivation, counterfactual predictions and constraints: normativity of predictive mechanisms. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03837-1.

Coren:2022:IBP

[2132] Daniel Coren. Indecision and Buridan’s principle. Synthese, 200 (5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03843-3.

Sterner:2022:EAS

[2133] Beckett Sterner. Explaining ambiguity in scientific language. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03792-x.

Collins:2022:NLQ

[2134] John Collins. Natural language quantification is not polysemous. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03773-0.

Levy:2022:DYO

[2135] Neil Levy. Do your own research! Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03793-w.

Steel:2022:PDB

[2136] Daniel Steel and Karoline Paier. Pro-diversity beliefs and the diverse person’s burden. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03785-w.

Hess:2022:ISL

[2137] Leopold Hess. Inferentialist semantics for lexicalized social meanings. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03817-5.

Cartlidge:2022:LDP

[2138] James Cartlidge. Lyotard, the differend and the philosophy of deep disagreement. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03841-5.

Yu:2022:DEP

[2139] Li an Yu. Distributing epistemic and practical risks: a comparative study of communicating earthquake damages. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03838-0.

Andreoletti:2022:SR

[2140] Giacomo Andreoletti and Louis Vervoort. Superdeterminism: a reappraisal. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03832-6.

Coffman:2022:II

[2141] E. J. Coffman. Involuntarism impugned? Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03849-x.

Pfister:2022:CTT

[2142] Rolf Pfister. Correction: Towards a theory of abduction based on conditionals. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03790-z. See [1985].

Ho:2022:OCS

[2143] Eugene Ho. An object-centric solution to Edelberg’s puzzles of intentional identity. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03834-4.

Messerli:2022:VRW

[2144] Michael Messerli, Julian Fink, and Kevin Reuter. The varying rationality of weakness of the will: an empirical investigation and its challenges for a unified theory of rationality. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03807-7.

Nicolai:2022:GGT

[2145] Carlo Nicolai. Gaps, gluts, and theoretical equivalence. Synthese, 200 (5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03816-6.

Veigl:2022:RHR

[2146] Sophie J. Veigl, Javier Suárez, and Adrian Stencel. Rethinking hereditary relations: the reconstitutor as the evolutionary unit of heredity. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03810-y.

Hlobil:2022:TMS

[2147] Ulf Hlobil. A truth-maker semantics for ST: refusing to climb the strict/ tolerant hierarchy. Synthese, 200(5):??, October 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03820-w.

Gardiner:2022:SSS

[2148] Georgi Gardiner and Brian Zaharatos. The safe, the sensitive, and the severely tested: a unified account. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03731-w.

Kuokkanen:2022:NCI

[2149] Jesse Kuokkanen. No computation without implementation? A potential problem for the single hierarchy view of physical computation. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03696-w.

Kimpton-Nye:2022:PMP

[2150] Samuel Kimpton-Nye. Pandispositionalism and the metaphysics of powers. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03857-x.

Pettigrew:2022:AAO

[2151] Richard Pettigrew. Aggregating agents with opinions about different propositions. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03850-4.

McKenna:2022:SAM

[2152] Travis McKenna. Structure and applied mathematics. Synthese, 200 (5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03873-x.

Vassallo:2022:PMS

[2153] Antonio Vassallo, Pedro Naranjo, and Tim Koslowski. A proposal for a metaphysics of self-subsisting structures. I. Classical physics. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03865-x.

Pavese:2022:PKF

[2154] Carlotta Pavese. Practical knowledge first. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03848-y.

Hoerl:2022:KFA

[2155] Christoph Hoerl. A knowledge-first approach to episodic memory. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03702-1.

Keck:2022:MMM

[2156] John W. Keck. Maximal motion and minimal matter: Aristotelian physics and special relativity. Synthese, 200(5):??, October 2022. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03771-

2. See correction [2275].

Francken:2022:COS

[2157] Jolien C. Francken, Marc Slors, and Carl F. Craver. Cognitive ontology and the search for neural mechanisms: three foundational problems. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03701-2.

Baysan:2022:TPO

[2158] Umut Baysan. Truthmaker puzzles for one-level physicalists. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03734-7.

Sebens:2022:FF

[2159] Charles T. Sebens. The fundamentality of fields. Synthese, 200 (5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03844-2.

Cangelosi:2022:MER

[2160] Josh Cangelosi. From moral to epistemic responsibility. Synthese, 200 (5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03738-3.

Polzler:2022:TEB

[2161] Thomas Pölzler and Ivar R. Hannikainen. The typicality effect in basic needs. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03859-9.

Nencha:2022:TQM

[2162] Cristina Nencha. On the translation from quantified modal logic to counterpart theory. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03813-9.

Azzano:2022:NFD

[2163] Lorenzo Azzano and Andrea Raimondi. New foundations of dispositionalism — introduction. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03847-z.

Pruss:2022:CNC

[2164] Alexander R. Pruss. Correction: Non-classical probabilities invariant under symmetries. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03878-6. See [1546].

Price:2022:GEA

[2165] Huw Price. Global expressivism and alethic pluralism. Synthese, 200 (5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03874-w.

Taliga:2022:HMC

[2166] Milos Taliga. How to make critical rationalism comprehensive and nonparadoxical. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03881-x.

Lange:2022:DRS

[2167] Marc Lange. In defense of really statistical explanations. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03868-8.

Fernandes:2022:HED

[2168] Alison Fernandes. How to explain the direction of time. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03818-4.

Simonelli:2022:CEF

[2169] Ryan Simonelli. Considering the exceptions: on the failure of cumulative transitivity for indicative conditionals. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03854-0.

McMyler:2022:EKK

[2170] Benjamin McMyler. Encultured knowing: knowledge transmission and varieties of cultural learning. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03876-8.

Fard:2022:SID

[2171] Sahar Heydari Fard. Strategic injustice, dynamic network formation, and social movements. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03576-3.

vantHoff:2022:DCE

[2172] Alice van’t Hoff. In defense of causal eliminativism. Synthese, 200 (5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03875-9.

Mukerji:2022:WHP

[2173] Nikil Mukerji and Edzard Ernst. Why homoeopathy is pseudoscience. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03882-w.

Sytsma:2022:SEE

[2174] Justin Sytsma, Robert Bishop, and John Schwenkler. Has the sideeffect effect been cancelled? (no, not yet.). Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03864-y.

Bonatti:2022:RCA

[2175] Nicola Bonatti. A reassessment of Cantorian abstraction based on the ε-operator. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03840-6.

Gamester:2022:TPD

[2176] Will Gamester. Truth pluralism without domains. Synthese, 200 (5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03828-2.

Moruzzi:2022:DPT

[2177] Sebastiano Moruzzi. Diaphonic pluralism: from truth pluralism to disagreement pluralism. Synthese, 200(5):??, October 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03852-2.

Mucino:2022:ARQ

[2178] Ricardo Muciño, Elias Okon, and Daniel Sudarsky. Assessing relational quantum mechanics. Synthese, 200(5):??, October 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03886-6.

Ichikawa:2022:YOK

[2179] Jonathan Jenkins Ichikawa. You ought to have known: positive epistemic norms in a knowledge-first framework. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03872-y.

Holik:2022:QST

[2180] Federico Holik, Juan Pablo Jorge, Décio Krause, and Olimpia Lombardi. Quasi-set theory: a formal approach to a quantum ontology of properties. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03884-8.

Wajnerman-Paz:2022:RCF

[2181] Abel Wajnerman-Paz and Daniel Rojas-Líbano. On the role of contextual factors in cognitive neuroscience experiments: a mechanistic approach. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03870-0.

Ferrari:2022:AAN

[2182] Francesco Maria Ferrari. An argument against nominalism. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03520-5.

Eriksson:2022:DIP

[2183] John Eriksson. Disagreement and inconsistency: a problem for orthodox expressivism. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03867-9.

VanSchoelandt:2022:TVO

[2184] Chad Van Schoelandt. Three vulnerability objections to justice as mutual advantage. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03871-z.

Hickson:2022:ICW

[2185] Michael Hickson. Illusory checkmates: why chess is not a game. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03855-z.

Ippoliti:2022:HPM

[2186] Emiliano Ippoliti. On the heuristic power of mathematical representations. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03883-9.

Watson:2022:ENT

[2187] Jamie Carlin Watson. Epistemic neighbors: trespassing and the range of expert authority. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03709-8.

Masterman:2022:PCP

[2188] Christopher James Masterman. Propositional contingentism and possible worlds. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03879-5.

Thorstad:2022:TNE

[2189] David Thorstad. There are no epistemic norms of inquiry. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03896-4.

Jellema:2022:EET

[2190] Hylke Jellema. Eyewitness evaluation through inference to the best explanation. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03858-w.

Baumgaertner:2022:PBI

[2191] Bert Baumgaertner and Florian Justwan. The preference for belief, issue polarization, and echo chambers. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03880-y.

Kang:2022:SCA

[2192] Shao-Pu Kang. Shared consciousness and asymmetry. Synthese, 200 (5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03890-w.

Rushing:2022:NFT

[2193] Bruce Rushing. No free theory choice from machine learning. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03901-w.

Burri:2022:CMP

[2194] Susanne Burri. Conceptualising morally permissible risk imposition without quantified individual risks. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03888-4.

Rubio:2022:II

[2195] Daniel Rubio. Ideological innocence. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03898-2.

Park:2022:CIM

[2196] John H. Park. Current issues in medical epistemology and statistics: a view from the frontline of medicine. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03765-0.

Choi:2022:TIP

[2197] Seungrak Choi. Is there an inconsistent primitive recursive relation? Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03900-x.

Khalaj:2022:CKH

[2198] M. Hosein M. A. Khalaj. Correction: Know-how-first anti-intellectualism: Williamson against Williamson. Synthese, 200(5):??, October 2022. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03851-

3. See [2119].

MelendezGutierrez:2022:ARS

[2199] Sofía Meléndez Gutiérrez. Ante rem structuralism and the semantics of instantial terms. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03856-y.

Greco:2022:SVT

[2200] John Greco. Skepticism, virtue and transmission in the theory of knowledge: an anti-reductionist and anti-individualist account. Synthese, 200 (5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03893-7.

Adlam:2022:TRR

[2201] Emily Adlam. Two roads to retrocausality. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03919-0.

Schwenkenbecher:2022:CIO

[2202] Anne Schwenkenbecher. Collective inaction, omission, and non-action: when not acting is indeed on ‘us’. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03827-3.

Lorenz:2022:QCM

[2203] Robin Lorenz. Quantum causal models: the merits of the spirit of Reichenbach’s principle for understanding quantum causal structure. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03887-5.

Schuster:2022:PLJ

[2204] Daniela Schuster and Leon Horsten. On the pure logic of justified belief. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03905-6.

Pritchard:2022:PLM

[2205] Tim Pritchard. Proprietary linguistic meaning. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03776-x.

Acquaviva:2022:WML

[2206] Paolo Acquaviva. Word meaning: a linguistic dimension of conceptualization. Synthese, 200(5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03910-9.

Trivino:2022:TCM

[2207] Vanesa Triviño. Towards a characterization of metaphysics of biology: metaphysics for and metaphysics in biology. Synthese, 200 (5):??, October 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03897-3.

Ackermans:2022:CBM

[2208] Lennart B. Ackermans. Causal bias in measures of inequality of opportunity. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03899-1.

Carston:2022:WSS

[2209] Robyn Carston. Words: Syntactic structures and pragmatic meanings. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03861-1. See correction [2257].

Tarnowski:2022:GKE

[2210] Maciej Tarnowski, Adrian Ziółkowski, and Mieszko Tałasiewicz. The Group Knobe Effect revisited: epistemic and doxastic side-effect effects in intuitive judgments concerning group agents. Synthese, 200 (6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03917-2.

Rice:2022:MMP

[2211] Collin Rice. Modeling multiscale patterns: active matter, minimal models, and explanatory autonomy. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03885-7.

Rattan:2022:MH

[2212] Gurpreet Rattan and Marion Durand. Mates and the hierarchy. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03915-4.

ElSkaf:2022:WCW

[2213] Rawad El Skaf and Patricia Palacios. What can we learn (and not learn) from thought experiments in black hole thermodynamics? Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03927-0.

Goh:2022:ATR

[2214] Esther Goh and Frederick Choo. Addressing two recent challenges to the factive account of knowledge. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03916-3.

Fisher:2022:TIF

[2215] Alex Fisher. Truth in interactive fiction. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03935-0.

Suarez:2022:TDS

[2216] Javier Suárez and Federico Boem. Technology-driven surrogates and the perils of epistemic misalignment: an analysis in contemporary microbiome science. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03912-7.

Kapsner:2022:LNL

[2217] Andreas Kapsner. Lord nicholls on the likelihood of crimes. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03918-1.

Nappo:2022:EAA

[2218] Francesco Nappo, Giulio Magli, and Giovanni Valente. Evidence and analogy in archaeoastronomy. Synthese, 200(6):??, December 2022. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03863-
z.

Gauld:2022:PKH

[2219] Christophe Gauld. From psychiatric kinds to harmful symptoms. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03922-5.

Bschir:2022:PPP

[2220] Karim Bschir and Simon Lohse. Pandemics, policy, and pluralism: a Feyerabend-inspired perspective on COVID-19. Synthese, 200(6): ??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03923-4.

Heersmink:2022:HUU

[2221] Richard Heersmink. Human uniqueness in using tools and artifacts: flexibility, variety, complexity. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03892-8.

Sogaard:2022:UMU

[2222] Anders Søgaard. Understanding models understanding language. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03931-4.

Pruss:2022:APB

[2223] Alexander R. Pruss. Accuracy, probabilism and Bayesian update in infinite domains. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03938-x.

Cintula:2022:DGE

[2224] Petr Cintula, Berta Grimau, Carles Noguera, and Nicholas J. J. Smith. These degrees go to eleven: fuzzy logics and gradable predicates. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03909-2.

Wainer:2022:UDU

[2225] Juan Manuel Garrido Wainer. Understanding the development and use of tools in neuroscience: the case of the tungsten micro-electrode. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03934-1.

Harris:2022:SPP

[2226] Keith Raymond Harris. Some problems with particularism. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03948-9.

Tan:2022:TEC

[2227] Peter Tan. Two epistemological challenges regarding hypothetical modeling. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03928-z.

Arenhart:2022:IPI

[2228] Jonas R. Becker Arenhart. Interpreting philosophical interpretations of paraconsistency. Synthese, 200(6):??, December 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03941-2.

Talasiewicz:2022:FMP

[2229] Mieszko Tałasiewicz. Files are mental particulars: a rejoinder to the relationist challenge. Synthese, 200(6):??, December 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03932-3.

Stuart:2022:STI

[2230] Michael T. Stuart. Sharpening the tools of imagination. Synthese, 200 (6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03939-w.

Mulder:2022:CDP

[2231] Jesse M. Mulder, Thomas Müller, Dawa Ometto, and Niels van Miltenburg. Causality and determination, powers and agency: Anscombean perspectives. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03853-1.

Colaco:2022:WSP

[2232] David Colaço. Why studying plant cognition is valuable, even if plants aren’t cognitive. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03869-7.

Blancke:2022:SMS

[2233] Stefaan Blancke. Science as a moral system. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03877-7.

Esfeld:2022:NRM

[2234] Michael Esfeld and Guillaume Köstner. Normative relations, mind points and social ontology. Synthese, 200(6):??, December 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03889-3.

Ivy:2022:RCE

[2235] Spencer Ivy. The role of creativity in expertise and skilled action. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03822-8.

Mollica:2022:AVT

[2236] Leona Mollica. Argle victorious: a theory of holes as hole-linings. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03952-z.

Kim:2022:SBC

[2237] Namjoong Kim. Sleeping beauty and the current chance evidential immodest dominance axiom. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03926-1.

Torza:2022:LNT

[2238] Alessandro Torza. Laws of nature and theory choice. Synthese, 200 (6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03950-1.

Christensen:2022:CCI

[2239] Wayne Christensen and Kath Bicknell. Cognitive control, intentions, and problem solving in skill learning. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03920-7.

Cornelissen:2022:USC

[2240] Milo D. Cornelissen and Henk W. de Regt. Understanding in synthetic chemistry: the case of periplanone B. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03929-y.

Milkowski:2022:TBT

[2241] Marcin Miłkowski and Piotr Litwin. Testable or bust: theoretical lessons for predictive processing. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03891-9.

Clark:2022:DDH

[2242] Elinor Clark and Donal Khosrowi. Decentring the discoverer: how AI helps us rethink scientific discovery. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03902-9.

Eisenthal:2022:REA

[2243] Joshua Eisenthal and Lydia Patton. Reichenbach’s empirical axiomatization of relativity. Synthese, 200(6):??, December 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03936-z.

Fehr:2022:CEE

[2244] Carla Fehr and Janet Minji Jones. Culture, exploitation, and epistemic approaches to diversity. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03787-8.

Read:2022:ESF

[2245] James Read and Bryan Cheng. Euclidean spacetime functionalism. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03951-0.

Ciecierski:2022:DA

[2246] Tadeusz Ciecierski and Piotr Tomasz Makowski. Demonstrations as actions. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03942-1.

Kovacs:2022:ASR

[2247] David Mark Kovacs. Against the status response to the argument from Vagueness. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03945-y.

Desai:2022:EFD

[2248] Jules Desai, David Watson, Vincent Wang, Mariarosaria Taddeo, and Luciano Floridi. The epistemological foundations of data science: a critical review. Synthese, 200(6):??, December 2022. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03933-2.

Bumatay:2022:FSI

[2249] Vlademire Kevin D. Bumatay. Formal schemas of induction as models. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03921-6.

Caterina:2022:NDS

[2250] Gianluca Caterina, Rocco Gangle, and Fernando Tohmé. Native diagrammatic soundness and completeness proofs for Peirce’s Existential Graphs (Alpha). Synthese, 200(6):??, December 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03903-8.

Fried:2022:SEC

[2251] Avital Fried. Statistical evidence and the criminal verdict asymmetry. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03947-w.

Iranzo-Ribera:2022:SCM

[2252] Noelia Iranzo-Ribera. Scientific counterfactuals as make-believe. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03949-8.

Bergman:2022:FFD

[2253] Karl Gustav Bergman and Nils Franzén. The force of fictional discourse. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03955-w.

Smith:2022:REB

[2254] Nicholas J. J. Smith. Respecting evidence: belief functions not imprecise probabilities. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03937-y.

Garbacz:2022:NPC

[2255] Paweł Garbacz. A new perspective on criteria of identity. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03954-x.

Dahlman:2022:CMV

[2256] Christian Dahlman and Eivind Kolflaath. Causal models versus reason models in Bayesian networks for legal evidence. Synthese, 200 (6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03953-y.

Carston:2022:CWS

[2257] Robyn Carston. Correction to: Words: Syntactic structures and pragmatic meanings. Synthese, 200(6):??, December 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03946-x. See

[2209].

Williamson:2022:AKH

[2258] Timothy Williamson. Acting on knowledge-how. Synthese, 200(6): ??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03677-z.

Wheeler:2022:HRI

[2259] Billy Wheeler. How realist is informational structural realism? Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03911-8.

Ghiara:2022:TRW

[2260] Virginia Ghiara. Taking the Russo–Williamson thesis seriously in the social sciences. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03924-3.

Blanchard:2022:SAE

[2261] Thomas Blanchard. Specificity of association in epidemiology. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03944-z.

Koskinen:2022:HIS

[2262] Inkeri Koskinen. How institutional solutions meant to increase diversity in science fail. Synthese, 200(6):??, December 2022. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03959-6.

Zagzebski:2022:IRT

[2263] Linda Zagzebski. Intersubjective reasons and the transmission of religious knowledge. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03914-5.

Dixon:2022:IPS

[2264] Jonathan Dixon. The insignificance of philosophical skepticism. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03957-8.

Cote:2022:MRF

[2265] Nicolas Côté. Measuring republican freedom. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03964-9.

Newall:2022:SB

[2266] Michael Newall. A study in brown. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03956-9.

Zouhar:2022:PPT

[2267] Marián Zouhar. Predicates of personal taste and normative meaning. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03894-6.

Khawaja:2022:IZ

[2268] Jake Khawaja. Interactionist zombies. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03970-x.

Shin:2022:SED

[2269] Justin Shin. Statistical evidence, discrimination, and causation. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03958-7.

Duede:2022:IAA

[2270] Eamon Duede. Instruments, agents, and artificial intelligence: novel epistemic categories of reliability. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03975-
6.

Soe:2022:DIP

[2271] Sille Obelitz Søe and Jens-Erik Mai. Data identity: privacy and the construction of self. Synthese, 200(6):??, December 2022. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03968-5.

Gonzalez-Cabrera:2022:LEH

[2272] Ivan Gonzalez-Cabrera. A lineage explanation of human normative guidance: the coadaptive model of instrumental rationality and shared intentionality. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03925-2.

Montero:2022:MPC

[2273] Barbara Gail Montero. Mathematical Platonism and the causal relevance of abstracta. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03962-x.

Bowker:2022:INW

[2274] Mark Bowker. Introducing new work on indeterminacy and underdetermination. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03982-7.

Keck:2022:CMM

[2275] John W. Keck. Correction to: Maximal motion and minimal matter: Aristotelian physics and special relativity. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-03961-y. See [2156].

Saad:2022:HPW

[2276] Bradford Saad. Harmony in a panpsychist world. Synthese, 200(6): ??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03974-7.

Archer:2022:QAA

[2277] Avery Archer. The questioning-attitude account of agnosticism. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03971-w.

Jobczyk:2022:PMT

[2278] Krystian Jobczyk. Putnam’s model-theoretic argument (meta)reconstructed. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03960-z.

Lissia:2022:CCP

[2279] Lina Maria Lissia. Correction to: Cut-off points for the rational believer. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03979-2. See [1877].

Porpora:2022:MIS

[2280] Douglas V. Porpora. The metaphysical issues in the social sciences and how social scientists debate them. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03978-3.

Dewey:2022:ANR

[2281] Aliya R. Dewey. Arbitrating norms for reasoning tasks. Synthese, 200 (6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03981-8.

Peterson:2022:GFB

[2282] Jared Peterson. Groups that fly blind. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03973-8.

Knuuttila:2022:CIE

[2283] Tarja Knuuttila. Correction to: Imagination extended and embedded: artifactual versus fictional accounts of models. Synthese, 200 (6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03866-w. See [525].

Sharp:2022:NRS

[2284] William A. Sharp. Naïve realism and supersaturated hue. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03994-3.

Cao:2022:MRS

[2285] Rosa Cao. Multiple realizability and the spirit of functionalism. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03524-1.

Farina:2022:MEE

[2286] Mirko Farina and Andrea Lavazza. Mind embedded or extended: transhumanist and posthumanist reflections in support of the extended mind thesis. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03963-w.

Wronski:2022:WCC

[2287] Leszek Wroński, Zalán Gyenis, and Mariangela Zoe Cocchiaro. What chance-credence norms should be. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-03996-
1.

Papineau:2022:STK

[2288] David Papineau. Swampman, teleosemantics and kind essences. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03966-7.

Lorenzano:2022:SEA

[2289] Pablo Lorenzano and José Díez. Scientific explanation as ampliative, specialized embedding: the case of classical genetics. Synthese, 200 (6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03983-6.

King:2022:SSA

[2290] Zoë A. Johnson King. Sensitivity, safety, and admissibility. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03972-9.

Baltag:2022:JBK

[2291] Alexandru Baltag, Nick Bezhanishvili, Aybüke Özgün, and Sonja Smets. Justified belief, knowledge, and the topology of evidence. Synthese, 200 (6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03967-6.

Ehmann:2022:RTB

[2292] Alexander Ehmann. Reducing thermodynamics to Boltzmannian statistical mechanics: the case of macro values. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03985-4.

Joshi:2022:DCB

[2293] Hrishikesh Joshi. Debunking creedal beliefs. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03991-6.

Alter:2022:PSM

[2294] Torin Alter and Robert J. Howell. Physicalism, supervenience, and monism. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03965-8.

Harris:2022:OFE

[2295] Keith Raymond Harris. Outward-facing epistemic vice. Synthese, 200 (6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03995-2.

Dybowski:2022:GAL

[2296] Maciej Dybowski, Weronika Dziegielewska, and Wojciech Rzepiński. Game analogy in law reconsidered: is evidence at stake? Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03997-0.

Atencia-Linares:2022:DSE

[2297] Paloma Atencia-Linares and Marc Artiga. Deepfakes, shallow epistemic graves. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-04003-3.

Dietrich:2022:RA

[2298] Franz Dietrich and Antonios Staras. Reasoning in attitudes. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03986-3.

Goffin:2022:FBR

[2299] Kris Goffin. Feeling is believing: recalcitrant emotion & Spinozan belief formation. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03980-9.

Bar-On:2022:PTS

[2300] Kati Kish Bar-On. From philosophical traditions to scientific developments: reconsidering the response to Brouwer’s intuitionism. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03908-3.

Adlam:2022:DSN

[2301] Emily Adlam. Does science need intersubjectivity? The problem of confirmation in orthodox interpretations of quantum mechanics. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03989-0.

Wolf:2022:IPP

[2302] Julia Wolf. Implications of pretend play for theory of mind research. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03984-5.

Carl:2022:UMT

[2303] Merlin Carl. Understanding mathematical texts: a hermeneutical approach. Synthese, 200(6):??, December 2022. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-03930-5.

Kenan:2022:STC

[2304] Avraham Max Kenan. Simply too complex: against non-conceptual representation of (most) complex properties. Synthese, 200(6):??, December 2022. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03992-5.

Thompson:2023:ERM

[2305] Morgan Thompson. Epistemic risk in methodological triangulation: the case of implicit attitudes. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03943-0.

Bieber:2023:PBE

[2306] Friedemann Bieber and Juri Viehoff. A paradigm-based explanation of trust. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03993-4.

Veit:2023:HCE

[2307] Walter Veit. Health, consciousness, and the evolution of subjects. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03998-z.

Sekowski:2023:RUL

[2308] Krzysztof Sekowski. Reference the untouchable. On the limits of revising concepts using the method of cases. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03990-7.

Warren:2023:IP

[2309] Jared Warren. Imagination and the a priori. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03988-1.

Battaly:2023:ADE

[2310] Heather Battaly. Access denied: epistemic obstruction and the distribution of knowledge. Synthese, 201(1):??, January 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-04007-z.

Vidal:2023:ABS

[2311] Oriol Vidal and David Teira. Are animal breeds social kinds? Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-04016-y.

Megger:2023:CVE

[2312] Dawid Megger and Igor Wysocki. Coercion, voluntary exchange, and the Austrian School of Economics. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-04005-1.

Kastner:2023:MPM

[2313] Lena Kästner. Modeling psychopathology: 4D multiplexes to the rescue. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-04008-y.

Brown:2023:QCU

[2314] Christopher Devlin Brown. Quantum computation and the untenability of a “No fundamental mentality” constraint on physicalism. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-04015-z.

Malinowska:2023:TMP

[2315] Joanna Karolina Malinowska and Tomasz Zuradzki. Towards the multileveled and processual conceptualisation of racialised individuals in biomedical research. Synthese, 201(1):??, January 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-04004-2.

Kelley:2023:AID

[2316] Mikayla Kelley and Sven Neth. Accuracy and infinity: a dilemma for subjective Bayesians. Synthese, 201(1):??, January 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-04019-9.

Fan:2023:CBC

[2317] Da Fan. From contrastivism back to contextualism. Synthese, 201 (1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-04010-4.

Chellappoo:2023:WSM

[2318] Azita Chellappoo and Jan Baedke. Where the social meets the biological: new ontologies of biosocial race. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-04006-0.

Perrett:2023:UW

[2319] Roy W. Perrett. The unreality of words. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03904-7.

Tambolo:2023:RAT

[2320] Luca Tambolo and Gustavo Cevolani. Realism, antirealism, and theoretical conservatism. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-04022-0.

Egre:2023:TFB

[2321] Paul Égré. Truth and falsity in Buridan’s Bridge. Synthese, 201 (1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03907-4.

Munroe:2023:CWI

[2322] Wade Munroe. Correction to: What it takes to make a word (token). Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-04024-y. See [2066].

Schwengerer:2023:CVC

[2323] Lukas Schwengerer. Collective vice and collective self-knowledge. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-03969-4.

Benitez:2023:PFO

[2324] Federico Benitez. Particles, fields, and the ontology of the standard model. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-04035-9.

Hindriks:2023:RAM

[2325] Frank Hindriks and Herman Veluwenkamp. The risks of autonomous machines: from responsibility gaps to control gaps. Synthese, 201 (1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-04001-5.

Pettigrew:2023:NCS

[2326] Richard Pettigrew. Nudging for changing selves. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-04020-2.

McCarroll:2023:CTF

[2327] Christopher Jude McCarroll and André Sant’Anna. Cryptomnesia: a three-factor account. Synthese, 201(1):??, January 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-04002-4.

Doroudi:2023:WRW

[2328] Shayan Doroudi. What is a related work? A typology of relationships in research literature. Synthese, 201(1):??, January 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03976-5.

Bagwell:2023:DIT

[2329] Jeffrey N. Bagwell. Debunking interface theory: why Hoffman’s skepticism (really) is self-defeating. Synthese, 201(1):??, January 2023. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-022-04021-
1.

Skoczen:2023:RAV

[2330] Izabela Skoczeń. Renouncing the attempt versus perpetration distinction. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-04000-6.

McLoone:2023:CSE

[2331] Brian McLoone, Cassandra Grützner, and Michael T. Stuart. Counterpossibles in science: an experimental study. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-04014-0.

Litland:2023:GFA

[2332] Jon Erling Litland. Ground first: against the proof-theoretic definition of ground. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-04028-8.

Lopez:2023:MUT

[2333] Cristian López. The metaphysical underdetermination of time-reversal invariance. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04039-z.

Viola:2023:DAD

[2334] Marco Viola and Cristina Voto. Designed to abuse? Deepfakes and the non-consensual diffusion of intimate images. Synthese, 201(1):??, January 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-04012-2.

Arcangeli:2023:AD

[2335] Margherita Arcangeli. Aphantasia demystified. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-04027-9.

Hanschmann:2023:MSG

[2336] Jason Hanschmann. Memberless social groups. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-022-04034-w.

Nizzardo:2023:WAH

[2337] Matteo Nizzardo. Why I am not an anti-Haecceitist. Synthese, 201 (2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-04029-7.

Sartori:2023:PEB

[2338] Lorenzo Sartori. Putting the ‘experiment’ back into the ‘thought experiment’. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-022-04011-3.

Hansen:2023:IEK

[2339] Jens Ulrik Hansen and Paula Quinon. The importance of expert knowledge in big data and machine learning. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04041-5.

Baron:2023:PRS

[2340] Sam Baron, Kristie Miller, and Jonathan Tallant. Presentism and representation: saying it without words. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03987-2.

Wu:2023:HSW

[2341] Jingyi Wu and Cailin O’Connor. How should we promote transient diversity in science? Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04037-1.

Panahy:2023:SP

[2342] Salman Panahy. Synthetic proofs. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-04026-w.

Rumfitt:2023:DP

[2343] Ian Rumfitt. In defence of PKF. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-04009-x.

Binney:2023:LFR

[2344] Nicholas Binney. Ludwik Fleck’s reasonable relativism about science. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-04018-w.

Matthews:2023:DFB

[2345] Taylor Matthews. Deepfakes, fake barns, and knowledge from videos. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-04033-x.

Herok:2023:INU

[2346] Tomasz Herok. Intuitions are never used as evidence in ethics. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-04031-z.

Roberts:2023:HDT

[2347] Tom Roberts. How to do things with deepfakes. Synthese, 201 (2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04044-2.

Jacoby:2023:EJE

[2348] Franklin Jacoby. Evidence, justification, and epistemic standards. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04040-6.

Keren:2023:DPG

[2349] Arnon Keren. Disagreement, progress, and the goal of philosophy. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-022-04030-0.

Ding:2023:SKL

[2350] Yifeng Ding, Jixin Liu, and Yanjing Wang. Someone knows that local reasoning on hypergraphs is a weakly aggregative modal logic. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-04032-y.

Andow:2023:EDE

[2351] James Andow. Epistemic deontology, epistemic trade-offs, and Kant’s formula of humanity. Synthese, 201(2):??, February 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-023-04043-3.

Williams:2023:AME

[2352] J. Robert G. Williams. Aptness and means-end coherence: a dominance argument for causal decision theory. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/

s11229-022-04017-x. See correction [2484].

Brunet:2023:CMR

[2353] Tyler D. P. Brunet. Categorial modal realism. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-023-04042-4.

Plunkett:2023:GMN

[2354] David Plunkett, Rachel Katharine Sterken, and Timothy Sundell. Generics and metalinguistic negotiation. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03862-0.

Tamir:2023:MUD

[2355] Michael Tamir and Elay Shech. Machine understanding and deep learning representation. Synthese, 201(2):??, February 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-022-03999-y.

Tourville:2023:CCC

[2356] Nicholas Tourville and Roy T. Cook. Conditionals, curry, and consequence: embracing deduction. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-04013-1.

Glass:2023:HGE

[2357] David H. Glass. How good is an explanation? Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-04025-x.

Brenner:2023:TVM

[2358] Andrew Brenner. Theoretical virtues and the methodological analogy between science and metaphysics. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04047-z.

Bazhenov:2023:HAF

[2359] Nikolay Bazhenov, Manat Mustafa, Sergei Ospichev, and Luca San Mauro. How to approximate fuzzy sets: mind-changes and the Ershov Hierarchy. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04056-y.

Wilcox:2023:AHA

[2360] John Wilcox. Ad hocness, accommodation and consilience: a Bayesian account. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04036-2.

Copeland:2023:CAT

[2361] B. Jack Copeland and Diane Proudfoot. Correction to: On Alan Turing’s anticipation of connectionism. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-04023-z. See [1].

Bonk:2023:FSC

[2362] Thomas Bonk. Functionspaces, simplicity and curve fitting. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-022-03906-5.

Tiefensee:2023:TTT

[2363] Christine Tiefensee. Through thick and thin: seamless metaconceptualism. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03913-6.

Henne:2023:FFI

[2364] Céline Henne. Framed and framing inquiry: a pragmatist proposal. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04059-9.

Callahan:2023:JAV

[2365] Laura Frances Callahan. (Joint) achievements and the value problem. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-022-03860-2.

Ferrari:2023:AEA

[2366] Filippo Ferrari, Ben Martin, and Maria Paola Fogliani Sforza. Antiexceptionalism about logic: an overview. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04082-w.

Chen:2023:MNC

[2367] Kuei-Chen Chen and Jeff Yoshimi. The metaphysical neutrality of cognitive science. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04046-0.

Singleton:2023:EIE

[2368] Joseph Singleton and Richard Booth. Expertise and information: an epistemic logic perspective. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04064-y.

Tan:2023:IMT

[2369] Peter Tan. Interdisciplinary model transfer and realism about physical analogy. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04065-x.

Michaelson:2023:UN

[2370] Eliot Michaelson. Unspeakable names. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04038-0.

Nesic:2023:EEA

[2371] Janko Nesić. Ecological-enactive account of autism spectrum disorder. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04073-x.

Kuipers:2023:CTD

[2372] Theo A. F. Kuipers. A coherent trio of, distance and size based, measures for nomic and actual truthlikeness. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-022-03977-4.

Teixeira:2023:AP

[2373] Célia Teixeira. Abducting the a priori. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04050-4.

Stojnic:2023:BC

[2374] Una Stojnić. Being called. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-022-03940-3.

Saarinen:2023:EEF

[2375] Jussi A. Saarinen. Extending existential feeling through sensory substitution. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04049-x.

Indrzejczak:2023:WIM

[2376] Andrzej Indrzejczak and Michał Zawidzki. When iota meets lambda. Synthese, 201(2):??, February 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04048-y.

Jiang:2023:TCD

[2377] Asher Jiang. Truthmaking cannot be done afar. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04052-2.

Baratella:2023:PTM

[2378] Riccardo Baratella. Processes and their modal profile. Synthese, 201 (3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04051-3.

Tahko:2023:MBS

[2379] Tuomas E. Tahko. The modal basis of scientific modelling. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04063-z.

Strandberg:2023:RDI

[2380] Caj Strandberg. Realist dependence and irrealist butterflies. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04072-y.

Naumov:2023:ELP

[2381] Pavel Naumov and Anna Ovchinnikova. An epistemic logic of preferences. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04060-2.

Wolff:2023:COR

[2382] J. E. Wolff. Coordination and obsolescence: a response on behalf of measurement realism. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04061-1.

Houkes:2023:ECT

[2383] Wybo Houkes. Embedding and customizing templates in crossdisciplinary modeling. Synthese, 201(3):??, March 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-023-04076-8.

Bernecker:2023:GES

[2384] Sven Bernecker. Greco’s explanatory salience contextualism revisited. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04067-9.

Golemon:2023:DVN

[2385] Luke Golemon and Abraham Graber. A deductive variation on the no miracles argument. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04078-6.

Donhauser:2023:FLT

[2386] Justin Donhauser. Five lessons from teleology-neutrality and metaphor in ecology: bottom-up and top-down all at once. Synthese, 201(3): ??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-023-04079-5.

Baz:2023:PIO

[2387] Avner Baz. On philosophical idling: the ordinary language philosophy critique of the philosophical method of cases. Synthese, 201 (3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04070-0.

Fischer:2023:FCF

[2388] Eugen Fischer, Keith Allen, and Paul E. Engelhardt. Fragmented and conflicted: folk beliefs about vision. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-023-04066-
w.

vanLier:2023:IFF

[2389] Maud van Lier. Introducing a four-fold way to conceptualize artificial agency. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04083-9.

Dixon:2023:MF

[2390] Scott Dixon. Metaphysical foundherentism. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04053-1.

Villiger:2023:RTD

[2391] Daniel Villiger. Rational transformative decision-making. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04075-9.

Tunc:2023:SKC

[2392] Duygu Uygun Tunç. The subject of knowledge in collaborative science. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04080-y.

Brun:2023:ITC

[2393] Georg Brun, Catherine Herfeld, and Kevin Reuter. Introduction to the topical collection: Concept formation in the natural and social sciences: empirical and normative aspects. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04094-6.

Langsam:2023:TKP

[2394] Harold Langsam. Two kinds of a priori justification. Synthese, 201 (3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04087-5.

Budzynska:2023:AAA

[2395] Katarzyna Budzynska and Chris Reed. Argumentum Ad Alia: argument structure of arguing about what others have said. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04058-w.

Weckstrom:2023:NLM

[2396] Markus Mikael Weckström. Natural law, the modeling relation, and two roots of perspectivism. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04092-8.

Latham:2023:ANA

[2397] Andrew J. Latham, Kristie Miller, and James Norton. Against a normative asymmetry between near- and future-bias. Synthese, 201 (3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04045-1. See correction [2483].

FernandezVelasco:2023:EPC

[2398] Pablo Fernandez Velasco. Entropy, prediction and the cultural ecosystem of human cognition. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04089-3.

Gillespie:2023:ACE

[2399] James Gillespie. Agent-centered epistemic rationality. Synthese, 201 (3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04088-4.

Hyska:2023:LVC

[2400] Megan Hyska. Luck and the value of communication. Synthese, 201 (3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04077-7.

Pierini:2023:DDE

[2401] Francesco Pierini. Deepfakes and depiction: from evidence to communication. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04093-7.

Armour-Garb:2023:ACC

[2402] Bradley Armour-Garb and James A. Woodbridge. Answering the conceptual challenge: three strategies for deflationists. Synthese, 201 (3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04071-z.

Skyrms:2023:QC

[2403] Brian Skyrms. Quasi-conventions. Synthese, 201(3):??, March 2023. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-023-04103-
8.

Aumann:2023:AIE

[2404] Antony Aumann. Art, imagination, and experiential knowledge. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04109-2.

Grant:2023:EOR

[2405] David Gray Grant. Equalized odds is a requirement of algorithmic fairness. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04054-0.

Fischer:2023:COL

[2406] Eugen Fischer. Critical ordinary language philosophy: A new project in experimental philosophy. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04068-8.

Habgood-Coote:2023:DEA

[2407] Joshua Habgood-Coote. Deepfakes and the epistemic apocalypse. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04097-3.

Sarkia:2023:TAN

[2408] Matti Sarkia and Tuukka Kaidesoja. Two approaches to naturalistic social ontology. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04105-6.

Piovarchy:2023:VEN

[2409] Adam Piovarchy. Vice epistemology, norm-maintenance and epistemic evasiveness. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04086-6.

Calosi:2023:UER

[2410] Claudio Calosi and Alessandro Giordani. Universalism and extensionalism revisited. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04091-9.

Noichl:2023:HLC

[2411] Maximilian Noichl. How localized are computational templates? A machine learning approach. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04057-x.

Ziolkowski:2023:TCS

[2412] Adrian Ziółkowski, Alex Wiegmann, Joachim Horvath, and Edouard Machery. Truetemp cooled down: the stability of Truetemp intuitions. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04055-z.

BarrancoLopez:2023:HCA

[2413] Aliosha Barranco Lopez. Hinge commitments as arational beliefs. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04090-w.

Collins:2023:IPP

[2414] John Collins. Internalist priorities in a philosophy of words. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04074-w.

Buchholz:2023:DNN

[2415] Oliver Buchholz. The deep neural network approach to the reference class problem. Synthese, 201(3):??, March 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04110-9.

Babcock:2023:TFN

[2416] Gunnar Babcock. Teleology and function in non-living nature. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04099-1.

Poslajko:2023:MMG

[2417] Krzysztof Poslajko. The meta-metaphysics of group beliefs: in search of alternatives. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04108-3.

Haziza:2023:QAK

[2418] Eliran Haziza. Questioning and addressee knowledge. Synthese, 201 (4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-023-04113-6.

Petracca:2023:HCE

[2419] Enrico Petracca and James Grayot. How can embodied cognition naturalize bounded rationality? Synthese, 201(4):??, April 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-023-04124-3. See cor-

rection [2474].

Boge:2023:ISC

[2420] Florian J. Boge, Miguel-Ángel Carretero-Sahuquillo, Paul Grünke, and Martin King. Introduction: Simplicity out of complexity? Physics and the aims of science. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04126-1.

Hill:2023:WWV

[2421] Jesse Hill and James Fanciullo. What’s wrong with virtue signaling? Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04131-4.

Scott:2023:SEI

[2422] Jordan Scott. Should epistemic instrumentalists be more social? Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04118-1.

Dewar:2023:IEE

[2423] Neil Dewar. Interpretation and equivalence; or, equivalence and interpretation. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04102-9.

Kao:2023:IRI

[2424] Molly Kao, Deborah G. Mayo, and Elay Shech. Introduction to recent issues in philosophy of statistics: evidence, testing, and applications. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04128-z.

Sanchez-Canizares:2023:MEP

[2425] Javier Sánchez-Cañizares. Is the maximum entropy production just a heuristic principle? Metaphysics on natural determination. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04129-y.

Turyn:2023:GFP

[2426] Gus Turyn. Gender and first-person authority. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04125-2.

Wang:2023:CTB

[2427] Xun Wang. Completeness theorems for ∃□-bundled fragment of firstorder modal logic. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04104-7.

Strollo:2023:TBP

[2428] Andrea Strollo. Truth (bearers) pluralism. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04081-x.

Boumans:2023:MSE

[2429] Marcel Boumans. Materials selection in economic modeling. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04120-7.

Pisano:2023:ITM

[2430] Niccolò Aimone Pisano. An instrumentalist take on the models of the Free-Energy Principle. Synthese, 201(4):??, April 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-023-04111-8.

Pearson:2023:JEA

[2431] Phyllis Pearson. Justification and epistemic agency. Synthese, 201(4): ??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-023-04115-4. See correction [2481].

Trappes:2023:HTT

[2432] Rose Trappes. How tracking technology is transforming animal ecology: epistemic values, interdisciplinarity, and technology-driven scientific change. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04122-5.

Lenhard:2023:TTP

[2433] Johannes Lenhard and Hans Hasse. Traveling with TARDIS. Parameterization and transferability in molecular modeling and simulation. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04116-3.

Barrantes:2023:SEI

[2434] Manuel Barrantes. Structural explanations: impossibilities vs failures. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04121-6.

Li:2023:CMA

[2435] Yanjun Li. Complexity of multi-agent conformant planning with group knowledge. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04095-5.

Owesen:2023:WLT

[2436] Erlend Owesen. What-it’s-like talk is technical talk. Synthese, 201 (4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-023-04123-4.

Bertea:2023:AWP

[2437] Stefano Bertea. Is the autonomy of the will a paradoxical idea? Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04112-7.

Schmidt:2023:BSE

[2438] Eva Schmidt, Andreas Sesing-Wagenpfeil, and Maximilian A. Köhl. Bare statistical evidence and the legitimacy of software-based judicial decisions. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04141-2.

Ortmann:2023:ERE

[2439] Andreas Ortmann and Leonidas Spiliopoulos. Ecological rationality and economics: where the Twain shall meet. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04136-z.

Turyn:2023:CDM

[2440] Gus Turyn. Correction to: On dispositional masks. Synthese, 201(4): ??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-023-04139-w. See [1684].

Sandmann:2023:II

[2441] Jan-Paul Sandmann. Irrationality and indecision. Synthese, 201(4): ??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-023-04107-4.

Morales:2023:EAC

[2442] Juan Morales. Emergent agent causation. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04117-2.

Schoenmakers:2023:SPN

[2443] Ludo L. J. Schoenmakers. Scientific progress, normative discussions, and the pragmatic account of definitions of life. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04085-7.

Zhang:2023:CCD

[2444] Xiaoxing Zhang. Correction to: Closure, deduction and hinge commitments. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04132-3. See [451].

Kiani:2023:SPS

[2445] Amirhossein Kiani. Structured propositions and a semantics for unrestricted impure logics of ground. Synthese, 201(4):??, April 2023. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-023-04114-
5.

Nahas:2023:WSD

[2446] Auguste Nahas and Carl Sachs. What’s at stake in the debate over naturalizing teleology? An overlooked metatheoretical debate. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04147-w.

Robertson:2023:AIA

[2447] Ian Robertson and Daniel D. Hutto. Against intellectualism about skill. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04096-4.

McAllister:2023:RTA

[2448] Blake McAllister. Rescuing a traditional argument for internalism. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04119-0.

Guala:2023:NSS

[2449] Francesco Guala and Frank Hindriks. The nature and significance of social ontology. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04142-1.

Chapman:2023:WAM

[2450] Siobhan Chapman. When Arne met J. L.: attitudes to scientific method in empirical semantics, ordinary language philosophy and linguistics. Synthese, 201(4):??, April 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04106-5.

Garcia:2023:EIN

[2451] Claudia Lorena García and Alejandro Vázquez del Mercado. Epistemological issues in neurodivergence and atypical cognition: introduction. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04100-x.

Meyer:2023:EVC

[2452] Marco Meyer. The epistemic vices of corporations. Synthese, 201(5): ??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04133-2.

Coliva:2023:HKE

[2453] Annalisa Coliva. Hinges in the knowledge economy. On Greco’s common and procedural knowledge. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04130-5.

Bogucki:2023:DAV

[2454] Krystian Bogucki. A defence of the austere view of nonsense. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04138-x.

Elkind:2023:CCV

[2455] Landon D. C. Elkind. Correction to: Computer verification for historians of philosophy. Synthese, 201(5):??, May 2023. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04151-0. See [1977].

Adlam:2023:TCF

[2456] Emily Adlam. Is there causation in fundamental physics? New insights from process matrices and quantum causal modelling. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04160-z.

Pruss:2023:DAA

[2457] Alexander R. Pruss. The dialectics of accuracy arguments for probabilism. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04145-y.

Marfori:2023:EHP

[2458] Marianna Antonutti Marfori, Sorin Bangu, and Emiliano Ippoliti. The explanatory and heuristic power of mathematics. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04127-0.

Gambarotto:2023:TMD

[2459] Andrea Gambarotto. Teleology and mechanism: a dialectical approach. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04137-y.

Smith:2023:OMA

[2460] Martin Smith. An objection to the modal account of risk. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04134-1.

Riccardi:2023:SBN

[2461] Mattia Riccardi. Sameness beyond numerical identity. A defence of the One Object View of Kant’s transcendental idealism. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04159-6.

Brown:2023:CBM

[2462] Rebecca C. H. Brown, Mícheál de Barra, and Brian D. Earp. Correction to: Broad Medical Uncertainty and the ethical obligation for openness. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04155-w. See [1900].

Yan:2023:CCA

[2463] Karen Yan and Chuan-Ya Liao. A co-citation analysis of crossdisciplinarity in the empirically-informed philosophy of mind. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04146-x.

Drozdzowicz:2023:LMS

[2464] Anna Drozdzowicz. Linguistic modalities and the sources of linguistic utterances. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04062-0.

Dohrn:2023:MMM

[2465] Daniel Dohrn. Modals model models: scientific modeling and counterfactual reasoning. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04135-0.

Olafsson:2023:UKT

[2466] Ísak Andri Ólafsson. Unwanted knowledge transmission. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04140-3.

Das:2023:IQM

[2467] Ramit Das, Anantha Padmanabha, and R. Ramanujam. Implicit quantification for modal reasoning in large games. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04156-9.

Rybarikova:2023:STT

[2468] Zuzana Rybaríková. Specification of time in Tichý’s transparent intensional logic and Prior’s temporal logic. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04098-2.

Pendoley:2023:SES

[2469] Kate Pendoley. Stubborn emotions, stubborn beliefs. Synthese, 201 (5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-023-04165-8.

Liu:2023:XPC

[2470] Michelle Liu. X-Phi and the challenge from ad hoc concepts. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04149-8.

Bonard:2023:NMP

[2471] Constant Bonard. Natural meaning, probabilistic meaning, and the interpretation of emotional signs. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04144-z.

Levy:2023:EDA

[2472] Arnon Levy and Itamar Weinshtock Saadon. Evolutionary debunking of (arguments for) moral realism. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04157-8.

Kelp:2023:HC

[2473] Christoph Kelp. How to be a capacitist. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04143-0.

Petracca:2023:CHC

[2474] Enrico Petracca and James Grayot. Correction to: How can embodied cognition naturalize bounded rationality? Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04152-z. See [2419].

Verreault-Julien:2023:TMD

[2475] Philippe Verreault-Julien. Toy models, dispositions, and the power to explain. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04084-8.

Kroedel:2023:NEN

[2476] Thomas Kroedel. Norms, epistemic norms, context, and counterfactuals. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04162-x.

Serrahima:2023:ED

[2477] Carlota Serrahima and Manolo Martínez. The experience of dysmenorrhea. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04148-9.

Gasparri:2023:FWE

[2478] Luca Gasparri. The first words ever spoken. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04166-7.

Brandenburger:2023:RBS

[2479] Adam Brandenburger, Amanda Friedenberg, and H. Jerome Keisler. The relationship between strong belief and assumption. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04167-6.

Dziurosz-Serafinowicz:2023:EDA

[2480] Patryk Dziurosz-Serafinowicz. Expert deference and Adams conditionalization. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04175-6.

Pearson:2023:CJE

[2481] Phyllis Pearson. Correction to: Justification and epistemic agency. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04170-x. See [2431].

McShea:2023:ETG

[2482] Daniel W. McShea. Evolutionary trends and goal directedness. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04164-9.

Latham:2023:CAN

[2483] Andrew J. Latham, Kristie Miller, and James Norton. Correction to: Against a normative asymmetry between near- and future-bias. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04171-w. See [2397].

Williams:2023:CAM

[2484] J. Robert G. Williams. Correction to: Aptness and means-end coherence: a dominance argument for causal decision theory. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04154-x. See [2352].

Shech:2023:INS

[2485] Elay Shech and Alison A. Springle. Inductive neutrality and scientific representation. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04161-y.

FernandezPinto:2023:EDI

[2486] Manuela Fernández Pinto and Daniel Fernández Pinto. Epistemic diversity and industrial selection bias. Synthese, 201(5):??, May 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04158-7. See

correction [2507].

Gordon:2023:RHD

[2487] Noah Gordon. Resurrecting the Hume’s Dictum argument against metaethical non-naturalism. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04173-8.

Campi:2023:IKU

[2488] Marco C. Campi. Inductive knowledge under dominance. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04172-9.

Keeling:2023:ADA

[2489] Sophie Keeling. Accounting for doxastic agency: Mental action and self-awareness. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04069-7.

Loi:2023:HMD

[2490] Michele Loi, Andrea Ferrario, and Eleonora Viganò. How much do you trust me? A logico-mathematical analysis of the concept of the intensity of trust. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04169-4.

Caponetto:2023:PSR

[2491] Laura Caponetto. The pragmatic structure of refusal. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/

10.1007/s11229-023-04177-4. See publisher correction [2522].

Nes:2023:PNM

[2492] Anders Nes. Perception needs modular stimulus-control. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04181-8.

Bailey:2023:WML

[2493] Olivia Bailey. What must be lost: on retrospection, authenticity, and some neglected costs of transformation. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04179-2.

Battaly:2023:CFL

[2494] Heather Battaly. Can fanaticism be a liberatory virtue? Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04174-7.

Barrett:2023:ACC

[2495] Thomas William Barrett, J. B. Manchak, and James Owen Weatherall. On automorphism criteria for comparing amounts of mathematical structure. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04186-3.

Rudnicki:2023:DDB

[2496] Jakub Rudnicki and Joanna Odrowaz-Sypniewska. Don’t be deceived: bald-faced lies are deceitful assertions. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04180-9.

Schirn:2023:FIR

[2497] Matthias Schirn. Frege on the introduction of real and complex numbers by abstraction and cross-sortal identity claims. Synthese, 201(6): ??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04150-1.

Smart:2023:CBM

[2498] Benjamin T. H. Smart. The core business of medicine: a defence of the best available intervention thesis. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/

s11229-023-04176-5. See publisher correction [2621].

Tarsney:2023:ECL

[2499] Christian Tarsney. The epistemic challenge to longtermism. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04153-y.

Koskinen:2023:KMM

[2500] Rami Koskinen. Kinds of modalities and modeling practices. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04185-4.

Wysocki:2023:CC

[2501] Tomasz Wysocki. Conjoined cases. Synthese, 201(6):??, June 2023. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-023-04101-
w.

Gordon:2023:UNP

[2502] Emma C. Gordon. Understanding of the norm of political discourse. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04183-6.

Jaksland:2023:NMD

[2503] Rasmus Jaksland. Naturalized metaphysics or displacing metaphysicians to save metaphysics. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04207-1.

Knuuttila:2023:MTT

[2504] Tarja Knuuttila and Andrea Loettgers. Model templates: transdisciplinary application and entanglement. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-023-04178-
3.

Corbi:2023:DSS

[2505] Josep E. Corbí. Distinctive substantial self-knowledge and the possibility of self-improvement. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04195-2.

Elzinga:2023:KHW

[2506] Benjamin Elzinga. Know-how and why self-regulation will not go away. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04168-5.

FernandezPinto:2023:CED

[2507] Manuela Fernández Pinto and Daniel Fernández Pinto. Correction to: Epistemic diversity and industrial selection bias. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04205-3. See [2486].

Richardson:2023:CSO

[2508] Kevin Richardson. Critical social ontology. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04197-0.

Hricko:2023:CSE

[2509] Jonathon Hricko. Comparative success and empirical progress without approximate truth. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04204-4.

Hung:2023:PPA

[2510] Tzu-Wei Hung and Chun-Ping Yen. Predictive policing and algorithmic fairness. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04189-0.

Larese:2023:HCS

[2511] Costanza Larese. Hintikka’s conception of syntheticity as the introduction of new individuals. Synthese, 201(6):??, June 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-023-04200-8.

Wirling:2023:IST

[2512] Ylwa Sjölin Wirling and Till Grüne-Yanoff. Introduction to the Synthese topical collection ‘modal modeling in science: modal epistemology meets philosophy of science’. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04188-1.

Noh:2023:IOU

[2513] Hyungrae Noh. Interpreting ordinary uses of psychological and moral terms in the AI domain. Synthese, 201(6):??, June 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-023-04194-3.

Sytsma:2023:EOP

[2514] Justin Sytsma and Eugen Fischer. ‘Experience’, ordinary and philosophical: a corpus study. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04190-7.

Grim:2023:IT

[2515] Patrick Grim and Nicholas Rescher. Influence theory. Synthese, 201 (6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-023-04163-w.

Lipski:2023:EEF

[2516] Jonas Lipski. Explanatory efficiency: a framework for analyzing the dynamic properties of explanatory games. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04208-0.

Degerman:2023:EIN

[2517] Dan Degerman. Epistemic injustice, naturalism, and mental disorder: on the epistemic benefits of obscuring social factors. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04210-6.

Reijula:2023:DCL

[2518] Samuli Reijula, Jaakko Kuorikoski, and Miles MacLeod. The division of cognitive labor and the structure of interdisciplinary problems. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04193-4.

Mehl:2023:MMO

[2519] Kayla R. Mehl. The medical model of “obesity” and the values behind the guise of health. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04209-z.

Engler:2023:WWL

[2520] Jann Paul Engler. Wittgenstein on Weyl: the law of the excluded middle and the natural numbers. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04199-y.

Horvath:2023:AAT

[2521] Joachim Horvath and Alex Wiegmann. Arguing about thought experiments. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04182-7.

Caponetto:2023:PCP

[2522] Laura Caponetto. Publisher correction to: The pragmatic structure of refusal. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04214-2. See [2491].

dAragona:2023:PTS

[2523] Antonio Piccolomini d’Aragona. The proof-theoretic square. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04203-5.

ChoGlueck:2023:VHC

[2524] Christopher ChoGlueck and Elisabeth A. Lloyd. Values as heuristics: a contextual empiricist account of assessing values scientifically. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04198-z.

Brancazio:2023:IAD

[2525] Nick Brancazio. Interactive agential dynamics. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04192-5.

Rowson:2023:EEP

[2526] Rebecca Rowson. Expressions of emotion as perceptual media. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04212-4.

Settegast:2023:FGH

[2527] Sascha Settegast. Fineness of grain and the hylomorphism of experience. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04201-7.

Baave:2023:CCP

[2528] Arvid Båve. Correctness conditions for property nominalists. Synthese, 201(6):??, June 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04206-2.

Porter:2023:QHO

[2529] J. D. Porter and Nat Hansen. A quantitative history of ordinary language philosophy. Synthese, 201(6):??, June 2023. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04187-2.

Puddifoot:2023:EAG

[2530] Katherine Puddifoot and Marina Trakas. Epistemic agency and the generalisation of fear. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04222-2.

Martinez:2023:CPA

[2531] Joseph Martinez. Causal pluralism: agent causation without the panicky metaphysics. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04216-0.

Gustafsson:2023:TTS

[2532] Martin Gustafsson. Taking truth seriously: the case of generics. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04213-3.

deOliveira:2023:SPE

[2533] Guilherme Sanches de Oliveira, Thomas van Es, and Inês Hipólito. Scientific practice as ecological-enactive co-construction. Synthese, 202(1): ??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04215-1.

Datteri:2023:CSS

[2534] Edoardo Datteri and Viola Schiaffonati. Computer simulations and surrogative reasoning for the design of new robots. Synthese, 202(1): ??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04220-4.

Tanesini:2023:IAI

[2535] Alessandra Tanesini. Intellectual arrogance: individual, group-based, and corporate. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04224-0.

Wanderer:2023:IE

[2536] Jeremy Wanderer and Leo Townsend. Illocution by example. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04229-9.

Schumann:2023:FFL

[2537] Andrew Schumann and Jens Lemanski. Fichte’s formal logic. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04218-y.

Rudnicki:2023:SCR

[2538] Jakub Rudnicki. Semantic conventions and referential intentions. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04227-x.

Siu:2023:IBD

[2539] Charlie Siu. Interval-based dynamics of loose talk. Synthese, 202(1): ??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04231-1.

Chung:2023:MEU

[2540] WooJin Chung and Salvador Mascarenhas. Modality, expected utility, and hypothesis testing. Synthese, 202(1):??, July 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-023-04191-6.

Rohr:2023:FHC

[2541] Tabea Rohr. The Frege–Hilbert controversy in context. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04196-1.

Barnett:2023:CM

[2542] David James Barnett. Cogito and Moore. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04230-2.

Krickel:2023:HWT

[2543] Beate Krickel, Leon de Bruin, and Linda Douw. How and when are topological explanations complete mechanistic explanations? The case of multilayer network models. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04241-z.

Paulson:2023:GRA

[2544] Spencer Paulson. Good reasons are apparent to the knowing subject. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04242-y.

Jahn:2023:DYB

[2545] Marcel Jahn and Lukas Beck. Do you believe in Deep Down? On two conceptions of valuing. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04221-3.

Wilson:2023:PAT

[2546] Joseph Wilson. Paleoclimate analogues and the threshold problem. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04202-6.

Jago:2023:WBT

[2547] Mark Jago. On ways of being true. Synthese, 202(1):??, July 2023. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-023-04234-

y. See correction [2842].

Peralta:2023:SP

[2548] Esteban Peralta and Fernando Tohmé. The scope of provability. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04238-8.

Rosenqvist:2023:SCP

[2549] Tiina Carita Rosenqvist. Seeing with color: Psychophysics and the function of color vision. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04226-y.

Dimitrakos:2023:DKA

[2550] Thodoris Dimitrakos. Do Kuhnians have to be anti-realists? Towards a realist reconception of Kuhn’s historiography. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04225-z.

Miyazono:2023:GIA

[2551] Kengo Miyazono and Rie Iizuka. A group identification account of collective epistemic vices. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04235-x.

Coumans:2023:DPI

[2552] V. J. W. Coumans and L. Consoli. Definitions in practice: an interview study. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04240-0.

Zhang:2023:HUG

[2553] Wenjun Zhang. How to unify grounding and causation. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04237-9.

Glick:2023:PLA

[2554] David Glick. The principle of least action and teleological explanation in physics. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04251-x.

Crippen:2023:CMA

[2555] Matthew Crippen. Conceptual and moral ambiguities of deepfakes: a decidedly old turn. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04250-y.

Hsiung:2023:NPV

[2556] Ming Hsiung. Necessity predicate versus truth predicate from the perspective of paradox. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04244-w.

Holm:2023:SEA

[2557] Sune Holm. Statistical evidence and algorithmic decision-making. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04246-8.

Lorenat:2023:RMC

[2558] Jemma Lorenat. Relocating mathematics: a case of moving texts between the front and back of mathematics. Synthese, 202(1):??, July 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04211-5.

Francis:2023:RCK

[2559] Kathryn B. Francis and C. Philip Beaman. The role of confidence in knowledge ascriptions: an evidence-seeking approach. Synthese, 202 (2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04236-w.

Lopez:2023:HTR

[2560] Cristian López and Michael Esfeld. Humean time-reversal symmetry. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04247-7.

Roe:2023:CAN

[2561] Henry Roe. Collective arrogance: a norms-based account. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04259-3.

Buchholz:2023:MEA

[2562] Oliver Buchholz. A means–end account of explainable artificial intelligence. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04260-w.

Weber:2023:ICS

[2563] Keith Weber. Instructions and constructions in set theory proofs. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04239-7.

Ribeiro:2023:AMM

[2564] Max Minden Ribeiro. An assembled message: Matthen on the content of perceptual experience. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04232-0.

Rolin:2023:SCD

[2565] Kristina Rolin, Inkeri Koskinen, Jaakko Kuorikoski, and Samuli Reijula. Social and cognitive diversity in science: introduction. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04261-9.

Sinhababu:2023:QST

[2566] Neil Sinhababu. Qualia share their correlates’ locations. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04256-6.

Arana:2023:DDM

[2567] Andrew Arana and Will Stafford. On the difficulty of discovering mathematical proofs. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04184-5.

Turska:2023:BPD

[2568] Julia J. Turska and David Ludwig. Back by popular demand, ontology. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04243-x.

Barber:2023:FEM

[2569] Alex Barber. Freedom of expression meets deepfakes. Synthese, 202 (2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04266-4.

Zavaleta:2023:PU

[2570] Martín Abreu Zavaleta. Partial understanding. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04268-2.

Eker:2023:PAT

[2571] Bahadir Eker. Perspectivalism about temporal reality. Synthese, 202 (2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04269-1.

James:2023:MRO

[2572] Aaron James. Money, recognition, and the outer limits of obliviousness. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04252-w.

Rossi:2023:HAP

[2573] Niccolò Rossi and Aybüke Özgün. A hyperintensional approach to positive epistemic possibility. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04219-x. See

correction [2628].

Randell:2023:FTE

[2574] Petronella Randell. Familiar transformative experiences. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04275-3.

Wagner:2023:PCP

[2575] Roy Wagner. Paul Cohen’s philosophy of mathematics and its reflection in his mathematical practice. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04273-5.

Symonds:2023:TDE

[2576] Hezki Symonds. Thomasson’s defence of easy arguments. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04217-z.

Shangin:2023:CFO

[2577] Vasily Shangin. A classical first-order normalization procedure with ∀ and ∃ based on the Milne–Kürbis approach. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04276-2.

Baron:2023:NDT

[2578] Sam Baron, Andrew J. Latham, and Somogy Varga. Non-dynamism and temporal disturbances. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04253-9.

Biebel:2023:PEJ

[2579] Nathan Biebel. Pragmatic encroachment and justified group belief. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04271-7.

Larsson:2023:RR

[2580] Felix Larsson and Martin Kaså. Re: the rhetic. Synthese, 202(2): ??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-023-04277-1.

Wiegmann:2023:DLR

[2581] Alex Wiegmann. Does lying require objective falsity? Synthese, 202 (2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04291-3.

Buckareff:2023:EMP

[2582] Andrei A. Buckareff and Jessica Hawkins. Emergent mental properties are not just double-preventers. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04282-4.

Schoenegger:2023:SSC

[2583] Philipp Schoenegger and Raimund Pils. Social sciences in crisis: on the proposed elimination of the discussion section. Synthese, 202 (2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04267-3.

Schiemer:2023:SIP

[2584] Georg Schiemer and John Wigglesworth. Structuralism and informal provability. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04274-4.

Ross:2023:ENC

[2585] Lauren N. Ross. The explanatory nature of constraints: Law-based, mathematical, and causal. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04281-5.

Assadian:2023:PR

[2586] Bahram Assadian and Giorgio Sbardolini. Performative reference. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04284-2.

Piekarski:2023:IVF

[2587] Michał Piekarski. Incorporating (variational) free energy models into mechanisms: the case of predictive processing under the free energy principle. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04292-2.

Kearl:2023:KHL

[2588] Timothy R. Kearl. Knowledge-how and the limits of defeat. Synthese, 202(2):??, August 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04280-6.

Koslicki:2023:PDS

[2589] Kathrin Koslicki and Olivier Massin. A plea for descriptive social ontology. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04263-7.

Gibbons:2023:PIB

[2590] Adam F. Gibbons. Political ignorance is both rational and radical. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04293-1.

Clinkenbeard:2023:CDT

[2591] Cayla Clinkenbeard. Collective deception: toward a network model of epistemic responsibility. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04299-9.

Chirimuuta:2023:HRN

[2592] M. Chirimuuta. Haptic realism for neuroscience. Synthese, 202(3): ??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04295-z.

Johnson:2023:IR

[2593] Casey Rebecca Johnson. Illocutionary relativism. Synthese, 202(3): ??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04289-x.

Egerton:2023:TEP

[2594] Karl Egerton and Helen Capitelli-McMahon. Transformative experience and the principle of informed consent in medicine. Synthese, 202 (3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04258-4.

Kuukkanen:2023:PRH

[2595] Jouni-Matti Kuukkanen. Pluralism in reasoning: how to legitimate material inferences. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04287-z.

Ross:2023:WDC

[2596] Ryan Ross. When to dismiss conspiracy theories out of hand. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04257-5.

Linquist:2023:EEC

[2597] Stefan Linquist, Brady Fullerton, and Akashdeep Grewal. Epigenetic this, epigenetic that: comparing two digital humanities methods for analyzing a slippery scientific term. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04265-5.

Halperin:2023:NEN

[2598] Topaz Halperin. Niches and ecological neutrality. Synthese, 202(3): ??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04278-0.

Noonan:2023:EAW

[2599] Christopher Noonan. An explanationist account of what goes wrong with third-factor replies to debunking arguments. Synthese, 202(3): ??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04270-8.

Lam:2023:LBS

[2600] Vincent Lam and Christian Wüthrich. Laws beyond spacetime. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04305-0.

Bourrat:2023:MSM

[2601] Pierrick Bourrat. Multilevel selection 1, multilevel selection 2, and the Price equation: a reappraisal. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04285-1.

Mendonca:2023:DGD

[2602] Bruno Ramos Mendonça. Dialogue games and deductive information: a dialogical account of the concept of virtual information. Synthese, 202 (3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04303-2.

Muller:2023:ECI

[2603] Julian F. Müller and Amin Ebrahimi Afrouzi. An egalitarian challenge to increasing epistemic value in democracy. Synthese, 202(3): ??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04283-3.

Hildebrandt:2023:RHC

[2604] Frauke Hildebrandt, Ramiro Glauer, and Richard Moore. Rethinking how children individuate objects: spatial indexicals in early development. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04300-5.

Hochstein:2023:IIM

[2605] Eric Hochstein. Integration without integrated models or theories. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04298-w.

Gottlieb:2023:DGD

[2606] Joseph Gottlieb and Saja Parvizian. Descartes’ God is a deceiver, and that’s OK. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04290-4.

Oleksowicz:2023:NDE

[2607] Michał Oleksowicz and Leandro Gaitán. Near-death experiences: feasibility and advantages of the mechanistic explanation. Synthese, 202 (3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04307-y.

Pedersen:2023:DBT

[2608] Mina Young Pedersen, Marija Slavkovik, and Sonja Smets. Detecting bots with temporal logic. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04264-6.

Phillips:2023:VBI

[2609] Callie K. Phillips. Value-based interpretationism. Synthese, 202(3): ??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04272-6.

Parkinson-Coombs:2023:RPI

[2610] Oisín Parkinson-Coombs and Rafael Núñez. Realism and the point at infinity: The end of the line? Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04228-w.

Grossi:2023:QIT

[2611] Marco Grossi. Quantification in the interpretational theory of validity. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04302-3.

Giraud:2023:OMT

[2612] Thibaut Giraud, Maicol Neves Leal, and Florian Cova. “One more time”: time loops as a tool to investigate folk conceptions of moral responsibility and human agency. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04245-9.

Allen:2023:MWD

[2613] Nicholas D. C. Allen. The misappropriation of “woke”: discriminatory social media practices, contributory injustice and context collapse. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04249-5.

Cappuccio:2023:DRK

[2614] Massimiliano L. Cappuccio. Dreyfus is right: knowledge-that limits your skill. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04248-6.

Dings:2023:SAE

[2615] Roy Dings, Christopher J. McCarroll, and Albert Newen. Situated authenticity in episodic memory. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04309-w.

Schwengerer:2023:VRP

[2616] Lukas Schwengerer. Virtual reality as a path to self-knowledge. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04317-w.

Kilov:2023:BVB

[2617] Daniel Kilov. Brittle virtue or bust: a new challenge to virtue-as-skill theories. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04306-z.

Williams:2023:CPM

[2618] Daniel Williams. The case for partisan motivated reasoning. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04223-1.

Rudolph:2023:CMN

[2619] Rachel Etta Rudolph. Contested metalinguistic negotiation. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04288-y.

Kazemi:2023:RC

[2620] Alireza Kazemi. Reasoning and commitment. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-023-04308-x.

Smart:2023:PCC

[2621] Benjamin T. H. Smart. Publisher correction to: The core business of medicine: a defence of the best available intervention thesis. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04313-0. See [2498].

Intemann:2023:MME

[2622] Kristen Intemann and Inmaculada de Melo-Martín. On masks and masking: epistemic harms and science communication. Synthese, 202 (3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04322-z.

Baron:2023:MEV

[2623] Sam Baron, Brigitte C. Everett, Andrew J. Latham, Kristie Miller, Hannah Tierney, and Jordan Veng Thang Oh. Moving ego versus moving time: investigating the shared source of future-bias and near-bias. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04286-0.

Farshi:2023:KTR

[2624] Farhang Hadad Farshi. On knots and temporality: a relational view of time. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04321-0.

Corti:2023:ONT

[2625] Luca Corti. Organizational normativity and teleology: a critique. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04255-7.

Muhlebach:2023:NIA

[2626] Deborah Mühlebach. A non-ideal approach to slurs. Synthese, 202 (3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04315-y.

Indrzejczak:2023:DDH

[2627] Andrzej Indrzejczak and Michał Zawidzki. Definite descriptions and hybrid tense logic. Synthese, 202(3):??, September 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-023-04319-8.

Rossi:2023:CHA

[2628] Niccolò Rossi and Aybüke Özgün. Correction to: A hyperintensional approach to positive epistemic possibility. Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04312-1. See [2573].

Azzouni:2023:WPP

[2629] Jody Azzouni. What is PA + con(PA) about, and where? Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04297-x.

Andler:2023:WM

[2630] Matthew Andler. What is masculinity? Synthese, 202(3):??, September 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04296-y.

Wrenn:2023:BBA

[2631] Chase B. Wrenn. Blindspots and brightspots for alethic pluralism. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04320-1.

Gert:2023:NST

[2632] Joshua Gert. Neopragmatism as a solution to Twin Earth problems. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04328-7.

Grimmer:2023:PQM

[2633] Daniel Grimmer. The pragmatic QFT measurement problem and the need for a Heisenberg-like cut in QFT. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04301-4.

Eldridge-Smith:2023:IHL

[2634] Peter Eldridge-Smith. The import of hypodoxes for the Liar and Russell’s paradoxes. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04326-9.

Anderau:2023:FNE

[2635] Glenn Anderau. Fake news and epistemic flooding. Synthese, 202 (4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04336-7.

Begley:2023:KGY

[2636] Keith Begley. Katz got your tongue? The metaphysics of words. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04324-x.

Fisseni:2023:HFI

[2637] Bernhard Fisseni, Deniz Sarikaya, and Bernhard Schröder. How to frame innovation in mathematics. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04310-3.

Freiesleben:2023:BGT

[2638] Timo Freiesleben and Thomas Grote. Beyond generalization: a theory of robustness in machine learning. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04334-9.

Arana:2023:MH

[2639] Andrew Arana and Heather Burnett. Mathematical hygiene. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04254-8.

Dandelet:2023:ASC

[2640] Sophia Dandelet. Arbitrary switching and concern for truth. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04337-6.

Kostic:2023:MEL

[2641] Daniel Kostić and Willem Halffman. Mapping explanatory language in neuroscience. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04329-6.

Mackiewicz:2023:IPT

[2642] Bartosz Maćkiewicz, Katarzyna Kuś, and Witold M. Hensel. The influence of philosophical training on the evaluation of philosophical cases: a controlled longitudinal study. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04316-x.

Vivanco:2023:NP

[2643] Melisa Vivanco. Numbers as properties. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04330-z.

Shimamura:2023:NMI

[2644] Shuhei Shimamura and Tuomo Tiisala. Normativity of meaning: an inferentialist argument. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04262-8.

Blumson:2023:DDS

[2645] Ben Blumson. Dialetheism and distributed sorites. Synthese, 202 (4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04346-5.

Sims:2023:BOT

[2646] Ric Sims. Boundary objects, trading zones, and stigmergy: the social and the cognitive in science. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04340-x.

Taylor:2023:CPQ

[2647] Henry Taylor. Compound powerful qualities: properties as compounds of distinct powers and qualities. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04339-4.

Park:2023:EEB

[2648] Ilho Park. Evidence and the epistemic betterness. Synthese, 202 (4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04350-9.

Mooney:2023:OUP

[2649] Justin Mooney. Ordinary undetached parts. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/

s11229-023-04343-8. See correction [2718].

Jacobs:2023:MOT

[2650] Caspar Jacobs. Are models our tools not our masters? Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04352-7.

Coleman:2023:UTE

[2651] Sam Coleman and Barbara Gail Montero. Unconscious transformative experience. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04332-x.

Moon:2023:DDG

[2652] Stella S. Moon. Demarcating Descartes’s geometry with clarity and distinctness. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04356-3.

Chao:2023:TKL

[2653] Hsiang-Ke Chao. Three kinds of the Lotka–Volterra model transfer from biology to economics. Synthese, 202(4):??, October 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-023-04341-w.

Scarpati:2023:EAW

[2654] Maria Scarpati and Claudio Calosi. By(e) enduring? An answer to Wasserman. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04351-8.

Tricard:2023:SGI

[2655] Julien Tricard. Symmetries as grounds for induction: the case of the Ω^{−} baryon. Synthese, 202(4):??, October 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04344-7.

Caponetto:2023:HMM

[2656] Laura Caponetto, Luca San Mauro, and Giorgio Venturi. How to make (mathematical) assertions with directives. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-023-04360-7.

Hurshman:2023:AIP

[2657] Clint Hurshman. Artifacts and intervention: a persistence theory of artifact functions. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04347-4.

Asma:2023:IBU

[2658] Lieke Joske Franci Asma. Implicit bias as unintentional discrimination. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04364-3.

Kaspers:2023:TUD

[2659] Tom Kaspers. Truth and its uses: deflationism and alethic pluralism. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04362-5.

Eker:2023:RAT

[2660] Bahadir Eker. Realism about tense and atemporality. Synthese, 202 (5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04335-8.

Laas:2023:DTT

[2661] Oliver Laas. Deepfakes and trust in technology. Synthese, 202(5): ??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04363-4.

Ambrosj:2023:VFI

[2662] Jacopo Ambrosj, Hugh Desmond, and Kris Dierickx. The value-free ideal in codes of conduct for research integrity. Synthese, 202(5): ??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04377-y.

Elliott:2023:NDM

[2663] Kevin C. Elliott. Navigating dissent by managing value judgments: the case of Lyme disease. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04368-z.

Hilligardt:2023:PSD

[2664] Hannah Hilligardt. Partisan science and the democratic legitimacy ideal. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04370-5.

Herzog:2023:WCC

[2665] Michael H. Herzog, Adrien Doerig, and Christian Sachse. Why computational complexity may set impenetrable barriers for epistemic reductionism. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04366-1.

Almagro:2023:RPS

[2666] Manuel Almagro, David Bordonaba-Plou, and Neftalí Villanueva. Retraction in public settings. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04348-3.

Dung:2023:CCA

[2667] Leonard Dung. Current cases of AI misalignment and their implications for future risks. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04367-0.

Yarandi:2023:ZSP

[2668] Seyed Mohammad Yarandi. On the zetetic significance of peer disagreement. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04331-y.

Ratoff:2023:HSP

[2669] William Ratoff. How should predictive processors conceive of practical reason? Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04354-5.

SantAnna:2023:RCI

[2670] André Sant’Anna. Is remembering constructive imagining? Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04338-5.

Skrzypulec:2023:BSS

[2671] Błazej Skrzypulec. Bodily sense and structural content. Synthese, 202 (5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04371-4.

Chislenko:2023:VW

[2672] Eugene Chislenko. Virtues of willpower. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04369-y.

Lauer:2023:SOR

[2673] Richard Lauer and Kareem Khalifa. Should ordinary race talk be ontologically privileged? Moving social science into the philosophical mainstream. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04361-6.

Collins:2023:SCE

[2674] Harry Collins. Science as a counter to the erosion of truth in society. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04357-2.

Ross:2023:SMS

[2675] Don Ross. Scientific metaphysics and social science. Synthese, 202 (5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04358-1.

Arnaud:2023:FPP

[2676] Sarah Arnaud. First-person perspectives and scientific inquiry of autism: towards an integrative approach. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04384-z.

Hermida:2023:TES

[2677] Margarida Hermida. Thought experiments, sentience, and animalism. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04349-2.

Pritchard:2023:HCT

[2678] Duncan Pritchard. Hinge commitments and trust. Synthese, 202 (5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04378-x.

Hirvonen:2023:HTE

[2679] Ilpo Hirvonen. On Husserl’s Twin Earth. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04294-0.

Helton:2023:HCE

[2680] Grace Helton and Christopher Register. Hot–cold empathy gaps and the grounds of authenticity. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04311-2.

Todd:2023:AMI

[2681] Cain Todd. Affective memory, imagined emotion, and bodily imagery. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04372-3.

Pokropski:2023:FPC

[2682] Marek Pokropski and Piotr Suffczynski. First-person constraints on dynamic-mechanistic explanations in neuroscience: The case of migraine and epilepsy models. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04396-9.

Wawer:2023:PFP

[2683] Jacek Wawer. A problem with the fixed past fixed. Synthese, 202 (5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04345-6.

Cassinadri:2023:REC

[2684] Guido Cassinadri and Marco Fasoli. Rejecting the extended cognition moral narrative: a critique of two normative arguments for extended cognition. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04397-8. See correction

[2720].

deVignemont:2023:EP

[2685] Frederique de Vignemont. Expecting pain. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04394-x.

Darling:2023:SBP

[2686] Tom Darling. Synthesising boredom: a predictive processing approach. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04380-3.

Hirvela:2023:VRS

[2687] Jaakko Hirvelä. A virtue reliabilist solution to Moore’s Paradox. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04392-z.

Kietzmann:2023:WRR

[2688] Christian Kietzmann. Why reasons and reasoning don’t come apart. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04355-4.

Wuthrich:2023:CCC

[2689] Adrian Wüthrich. Characterizing a collaboration by its communication structure. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04376-z.

Thresher:2023:CSP

[2690] Ann C. Thresher. Compact spaces and privileged times; what the video game asteroids can teach us about the present. Synthese, 202 (5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04325-w.

Brauner:2023:TF

[2691] Torben Braüner. The true futures. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04386-x.

Simonelli:2023:HHI

[2692] Ryan Simonelli. How to be a hyper-inferentialist. Synthese, 202(5): ??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04382-1.

Birgul:2023:UPF

[2693] Osman Gazi Birgül. Unveiling the philosophical foundations: On Cantor’s transfinite infinites and the metaphorical accounts of infinity. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04379-w.

Meadows:2023:RTE

[2694] Toby Meadows. Risk and theoretical equivalence in mathematical foundations. Synthese, 202(5):??, November 2023. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04374-1.

Dawson:2023:MFA

[2695] Patrick Dawson. Meta-fictionalism about the non-present. Synthese, 202 (5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04390-1.

Dover:2023:II

[2696] Daniela Dover. Identity and influence. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04279-z.

Mussavifard:2023:MC

[2697] Nima Mussavifard. Metarepresenting in communication. Synthese, 202 (5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04420-y.

Sher:2023:CP

[2698] Gila Sher. Correspondence pluralism. Synthese, 202(5):??, November 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04318-9.

Aikin:2023:ACD

[2699] Scott F. Aikin. The academic at the crossroads: a dialectical assessment of Augustinian pragmatic anti-skepticism. Synthese, 202(6): ??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04388-9.

Freivogel:2023:DRE

[2700] Andreas Freivogel. Does reflective equilibrium help us converge? Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04375-0.

Karadimas:2023:EIE

[2701] Panagiotis Karadimas. The epistemic impossibility of economic calculation. Synthese, 202(6):??, December 2023. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04381-2.

Robertson:2023:URS

[2702] Ian Robertson. The unbearable rightness of seeing? Conceptualism, enactivism, and skilled engagement. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04385-y.

Matey:2023:MGY

[2703] Jennifer Matey. Meaningfulness and grief: you don’t know what you got till it’s gone. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04333-w.

Shea:2023:ORF

[2704] Nicholas Shea. Organized representations forming a computationally useful processing structure. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04373-2.

Lee:2023:PPA

[2705] Jonny Lee and Paco Calvo. The potential of plant action potentials. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04398-7.

Kingsbury:2023:NCS

[2706] Justine Kingsbury. New challenges to the selected effects account of biological function. Synthese, 202(6):??, December 2023. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-023-04404-y.

Fellowes:2023:IIE

[2707] Sam Fellowes. The importance of involving experts-by-experience with different psychiatric diagnoses when revising diagnostic criteria. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04399-6.

Stanton:2023:CWN

[2708] Kate Hazel Stanton. Composing words and non-words. Synthese, 202 (6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04233-z.

Lewis:2023:SSL

[2709] Peter J. Lewis and Don Fallis. Simulation and self-location. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04413-x.

Baratella:2023:PER

[2710] Riccardo Baratella. Processes and events as rigid embodiments. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04304-1.

Syrjanen:2023:NPP

[2711] Pekka Syrjänen. Novel prediction and the problem of low-quality accommodation. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04400-2.

Cato:2023:PRC

[2712] Susumu Cato. Potential rationality in collective decision-making. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04365-2.

Bhakthavatsalam:2023:RSV

[2713] Sindhuja Bhakthavatsalam. Revisiting stance voluntarism: in defense of an active stance pluralism. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04410-0.

Gamester:2023:TNM

[2714] Will Gamester. Truth as none and many. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04353-6.

Sneddon:2023:TA

[2715] Andrew Sneddon. A theory of advice. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04389-8.

Cousens:2023:TAQ

[2716] Chris Cousens. Tweet acts and quote-tweetable acts. Synthese, 202 (6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04395-w.

deBruin:2023:RPJ

[2717] Boudewijn de Bruin. Ranking philosophy journals: a meta-ranking and a new survey ranking. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04342-9.

Mooney:2023:COU

[2718] Justin Mooney. Correction to: Ordinary undetached parts. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04424-8. See [2649].

Skowron:2023:TPS

[2719] Bartłomiej Skowron, Tomasz Kakol, and Zbigniew Król. The topology of persons, and surviving to some degree. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-023-04406-w.

Cassinadri:2023:CRE

[2720] Guido Cassinadri and Marco Fasoli. Correction to: Rejecting the extended cognition moral narrative: a critique of two normative arguments for extended cognition. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04431-9. See

[2684].

Boardman:2023:EMW

[2721] Samuel Boardman and Tom Schoonen. Epistemologists of modality wanted. Synthese, 202(6):??, December 2023. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04416-8. See correction

[2739].

Graziani:2023:CST

[2722] Ernesto Graziani, Francesco Orilia, Elena Capitani, and Roberto Burro. Common-sense temporal ontology: an experimental study. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04323-y.

Dominguez:2023:UZT

[2723] David Domínguez. Unzipping the zetetic turn. Synthese, 202(6): ??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04407-9.

Sridharan:2023:POR

[2724] Vishnu Sridharan. Praise, objective rightness and extended action. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04414-w.

Czerkawski:2023:WBF

[2725] Maciej Czerkawski. Why being fragments. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04403-z.

Yaghmaie:2023:GSR

[2726] Aboutorab Yaghmaie. Grounding scientific representation. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04423-9.

Wegrecki:2023:OAC

[2727] Jakub Wegrecki. The operator argument and the case of timestamp semantics. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04411-z.

Gularte:2023:SBE

[2728] Louis Gularte. Suffering is bad: experiential understanding and the impossibility of intrinsically valuing suffering. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04405-x.

Yao:2023:EA

[2729] Vida Yao. Eros and anxiety. Synthese, 202(6):??, December 2023. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-023-04391-
0.

Taieb:2023:REC

[2730] Hamid Taieb. Reinach on the essence of colours. Synthese, 202(6): ??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04327-8.

Buijsman:2023:CSE

[2731] Stefan Buijsman. Causal scientific explanations from machine learning. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04429-3.

Mortini:2023:EPF

[2732] Dario Mortini. Etiological proper function and the safety condition. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04412-y.

Sass:2023:NRB

[2733] Reuben Sass. Normative realism and Brentanian accounts of fittingness. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04419-5.

Cepollaro:2023:SR

[2734] Bianca Cepollaro and Dan López de Sa. The successes of reclamation. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04393-y.

Knudstorp:2023:LTS

[2735] Søren Brinck Knudstorp. Logics of truthmaker semantics: comparison, compactness and decidability. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04401-1.

Hughes:2023:PWG

[2736] Thomas J. Hughes. Possible words: generativity, instantiation, and individuation. Synthese, 202(6):??, December 2023. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04428-4.

Eyre:2024:RMP

[2737] John Eyre. Resolution of the Miller–Popper paradox. Synthese, 203 (1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04425-7.

Kornmesser:2024:CCC

[2738] Stephan Kornmesser. Conceptual change and conceptual enrichment: a frame-based reconstruction of Austin’s theory of speech acts. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04417-7.

Boardman:2024:CEM

[2739] Samuel Boardman and Tom Schoonen. Correction to: Epistemologists of modality wanted. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04454-2. See [2721].

Mikalonyte:2024:MWM

[2740] Elze Sigute Mikalonyte. Musical works are mind-independent artifacts. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04402-0.

Gjesdal:2024:LPA

[2741] Adam Gjesdal. Liberalism, polarization, and the aggregation problem. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04441-7.

Fischer:2024:PS

[2742] Enno Fischer. The promise of supersymmetry. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04447-1.

Facchin:2024:NRU

[2743] Marco Facchin. Neural representations unobserved-or: a dilemma for the cognitive neuroscience revolution. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04418-6.

Mai:2024:RI

[2744] Jonathan Mai. Realizing impossibilities. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04422-w.

Schwenkler:2024:CC

[2745] John Schwenkler. The categories of causation. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04409-7.

Mayne:2024:CFP

[2746] Zachary J. Mayne. The conceptual foundation of the propensity interpretation of fitness. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04437-3.

Lohse:2024:OCP

[2747] Richard Lohse. Overlapping consensus in pluralist societies: simulating Rawlsian full reflective equilibrium. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04415-9.

Redei:2024:ETL

[2748] Miklós Rédei and Márton Gömöri. Entropic taming of the Look Elsewhere Effect. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04434-6.

Paulson:2024:RN

[2749] Spencer Paulson. Reflective naturalism. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04430-w.

Lauer:2024:HDT

[2750] Helen Lauer. How to do things with insecure extensions. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04387-w.

Buder-Grondahl:2024:ABW

[2751] Tommi Buder-Gröndahl. The ambiguity of BERTology: what do large language models represent? Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04435-5.

Boge:2024:FCP

[2752] Florian J. Boge. Functional concept proxies and the Actually Smart Hans Problem: What’s special about deep neural networks in science. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04440-8.

Dudley:2024:KCK

[2753] Rachel Dudley and Christopher Vogel. Knowledge, the concept know, and the word know: considerations from polysemy and pragmatics. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04439-1.

Kedar:2024:EPQ

[2754] Yael Kedar. Propter quid demonstrations: Roger Bacon on geometrical causes in natural philosophy. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04433-7.

Bochman:2024:CRA

[2755] Alexander Bochman. Causal reasoning from almost first principles. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04442-6.

Stamenkovic:2024:SVL

[2756] Philippe Stamenkovic. Straightening the ‘value-laden turn’: minimising the influence of extra-scientific values in science. Synthese, 203 (1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04446-2.

Bacharach:2024:TTT

[2757] Julian Bacharach. Tensed truth, temporal particularity, and the fixity of the past. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04470-2.

Devanesan:2024:PTC

[2758] Arjun Devanesan. Pregnancy, a test case for immunology. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04444-4.

Puczylowski:2024:WGS

[2759] Tomasz Puczyłowski. Why a Gricean-style defense of the vacuous truth of counterpossibles won’t work, but a defense based on heuristics just might. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04436-4.

dePrado-Gordillo:2024:BWI

[2760] Miguel Núñez de Prado-Gordillo. Broken wills and ill beliefs: Szaszianism, expressivism, and the doubly value-laden nature of mental disorder. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04427-5.

Atmanspacher:2024:PND

[2761] Harald Atmanspacher. Psychophysical neutrality and its descendants: a brief primer for dual-aspect monism. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04449-z.

Fischer:2024:DTD

[2762] Martin Fischer and Matteo Zicchetti. Definite totalities and determinate truth in conceptual structuralism. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04443-5.

Kincaid:2024:NAS

[2763] Harold Kincaid. A naturalist approach to social ontology. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04432-8.

Matteoli:2024:OAP

[2764] Kory Matteoli. The once and always possible. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04426-6.

Kalewold:2024:HAA

[2765] Kalewold Hailu Kalewold. The hybrid account of activities. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04452-4.

Grujicic:2024:DCN

[2766] Bojana Grujicić. Deep convolutional neural networks are not mechanistic explanations of object recognition. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04461-3.

Krifka:2024:PUM

[2767] Manfred Krifka. Performative updates and the modeling of speech acts. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04359-0.

Colclough:2024:CAT

[2768] Thomas M. Colclough. Components of arithmetic theory acceptance. Synthese, 203(1):??, January 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04465-z.

Conner:2024:RET

[2769] William Conner. Radical epistemology, theory choice, and the priority of the epistemic. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04448-0.

Klemick:2024:PJD

[2770] Griffin Klemick. Perceptual justification and the demands of effective agency. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04453-3.

Costa:2024:AUS

[2771] Damiano Costa and Alessandro Giordani. Aristotelian universals, strong immanence, and construction. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04421-x.

Glasman:2024:ICS

[2772] Thomas Glasman. The iterative conception of set does not justify ZFC. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04408-8.

Berg:2024:WWR

[2773] Ásgeir Berg. Was Wittgenstein a radical conventionalist? Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04457-z.

Kadic:2024:MP

[2774] Nino Kadić. Monadic panpsychism. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04464-0.

Rolin:2024:EET

[2775] Kristina H. Rolin. Earning epistemic trustworthiness: an impact assessment model. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04472-0.

Douglas:2024:SCS

[2776] Heather Douglas and T. Y. Branch. The social contract for science and the value-free ideal. Synthese, 203(2):??, February 2024. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-023-04477-9.

Fernandez:2024:MTD

[2777] Jordi Fernández. Memory in two dimensions. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-023-04462-2.

Intemann:2024:VTP

[2778] Kristen Intemann. Value transparency and promoting warranted trust in science communication. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04471-1.

White:2024:PAA

[2779] Dylan J. White. Paying attention to attention: psychological realism and the attention economy. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04460-4.

Saad:2024:SDL

[2780] Bradford Saad. Should dualists locate the physical basis of experience in the head? Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04456-0.

Dellsen:2024:PRE

[2781] Finnur Dellsén. Probabilifying reflective equilibrium. Synthese, 203 (2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04466-y.

Freeborn:2024:RFA

[2782] David Peter Wallis Freeborn. Rational factionalization for agents with probabilistically related beliefs. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04491-5.

Shiller:2024:FID

[2783] Derek Shiller. Functionalism, integrity, and digital consciousness. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04473-z.

Yu:2024:DGI

[2784] Haomiao Yu and Stefan Petkov. Don’t get it wrong! On understanding and its negative phenomena. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/

s11229-023-04445-3. See correction [2831].

Sorensen:2024:WEM

[2785] Henrik Kragh Sørensen, Sophie Kjeldbjerg Mathiasen, and Mikkel Willum Johansen. What is an experiment in mathematical practice? New evidence from mining the Mathematical Reviews. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04475-x.

Rescorla:2024:BDC

[2786] Michael Rescorla. Bayesian defeat of certainties. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04383-0.

Schmid:2024:BPP

[2787] Joseph C. Schmid. Benardete paradoxes, patchwork principles, and the infinite past. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04481-z.

vanderBurg:2024:PRE

[2788] Wibren van der Burg. Pragmatist reflective equilibrium. Synthese, 203 (2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04455-1.

Murday:2024:IRC

[2789] Brendan Murday. Infinitism: rival or common ground in answering the epistemic regress? Synthese, 203(2):??, February 2024. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-023-04476-w.

Dopico:2024:DIT

[2790] Pablo Dopico. A defense of Isaacson’s thesis, or how to make sense of the boundaries of finite mathematics. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04488-0.

Fuentes:2024:CSH

[2791] Jorge Ignacio Fuentes. Computational systems as higher-order mechanisms. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04482-y.

Saborido:2024:DSP

[2792] Cristian Saborido and Jesús Zamora-Bonilla. Diseases as social problems. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04468-w.

Schafer:2024:HIS

[2793] Leon-Philip Schäfer. How (not) to integrate scientific and moral realism. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04467-x.

Martin:2024:REL

[2794] Ben Martin. Reflective equilibrium in logic. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04480-0.

Palazzolo:2024:CAR

[2795] Giulia Palazzolo. A case for animal reference: beyond functional reference and meaning attribution. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-023-04469-9.

Metzen:2024:OSV

[2796] Hanna Metzen. Objectivity, shared values, and trust. Synthese, 203 (2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04493-3.

Greene:2024:IDT

[2797] Preston Greene, Andrew J. Latham, Kristie Miller, and Michael Nielsen. The implicit decision theory of non-philosophers. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/

s11229-023-04478-8. See correction [2828].

Zoh:2024:HEB

[2798] Yoonseo Zoh, L. A. Paul, and M. J. Crockett. How the evaluability bias shapes transformative decisions. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04474-y.

Muller:2024:BSN

[2799] Basil Müller. Bad social norms rather than bad believers: examining the role of social norms in bad beliefs. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04483-5.

Garcia-Godinez:2024:DAL

[2800] Miguel Garcia-Godinez. A deflationary approach to legal ontology. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04487-1.

Domoslawski:2024:OSP

[2801] Aleksander Domosławski. The omniscient speaker puzzle. Synthese, 203 (2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04450-6.

Lorenzetti:2024:TFF

[2802] Lorenzo Lorenzetti. Two forms of functional reductionism in physics. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04507-0.

Hessbruggen-Walter:2024:ICC

[2803] Stefan Heßbrüggen-Walter. Interdisciplinarity in the 17th century? A cooccurrence analysis of early modern German dissertation titles. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04494-2.

Molinari:2024:ACA

[2804] Giacomo Molinari. An accuracy characterisation of approximate coherence. Synthese, 203(2):??, February 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04490-6.

Freitag:2024:LSD

[2805] Wolfgang Freitag. Lying and self-defeating prophecies. Synthese, 203 (3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04484-4.

Klinedinst:2024:CIT

[2806] Nathan Klinedinst. Coordinated if ’s and theories of conditionals. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-023-04458-y.

Lockhart:2024:LFR

[2807] Jennifer Ryan Lockhart and Micah Lott. Are life forms real? Aristotelian naturalism and biological science. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-023-04459-x.

Raineri:2024:PCE

[2808] Carlo Raineri. Phenomenal character and the epistemic role of perception. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-023-04479-7.

Kortabarria:2024:SEG

[2809] Markel Kortabarria and Joaquim Giannotti. Scientific explanation as a guide to ground. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04492-4.

Springle:2024:PPR

[2810] Alison A. Springle and Alessandra Buccella. Practical perceptual representations: a contemporary defense of an old idea. Synthese, 203 (3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-023-04463-1.

Orilia:2024:ESO

[2811] Francesco Orilia and Michele Paolini Paoletti. An experimental study on the ontology of relations. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04485-3.

Pharr:2024:DZS

[2812] Quentin Pharr. The distinctly zetetic significance of disagreement. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04496-0.

Stovall:2024:TMP

[2813] Preston Stovall. The teleological modal profile and subjunctive background of organic generation and growth. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/

s11229-023-04438-2. See correction [3006].

Pope:2024:ASD

[2814] Sara Pope. Authenticity as self-discovery and interpretation of value. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04489-z.

Beisbart:2024:TDC

[2815] Claus Beisbart and Georg Brun. Is there a defensible conception of reflective equilibrium? Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04495-1.

Gaszczyk:2024:LEE

[2816] Grzegorz Gaszczyk and Aleksandra Krogulska. Lying by explaining: an experimental study. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-023-04451-5. See correction

[2854].

Amico-Korby:2024:EE

[2817] Dallas Amico-Korby, Maralee Harrell, and David Danks. Environmental epistemology. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04504-3.

Zhao:2024:WG

[2818] Yingying Zhao and Zhiqiang Hu. What is gullibility? Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04497-z.

McLarty:2024:SDE

[2819] Colin McLarty. Structuralism in differential equations. Synthese, 203 (3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04522-1.

Simpson:2024:NWG

[2820] James Simpson. A note on Williamson’s Gettier cases in epistemic logic. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04506-1.

Asay:2024:IT

[2821] Jamin Asay. Irreplaceable truth. Synthese, 203(3):??, March 2024. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-024-04498-
y.

Barkasi:2024:CSR

[2822] Michael Barkasi. Consumer-side reference through promiscuous memory traces. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04509-y.

Pawlowski:2024:RF

[2823] Pawel Pawlowski and Karim Zahidi. Rigor and formalization. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04521-2.

Baumgaertner:2024:PRS

[2824] Bert Baumgaertner and Charles Lassiter. Precedent and rest stop convergence in reflective equilibrium. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04499-x.

Kleiner:2024:WMS

[2825] Johannes Kleiner and Tim Ludwig. What is a mathematical structure of conscious experience? Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04503-4.

Openshaw:2024:RRT

[2826] James Openshaw and Kourken Michaelian. Reference in remembering: towards a simulationist account. Synthese, 203(3):??, March 2024. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-024-04508-
z.

Rellihan:2024:FIH

[2827] Matthew Rellihan. Functionalism, interventionism, and higher-order causation. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04500-7.

Greene:2024:CID

[2828] Preston Greene, Andrew J. Latham, Kristie Miller, and Michael Nielsen. Correction to: The implicit decision theory of non-philosophers. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04538-7. See [2797].

Kappes:2024:BTL

[2829] Yannic Kappes. Bolzano’s Tortoise and a loophole for Achilles. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04515-0.

Subotic:2024:EEE

[2830] Vanja Subotić. Exploring, expounding & ersatzing: a three-level account of deep learning models in cognitive neuroscience. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04514-1.

Yu:2024:CDG

[2831] Haomiao Yu and Stefan Petkov. Correction to: Don’t get it wrong! On understanding and its negative phenomena. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04510-5. See [2784].

Coffman:2024:CAO

[2832] E. J. Coffman. The consequence argument and ordinary human agency. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04517-y.

vanRemmen:2024:IRN

[2833] Maximilian van Remmen. Inquiry, reasoning and the normativity of logic. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04533-y.

Berger:2024:HHH

[2834] Jacob Berger and Myrto Mylopoulos. HOTT and heavy: higher-order thought theory and the theory-heavy approach to animal consciousness. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04529-8.

Kersten:2024:IAM

[2835] Luke Kersten. An idealised account of mechanistic computation. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04526-x.

Briciu:2024:ICA

[2836] Adrian Briciu. Indexicals and communicative affordances. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04542-x.

Luc:2024:HAN

[2837] Joanna Luc. The Hole Argument without the notion of isomorphism. Synthese, 203(3):??, March 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04501-6.

Weber:2024:DGE

[2838] Erik Weber, Kristian González Barman, and Thijs De Coninck. Distinctively generic explanations of physical facts. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04486-2.

Rinner:2024:AAN

[2839] Stefan Rinner. Attitude ascriptions: a new old problem for Russell’s theory of descriptions. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04519-w.

Guarino:2024:PVE

[2840] Nicola Guarino and Giancarlo Guizzardi. Processes as variable embodiments. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04505-2.

ElSkaf:2024:SRS

[2841] Rawad El Skaf, Laura Felline, Patricia Palacios, and Giovanni Valente. Surrogative reasoning in the sciences. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04518-x.

Jago:2024:CWB

[2842] Mark Jago. Correction to: On ways of being true. Synthese, 203(4): ??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-024-04548-5. See [2547].

Tobia:2024:LCL

[2843] Kevin Tobia. Legal concepts and legal expertise. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/

s11229-024-04512-3. See correction [2882].

Schindler:2024:NSU

[2844] Samuel Schindler. Normal science: not uncritical or dogmatic. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04527-w.

Garancini:2024:EDR

[2845] Daniele Bruno Garancini. Experiment-driven rationalism. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04513-2.

Gausselin:2024:CPU

[2846] Kevin Gausselin. Conciliationism and the Peer-undermining Problem. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04511-4.

Boucher:2024:PTS

[2847] Sandy C. Boucher and Curtis Forbes. The pragmatic turn in the scientific realism debate. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04528-9.

Orlando:2024:TAP

[2848] Eleonora Orlando and Ramiro Caso. Are thick aesthetic predicates assessment-sensitive? Synthese, 203(4):??, April 2024. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-024-04520-3.

Munro:2024:CTE

[2849] Daniel Munro. Conspiracy theories, epistemic self-identity, and epistemic territory. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04541-y.

Nguyen:2024:CCM

[2850] James Nguyen. Confidence in Covid-19 models. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04530-1.

Dusi:2024:REM

[2851] Giovanni Dusi. Reliabilist epistemology meets bounded rationality. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04525-y.

Recanati:2024:MBM

[2852] François Recanati. Memory-based modes of presentation. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04531-0.

Berto:2024:HO

[2853] Francesco Berto. Hyperintensionality and overfitting. Synthese, 203 (4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-024-04556-5.

Gaszczyk:2024:CLE

[2854] Grzegorz Gaszczyk and Aleksandra Krogulska. Correction to: Lying by explaining: an experimental study. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04570-7. See [2816].

Gadsby:2024:MPM

[2855] Stephen Gadsby. Motivational pessimism and motivated cognition. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04546-7. See correction [2931].

Kind:2024:IP

[2856] Amy Kind. The impoverishment problem. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04551-w.

Krstic:2024:BFL

[2857] Vladimir Krstić. Bald-faced lying to institutions: deception or manipulation. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04547-6.

Cordes:2024:PFE

[2858] Manuel Cordes. The playing field of empirical facts: on the interrelations between moral and empirical beliefs in reflective equilibrium. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04553-8.

Stotts:2024:MMB

[2859] Megan Henricks Stotts. Moving from the mental to the behavioral in the metaphysics of social institutions. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04532-z.

Grodniewicz:2024:BRP

[2860] J. P. Grodniewicz. Belief revision in psychotherapy. Synthese, 203 (4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-024-04523-0.

Ruyant:2024:LP

[2861] Quentin Ruyant and Alexandre Guay. Lagrangian possibilities. Synthese, 203(4):??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04550-x.

Bradley:2024:HLY

[2862] Darren Bradley. How to lose your memory without losing your money: shifty epistemology and Dutch strategies. Synthese, 203(4): ??, April 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-024-04516-z.

Trinh:2024:LPT

[2863] Tue Trinh. Logicality and the picture theory of language. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04549-4.

Overgaard:2024:POD

[2864] Søren Overgaard. Perceptual occlusion and the differentiation condition. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04574-3.

Snyder:2024:RNA

[2865] Eric Snyder, Richard Samuels, and Stewart Shapiro. Restricted nominalism about number and its problems. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04564-5.

Miller:2024:MP

[2866] JTM Miller. The metaphysics of puns. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-024-04583-
2.

Reinikainen:2024:DPC

[2867] Jaakko Tapio Reinikainen. Defending the pure causal-historical theory of reference fixing for natural kind terms. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04568-1.

Keven:2024:CEM

[2868] Nazim Keven. Can episodic memory deter cheating and promote altruism? Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04560-9.

Milevski:2024:HTM

[2869] Voin Milevski. The Humean theory of motivation: much ado about nothing? Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04575-2.

Colo:2024:TJU

[2870] Philippe Colo. Testimonial justification under epistemic conflict of interest. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04585-0.

Pan:2024:PEL

[2871] Yixin Pan and Meiyun Guo. Probabilistic epistemic logic based on neighborhood semantics. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04557-4.

Orlandi:2024:KAA

[2872] Martina Orlandi. ‘I knew all along’: making sense of post-self-deception judgments. Synthese, 203(5):??, May 2024. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-024-04563-6.

Dewey:2024:ARM

[2873] Aliya R. Dewey. Anatomy’s role in mechanistic explanations of organism behaviour. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04534-x.

Facchin:2024:EAC

[2874] Marco Facchin and Giulia Leonetti. Extended animal cognition. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04579-y.

Young:2024:GKH

[2875] Garry Young. On the gradability of knowledge how, and its relationship to motor representations and ability. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04543-w.

Inoue:2024:REP

[2876] Akira Inoue, Kazumi Shimizu, Daisuke Udagawa, and Yoshiki Wakamatsu. Reflective equilibrium in practice and model selection: a methodological proposal from a survey experiment on the theories of distributive justice. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04544-9.

Jalloh:2024:BNI

[2877] Mahmoud Jalloh. Bridgman and the normative independence of science: an individual physicist in the shadow of the bomb. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04567-2.

Davis:2024:RRU

[2878] Isaac Davis. Rational representations of uncertainty: a pluralistic approach to bounded rationality. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04586-z.

Villegas:2024:VPD

[2879] Cristina Villegas. Variational propensities: development and ultimate causes. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04576-1.

Spiegel:2024:CVG

[2880] Thomas J. Spiegel. Can video games be philosophical? Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04580-5.

Bocchi:2024:MBC

[2881] Federica Bocchi. Metrics in biodiversity conservation and the value-free ideal. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04561-8.

Tobia:2024:CLC

[2882] Kevin Tobia. Correction: Legal concepts and legal expertise. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04598-9. See [2843].

Elabbar:2024:EMS

[2883] Ahmad Elabbar. Expertise, moral subversion, and climate deregulation. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04593-0.

Dixon:2024:NHC

[2884] Jonathan Dixon. No hope for conciliationism. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04590-3.

Friedman:2024:MMB

[2885] Michael Friedman. On metaphors of mathematics: Between Blumenberg’s nonconceptuality and Grothendieck’s waves. Synthese, 203(5): ??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04558-3.

Hoerl:2024:STT

[2886] Christoph Hoerl. Singular thought without temporal representation? Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04562-7.

Robins:2024:ITC

[2887] Sarah K. Robins and Felipe De Brigard. Introduction to the topical collection ‘locating representations in the brain: interdisciplinary perspectives’. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04565-4.

Werner:2024:EIP

[2888] Konrad Werner. Enacted institutions, participatory sense-making and social norms. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04569-0.

Carranza-Pinedo:2024:RCA

[2889] Víctor Carranza-Pinedo. Rethinking core affect: the role of dominance in animal behaviour and welfare research. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04591-2.

Murphy:2024:NLS

[2890] Elliot Murphy, Emma Holmes, and Karl Friston. Natural language syntax complies with the free-energy principle. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04566-3.

Forbes:2024:MLD

[2891] Owen Forbes. Metaphysical laws and the directionality of grounding. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04552-9.

Peet:2024:PPD

[2892] Andrew Peet. The puzzle of plausible deniability. Synthese, 203(5): ??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04600-4.

Moosavi:2024:FAA

[2893] Parisa Moosavi. The function argument for ascribing interests. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04592-1.

Esanu:2024:SFC

[2894] Andreea Esanu. Scrutinizing the foundations: could large language models be solipsistic? Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04589-w.

Watkins:2024:APD

[2895] Aja Watkins. The adequacy of purposes for data: a paleoecological case study. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04588-x.

Desmond:2024:GSC

[2896] Hugh Desmond. Gatekeeping should be conserved in the open science era. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04559-2.

Bielecka:2024:RRW

[2897] Krystyna Bielecka and Marcin Miłkowski. Representationalism and rationality: why mental representation is real. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04540-z.

Varga:2024:WMH

[2898] Somogy Varga and Andrew J. Latham. What is mental health and disorder? Philosophical implications from lay judgments. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04555-6.

Navarro:2024:EOP

[2899] Jesús Navarro. Epistemic ownership and the practical/epistemic parallelism. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04578-z.

Sikorski:2024:VBR

[2900] Michał Sikorski. Values, bias and replicability. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04573-4.

Martinis:2024:PJO

[2901] Valentina Martinis. Perceptual justification and objectual attitudes. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04597-w.

Lyons:2024:FHJ

[2902] Jack C. Lyons. Factivity, hallucination, and justification. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04587-y.

Kourany:2024:TIS

[2903] Janet A. Kourany. The two ideals shaping the content of modern science. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04554-7.

Guillon:2024:IPF

[2904] Jean-Baptiste Guillon. If presentism is false, then I don’t exist. On common-sense presentism. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04545-8.

Marvan:2024:DIA

[2905] Tomás Marvan and Sam Coleman. Defending internalism about unconscious phenomenal character. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04582-3.

Schmidt:2024:DMR

[2906] Michael W. Schmidt. Defining the method of reflective equilibrium. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04571-6.

deWeerd:2024:CBT

[2907] C. R. de Weerd. A credence-based theory-heavy approach to non-human consciousness. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04539-6.

Melchior:2024:SLP

[2908] Guido Melchior. The sensitivity of legal proof. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04609-9.

Rowbottom:2024:DNM

[2909] Darrell P. Rowbottom, William Peden, and André Curtis-Trudel. Does the no miracles argument apply to AI? Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04524-z.

Marques:2024:HSE

[2910] Teresa Marques. How slurs enact norms, and how to retract them. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04595-y.

Bowden:2024:IVS

[2911] Chelsea Bowden. Inquiry, value, and some peculiarities of the Pyrrhonist’s psychology. Synthese, 203(5):??, May 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04601-3.

Nawar:2024:TER

[2912] Tamer Nawar. Temporalism and eternalism reconsidered: perceptual experience, memory, and knowledge. Synthese, 203(6):??, June 2024. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-024-04535-
w.

Massin:2024:NDP

[2913] Olivier Massin. The nature and difficulty of physical efforts. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04572-5.

Kee:2024:HAC

[2914] Hayden Kee. “Humanity is another corporeity”: The evolution of human bodily appearance and sociality. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04581-4.

Gaultier:2024:EDA

[2915] Benoit Gaultier. An epistemic distinction among essences, its metaphysical ground, and the role of philosophy. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04619-7.

Carter:2024:SDS

[2916] J. Adam Carter and Robert Cowan. Safety and dream scepticism in Sosa’s epistemology. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04577-0.

Vollmer:2024:AZE

[2917] Michael Vollmer. Against zetetic encroachment. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04615-x.

Suikkanen:2024:ALA

[2918] Jussi Suikkanen. Anti-luminosity and anti-realism in metaethics. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04616-w.

Stroppa:2024:MC

[2919] Luca Stroppa. The monstrous conclusion. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04603-1.

Gasparri:2024:MCY

[2920] Luca Gasparri. Meaning change you can make. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04604-0.

Neijzen:2024:EVD

[2921] Mara Neijzen. The epistemic value of doombehaviour: beyond the prudential consequences of doomscrolling, doomchecking, and doomsurfing. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04602-2.

Dimitrijevic:2024:PCI

[2922] Dejan R. Dimitrijević. On the probabilistic character of irreducible mental causation. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04638-4.

Oest:2024:LMC

[2923] Sebastian Sander Oest. Life-mind continuity: untangling categorical, extensional, and systematic aspects. Synthese, 203(6):??, June 2024. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-024-04645-
5.

Mancini:2024:LAA

[2924] Filippo Mancini. Logical abductivism on abductive logic. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04626-8.

Sebastian:2024:CEP

[2925] Miguel Ángel Sebastián. Consciousness and Perspectival De Se content. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04611-1.

Sass:2024:NVE

[2926] Reuben Sass. Naturalness, veritism, and epistemic significance. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04629-5.

Elkin:2024:NUM

[2927] Lee Elkin. Normative uncertainty meets voting theory. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04644-6.

Montessori:2024:ORB

[2928] Auke Montessori. Overlap: on the relation between perceiving and believing. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04610-2.

Mezzadri:2024:RSP

[2929] Angelica Mezzadri. The role of scenarios in paradoxes. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04617-9.

Bagwala:2024:IIE

[2930] Abbas Bagwala. On informational injustice and epistemic exclusions. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04636-6.

Gadsby:2024:CMP

[2931] Stephen Gadsby. Correction to: Motivational pessimism and motivated cognition. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04635-7. See [2855].

Imaguire:2024:PEG

[2932] Guido Imaguire. The problem of extrinsic grounding. Synthese, 203 (6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-024-04614-y.

Wajzer:2024:GIP

[2933] Mateusz Wajzer. Gradual de-idealisation and progress in political science: a case study. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04627-7.

Heil:2024:HAW

[2934] Roman Heil. How to act on what you know. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04628-6.

Yon:2024:SLM

[2935] Daniel Yon and Cecilia Heyes. Social learning in models and minds. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04632-w.

Popa:2024:VPH

[2936] Elena Popa. Values in public health: an argument from trust. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04650-8.

Dranseika:2024:TST

[2937] Vilius Dranseika. Two ships of Theseus. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04623-x.

Duzi:2024:KWO

[2938] Marie Duzí and Martina Cíhalová. Knowing who occupies an office: purely contingent, necessary and impossible offices. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04596-x.

Torres:2024:TTT

[2939] Christopher A. Torres. Types of testimony and their reliability. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04641-9.

Tang:2024:LPS

[2940] Liping Tang. Linguistic politeness in social networks. Synthese, 203 (6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-024-04642-8.

Katsafanas:2024:WBA

[2941] Paul Katsafanas. What’s so bad about fanaticism? Synthese, 203 (6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-024-04647-3.

Gebharter:2024:MCI

[2942] Alexander Gebharter and Maria Sekatskaya. Mental causation, interventionism, and probabilistic supervenience. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04608-w.

Dohrn:2024:DIJ

[2943] Daniel Dohrn. Don’t imagine junk! Positive conceivability and modal illusion in mereology. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04658-0.

Gonzalez:2024:SCB

[2944] Stéphane Gonzalez and Nikolaos Pnevmatikos. A story of consistency: bridging the gap between Bentham and Rawls foundations. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04612-0.

Torsell:2024:LHL

[2945] Christian Torsell and Jeffrey A. Barrett. Learning how to learn by selftuning reinforcement. Synthese, 203(6):??, June 2024. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-024-04649-1.

Jiang:2024:LHO

[2946] Junli Jiang and Pavel Naumov. A logic of higher-order preferences. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04655-3.

Bos:2024:LCC

[2947] Gregor E. Bös. A logical challenge to correlationism: the Church–Fitch paradox in Husserl’s account of fulfilment, truth, and meaning. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04584-1.

Rivers:2024:OUT

[2948] Imogen Lucy Grace Rivers. Once upon a time in superspace: the diegetic ideal for the interpretation of physical theories. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04618-8.

Peters:2024:GNA

[2949] Uwe Peters and Ignacio Ojea Quintana. Are generics and negativity about social groups common on social media? A comparative analysis of Twitter (X) data. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04639-3.

Stradis:2024:PRP

[2950] Athamos Stradis. Present records of the Past Hypothesis. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04643-7.

Nimtz:2024:PSN

[2951] Christian Nimtz. The power of social norms: Why conceptual engineers should care about implementation. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04625-9.

Ke:2024:DAE

[2952] Xiaoyu Ke. Developing appropriate emotions. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04630-y.

Percy:2024:YRI

[2953] Chris Percy. Your red isn’t my red! Connectionist Structuralism and the puzzle of abstract objects. Synthese, 203(6):??, June 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04652-6.

Durr:2024:ICP

[2954] Patrick Dürr and James Read. An invitation to conventionalism: a philosophy for modern (space-)times. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04605-z.

Hambly:2024:INT

[2955] Jesse Hambly. The intrapersonal normative twin earth argument. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04651-7.

Goff:2024:FTE

[2956] Philip Goff. Is the fine-tuning evidence for a multiverse? Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04621-z.

Cox:2024:ARB

[2957] Ryan Cox. Actions, reasons, and becauses. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04656-2.

Karakas:2024:MV

[2958] Alexandra Karakas and Adam Tamas Tuboly. Materializing values. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04663-3.

Avni:2024:ICP

[2959] Ron Avni. Indistinguishability as a constraint on priors. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04599-8.

Weinstein:2024:SBE

[2960] Zachary Weinstein. Sounds are broad events. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04662-4.

Bromand:2024:WL

[2961] Joachim Bromand. Wittgenstein and the liar. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04620-0.

Finley:2024:TPE

[2962] Robby Finley. Towards a pragmatist epistemology for theory choice in logic. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04654-4.

Drobnak:2024:CCC

[2963] Matej Drobnák. Compositionality, communication, and commitments. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04594-z.

Beni:2024:CTJ

[2964] Majid D. Beni and Karl Friston. Carving teleology at its joints. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04660-6.

Goodman:2024:MFC

[2965] Rachel Goodman and Aidan Gray. Mental filing, continued. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04670-4.

Bonardi:2024:SSF

[2966] Paolo Bonardi. Salmón, Schiffer and Frege’s constraint. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04502-5.

Cavedon-Taylor:2024:DSI

[2967] Dan Cavedon-Taylor. Deepfakes: a survey and introduction to the topical collection. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04634-8.

Thorvaldsen:2024:RQN

[2968] Steinar Thorvaldsen, Peter Øhrstrøm, and Ola Hössjer. The representation, quantification, and nature of genetic information. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04613-z.

Araya:2024:EAR

[2969] J. M. Araya. Emotion against reason? Self-control conflict as selfmodelling rivalry. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04672-2.

Kirkpatrick:2024:GQ

[2970] James Ravi Kirkpatrick. Are generics quantificational? Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04653-5.

Winokur:2024:EMA

[2971] Benjamin Winokur. Extended mentality and ascriptive authority. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04637-5.

Cohen:2024:IG

[2972] Shlomit Wygoda Cohen. Indeterministic grounding. Synthese, 204(1): ??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04648-2.

Salmon:2024:SC

[2973] Nathan Salmón. Singular concepts. Synthese, 204(1):??, July 2024. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-024-04624-
w.

Lehrer:2024:IIK

[2974] Ehud Lehrer and Dov Samet. Interpersonal independence of knowledge and belief. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04607-x.

Grandjean:2024:ITR

[2975] Vincent Grandjean and Matteo Pascucci. Introduction temporal reasoning and tensed truths. Synthese, 204(1):??, July 2024. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-024-04682-0.

Adelstein:2024:ASI

[2976] Matthew Adelstein. Alternatives to the self-indication assumption are doomed. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04686-w.

Politi:2024:VFI

[2977] Vincenzo Politi. “The value-free ideal, the autonomy thesis, and cognitive diversity”. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04673-1.

Ferretti:2024:APF

[2978] Gabriele Ferretti and Silvano Zipoli Caiani. An all-purpose framework for affordances. Reconciling the behavioral and the neuroscientific stories. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04646-4.

Gupte:2024:DHN

[2979] Firdaus Gupte. In defense of Humean non-universal laws. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04678-w.

Tamborini:2024:ETB

[2980] Marco Tamborini. Exploring the transition: biology, technology, and epistemic activities. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04675-z.

Moehler:2024:SJC

[2981] Michael Moehler and John Thrasher. Strategic justice, conventions, and game theory: introduction to a Synthese topical collection. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04680-2.

Astola:2024:MV

[2982] Mandi Astola, Steven Bland, and Mark Alfano. Mandevillian vices. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04676-y.

Hung:2024:RAR

[2983] Jenny Hung. Reflexive awareness and reflexivity: an identity model of reflexive awareness with Korta and Perry’s reflexive-referential theory of content (RRT). Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04688-8.

Gersel:2024:MOA

[2984] Johan Gersel. McDowell’s overlooked argument for disjunctivism — realism, self-consciousness, and knowledge. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04695-9.

Jiang:2024:MMO

[2985] Yihan Jiang. The metaphysics of mechanisms: an ontic structural realist perspective. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04684-y.

Andreoletti:2024:REP

[2986] Mattia Andreoletti and Saúl Pérez-González. The reliability of evidential pluralism in drug regulation. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04677-x.

Benito-Monsalvo:2024:LSP

[2987] Carlos Benito-Monsalvo and José Martínez-Fernández. A localist solution to the problem of mixed inferences by juxtaposition. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04690-0.

Baston:2024:BPN

[2988] René Baston. Beyond prediction: a new paradigm for understanding suicide risk. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04689-7.

Thobani:2024:TWI

[2989] Imran Thobani. A triviality worry for the internal model principle. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04693-x.

Hatchwell:2024:IVI

[2990] Rivkah Hatchwell. Individual values and inductive risk: remotivating the Bayesian alternative. Synthese, 204(1):??, July 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04674-0.

Goldwasser:2024:ISM

[2991] Seth Goldwasser. Imagining as a skillful mental action. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04681-1.

Reuter:2024:SS

[2992] Kevin Reuter. Salient semantics. Synthese, 204(2):??, August 2024. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-024-04669-
x.

Pritchard:2024:BDH

[2993] Duncan Pritchard. Beliefs, delusions, hinge commitments. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04703-y.

DellUtri:2024:PAP

[2994] Massimo Dell’Utri. Plain alethic pluralism: the human face of truth. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04700-1.

Dagher:2024:PPM

[2995] Ibrahim Dagher. Powered properties, modal continuity, and the patchwork principle. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04696-8.

Spagnesi:2024:NIE

[2996] Lorenzo Spagnesi. Natures, ideas, and essentialism in Kant. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04659-z.

Fletcher:2024:RDS

[2997] Samuel C. Fletcher and David E. Taylor. The representation and determinable structure of quantum properties. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04631-x.

Brewer:2024:MEC

[2998] Bill Brewer. The metaphysics and epistemology of conscious perception. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04698-6.

Jiang:2024:LTB

[2999] Junli Jiang and Pavel Naumov. A logic of trust-based beliefs. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04671-3.

Lusk:2024:AGA

[3000] Greg Lusk and Kevin C. Elliott. Against global aims for science: values, epistemic priority, and a local aims approach. Synthese, 204 (2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04633-9.

White:2024:ASE

[3001] Ben White and Mark Miller. Ambient smart environments: affordances, allostasis, and wellbeing. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04679-9.

Glass:2024:CEW

[3002] David H. Glass and Jonah N. Schupbach. Conjunctive explanations: when are two explanations better than one? Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04683-z.

Alter:2024:PFM

[3003] Torin Alter. Physicalism and fundamental mentality. Synthese, 204 (2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04697-7.

Baz:2024:CCC

[3004] Avner Baz. Can contemporary cognitive science coherently accommodate itself? Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04704-x.

Koiv:2024:REA

[3005] Riin Kõiv. To reform or to eliminate an attractor? Synthese, 204 (2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04685-x.

Stovall:2024:CTM

[3006] Preston Stovall. Correction to: The teleological modal profile and subjunctive background of organic generation and growth. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04692-y. See [2813].

Kriegel:2024:EP

[3007] Uriah Kriegel. Event plenitude. Synthese, 204(2):??, August 2024. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-024-04640-
w.

Davis:2024:NSM

[3008] Cruz Austin Davis. Non-spatial matters on the possibility of non-spatial material objects. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04606-y.

Schipani:2024:JPT

[3009] Vanessa Schipani. Journalism and public trust in science. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04701-0.

Veigl:2024:MTJ

[3010] Sophie Juliane Veigl. More than just principles: revisiting epistemic systems. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04708-7.

Logins:2024:IR

[3011] Arturs Logins. Inquiry and reasons. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04687-9.

Kijania-Placek:2024:DRM

[3012] Katarzyna Kijania-Placek. Deferred reference, meaning transfer or coercion? Toward a new principle of accounting for systematic uses of proper names. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04667-z.

Wei:2024:WT

[3013] Xintong Wei. The weight of truth. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04709-6.

Deigan:2024:HCC

[3014] Michael Deigan. Having a concept has a cost. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04661-5.

Wenmackers:2024:RPB

[3015] Sylvia Wenmackers. Remote possibilities in branching time structures. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04622-y.

Lohr:2024:WDI

[3016] Guido Löhr. What does it mean for a duty to be directed in joint action? Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04713-w.

Athea:2024:PLB

[3017] Héloïse Athéa, Nicolas Heck, and Denis Forest. The private life of the brain: issues and promises in the neuroscientific study of internal states.

Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04717-6.

Sorensen:2024:TTC

[3018] Roy Sorensen. Time traveler confirms five minute hypothesis! Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04719-4.

Varga:2024:SUB

[3019] Somogy Varga. Scientific understanding in biomedical research. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04694-w.

Carter:2024:IAK

[3020] J. Adam Carter and Gloria Andrada. Intentional action, knowledge, and cognitive extension. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04691-z.

Skrzypulec:2024:SDO

[3021] Błazej Skrzypulec. Smellscapes and diachronic olfaction. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04715-8.

Prentner:2024:CTC

[3022] Robert Prentner. Category theory in consciousness science: going beyond the correlational project. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04718-5.

Garcia-Carpintero:2024:MBR

[3023] Manuel García-Carpintero. Memory-based reference and immunity to error through misidentification. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04664-2.

Grindrod:2024:LLM

[3024] Jumbly Grindrod. Large language models and linguistic intentionality. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04723-8.

Lane:2024:DDA

[3025] Benedict Lane. Deep disagreement across moral revolutions. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04710-z.

Ryan:2024:CUC

[3026] Liam D. Ryan. Credo in unam credentiam: religious beliefs are standard beliefs. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04711-y.

Spelda:2024:WHC

[3027] Petr Spelda and Vit Stritecky. Why and how to construct an epistemic justification of machine learning? Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04702-z.

Garcia-Valdecasas:2024:OBT

[3028] Miguel García-Valdecasas and Terrence W. Deacon. Origins of biological teleology: how constraints represent ends. Synthese, 204(2): ??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-024-04705-w.

Barker:2024:KHB

[3029] Beth Barker. Knowing how and being able. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04707-8.

Smith:2024:PVN

[3030] George E. Smith. Physical vs. numerical approximation in Isaac Newton’s Principia. Synthese, 204(2):??, August 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04665-1.

Murez:2024:ECT

[3031] Michael Murez and Brent Strickland. Event completion: a test case for theories of reference in memory. Synthese, 204(3):??, September 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04722-9.

Cavedon-Taylor:2024:MDI

[3032] Dan Cavedon-Taylor. Must depression be irrational? Synthese, 204 (3):??, September 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04714-9.

Adlam:2024:QFT

[3033] Emily Adlam. Quantum field theory and the limits of reductionism. Synthese, 204(3):??, September 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04734-5.

Zhan:2024:NAO

[3034] Xinyi Zhan. I am no abstract object: a novel challenge to mind uploading. Synthese, 204(3):??, September 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04744-3.

Vaynberg:2024:RDD

[3035] Eugene Vaynberg. Realism and the detection of dark matter. Synthese, 204(3):??, September 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04728-3.

Popa:2024:JPT

[3036] Elena Popa. Just pluralism: thinking about concepts of mental disorder in global context. Synthese, 204(3):??, September 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04724-7.

Legault:2024:BSA

[3037] Mylène Legault, Amandine Catala, and Pierre Poirier. Breaking the stigma around autism: moving away from neuronormativity using epistemic justice and 4E cognition. Synthese, 204(3):??, September 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04730-9.

Lopez-Silva:2024:MSD

[3038] Pablo López-Silva. Making sense of the doxastic approach to thought insertion. Synthese, 204(3):??, September 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04721-w.

Tersman:2024:SRE

[3039] Folke Tersman. Seeking a reflective equilibrium in the face of disagreement. Synthese, 204(3):??, September 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04729-2.

Westfall:2024:PEI

[3040] Mason Westfall. Polarization is epistemically innocuous. Synthese, 204 (3):??, September 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04739-0.

Genin:2024:CPE

[3041] Konstantin Genin, Thomas Grote, and Thomas Wolfers. Computational psychiatry and the evolving concept of a mental disorder. Synthese, 204 (3):??, September 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04741-6.

Diez:2024:NCA

[3042] José Díez and David Pineda. A non-causalist account of the explanatory autonomy in the psychological sciences. Synthese, 204(3): ??, September 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04657-1.

Franken:2024:PCW

[3043] Dirk Franken. Persistence conditions: what they are, and what it is to have them. Synthese, 204(3):??, September 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04755-0.

Brown:2024:VSA

[3044] Matthew J. Brown. For values in science: Assessing recent arguments for the ideal of value-free science. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04762-1.

Richardson:2024:SET

[3045] Louise Richardson. Shared emotion without togetherness: the case of shared grief. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04760-3.

Jonsson:2024:WWM

[3046] Martin L. Jönsson and Hubert Hågemark. What we mean as what we said or would have said. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04761-2.

Placek:2024:TSW

[3047] Tomasz Placek. On testimony in scenarios with Wigner and Friend. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04754-1. See correction [3170].

Hunt:2024:SSM

[3048] Josh Hunt. Staying on-shell: manifest properties and reformulations in particle physics. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04750-5.

Fulmer:2024:VC

[3049] Everett C. Fulmer. Is validity circular? Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04666-0.

Cerezo:2024:MEL

[3050] María Cerezo and Vanessa Triviño. 3D/4D metaphysical equivalence: lessons from the species debate for the metaphysics of change and persistence. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04747-0.

Declos:2024:OV

[3051] Alexandre Declos. The ontology of videogames. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04774-x.

Britten-Neish:2024:NPA

[3052] George Britten-Neish. Neuroidealism, perceptual acquaintance and the Kantian roots of predictive processing. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04778-7.

Lawry:2024:MBS

[3053] Jonathan Lawry. Multiple belief states in social learning: an evidence tokens model. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04770-1.

Ruetsche:2024:PAO

[3054] Laura Ruetsche. The perils of approximate ontology. Synthese, 204 (4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04777-8.

Jago:2024:BTS

[3055] Mark Jago. Belief in truthmaker semantics. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04748-z.

Ivani:2024:TPS

[3056] Silvia Ivani and Alfred Archer. Taking the public seriously: the role of respect in interactions between scientific experts and lay publics. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04772-z.

Viglione:2024:TIC

[3057] Federico Viglione. The traversal of the infinite: considering a beginning for an infinite past. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04735-4.

Mulder:2024:SCA

[3058] Ruward Mulder and James Read. Is spacetime curved? Assessing the underdetermination of general relativity and teleparallel gravity. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04773-y.

Harris:2024:HOM

[3059] Keith Raymond Harris. Higher-order misinformation. Synthese, 204 (4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04763-0.

Lane:2024:DTK

[3060] Elek Lane. Distinguishing two kinds of fictionalism: metaphor, autism, and the imagination. Synthese, 204(4):??, October 2024. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-024-04742-5.

Esfeld:2024:MMM

[3061] Michael Esfeld and Cristian López. Minimal metaphysics in moral and political philosophy. Synthese, 204(4):??, October 2024. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-024-04775-w.

Zuradzki:2024:ERC

[3062] Tomasz Zuradzki and Joanna Karolina Malinowska. Ethno-racial categorisations for biomedical studies: the fair selection of research participants and population stratification. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04769-8.

Hodroj:2024:RCM

[3063] Batoul Hodroj, Andrew J. Latham, Kristie Miller, Rasmus Pedersen, and Danqi Wang. The role of causal manipulability in the manifestation of time biases. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04740-7.

Iikawa:2024:TED

[3064] Haruka Iikawa and Go Sasaki. Temporal externalist descriptivism on natural kind terms: beyond the causal-historical analysis. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04765-y.

Moody:2024:MTM

[3065] Kristoffer Moody. The Manipulationist Threat to moral responsibility. Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04771-0.

Aviles-Meza:2024:ND

[3066] Edvard Aviles-Meza. The neurotic and the dogmatist. Synthese, 204 (4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04781-y. See correction [3108].

Wiegman:2024:CEV

[3067] Isaac Wiegman and Michael T. Dale. The co-evolution of virtue and desert: debunking intuitions about intrinsic value. Synthese, 204 (4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04751-4.

Akhlaghi:2024:HET

[3068] Farbod Akhlaghi. How are ethical theories explanatory? Synthese, 204(4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04784-9.

Klein:2024:HST

[3069] Peter D. Klein. How sceptics teach us to know. Synthese, 204 (4):??, October 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04727-4. See correction [3100].

SantosGenta:2024:FMJ

[3070] Benjamin Santos Genta. Formal models and justifications of democracy. Synthese, 204(5):??, November 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04783-w.

Restall:2024:QJR

[3071] Greg Restall. Questions, justification requests, inference, and definition. Synthese, 204(5):??, November 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04752-3.

Nelson:2024:SAN

[3072] Erik Nelson. A Sellarsian argument for nonlinguistic conceptual capabilities. Synthese, 204(5):??, November 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04785-8.

Belardinelli:2024:IKU

[3073] Gaia Belardinelli and Burkhard C. Schipper. Implicit knowledge in unawareness structures. Synthese, 204(5):??, November 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04706-9.

Lehtinen:2024:IFM

[3074] Aki Lehtinen. Imagination and fiction in modelling; an epistemic critique. Synthese, 204(5):??, November 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04780-z.

Semeijn:2024:MSM

[3075] Merel Semeijn. Matraversian skepticism and models of memory. Synthese, 204(5):??, November 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04776-9.

Bianchi:2024:USA

[3076] Claudia Bianchi. Unfinished speech acts. Synthese, 204(5):??, November 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04782-x.

Cova:2024:DDE

[3077] Florian Cova, Céline Schöpfer, and Morgane Myling Bezat. Delving into depth: an empirical investigation of the ordinary concepts of depth and profundity. Synthese, 204(5):??, November 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04786-7.

Gough:2024:MME

[3078] Joe Gough. ‘Mind’ and ‘mental’: extended, pluralistic, eliminated. Synthese, 204(5):??, November 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04788-5.

Vogrin:2024:RLN

[3079] Michael Vogrin and Jan Willem Koten. RUM leads to noise: the significance of finding the sources of variability between experimental runs. Synthese, 204(5):??, November 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04798-3.

Farina:2024:NNI

[3080] Mirko Farina and Andrea Lavazza. The ‘NeuroGate’: neuromorphic intelligence, extended mind, and neurorights. Synthese, 204(5): ??, November 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04795-6.

Palmer:2024:AFW

[3081] David Palmer and Ning Fan. Achievements, free will, and meaning in life. Synthese, 204(5):??, November 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04791-w.

Bussiere-Caraes:2024:NAS

[3082] Lwenn Bussière-Caraes, Luca Incurvati, Giorgio Sbardolini, and Julian J. Schloeder. Non-assertoric speech acts: Introduction to the topical collection. Synthese, 204(5):??, November 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04799-2.

Molto:2024:RAA

[3083] Daniel Molto. Revisiting an argument against identity. Synthese, 204 (5):??, November 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04807-5.

Cosgrove:2024:OOR

[3084] Joseph K. Cosgrove. Order, organization, and randomness: on the mathematical formulation of life. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04806-6.

Elohim:2024:CGC

[3085] David Elohim. Correction to: Grounding, conceivability, and the mindbody problem. Synthese, 204(6):??, December 2024. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04712-x. See [5].

Andonovski:2024:EMF

[3086] Nikola Andonovski. Engrams as mental files. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04801-x.

Wolf:2024:PSD

[3087] William J. Wolf and Patrick M. Duerr. Promising stabs in the Dark: theory virtues and pursuit-worthiness in the Dark Energy problem. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04796-5.

Perrin:2024:RR

[3088] Denis Perrin. Re-remembering. Synthese, 204(6):??, December 2024. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-024-04794-
7.

Sullivan-Bissett:2024:MDM

[3089] Ema Sullivan-Bissett. Monothematic delusions are misfunctioning beliefs. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04803-9.

Karagiannopoulos:2024:ACW

[3090] Yorgos Karagiannopoulos and Alexios Stamatiadis-Bréhier. Against cross-world anchoring. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04793-8.

Roffe:2024:OTF

[3091] Ariel Jonathan Roffé, Karina Alleva, Santiago Ginnobili, and Sergio Barberis. The open texture of functions: a framework for analyzing functional concepts in molecular biology. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04821-7.

Antilici:2024:KFA

[3092] Francesco Antilici. Knowledge first: the argument from development. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04809-3.

Sares:2024:RDP

[3093] James Sares. The reasonableness of doubt: phenomenology and scientific realism. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04819-1.

Fabregas-Tejeda:2024:DNC

[3094] Alejandro Fábregas-Tejeda and Grant Ramsey. Driftability and niche construction. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04815-5.

Arslan:2024:CCS

[3095] Aran Arslan and Frank Zenker. Cohen’s convention, the seriousness of errors, and the body of knowledge in behavioral science. Synthese, 204 (6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04753-2.

Kallestrup:2024:DEK

[3096] Jesper Kallestrup. Digitally extended knowledge. Synthese, 204(6): ??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04818-2.

Beaute:2024:HEC

[3097] Robin Beauté. How is empirical content possible? Sellars after McDowell. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04790-x.

Wu:2024:ERC

[3098] Qiantong Wu. An enactivist response to the challenge from dreams. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04802-w.

Zhao:2024:CTC

[3099] Shimin Zhao. A causal theory of confirmation for Bayesians. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04779-6.

Klein:2024:CHS

[3100] Peter D. Klein. Correction to: How sceptics teach us to know. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04813-7. See [3069].

Kovacs:2024:PTM

[3101] David Mark Kovacs. The problem of too many mental tokens reconsidered. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04832-4.

Markunas:2024:AKL

[3102] Michael Markunas. Acquaintance, knowledge, and luck. Synthese, 204 (6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04822-6.

Bnefsi:2024:WIL

[3103] Sayid R. Bnefsi. Willful ignorance in law and epistemology. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04847-x.

Pedersen:2024:RPA

[3104] Rasmus Pedersen. Retrodictive and predictive attentional modulation in temporal binding. Synthese, 204(6):??, December 2024. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-024-04787-6.

Choi:2024:PP

[3105] Hong Hui Choi. Preregistration and predictivism. Synthese, 204(6): ??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04827-1.

Willems:2024:SIP

[3106] Ed Willems. Solving the inclusion problem: gender without representationalism. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04831-5.

Ylikoski:2024:GLL

[3107] Petri Ylikoski. Getting lost with levels: the sociological micro-macro problem. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04841-3.

Aviles-Meza:2024:CND

[3108] Edvard Aviles-Meza. Correction to: The neurotic and the dogmatist. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04804-8. See [3066].

Plourde:2024:TCT

[3109] Jimmy Plourde. Truth and correspondence in the Tractatus. Synthese, 204(6):??, December 2024. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04829-z.

Ferretti:2025:PID

[3110] Gabriele Ferretti. Philosophical implications of derealization disorder. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04808-4.

Mikalonyte:2025:FCA

[3111] Elze Sigute Mikalonyte and Markus Kneer. The folk concept of art. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04812-8.

Antoniou:2025:TMU

[3112] Antonis Antoniou and Karim P. Y. Thébault. Theories without models: uncontrolled idealizations in particle physics. Synthese, 205(1): ??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-024-04826-2.

Kinkaid:2025:HMM

[3113] James Kinkaid. Husserl and the marks of the mental. Synthese, 205 (1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04862-y.

Piccinini:2025:KKH

[3114] Gualtiero Piccinini and Stephen Hetherington. Knowing that as knowing how: a neurocognitive practicalism. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04843-1.

Assadian:2025:IMO

[3115] Bahram Assadian and Robert Fraser. The individuation of mathematical objects. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04814-6.

Evans:2025:MPA

[3116] Amanda Evans. Medical paternalism, anorexia nervosa, and the problem of pathological values. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04870-y.

Vollmer:2025:PDS

[3117] Michael Vollmer. The puzzle of defeated suspension. Synthese, 205 (1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04874-8.

Havlik:2025:MUL

[3118] Vladimír Havlík. Meaning and understanding in large language models. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04878-4.

Beck:2025:PDH

[3119] Pieter T. L. Beck. Patience, diligence, and humility: epistemic virtues and chemistry in the eighteenth century Dutch Republic. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04845-z.

Rivat:2025:OSA

[3120] Sébastien Rivat. Open systems across scales. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04817-3.

Mills:2025:LWD

[3121] Philip Mills. Lyotard’s Wittgenstein: from deep disagreement to Differend. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04872-w.

Kubiak:2025:PRF

[3122] Adam P. Kubiak. Perspectival realism and frequentist statistics: the case of Jerzy Neyman’s methodology and philosophy. Synthese, 205 (1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04842-2.

Richmond:2025:WTN

[3123] Andrew Richmond. What is a theory of neural representation for? Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04816-4.

Carrara:2025:NSL

[3124] Massimiliano Carrara and Andrea Strollo. On the non-substantiality of logic: a case study. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04846-y.

Baetu:2025:EPS

[3125] Tudor M. Baetu. The epistemic problem of singular causation. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04876-6.

Jackson:2025:FWP

[3126] Elizabeth Grace Jackson. Faith is weakly positive. Synthese, 205 (1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04811-9.

Klemick:2025:STR

[3127] Griffin Klemick. Sellars’s two responses to skepticism. Synthese, 205 (1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04833-3.

Schirn:2025:FPM

[3128] Matthias Schirn. Frege’s Platonism and mathematical creation: some new perspectives. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04764-z.

Mazzullo:2025:MSN

[3129] Giorgio Mazzullo. The Monist strategy: Naïve realism and the master argument from hallucination. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04834-2.

Stokke:2025:QL

[3130] Andreas Stokke and Derek Ball. Quotative be like. Synthese, 205 (1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04800-y.

Dimitrov:2025:EQT

[3131] Martin Dimitrov. The evaluative qualia theory of pleasure and displeasure. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04824-4.

Freeborn:2025:ETB

[3132] David Peter Wallis Freeborn. Effective theory building and manifold learning. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04844-0.

Faissner:2025:EOC

[3133] Mirjam Faissner, Esther Braun, and Christin Hempeler. Epistemic oppression and the concept of coercion in psychiatry. Synthese, 205 (1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04853-z.

Lee:2025:TCP

[3134] James Lee. Two can play at this game: a dual-individuation account of games. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04820-8.

Geiger:2025:NFR

[3135] Ido Geiger. Necessity for finite rational minds — Kant on empirical nomic necessity and the conceptual purposiveness of nature. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04861-z.

Masterman:2025:MTM

[3136] Christopher James Masterman. Modality, truth and mere picture thinking. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04840-4.

Zong:2025:TCL

[3137] Ning Zong. Transformative creativity in life: an epigenetic perspective based on biological autonomy. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04908-1.

Podosky:2025:WAS

[3138] Paul-Mikhail Catapang Podosky. What attitude should we take to conceptual engineering? Synthese, 205(1):??, January 2025. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-024-04885-5.

Crawford:2025:WB

[3139] Lindsay Crawford. Wronging in believing. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04855-x.

Benschop:2025:LPA

[3140] Cato Benschop and Annemarie Kalis. The life project account of eating disorders: agency in the pursuit of dietary goals. Synthese, 205 (1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04894-4.

Koch:2025:MVA

[3141] Steffen Koch. Merely verbal agreement, speaker-meaning, and defective context. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04888-2.

Entwistle:2025:GEM

[3142] Megan Entwistle. Generic episodic memories. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04869-5.

Geddes:2025:RPA

[3143] Alexander Geddes. Resolving a puzzle about moral responsibility and logical truth. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04838-y.

Vigiani:2025:REL

[3144] Pietro Vigiani. Relevant epistemic logic with state-sensitive topics. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04863-x.

Rehren:2025:PBM

[3145] Paul Rehren and Till Armbruster. Has philosophy become more ‘scientific’? A citation analysis. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04896-2.

Shan:2025:DPU

[3146] Yafeng Shan. The debate over proximate and ultimate causation in biology. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04890-8.

Barker:2025:EHE

[3147] Simon Barker. Epistemic health, epistemic self-trust, and bipolar disorder: a case study. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04867-7.

Pignocchi:2025:BBD

[3148] Tomaso Pignocchi. Beyond belief: deep disagreement and conversion in Wittgenstein’s philosophy. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04857-9.

Zhang:2025:YWL

[3149] Jiahe Zed Zhang. “You will live to regret this!”: Transformative choices and predicted regret. Synthese, 205(1):??, January 2025. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-024-04892-6.

Moore:2025:DIR

[3150] Graham Seth Moore. Deflationary and inflationary roles of truth. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04889-1.

Bykvist:2025:PME

[3151] Krister Bykvist and Joe Roussos. A plea for modelling in ethics. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04864-w.

Kohler:2025:GCM

[3152] Sebastian Köhler. Good classification matters: conceptual engineering in data science. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04850-2.

Jespersen:2025:LPD

[3153] Bjørn Jespersen. Lexical predicates do substitute in fine-grained attitudes. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04810-w.

Berghofer:2025:RBE

[3154] Philipp Berghofer. On the relationship between epistemology and science: synergies between experience-first epistemologies and agentcentered interpretations of quantum mechanics. Synthese, 205(1):??, January 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04875-7.

Cuffari:2025:IIE

[3155] Elena C. Cuffari and Nara M. Figueiredo. Intentions in interactions: an enactive reply to expressive communication proposals. Synthese, 205 (2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04836-0.

Britten-Neish:2025:COC

[3156] George Britten-Neish. Cognitive offloading and the causal structure of human action. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04887-3.

Siegel:2025:CBP

[3157] Gabriel Siegel. Can basic perceptual features be learned? Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04906-3.

Pritchard:2025:AHC

[3158] Duncan Pritchard. Axiological hinge commitments. Synthese, 205 (2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04898-0.

McKee:2025:DLM

[3159] Patrick McKee. Dualism leads to many minds. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04868-6.

Berg:2025:PSD

[3160] Ásgeir Berg. Problems for ‘standard’ dispositionalist accounts of semantic content. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04886-4.

Rijssenbeek:2025:SBS

[3161] Julia Rijssenbeek. Synthetic biology: supporting an anti-reductionist view of life. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-04913-y.

Garcia:2025:VNM

[3162] Enara García and Xabier E. Barandiaran. Varieties of normativity and mental health: an enactive approach. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04854-y.

Giraldo:2025:WEC

[3163] Andrés Giraldo. What is the experimentalist challenge to the method of cases? Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04877-5.

Ellefson:2025:AAI

[3164] Gretchen Ellefson. Authority, accommodation, and illocutionary success. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-04912-z.

Vacek:2025:HA

[3165] Daniela Vacek and Martin Vacek. The hyperintensionality of art. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04909-0.

Cangiotti:2025:AMV

[3166] Nicolò Cangiotti, Gianni Arioli, and Giovanni Valente. Approximations that matter: virtual particles as carriers of interactions. Synthese, 205 (2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04881-9.

Wee:2025:CFD

[3167] Michael Wee. Concept-formation and deep disagreements in theoretical and practical reasoning. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04884-6.

Beyermann:2025:BPS

[3168] Jasper Beyermann. Bending pharmaceutical science: epistemic diversity and regulation. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04910-7.

Vergouwen:2025:SSW

[3169] Sanne Vergouwen and Sebastian De Haro. Supersymmetry in the Seiberg–Witten theory: a window into quantum field theory. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04882-8.

Placek:2025:CTS

[3170] Tomasz Placek. Correction to: On testimony in scenarios with Wigner and Friend. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04849-9. See [3047].

Bourrat:2025:RRD

[3171] Pierrick Bourrat. Reproducees, reproducers, and Darwinian individuals. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04880-w.

Krueger:2025:HMA

[3172] Joel Krueger. Home as mind: AI extenders and affective ecologies in dementia care. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-04923-w.

Sakakibara:2025:OII

[3173] Eisuke Sakakibara. The oversight of implicature and implicational injustice in doctor–patient communication. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04865-9.

Nielsen:2025:DDH

[3174] Thor Hennelund Nielsen. Disorder or distress? The hermeneutical injustices of overdiagnosis within psychiatry. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04902-7.

Juvshik:2025:WMK

[3175] Tim Juvshik. What makes a kind an artifact kind? Synthese, 205 (2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-04914-x.

Ventura:2025:MQE

[3176] Rafael Ventura. Measuring the quality of experimental research. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04905-4.

Razavi:2025:WAT

[3177] Sepehr Razavi. What is active touch? Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-04918-7.

Bolinska:2025:EIM

[3178] Agnes Bolinska and Andrej Sali. Effective integration and models of information: lessons from integrative structure modeling. Synthese, 205 (2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-04911-0.

Roloff:2025:IRA

[3179] Jakob Roloff. The indispensability of relational, adapted, and derived proper functions. Synthese, 205(2):??, February 2025. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-025-04921-y.

Manhire:2025:TDA

[3180] Ryan Manhire and Camilla Kronqvist. “Tap-dancing around the conversation”: difficulties in an intimate deep moral disagreement. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-04920-z.

Lemanek:2025:IAL

[3181] Kamil Lemanek. An inferentialist account of lying. Synthese, 205 (2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-04932-9.

Roberts:2025:SCP

[3182] Tom Roberts and Sam Wilkinson. Status and constitution in psychiatric classification. Synthese, 205(2):??, February 2025. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04837-z.

Burnston:2025:ERC

[3183] Daniel C. Burnston. Epistemic reduction of the concept of ‘decision’. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-04917-8.

Lisciandra:2025:ENI

[3184] Chiara Lisciandra. Explanatory norms and interdisciplinary research. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04900-9.

Finnegan:2025:NNN

[3185] Colum Finnegan and Alexandra S. Ilieva. Naturalizing normativity: neopragmatist reflections on mental disorder. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04858-8.

Sul:2025:SSD

[3186] Jae Ryeong Sul. The structure of silence in depression. Synthese, 205 (2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-04933-8.

Kodaj:2025:TTF

[3187] Daniel Kodaj, László Bernáth, and Martin Pickup. Teleology for the twenty-first century: editorial introduction. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-04926-7.

Sapalski:2025:QRA

[3188] Szymon Sapalski. A quasi-realist approach to rules in inferentialism. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-04934-7.

Ortmann:2025:OOE

[3189] Jakob Ortmann. Of opaque oracles: epistemic dependence on AI in science poses no novel problems for social epistemology. Synthese, 205 (2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-04930-x.

White:2025:SBE

[3190] Ben White, Andy Clark, Avel Guènin-Carlut, Axel Constant, and Laura Desirée Di Paolo. Shifting boundaries, extended minds: ambient technology and extended allostatic control. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-04924-9.

Panjwani:2025:NSL

[3191] Raja Panjwani. A note on the semantics of linear regression. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04904-5.

Lee:2025:PEN

[3192] Wooram Lee. Pointless epistemic norms. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-04940-9.

Eyre:2025:CU

[3193] John Eyre. Corroboration and uncertainty. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04835-1.

White:2025:PIM

[3194] Ben White. The phenomenal intentionality of mental imagery and seeingas. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-04938-3.

Brun:2025:REC

[3195] Georg Brun, Gregor Betz, and Claus Beisbart. Reflective equilibrium: conception, formalization, application-introduction to the topical collection. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-04927-6.

Graham:2025:PEA

[3196] Peter J. Graham and Nikolaj Jang Lee Linding Pedersen. The phenomenal evidence argument. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04856-w.

Akiba:2025:FAT

[3197] Takeshi Akiba. A functionalist account of truth-grounding: refining Lynch’s view. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04879-3.

Picollo:2025:AGC

[3198] Lavinia Picollo. Against generalized comprehension. Synthese, 205 (2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04797-4.

Baron:2025:PME

[3199] Sam Baron and Jessica Pohlmann. Priority monism and the emergence of spacetime. Synthese, 205(2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04859-7.

Cocco:2025:BBE

[3200] Lorenzo Cocco and Joshua Babic. Burgess and the bucket: the emergence of spacetime in classical theories of gravitation. Synthese, 205 (2):??, February 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04823-5.

Cinti:2025:BSS

[3201] Enrico Cinti and Marco Sanchioni. Bridging supersymmetry and the spin-statistics theorem: a quest for emergence. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04895-3.

Konigs:2025:BHO

[3202] Peter Königs. Bracketing higher-order evidence in scholarly philosophical argumentation: why and which? Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-04935-6.

Irikefe:2025:CCP

[3203] Paul O. Irikefe. The calibration challenge to philosophical intuitions. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-024-04903-6.

Schwartz:2025:MCW

[3204] Arieh Schwartz and Nir Fresco. Mapping content: why cognitive maps are non-conceptual mental states. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-04922-x.

Pils:2025:SRD

[3205] Raimund Pils. The scientific realism debate and consensus reporting. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-025-04916-9.

Soria-Ruiz:2025:KHW

[3206] Andrés Soria-Ruiz. I know how to withstand the skeptic. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04873-9.

Hartzell:2025:RR

[3207] Robert Hartzell. Representing relevance. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-04943-6.

Heap:2025:MHP

[3208] Shaun Hargreaves Heap and Mehmet S. Ismail. Mill’s harm principle, rationality, and Pareto optimality in games. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04805-7.

Drozdzowicz:2025:CPW

[3209] Anna Drozdzowicz. Conversational pressures at work: professional roles and communication in mental healthcare settings. Synthese, 205 (3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04839-x.

Krupnik:2025:BMH

[3210] Valery Krupnik. Bayes meets Hegel: the dialectics of belief space and the active inference of suffering. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-04944-5.

Williams:2025:HRA

[3211] Danielle J. Williams. How to be a realist about computational neuroscience. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-04952-5.

Hermann:2025:MCD

[3212] Julia Hermann. Moral certainty, deep disagreement, and disruption. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-025-04948-1.

Lasonen-Aarnio:2025:KCD

[3213] Maria Lasonen-Aarnio. Knowledge-conducive dispositions: how to do consequentialist epistemology. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-024-04897-1.

Evans:2025:BBS

[3214] Andrew Evans. Beyond biological and social normativity: varieties of norm deviation and the justification for intervention. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-04945-4.

Clatterbuck:2025:LIC

[3215] Hayley Clatterbuck and Hunter Gentry. Learning incommensurate concepts. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04883-7.

Odrowaz-Sypniewska:2025:IPL

[3216] Joanna Odrowaz-Sypniewska. Inferential potentials as lexical meanings. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-025-04942-7.

McDonald:2025:CMC

[3217] Jennifer McDonald. Causal models and causal relativism. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04893-5.

Toon:2025:WNB

[3218] Adam Toon. We have never been Cartesian. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-024-04891-7.

Filipovic:2025:MFE

[3219] Nenad Filipović. Material foundations of ecological rationality. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-04939-2.

Hueck:2025:CVP

[3220] Christoph J. Hueck. Cognizing the vital principle of the organism by interpreting the four Aristotelian causes in a Kantian perspective: an outline. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-024-04860-0.

Voggenauer:2025:EOP

[3221] Martin Voggenauer. Explaining objective probabilities by physical symmetries. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-04951-6.

Elgin:2025:HTI

[3222] Samuel Z. Elgin. Haeccities and the triviality of the identity of indiscernibles. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04851-1.

Openshaw:2025:RRE

[3223] James Openshaw, Kourken Michaelian, and Denis Perrin. Reference and remembering: editorial introduction. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-04956-1.

Franzen:2025:UAS

[3224] Nils Franzén and Andreas Stokke. A unified account of semantic and pragmatic infelicity. Synthese, 205(3):??, March 2025. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-025-04954-3.

Calosi:2025:GQ

[3225] Claudio Calosi and Robert Michels. Graded qualities. Synthese, 205 (3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04848-w.

Greve:2025:MME

[3226] Sebastian Sunday Grève. Method and morality: Elenchus from Socrates to Wittgenstein. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-04919-6.

Sganzerla:2025:BEM

[3227] Lorena Sganzerla, Daniel D. Hutto, and Michael D. Kirchhoff. Beyond the extended mind: new arguments for extensive enactivism. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-04931-w.

Zakkou:2025:ECI

[3228] Julia Zakkou. Entailed conversational implicatures. Synthese, 205 (3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-024-04866-8.

Quinn:2025:CSV

[3229] Aleta Quinn. Community science and the value-free ideal. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-04955-2.

Xie:2025:DLB

[3230] Kaibo Xie and Qingyu He. Doxastic logic based on causal structures. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-025-04946-3.

Dardashti:2025:LCS

[3231] Radin Dardashti, Enno Fischer, and Robert Harlander. The life cycle of scientific principles-a template for characterizing physical principles. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-04972-1.

Ceragioli:2025:ISC

[3232] Leonardo Ceragioli. Inferential semantic contamination, harmony and realist pollution. Synthese, 205(3):??, March 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-024-04871-x.

Silins:2025:BBB

[3233] Nicholas Silins. Blindsight, blandsight, and blingsight: unconscious perception, attention, and the epistemology of perception. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05047-x.

Michelini:2025:BPS

[3234] Matteo Michelini, Eugenia Polizzi, and Daniele Vilone. Bad polarization in structurally cohesive communities. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-04977-w.

Beck:2025:WLC

[3235] Ori Beck and Assaf Weksler. Weber’s law as a challenge to inheritance. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05076-6.

Vredenburgh:2025:FRD

[3236] Kate Vredenburgh. Fairness and randomness in decision-making: the case of decision thresholds. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-05091-7.

Becattini:2025:ASE

[3237] Valeria Becattini and Liberty Severs. Are ambient smart environments sufficient for developing good habits? Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05060-0.

Arroyo:2025:NMP

[3238] Raoni Arroyo and Matteo Morganti. Naturalistic metaphysics and the parity thesis: why scientific realism doesn’t lead to realism about metaphysics. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05075-7.

Chandrasekharan:2025:MCM

[3239] Sanjay Chandrasekharan, Ravi Sinha, and Geetanjali Date. Making cognition: a mechanism account of the way humans develop the ability to build and alter their environments. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05081-9.

Negro:2025:TAA

[3240] Niccolò Negro. Three arguments against metaphysical structuralism in consciousness research. Synthese, 206(1):??, July 2025. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-025-05103-6.

Yoshida:2025:GWE

[3241] Yoshinari Yoshida. Generalizing while embracing differences: configurations of representations and cross-fertilization. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05083-7.

Franda:2025:WAC

[3242] Francesco Franda and Pauline Souman. Why art cannot (technically) malfunction. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05086-4.

Bentley:2025:FSC

[3243] Vanessa A. Bentley. A feminist standpoint for cognitive neuroscience. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-025-05106-3.

Greco:2025:RC

[3244] John Greco. Replies to critics. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-05082-8.

Maiese:2025:SEB

[3245] Michelle Maiese. Self-equilibration, borderline personality disorder, and dance-movement therapy. Synthese, 206(1):??, July 2025. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-025-05122-3.

Drayson:2025:RRL

[3246] Zoe Drayson. Rethinking the role of language in arguments for extended cognition. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05079-3.

Howard:2025:NFE

[3247] Stephen Howard. Noumenal forces? Essences, things in themselves, and dispositions in Kant. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05080-w.

Saad:2025:PFT

[3248] Bradford Saad. Planetary fine-tuning, cosmological fine-tuning, and the multiverse. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05071-x.

Sims:2025:TTS

[3249] Ric Sims. Traces of thinking: a stigmergic approach to 4E cognition. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05074-8.

Franks:2025:DSU

[3250] Curtis Franks. Disjunctive syllogism: the universal characterization of multiplicative “or”. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-04966-z.

Lee:2025:CN

[3251] Seungtaek Lee. Complex names. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-05100-9.

Hui:2025:HDD

[3252] Liyun Hui. How does the defense of closure fail again? Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05087-3.

Haber:2025:PME

[3253] Matthew H. Haber. Positively misleading errors. Synthese, 206(1): ??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05093-5.

Gomes:2025:RST

[3254] Henrique de A. Gomes. Representational schemes for theories with symmetry. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05045-z.

Camillo:2025:SAA

[3255] José Carlos Camillo. Success atomism: the accuracy of episodic memory and its relevant elements. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-05129-w.

Benito-Monsalvo:2025:BMM

[3256] Carlos Benito-Monsalvo. Beyond minimality: mixed inferences as an anti-collapse problem for standard combination mechanisms. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05051-1.

Cuesta:2025:QLC

[3257] José Alejandro Fernández Cuesta, Michele Piazzai, and Umberto Rivieccio. Quantum logics in cognition: a proposal. Synthese, 206(1): ??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05090-8.

Brown:2025:MIM

[3258] James L. D. Brown. Moral inferentialism and moral psychology. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05102-7.

Spagnesi:2025:TUN

[3259] Lorenzo Spagnesi. Truth, understanding, and normativity in scientific models. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05110-7.

Lukowski:2025:FCD

[3260] Piotr Lukowski. Formalizing the conditionals of Diodorus Cronus and Chrysippus: logic and potential applications in human thinking. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-04981-0.

Schettino:2025:IRS

[3261] Giulia Schettino and Jacopo Rosino Giraldo. Investigating the role of simulations in astrophysics: the case of the orbit determination problem. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05096-2.

Torres:2025:EH

[3262] Émile P. Torres. On the extinction of humanity. Synthese, 206(1): ??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05094-4.

Puy:2025:ADA

[3263] Nemesio G. C. Puy. The artifactual duality of architecture and its ontological implications for legal protection. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05124-1.

Rizza:2025:MTC

[3264] Davide Rizza. Measurement theory in the context of scientific enquiry. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-025-05109-0.

Elder-Vass:2025:SO

[3265] Dave Elder-Vass. Segmenting ontology. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-025-05119-
y.

Boucher:2025:FSB

[3266] Sandy C. Boucher. Functionalism and structuralism in biology: stances or explanatory strategies? Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-05128-x.

Castro:2025:SVN

[3267] Víctor Fernández Castro and Eduardo Pérez-Navarro. The semantics and virtues of nondescriptivist normativism about mental disorder. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05133-0.

Papandreou:2025:ATA

[3268] Marilù Papandreou. Artworks, technical artefacts, and a few other things: a fresh Aristotelian approach. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-025-05113-
4.

Liu:2025:WCU

[3269] Chang Liu. Why conventions are underestimated by intentionalism of illocutionary acts. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05141-0.

Fang:2025:SLC

[3270] Jingzhi Fang and Jiji Zhang. On Stalnakerian and Lewisian causal models. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05092-6.

Pailos:2025:NVN

[3271] Federico Pailos. Naïve non-substructural solutions to the Validity Paradox. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05097-1.

Lee:2025:CEG

[3272] Kok Yong Lee. A critical examination of Golemon and Graber’s deductive No-Miracle Argument. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-05144-x.

Li:2025:PHT

[3273] Yucheng Li. Plenitude and the Humean thesis. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05131-2.

Lauman-Lairson:2025:MRA

[3274] Jessica Lauman-Lairson. A middle-range account of analogical inference: Target Domain Salience. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-05136-x.

Donohue:2025:SCP

[3275] J. L. A. Donohue. The Scylla and Charybdis of paternalism and moral complicity. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05145-w.

Fu:2025:BLT

[3276] Lyu Fu. Beauty leads to truth: aesthetic induction on consistency. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05137-w.

Hong:2025:FK

[3277] Frank Hong. Fanaticism and knowledge. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-025-05114-

Matarese:2025:EVA

[3278] Vera Matarese. On the epistemic value of atmospheric retrieval models in exoplanetary science. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05149-6.

Greenstein:2025:DDM

[3279] Noah Greenstein. Disagreement, dogmatism, moral encroachment & Wittgensteinian etiquette. Synthese, 206(1):??, July 2025. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-025-05151-y.

Miller:2025:FWE

[3280] Kristie Miller and Challis Chair. Free will expressivism. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05152-x.

Donskov:2025:RCN

[3281] Jakob Donskov. Resisting counterevidence: no knowledge needed. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05155-8.

Englander:2025:NA

[3282] Alex Englander. Naivety and assurance. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-025-05143-
y.

Miragoli:2025:EIC

[3283] Martin Miragoli. The environmental image: the case of white ignorance for epistemic justice. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05135-y.

Malcolm:2025:FEC

[3284] Finlay Malcolm. Faith and epistemic coherence. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05117-0.

DeRizzo:2025:MP

[3285] Julio De Rizzo. Modal perspectivalism. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-025-05146-
9.

Wang:2025:ENA

[3286] Cuizhu Wang, Maciej Juzaszek, and Shane Ryan. Epistemic nudging for accurate normative expectations. Synthese, 206(1):??, July 2025. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-025-05127-
y.

Fatollahi:2025:DBI

[3287] Alireza Fatollahi and Kasra Alishahi. In defense of Bayesian Information Criterion. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05085-5.

Garcia-Arnaldos:2025:INA

[3288] Dolores García-Arnaldos and Concha Martínez-Vidal. Inferentialism naturalized and anti-exceptionalism about logic. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05104-5.

Fulda:2025:NBA

[3289] Fermín C. Fulda. Naturalizing biological agency: constitutive and dynamical strategies. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05132-1.

Marre:2025:BFG

[3290] Thomas Marré. Biological functions, generics, and explanation. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05126-z.

Thong:2025:SR

[3291] Joshua Thong. Not staying regular? Synthese, 206(1):??, July 2025. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-025-05115-

Rinner:2025:BVS

[3292] Stefan Rinner. Brains in vats and semantic externalism: a new response to the skeptical argument. Synthese, 206(1):??, July 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-05107-2.

Ryan:2025:BPP

[3293] Patrick J. Ryan. Bertrand’s Paradox is a Paradox of Infinity. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05140-1.

Meyer:2025:ROS

[3294] Aiden Meyer. Radical ontic structural realism as a structuralist existence monism. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05120-5.

LaCroix:2025:WDP

[3295] Travis LaCroix, Alexis Amero, and Benjamin Sidloski. What do philosophers talk about when they talk about autism? Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05116-1.

Russell:2025:PSS

[3296] Jodie Louise Russell. Prescriptive ‘selves’ and self-illness ambiguity. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05147-8.

Brandt:2025:RVI

[3297] Stefan Brandt. Ryle vs. intellectualism: a case study in philosophical method. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05112-5.

Bickle:2025:WKS

[3298] John Bickle. What kind of a scientific revolution was the discovery of mammalian adult neurogenesis? Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05138-9.

Kortenbach:2025:AGV

[3299] Bas Kortenbach. Appreciating global validity. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05089-1.

Duerr:2025:RWP

[3300] Patrick M. Duerr and Enno Fischer. Rationally warranted promise: the virtue-economic account of pursuit-worthiness. Synthese, 206(2): ??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-025-05077-5.

Raskoff:2025:TFM

[3301] Sarah Zoe Raskoff. The transparency fallacy in metaethics. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05162-9.

Blum:2025:PC

[3302] Alexander S. Blum and James D. Fraser. Perturbative causality. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-025-05099-z.

Belastegui:2025:CSC

[3303] Javier Belastegui. A conceptual space for classical concepts. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05148-7.

Corry:2025:FTA

[3304] Richard Corry. Finding temporal asymmetry within the many, many worlds of Everettian quantum mechanics. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05166-5.

McCullagh:2025:ICA

[3305] Mark McCullagh. Inferentialism and collaged assertions. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05167-4.

Overmark:2025:CPR

[3306] Rasmus Overmark. Competence and performance revised: pragmatic development may shape mentalizing development. Synthese, 206(2): ??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL https://link.springer.com/article/10.
1007/s11229-025-05171-8.

Heylen:2025:DAK

[3307] Jan Heylen. The dynamic approach to knowability. Synthese, 206 (2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05159-4.

Gyenis:2025:ESP

[3308] Balázs Gyenis. Empirical structure physicalism and realism, Hempel’s dilemma, and an optimistic meta-induction. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/

s11229-025-05160-x. See publisher correction [3329].

Wieczorkowski:2025:CDD

[3309] Michał Wieczorkowski. Context-dependence debate from an inferentialist perspective. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05156-7.

Jarczewski:2025:VET

[3310] Dominik Jarczewski. Virtue environmentalism: towards a robustly social virtue epistemology. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-05168-3.

Vincent:2025:RIT

[3311] Michael Vincent and Mark Hooper. Research integrity, thick and thin. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05157-6.

Frketich:2025:KPC

[3312] Elise Frketich. Kant on particular chemical ideas. Synthese, 206 (2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05134-z.

Lipshitz:2025:JGR

[3313] Nethanel Lipshitz and Tammy Harel Ben Shahar. Justice for grasshoppers: reassessing the effort principle in distributive justice. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05108-1.

Mangraviti:2025:FLB

[3314] Franci Mangraviti. Feminist logic, broadly speaking. Synthese, 206 (2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05176-3.

Fazio:2025:CEJ

[3315] Davide Fazio and Raffaele Mascella. Considerations on Everett J. Nelson’s connexive logic. Synthese, 206(2):??, August 2025. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-025-05121-4.

Choi:2025:PPT

[3316] Seungrak Choi. Problems of a proof-theoretic characterization of paradoxes. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05150-z.

Otsuka:2025:MCP

[3317] Jun Otsuka, Tomoyuki Hayashi, and Hayato Saigo. Modeling causal processes. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05084-6.

Allori:2025:MMT

[3318] Valia Allori. A minimal metaphysics of thin-objects and nomic structure. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05154-9.

Hohwy:2025:MPP

[3319] Jakob Hohwy. A metaphysics for predictive processing. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05169-2.

Marchiori:2025:CAH

[3320] Samuela Marchiori. Conceptual affordances: (how) should they inform conceptual engineering? Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-05161-w.

Kohler:2025:PFC

[3321] Sebastian Köhler and Herman Veluwenkamp. Proper foundations for conceptual ethics. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05175-4.

Gabriel:2025:ILE

[3322] Rami Gabriel. Inter-level explanation: the case of neuropsychoanalysis. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05158-5.

Zhou:2025:IRP

[3323] Peng Zhou. Indicating without revealing: a posteriori physicalism and the new challenge. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05163-8.

Witte:2025:ULO

[3324] Margot Witte. Using and losing our concepts. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05173-6.

Azhar:2025:SSI

[3325] Feraz Azhar and Mohammad Hossein Namjoo. Spacetime singularities and indeterminism. Synthese, 206(2):??, August 2025. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-025-05178-1.

He:2025:HOA

[3326] Jiacheng He. Higher-order accounts of agnosticism: belief vs. endorsement. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05179-0.

Snodgrass:2025:CIP

[3327] J. J. Snodgrass. Co-intensional properties and running afoul of Ockham’s Razor. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05153-w.

Ivy:2025:TGS

[3328] Spencer Ivy and Aleksandra Mroczko-Wasowicz. Three grades of subject-dependency in object perception. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05182-5.

Gyenis:2025:PCE

[3329] Balázs Gyenis. Publisher correction to: Empirical structure physicalism and realism, Hempel’s dilemma, and an optimistic meta-induction. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05195-0. See [3308].

Sushchin:2025:ICT

[3330] Mikhail Sushchin. On the idea of complexes of theories in the cognitive sciences. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05172-7.

Trueman:2025:MM

[3331] Robert Trueman and Simon Thunder. Manyism as mereologicism. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-025-05123-2.

vanderLinden:2025:SDF

[3332] R. R. van der Linden, M. H. N. Schermer, and N. R. Binney. From symptom to disease: a Fleckian comparative epistemology of the conceptualization of chronic pain. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-05180-7.

Li:2025:RUU

[3333] Dazhu Li, Sujata Ghosh, and Fenrong Liu. Reasoning under uncertainty in the game of Cops and Robbers. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05165-6.

Sneddon:2025:DAR

[3334] Andrew Sneddon. Doing and allowing, removing barriers: the problem of shared barriers. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05170-9.

Pruss:2025:IGC

[3335] Alexander R. Pruss. The impossibility of generating comparative probabilities from primitive conditional probabilities. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05177-2.

Adlam:2025:ROQ

[3336] Emily Adlam. Relational observables, quiddities, and structural realism. Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05190-5.

McKay:2025:WSE

[3337] Noah McKay. In what sense is evolution blind to moral truth? Synthese, 206(2):??, August 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05188-z.

Basu:2025:BP

[3338] Rima Basu. Bullshit philosophy. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05198-x.

Huneman:2025:WKE

[3339] Philippe Huneman. What kind of epistemology for network analyses? Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05105-4.

Kilinc:2025:SJT

[3340] Berna Kilinc. A skewed jury theorem: more theorems in search of the truth. Synthese, 206(3):??, September 2025. CODEN SYNTAE.

ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05125-0.

Morin-Martel:2025:VET

[3341] Alexis Morin-Martel. The value of epistemic trust in public experts. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05212-2.

Beauchemin:2025:MMU

[3342] Pierre-Hugues Beauchemin and Kent W. Staley. Making measurement useful: integrating measurement, uncertainty, and sensitivity. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05187-0.

Rubin:2025:PDI

[3343] Mark Rubin. Preregistration does not improve the transparent evaluation of severity in Popper’s philosophy of science or when deviations are allowed. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05191-4.

Hu:2025:GP

[3344] Xingming Hu. Grasping a proposition. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05224-y.

Dutta:2025:GG

[3345] Rajeev R. Dutta. Genetics and generics. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05230-0.

Du:2025:RBG

[3346] Shaokai Du. The relation between group epistemic structure and operative members in epistemic group. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05227-9.

List:2025:CAS

[3347] Christian List. Can AI systems have free will? Synthese, 206(3): ??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05209-x.

Freeborn:2025:CUS

[3348] David Peter Wallis Freeborn. Compositional understanding in signaling games. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05184-3.

Zeller:2025:AMF

[3349] Kaj André Zeller. Against metaphysics by fiat. Synthese, 206(3): ??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05192-3.

Reydon:2025:SSA

[3350] Thomas A. C. Reydon. Synchronic species aren’t diachronic species: why 3D/4D metaphysical equivalence fails for the case of biological species. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05206-0.

Schultz:2025:NBP

[3351] Michael Schultz. The natural-born posthuman: applying extended mind to post- and transhumanist discourse. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05202-4.

Milkowski:2025:TNP

[3352] Marcin Miłkowski and Przemysław Nowakowski. Tracking norms: a plea for a normative digital philosophy of science. Synthese, 206(3): ??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05200-6.

Butler:2025:HRM

[3353] Andrew P. Butler. Husserl and the reduction of modality to essence. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05205-1.

Duetz:2025:CCT

[3354] J. C. M. Duetz. Contrarianism, conspiracy theories, and epistemic forgeries. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05203-3.

DeDomenico:2025:OPS

[3355] Andrea De Domenico, Ali Farjami, and Xiaolong Wang. Obligations and permissions on selfextensional logics. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05207-z.

Rice:2025:EAT

[3356] Collin Rice. The experiment-applicability tradeoff in scientific model building. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05215-z.

Chomanski:2025:SRE

[3357] Bartlomiej Chomanski. Speculative risks of effectively combating misinformation: echo chambers and self-censorship. Synthese, 206(3): ??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05217-x.

Morrison:2025:MRK

[3358] Thomas Patrick Morrison. Minds and real kinds. Synthese, 206(3): ??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05221-1.

Felka:2025:BMA

[3359] Katharina Felka. Bad manners: Against a conversational implicature theory of slurs. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05118-z.

Kuan:2025:UAN

[3360] Ko-Hung Kuan. Updating by averaging: a network perspective. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-025-05225-x.

Lan:2025:BRA

[3361] Ying Lan. Belief, responsibility, and authenticity: reframing Heidegger’s existential ethics. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-05211-3.

Bangu:2025:WCP

[3362] Sorin Bangu and Jeffrey Schatz. Wittgenstein on Cantor’s Proof. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-025-05238-6.

Schulte:2025:MVM

[3363] Peter Schulte. Millikan vs. Millikan: biosemantics and the role of the consumer. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05241-x.

Malavolta:2025:EPU

[3364] Bruno Malavolta, Eros Moreira de Carvalho, and Rodrigo Sabadin Ferreira. Ecological perspectivism: understanding perspectival realism through ecological psychology. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05222-0.

Wang:2025:CKM

[3365] Ruoding Wang, Yiwen Ding, and Xiaolong Wang. Causal Kripke models. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05223-z.

Champion:2025:SNR

[3366] Heather Champion. Strong novelty regained: high-impact outcomes of machine learning for science. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05228-8.

Pillin:2025:FJ

[3367] Hannah Pillin. Fragmented justification. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05193-2.

Dandelet:2025:HDD

[3368] Sophia Dandelet. How do desires explain motivated reasoning? Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05214-0.

Ramirez-Vizcaya:2025:EEC

[3369] Susana Ramírez-Vizcaya. Extending the enactive concept of habit: from sensorimotor schemes to regional identities. Synthese, 206(3): ??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05237-7.

Song:2025:VAM

[3370] Ziming Song. Value-aligned but misguided: a dilemma in AI and AGI decision making. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05232-y.

Soles:2025:IVK

[3371] Donald E. Soles III. Inquiry & the value of knowledge. Synthese, 206 (3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05240-y.

Cheng:2025:ITA

[3372] Yong Cheng. Isaacson’s thesis on arithmetical truth. Synthese, 206 (3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05229-7.

Gamester:2025:EMV

[3373] Will Gamester. Expressivism and moral vagueness. Synthese, 206 (3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05196-z.

Ostos:2025:MFT

[3374] Elena Ostos. Mind the fungi: towards a philosophy of fungal cognition. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05063-x.

Blonde:2025:AGI

[3375] Ward Blondé. Adding genes and interaction to Smolin’s cosmological natural selection. Synthese, 206(3):??, September 2025. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-025-05189-y.

Herbert:2025:TMR

[3376] Jack L. Herbert. Testimonial monitoring and the redundancy challenge. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05186-1.

Ward:2025:WFF

[3377] Shane Ward. When factives fail to be factives. Synthese, 206(3): ??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05181-6.

Andersen:2025:EDE

[3378] Frederik J. Andersen. Engel’s dilemma and the epistemic significance of logical disagreement. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-05204-2.

Segura:2025:ETE

[3379] Diego Arana Segura. Equivalence and theory expansion. Synthese, 206 (3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05236-8.

Carter:2025:REE

[3380] Matilda Carter. Relational equality for extended minds. Synthese, 206 (3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05242-w.

Gill:2025:APL

[3381] H. Brigitte Gill. Agentive phenomenology and loss of control while eating. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05199-w.

Harder:2025:KFM

[3382] Rory Harder. Knowledge-first mentalizing: concept-based or merely state-based? Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05213-1.

Rosenthal:2025:DNA

[3383] David Rosenthal. Dennett, nonhuman animals, and consciousness. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-025-05216-y.

Dahl:2025:KWK

[3384] Niklas Dahl. Knowing who and knowing how: two ways to understand reference. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05208-y.

Vrahimis:2025:WTS

[3385] Andreas Vrahimis. The working of truths: Stebbing’s debate with Schiller. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05201-5.

Kaspers:2025:RR

[3386] Tom Kaspers. Reply to the reviewers. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05248-4.

Eskens:2025:ED

[3387] Floris Diederik Johannes Eskens. On effective dualities. Synthese, 206 (3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05185-2.

Gagnier:2025:VIT

[3388] Helena Gagnier. Value inheritance: the transmission of values through cognitive extenders. Synthese, 206(3):??, September 2025. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-025-05235-9.

Tejedor:2025:MDD

[3389] Chon Tejedor. Moral deep disagreement, self-stultification and forms of life. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05183-4.

Yee:2025:LEM

[3390] Adrian K. Yee. The limits of epidemiological models of misinformation. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05246-6.

Wells:2025:LFK

[3391] Aaron Wells. Laws first: Kant as a nomic essentialist. Synthese, 206 (3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05164-7.

Wallis:2025:REC

[3392] Jacqueline Mae Wallis. Relinquishing experimental control to improve translation. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05219-9.

Zahnoun:2025:WTG

[3393] Farid Zahnoun. What’s truth got to do with it? In defense of relational realism about perception. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-05251-9.

Tolbert:2025:WCI

[3394] Alexander Williams Tolbert. Why causal inference is necessary for algorithmic fairness. Synthese, 206(3):??, September 2025. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://

link.springer.com/article/10.1007/s11229-025-05044-0.

Rosa:2025:TNA

[3395] Luis Rosa. Is there a norm against inquiry into pseudo-questions? Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-025-05250-w.

Colyvan:2025:DCT

[3396] Mark Colyvan, Tim Smartt, and Hannah Tierney. Do conspiracytheory interventions rest on a mistake? Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05257-3.

Jovicevic:2025:SWQ

[3397] Andrej Jovi’cević. A semantics for weak, question-sensitive belief. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-025-05233-x.

Gunnemyr:2025:EDC

[3398] Mattias Gunnemyr. Estimating degrees of causal contribution. Synthese, 206(3):??, September 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05142-z.

Wehofsits:2025:COA

[3399] Anna Wehofsits. Convincing ourselves: accuracy motives and rationalization. Synthese, 206(4):??, October 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05259-1.

Grad:2025:CEB

[3400] Paweł Grad. Conceptual and epistemic bootstrapping of hinge commitments. Synthese, 206(4):??, October 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05273-3.

Balbiani:2025:IEL

[3401] Philippe Balbiani. Intuitionistic epistemic logic with two modal operators. Synthese, 206(4):??, October 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05226-w.

Hlobil:2025:IAM

[3402] Ulf Hlobil. An inferentialist approach to Mates’s Puzzle. Synthese, 206(4):??, October 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05275-1.

Calosi:2025:OSD

[3403] Claudio Calosi and Fabrice Correia. Omnipresence: secular and divine. Synthese, 206(4):??, October 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.

com/article/10.1007/s11229-025-05210-4.

Almagro:2025:WTM

[3404] Manuel Almagro and Neftalí Villanueva. A Wittgensteinian take on moral vs. deep disagreements. Synthese, 206(4):??, October 2025. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

https://link.springer.com/article/10.1007/s11229-025-05265-
3.

Dawid:2025:BMC

[3405] Richard Dawid and James D. Wells. A Bayesian model of credence in low energy supersymmetry. Synthese, 206(4):??, October 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https:

//link.springer.com/article/10.1007/s11229-025-05252-8.

Scarffe:2025:TJP

[3406] Eric Scarffe and Katherine Valde. Toward a jurisprudential philosophy of science: beyond the value-free ideal. Synthese, 206(4):??, October 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05262-6.

McIntyre:2025:CM

[3407] James H. McIntyre. Conceptual misinformation. Synthese, 206(4):??, October 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05274-2.

Butler:2025:HMP

[3408] Andrew P. Butler. Husserl and the metaphysics of perception: a unification story. Synthese, 206(4):??, October 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-025-05254-6.

Kohler:2025:CEE

[3409] Sebastian Köhler and Herman Veluwenkamp. Conceptual ethics and the ethics of implementation: case for divorce? Synthese, 206(4):??, October 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05231-z.

Kazankov:2025:UCF

[3410] Denis Kazankov. Unpacking conceptual function as tool function. Synthese, 206(4):??, October 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.springer.com/

article/10.1007/s11229-025-05263-5.

Ayhan:2025:QFL

[3411] Sara Ayhan. Queer feminist logic and contradictions: Or how logic and feminism can be relevant to each other. Synthese, 206(4):??, October 2025. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964

(electronic). URL https://link.springer.com/article/10.1007/
s11229-025-05220-2.

Leffler:2025:ONH

[3412] Olof Leffler. Outlines of a neo-Humean disjunctivism. Synthese, 206 (4):??, October 2025. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL https://link.springer.com/article/
10.1007/s11229-025-05255-5.